A Complete Bibliography of Publications in the Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences

halieutiques et aquatiques (2010–2019)

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
      WWW URL: http://www.math.utah.edu/~beebe/

31 August 2024 Version 1.08

Title word cross-reference

27 [376]. ^{+} [1539]. ^{13} [782, 287, 124, 1243, 56, 214, 1834]. ^{14} [962]. ^{15} [124, 1243, 755, 1834]. ^{18} [56, 110]. ^{86} [1712, 1232, 416]. ^{87} [1712, 1232, 416]. _{2} [89, 369, 1225, 1126]. a [70]. ≈ 500 [230]. β [399, 1566]. : [211, 207, 60, 657]. δ [782, 287, 124, 1243, 56, 340, 110, 1834]. F_{MSY} [479].

-ATPase [1539]. -year [376]. -year-old [230].

1 [426]. 100-year [1711]. 18-month [1367]. 1970s [527].

2014 [1071]. 20th [1505]. 21st [1505]. 2D [1233].

3D [1233].

51^{o} [1339]. 57^{o} [1339]. 57K [1526].

abalone [69, 1018, 40]. abbreviated [952]. ABC [1333]. ability [259, 735]. Abiotic [625, 846, 1689, 647, 1490, 1341]. ablation [842, 1232]. Absence [1161, 1496, 308, 1780]. Abundance [103, 1019, 882, 1173, 779, 1762, 1081, 334, 244, 1155, 144, 123, 18, 1680, 1480, 1202, 1546, 666, 273, 222, 1727, 1751, 1316, 1696, 507, 301, 280, 107, 372, 1299, 500, 450, 1353, 1412, 1067, 1429, 1367, 482, 654, 1706, 921, 1616, 1552, 1742, 1619, 682, 1157, 645, 614, 834, 389, 520, 1792, 174, 757, 862, 1647, 542, 736, 394, 565, 1055, 1395, 1403, 1555, 398, 387, 818, 78, 1383, 933, 1247, 726]. abundant [1108, 1433]. acanthias [792]. acceleration [1174]. acceptable [1333]. accessibility [666]. accessible [925, 1533]. accident [873, 1478]. acclimated [811]. acclimation [999]. acclimatized [1149]. account [1805, 1398]. Accounting [1503, 1457, 1401, 1022, 910, 1514, 315, 1395, 455, 587]. accounts [1462]. accumulation [873]. accuracy [623]. accurate [320, 677]. achieving [1455]. acid [1806, 440, 279, 64, 783, 1172, 1432, 800, 1539]. acid-impacted [783]. acidic [1434]. acidification [228, 235, 411, 1743]. acidified [199]. acids [287]. Acipenser [1464, 1506, 1640, 1731, 233, 1374, 362, 1330, 246, 991, 805, 839, 524, 186, 464, 497, 1632, 1634, 1494, 1769]. Acispenser [547]. Acoustic [1720, 1579, 820, 1111, 855, 779, 1664, 1586, 1681, 1457, 1007, 1603, 668, 301, 280, 1237, 557, 1429, 1262, 1609, 693, 1190, 1606, 1758, 1635, 1418, 426, 1755, 678]. acoustic-trawl [1429]. Acoustically [1112, 909]. acoustics [1262, 139]. acquisition [361, 541]. across [782, 1108, 1365, 1354, 1155, 1756, 1109, 1643, 149, 1545, 188, 1821, 753, 1167, 384, 1376, 745, 365, 1209, 1823, 1032, 265, 941, 902, 976, 1740, 1283]. Act [936, 585]. action [443, 1577]. actions [942]. Active [1064, 1663, 488, 1734, 1441]. actively [1285]. activities [878, 898, 1109]. activity [1639, 339, 1180, 1174, 668, 1186, 1723, 282, 1099, 641, 1726]. acute [811]. adaptation [652, 665, 1676, 1783]. adaptations [97]. adaption [1484]. adaptive [1639, 853, 1021, 1149, 1373]. added [634, 764]. addition [621, 1242, 214]. additional [1412]. additions [847, 1243]. additive [1059, 669, 848, 792, 828]. Addressing [679, 1776, 1777, 405]. Adfluvial [1458, 1672, 1618]. Adirondack [1806]. Adirondacks [228]. adjacent [314, 1236]. Adjusting [321]. admixture [291]. Adriatic [921]. Adult [1188, 882, 920, 1762, 65, 1311, 1549, 1731, 790, 465, 1603, 468, 1673, 708, 643, 803, 1454, 1353, 635, 1563, 15, 559, 310, 383, 811, 626, 330, 1154, 1792, 1097, 1608, 224, 599, 841]. advanced [1825]. Advances [1388, 1387, 1777, 1084]. Advancing [491, 1722, 1728]. advection [1666, 1406]. advice [1086, 1772]. aeglefinus [1459, 80, 632]. aeration [1099]. aestivalis [931, 975, 978]. affect [1827, 1422, 1323, 1317, 1001, 510, 1597, 731, 711, 988, 441, 324, 841]. affected [1762]. affecting [1751, 119, 916, 690]. affects [138, 661, 1035, 593, 1488, 1301, 868]. affinity [1823]. Africa [296]. African [589, 1247, 1402]. after [305, 782, 120, 235, 249, 1427, 308, 245, 1469, 146, 864, 435, 170, 972, 1825, 769, 344, 873]. again [1206]. against [1663, 1601, 616]. Age [1203, 1302, 967, 782, 1338, 728, 1507, 319, 987, 581, 1475, 667, 671, 1717, 263, 576, 75, 1443, 505, 538, 766, 893, 889, 531, 311, 672, 510, 502, 269, 1398, 487, 1824, 851, 664, 349, 392, 917, 1176, 1486, 80, 1534, 1134, 432, 1650, 1702, 704, 267, 433, 1467, 493, 768, 1328, 514, 772, 1189, 101, 1834, 1659, 1769, 1017]. age- [263, 349, 1659]. age-0 [269, 851, 392, 432, 1467, 1328, 1769]. age-2 [1650]. age-based [1398]. age-composition [768]. age-dependent [987, 1134]. age-frequency [917]. age-selection [267]. age-specific [667]. age-structured [576, 889, 311, 672, 664, 1176, 1486, 704, 493, 772, 1189, 101, 1659]. ageing [122, 1701]. agents [485]. aggregate [1647]. aggregated [1407]. aggregating [1007]. aggregation [1723, 183, 834]. aggregations [1733, 318, 383, 1632]. agreement [54]. agreements [54]. agricultural [454, 187, 1375]. aid [336, 1332]. aimed [697]. air [1692, 172, 461]. airborne [194]. al [1140, 686]. alalunga [215]. alarm [790, 1045]. Alaska [37, 1749, 600, 974, 1653, 789, 1775, 1470, 306, 669, 431, 908, 507, 515, 1532, 681, 778, 286, 516, 1304, 434, 181, 382, 1209, 204, 1649, 1009, 976, 815, 101]. Alaskan [1644, 1704, 333, 137, 787, 1357, 984, 1214, 396, 924]. albacares [343]. albacore [961, 215]. Albemarle [1336]. Alberta [454, 949]. albula [896]. aleuticus [871]. alevin [327]. alewife [1803, 1730, 1139, 1703, 975, 108]. algae [1164, 924]. Algal [475, 472]. algorithms [1641]. alien [1800]. aligned [1442]. aligns [89]. alive [1522]. alkylphenols [690]. Allee [183, 1652]. Allele [431]. alleviate [204]. Allier [64]. Allis [971]. allocation [335, 994, 361]. allochronic [1511]. allometric [1374]. allometry [1488, 326]. Allopatric [481]. allow [289]. allowing [1300]. allows [1027]. alone [676]. along [454, 606, 1244, 829, 1621, 67, 1325, 945, 303, 1534, 1368, 529, 70, 1291, 817, 351, 1094, 1366, 1472, 521, 570]. Alopias [699]. Alosa [1827, 77, 765, 92, 975, 108, 978]. alpha [1384]. alpine [298]. alpinus [895, 58, 495, 347, 415, 665, 1190, 811, 1533, 79]. also [1733]. alter [1264, 299, 187, 761]. alteration [906, 1100]. alterations [4, 517, 1149]. altered [1640, 590, 133, 1512, 653]. altering [955]. Alternative [1766, 1362, 1572, 166, 195, 136, 479, 1398, 1754, 1486, 757, 657, 1544, 1651, 412, 1719, 1232, 985]. alters [1213, 1136, 624, 1319, 1481, 1590]. aluminum [452, 249, 1539]. Ambient [920, 545, 208]. ambient-temperature [545]. Amblyraja [1816]. Ameiurus [1149]. America [468, 1499, 458, 1376, 116]. American [687, 1827, 855, 1605, 1582, 1201, 77, 1143, 1638, 263, 1666, 590, 765, 1505, 631, 1592, 92, 39, 696, 916, 691, 1779, 38, 693, 1194, 1610, 694, 1371, 1376, 1497, 1314, 692, 1700, 976, 788, 1625, 688, 1787, 655, 1444]. americanus [695, 1666, 1505, 916, 1779, 693, 1194, 1610, 1371, 1497, 689, 1688, 788, 828]. amid [1792]. amino [287]. among [1195, 1591, 1268, 830, 325, 375, 279, 1735, 1626, 534, 1415, 171, 1238, 895, 417, 1127, 708, 331, 218, 1596, 595, 1022, 38, 124, 485, 921, 1294, 1376, 1533, 50, 94, 607, 98]. among- [171]. among-population [830]. among-reach [595]. among-species [279]. amphibian [31]. amphibians [645]. amphidromous [1325]. amphipod [1436]. amplitude [880]. anadromous [1034, 1365, 77, 223, 584, 604, 1004, 946, 571, 1630, 960, 1483, 1479, 1842, 642, 971, 665, 1190, 1799, 1633, 522, 1154, 236, 341, 1192, 1296, 142, 108]. Anadromy [914, 79, 609, 308, 1842, 751, 827]. anaesthetic [208, 997]. analog [876, 621, 1135, 1243]. analyses [512, 1258, 939, 440, 1290, 562, 362, 938, 1282, 720, 1485, 260, 192, 40]. Analysis [158, 1644, 453, 112, 226, 238, 1278, 1503, 1802, 18, 630, 198, 700, 671, 1392, 1158, 1137, 1720, 1350, 849, 1369, 935, 1172, 119, 460, 67, 328, 1399, 1826, 1563, 1184, 1235, 991, 1053, 800, 104, 711, 988, 1529, 1411, 1308, 1818, 587, 139, 498, 131, 750, 976, 1670, 1698, 1040, 1791, 399, 863, 78, 497, 1232, 1834, 1710, 616, 1566, 1280, 729]. analytical [715]. analyze [12, 1374]. Anarhichas [1416]. ancestry [883]. anchoveta [844]. anchovy [1116, 465, 1313, 1079]. angle [1662]. angler [652, 929, 1716, 1430, 130, 161, 891, 1460, 384]. anglers [929, 753, 1724]. angles [1001]. Angling [1663, 906, 1430, 376, 885, 170, 1840, 1100, 1783, 241]. angling-induced [170]. Anguilla [263, 39, 1684, 1444]. anguillid [1326]. Animal [797, 1678]. animals [402]. Announcement [162, 549, 550, 1087, 1217]. Annual [1128, 854]. annually [197]. anomala [1132]. Anoplopoma [1274]. Antarctic [519, 103, 726]. Antarctica [473, 103]. Anthropogenic [777, 1109, 1620, 1558, 1809, 70]. Anticipated [472]. Anticipating [652]. antipredator [352]. apical [598]. Appalachian [888, 1328]. apparent [1349, 887, 1807, 147, 991]. applicability [1412, 1567]. Application [291, 849, 1045, 1810, 792, 1683, 675, 1272, 987, 878, 127, 876, 465, 1697, 491, 867, 1227, 1788, 1393, 1397, 284, 1176, 254, 131, 1105, 1000, 772, 94, 1247]. applications [520, 1293]. applied [1572, 1717, 1084, 292, 396]. Applying [903, 1577, 506]. Appointment [1581]. apports [469, 1206]. appreciation [518, 698, 1043, 1216, 1410, 1611, 1820]. appreciation/Note [518]. approach [652, 1203, 1338, 1656, 974, 191, 379, 545, 1767, 499, 12, 1664, 217, 1340, 781, 167, 122, 1471, 491, 1369, 551, 455, 1160, 1701, 1075, 994, 858, 1121, 1465, 706, 1537, 971, 1176, 1234, 282, 1488, 969, 1000, 902, 1530]. Approaches [899, 1102, 1104, 1398, 391, 715]. approaching [765, 385]. Approches [1102]. appropriateness [1656]. approximate [668]. approximations [1744]. aquaculture [305, 849, 661, 1727, 1750, 157, 1363]. aquariums [1776]. Aquatic [257, 469, 1621, 110, 1102, 369, 1058, 138, 1343, 452, 456, 494, 1104, 551, 345, 1722, 530, 574, 66, 881, 57, 323, 717, 623, 1225, 1347, 1141, 1064, 402, 936, 1566, 1478]. aquatique [1102]. aquatiques [257]. archaeological [230]. Archipelago [1166, 182, 1065]. archival [635, 1770]. archive [1711]. archived [498]. Arctic [1377, 1081, 600, 1704, 993, 1128, 903, 622, 667, 895, 58, 495, 347, 638, 553, 1637, 415, 960, 1483, 1842, 137, 679, 1266, 1689, 303, 665, 1190, 1633, 811, 596, 1533, 1804, 79, 616]. arcticus [1377, 1704]. Arctocephalus [320]. Arctoscopus [219]. Area [1552, 508, 1378, 1061, 1496, 1397, 1759, 438, 654, 866, 1616, 682, 758, 1595, 309]. area-under-the-curve [438, 654, 1616]. Areas [1204, 545, 1343, 322, 979, 43, 711, 289, 1732]. argus [368]. arid [816]. arid-land [816]. arising [702]. Arizona [1068]. Armstrong [177]. array [1480, 1681, 1457, 1526, 1606]. arrays [1546, 1758, 727]. Arripis [653]. arrival [202, 1045, 1175]. arrow [1027]. arrowtooth [379]. arsenic [154]. art [759]. artedi [4, 517, 1477, 1600]. arthropod [1551]. arthropods [1745]. artificial [36, 861]. asiaeorientalis [1382]. Asian [586, 1157]. ASLO [469]. asper [871]. Asphyxiation [404]. assay [965]. assemblage [558, 167, 411, 37, 1260, 59, 21, 1698]. assemblages [1806, 1109, 1248, 1502, 1594, 221, 890, 717, 776, 603, 1246, 611, 1440, 1829, 426, 794, 1375, 24]. assess [1051, 258, 216, 1223, 766, 706, 1282, 1826, 842, 772, 1307]. assessed [1560, 1492]. Assessing [1110, 211, 445, 443, 881, 139, 1562, 1079, 1834, 499, 1013, 1683, 1255]. Assessment [1388, 281, 700, 419, 1271, 752, 873, 508, 1310, 1387, 1258, 1837, 111, 939, 1085, 1272, 1273, 1394, 898, 375, 125, 961, 1535, 783, 1360, 427, 28, 263, 357, 271, 1014, 129, 1239, 935, 765, 651, 58, 253, 889, 672, 42, 1788, 197, 511, 1185, 1090, 1615, 880, 628, 664, 594, 151, 321, 1725, 1176, 1486, 1162, 289, 1684, 1372, 1407, 1618, 714, 1670, 1403, 1625, 101, 1405, 1455, 1659, 636]. assessments [1404, 388, 292, 1773, 910, 1514, 1218, 768, 1489, 173]. assignment [623]. associated [1549, 223, 584, 1556, 905, 618, 1220, 1542, 1564, 1084, 479, 158, 1573, 408, 1745, 758, 1163, 625, 816]. Association [469, 1549, 711, 1077]. associations [850, 861, 1823]. assumptions [136, 1486, 748]. asymmetric [256, 229]. asymmetrically [1545]. asymptotic [1160]. asynchronous [304]. asynchrony [382]. at-risk [1026]. Athabasca [780, 566]. Atheresthes [379]. Atlantic [1735, 1418, 1768, 256, 229, 1331, 1561, 640, 1762, 1230, 1838, 738, 894, 1358, 440, 1205, 982, 1549, 335, 545, 235, 1339, 1586, 1680, 613, 1502, 836, 452, 64, 961, 259, 648, 783, 1643, 588, 1728, 1249, 909, 171, 1257, 812, 149, 1603, 88, 386, 948, 938, 979, 93, 621, 579, 639, 1620, 553, 199, 1168, 141, 225, 950, 312, 1778, 1528, 670, 352, 218, 415, 869, 510, 11, 1409, 1468, 234, 61, 13, 539, 635, 746, 850, 1133, 1822, 147, 680, 1400, 1434, 1826, 1184, 555, 546, 693, 109, 663, 1617, 1706, 734, 1082, 866]. Atlantic [480, 593, 1652, 1416, 999, 1521, 522, 30, 383, 1534, 811, 9, 260, 626, 1167, 1732, 51, 1301, 1539, 868, 834, 536, 1598, 1276, 569, 641, 784, 19, 1645, 56, 1531, 1595, 526, 297, 1196, 1786, 1003, 1002, 437, 543, 366, 184, 1423, 1192, 134, 773, 1403, 1250, 1270, 525, 346, 1208, 399, 1632, 1634, 1036, 1695, 985, 608, 1224, 467]. atmospheric [298, 1225, 1705]. ATPase [1539]. attempt [1827]. attract [575]. attribute [1298]. attributes [857, 1016]. audax [76, 330]. auratus [1733, 466, 232]. auritus [1819]. austral [103]. Australia [305, 16, 1095, 1406, 1407]. Australian [421, 320, 484]. australis [1380, 289]. Austrovenus [1150]. Author [179]. Authorized [1265]. autocorrelated [951, 1659]. autocorrelative [23]. automata [379]. autonomous [1496, 1261]. auxiliary [1780]. Availability [1301, 1462, 125, 1129, 171, 137, 1035, 466, 1032]. available [520, 521, 570]. avancées [1387]. averaging [76]. avoid [234]. Avoidance [1715, 790, 708]. avoided [552].

back [1206, 1338, 455, 1534]. back-calculated [1534]. back-calculation [1338, 455]. background [1708]. backpack [949]. backscatter [127, 1262]. bacteria [924]. bacterial [364]. baffles [1642]. Bags [465]. Bags-of-Features [465]. balance [612, 28, 201, 772]. balanced [1237]. Balearic [820]. baleen [1077]. Balistes [1538]. ballast [1784, 993, 45, 32]. Baltic [700, 588, 1065, 1318, 664, 594, 1746, 1315, 1568, 441, 1685, 1292]. Banff [298]. Bank [75, 1617, 1569]. Banks [1816]. bans [711]. bar [288]. Barbara [1346]. Barents [1177, 1459, 119]. barotrauma [762]. barramundi [95]. Barrenlands [1377]. Barrier [1406, 1267, 1545, 1548, 504, 1157, 435]. Barriers [871, 256, 229, 1827, 807, 1577, 1675]. base [99, 1806, 721, 830, 817]. based [453, 1102, 1338, 1805, 749, 1101, 1362, 1520, 1572, 728, 1258, 705, 1155, 191, 217, 853, 259, 427, 364, 1728, 1415, 759, 362, 129, 332, 1104, 948, 1369, 1193, 360, 552, 292, 1227, 1120, 1346, 1121, 328, 1398, 706, 283, 664, 82, 1779, 806, 663, 742, 1661, 1463, 1721, 1407, 91, 220, 1514, 1584, 1057, 907, 56, 186, 1531, 533, 747, 1009, 637, 1776, 1313, 410, 1670, 788, 794, 514, 863, 1569, 1125]. basées [1102]. baseline [1737, 1759]. Basin [302, 261, 1632, 1634, 454, 889, 1160, 586, 1518, 583, 1599, 310, 1145, 637, 59, 21, 962, 1346, 1183, 152]. basin-specific [962]. basin-wide [586]. basins [1159]. basis [1195, 1464, 189, 187, 170]. bass [334, 376, 1343, 1344, 1586, 582, 1433, 1336, 1226, 1186, 378, 392, 237, 459, 1718]. batch [304]. batch-spawning [304]. bathymetric [940]. bathymetry [1414]. Bay [1585, 789, 1775, 1717, 1540, 37, 1366, 1572, 618, 530, 1226, 916, 378, 706, 414, 1079, 1651, 262]. Bayes [444]. Bayesian [1656, 912, 1381, 1503, 1767, 903, 136, 1592, 1160, 507, 292, 531, 867, 1525, 69, 328, 1824, 1725, 1682, 1742, 1661, 1308, 433, 377, 1780, 86, 1189, 1046, 309, 276]. bbe [1193, 810]. be [1582, 552, 1760, 80, 758, 1276]. beach [1604]. Beacham [27]. beaches [606]. beaked [1418]. Beamish [374]. bear [1097]. bears [1508]. Beaufort [1128]. Beaver [857, 1726, 618, 1099]. beaver-impounded [618]. beavers [1030]. bed [797]. bedload [62]. beds [986]. been [356, 1719]. before [245, 972]. behavior [652, 1805, 1604, 1311, 1731, 1681, 1374, 1470, 885, 891, 1624, 1285, 1822, 1609, 735, 1529, 805, 1305, 542, 224, 1761]. Behavioral [1050, 796, 822, 1157, 1501]. Behaviour [1007, 1100, 1420, 1762, 613, 909, 1459, 1512, 1748, 1436, 352, 130, 161, 680, 890, 1636, 1460, 1151, 232, 1417, 569, 25, 366, 1608, 678, 1441]. Behaviour-mediated [1100]. Behavioural [562, 1238, 1790, 675, 1343, 1626, 58, 1801, 1764, 232, 1466, 488, 1791]. behaviours [1716, 1220, 1755]. being [951]. Belgian [445]. bench [1784]. bench-scale [1784]. benchmarks [1477]. benefit [23]. benefits [1736, 238, 1245, 90, 1306, 1627]. Benthic [265, 956, 556, 381, 964, 1496, 875, 133, 395, 1737, 1689, 1759, 489, 776, 1610, 204, 1789, 163, 173, 317, 1472]. benthivores [1753]. benzoate [386]. Bergen [1071]. Bering [379, 400, 619, 1701, 301, 280, 311, 672, 548, 1262, 1804, 703, 1601, 1228]. Bertalanffy [2, 300]. best [1728]. beta [1384]. better [456, 1374, 1431, 1424, 917, 290, 1631]. Between [960, 1195, 256, 229, 1302, 1806, 1420, 29, 1557, 8, 5, 73, 242, 436, 334, 1640, 1303, 1587, 856, 481, 1124, 281, 1337, 1771, 28, 418, 1182, 812, 72, 754, 58, 1673, 777, 845, 552, 1708, 1545, 251, 312, 352, 1479, 1186, 1144, 702, 807, 66, 1038, 109, 791, 735, 665, 1215, 1130, 1041, 41, 389, 1417, 1097, 1008, 657, 1032, 904, 1538, 1229, 1589, 543, 25, 1423, 194, 192, 1077, 103, 678, 1834, 1695]. Between-watershed [960]. Beyond [532]. Bias [328, 1613, 122, 1005, 1654, 1813, 405, 321, 559]. biased [1805]. Bidirectional [954]. big [1665, 1352, 228]. bigeye [343, 850]. bigger [478]. bighead [1267, 1715]. Bight [1664, 1061]. billfish [932]. binding [1781]. bio [1572]. bio-economic [1572]. bioaccumulation [1331, 1561, 973, 1061, 1821]. Bioassessment [420]. biochemical [170]. Biocomplexity [391]. biodiversity [898, 1370, 220, 1566]. Bioeconomic [937, 130, 161, 891, 313]. Bioenergetic [133, 1692, 42, 1135, 58, 708, 806]. bioenergetic-based [806]. Bioenergetics [125, 250, 238, 1129, 495, 889, 1826, 1563, 96, 214]. Biofilm [876]. biofilms [204]. Biogeochemical [298, 922]. biogeochemistry [1312]. biogeography [577, 647]. bioindicators [1531]. bioinformatic [1360]. Biologging [1822]. biological [1349, 228, 20, 1401, 1001, 1759, 732, 785, 881, 1809, 1300, 1782, 1787, 1333]. biology [1297]. biomarkers [1013]. biomass [499, 1781, 1469, 573, 1648, 294, 139, 174, 1057, 907, 941, 1105, 486, 1544, 1079, 715, 832, 309]. biometric [174]. biophysical [1768]. biota [454, 1478]. Biotelemetry [1738, 544, 1822]. biotic [846, 1689, 1490, 625, 1341]. Bipartite [1430]. bird [1214]. Biscay [1572, 706, 414, 1079]. Bivalve [449, 159, 1150, 658]. Bivalve-enhanced [449, 658]. bivariate [1374]. black [1025]. blackside [752]. blade [1837]. blood [1556, 854]. bloom [501, 1148, 476, 1814]. Blooming [1164]. blooms [1355, 472]. Blowin [194]. blowout [397]. Blue [1425, 54, 1324, 1184, 911]. blueback [975, 978]. bluefin [823, 335, 973, 88, 938, 1409, 1447, 56, 1200, 1036]. bluefish [1061, 559]. bluegill [325, 725, 660, 241]. bluespine [1203]. bluntnose [1024]. boat [1790]. bocaccio [91]. bodied [931, 1753, 915]. body [1653, 1344, 1124, 1817, 1220, 1673, 1160, 1256, 1730, 1197, 423, 1231, 1453, 904, 486, 1056]. boldness [1448]. bomb [512, 1203, 1425, 1475, 122, 538, 962]. bomb-produced [962]. Bonavista [1629]. Bonneville [685]. Bootstrapping [768, 67]. Boreal [1698, 99, 780, 1161, 393, 719, 1295, 1753, 874, 1370, 1821, 371, 324, 317, 1319, 293, 1833]. Boreogadus [1266]. born [260, 802, 1276]. Bosmina [71]. both [596, 1041, 841]. bottleneck [685]. bottom [1815, 244, 1480, 1496, 1753, 1469, 1370, 557, 1429, 1262, 597, 762, 703, 565, 1442]. bottom-contact [1496]. bottom-moored [1480]. bottom-trawl [1429]. bottom-up [1469, 597]. boundaries [1142]. boundary [381]. bouvieri [42]. box [1642, 1370]. brachyurans [436]. brackish [1292]. brain [480, 841]. brave [692]. Brazilian [1510]. breadth [466]. break [188]. Breaking [564]. bream [1501]. breeders [171, 11]. Breeding [847, 1768, 242, 943, 1015]. brevirostrum [1464, 547, 464]. bridle [674]. brief [1575]. brings [1031]. Bristol [789, 1775, 1717, 1651]. British [1166, 26, 65, 1594, 182, 890, 927, 435, 6, 27, 113, 226, 1362, 1520, 1694, 1337, 1204, 373, 165, 1372]. broad [84, 902]. broadband [426]. broadcast [931, 183]. Brook [1066, 1277, 564, 1642, 445, 404, 1471, 1320, 1673, 777, 803, 1723, 428, 1818, 1598, 1008, 490, 1328, 625, 1012, 602]. Brosme [216]. Broughton [1166, 182]. Brown [1244, 1187, 256, 229, 1578, 1088, 1381, 35, 564, 1642, 1004, 445, 1471, 347, 1630, 885, 1696, 1277, 415, 1483, 1842, 1035, 190, 1723, 358, 1149, 1573, 642, 423, 596, 626, 1488, 1476, 483]. Brownie [1399]. browser [1203]. Brunswick [464, 1632, 1634]. buffer [918, 1601, 24]. bulk [1354]. bull [1187, 930, 808, 1450, 1421, 1767, 336, 913, 418, 537, 1599, 82, 949, 62, 1618, 625]. bullhead [1149]. burbot [708]. bycatch [1049, 1177, 705, 1339, 903, 117, 1147, 13, 1662, 1098, 1500, 1623]. Bythotrephes [14, 245, 422, 1314, 48].

C [782, 211, 551, 287, 124, 1419, 1243, 56, 214, 962, 1834]. Ca [60, 210]. cabled [890]. cadmium [159, 154]. cage [661, 157]. cages [305]. calamary [289]. Calanus [622]. calcarifer [95]. calcium [1622, 422, 691, 1093]. calcium-rich [1622]. calculated [1534]. Calculation [1397, 1338, 952, 1401, 455]. calculations [1135]. calibrated [1485]. Calibration [1830, 1174, 1680, 810]. California [1022, 102, 135, 291, 538, 560, 132, 1039, 1752, 1391, 275, 1619, 1215, 592, 1714, 542, 1843, 1588, 1494]. californicus [681]. call [1437]. camera [677]. campechanus [117, 1538]. camtschaticum [1804, 616]. camtschaticus [572]. Can [1487, 1170, 1480, 322, 1615, 822, 80, 1435, 959, 1349, 1422, 52, 478, 552, 1760, 1551, 1018, 983, 1575]. Canada [687, 26, 1838, 1664, 1594, 360, 890, 682, 1816, 688, 27, 977, 1694, 898, 235, 452, 578, 1265, 1727, 411, 1622, 1505, 579, 1016, 92, 373, 858, 1370, 1518, 182, 1235, 1026, 356, 1513, 407, 390, 1833, 30, 435, 165, 1372, 1093, 1355, 611, 1498, 409, 1533, 941, 1677, 543, 464, 398, 6, 1632, 1634, 936, 585]. canadensis [125, 1413, 1030]. Canadian [298, 1377, 1830, 993, 1245, 1128, 903, 1680, 494, 329, 1148, 1256, 422, 968, 1796, 1826, 303, 1190, 1675, 1416, 477, 1411, 1164, 1057, 907, 56, 657, 633, 1693, 354, 1428, 625, 48, 1036, 257]. canadien [257]. Canal [1323, 609, 1180]. Cancer [1307]. cannibalistic [44]. Cannibals [694]. Canopy [1469]. Canyon [1068, 511, 1067]. Cap [1655, 1294]. capabilities [1764]. capacities [1201, 860, 971]. Capacity [414, 356, 769]. Cape [188, 103]. capelin [119, 682, 1484]. capital [11]. capriscus [1538]. capsule [1167]. captive [928, 1335]. captivity [1801, 1614]. Capture [1200, 241, 1604, 1343, 544, 445, 1238, 867, 643, 1822, 1706, 1567, 727, 1776, 1608, 1403]. Capture-induced [1200]. captures [1371]. carbon [1173, 393, 47, 235, 1267, 364, 1199, 1715, 504, 1590, 82, 1467, 1032, 194, 755, 1657, 214]. Carcass [621, 876, 1135, 1243]. carcasses [299, 1144, 1208]. Carcharhinus [1557, 537, 82, 786]. Cardiac [1639]. Cardiorespiratory [811]. care [804]. Caretta [887]. Carey [457]. carnivorous [906, 653]. carp [1267, 1715, 586, 1157, 1386, 801]. carpio [1386]. carps [1437]. Carr [983]. case [705, 1738, 1086, 117, 825, 445, 844, 308, 1701, 679, 685, 664, 742, 114, 1407, 382, 1509, 1008, 752, 1023, 1341]. Castor [1030]. catadromous [206]. cataractae [859]. Catastrophic [135]. Catch [974, 1109, 909, 1839, 1835, 1613, 1572, 1049, 421, 8, 5, 73, 319, 1815, 705, 710, 878, 376, 581, 72, 75, 273, 222, 1239, 215, 505, 531, 1001, 679, 169, 774, 13, 1615, 1550, 461, 1359, 663, 407, 390, 921, 1463, 988, 1724, 1635, 868, 1840, 703, 1390, 716, 684, 1333, 276, 632]. Catch-and-release [909, 376, 1724, 1840]. catch-at-age [319, 581, 75, 505, 531]. Catch-per-unit-effort [1109, 215, 1463]. catchabilities [756]. Catchability [574, 703, 1809, 834, 86, 565, 716, 684, 276]. catches [558, 111, 1085, 1662, 1825]. Catchment [1003, 977, 138, 247]. Catchment-wide [1003]. catchments [1170, 1680]. categorical [131]. cations [99]. catostomid [709]. caught [1238, 786]. caurinus [155]. causal [1038]. Cause [1329, 1559, 991, 442, 1117]. caused [99, 918, 1065, 310, 471, 1698, 1375]. Causes [1249, 1179]. cavity [89]. Ceded [1189]. cellular [379]. Celtic [1077]. center [668]. Central [592, 102, 719, 1833, 610, 1619, 264, 949, 472, 1328, 1458, 774, 132, 664, 594]. centuries [1505]. century [1203, 1023]. cepedianum [258]. Cercopagis [245, 530]. CFRN [1497]. chain [1122, 968]. chalcogramma [379, 619, 301, 280, 311, 387]. chalcogrammus [1701, 1262, 815]. challenge [1669, 785, 383]. challenges [1421, 1543, 1777, 1579, 897, 1776]. challenging [1620]. Champlain [1461, 1703]. change [314, 918, 1830, 1547, 1110, 615, 1671, 16, 1475, 1020, 913, 1319, 239, 553, 298, 653, 17, 42, 169, 1090, 793, 1373, 991, 1676, 915, 1308, 1745, 472, 1714, 236, 1698, 1375, 1601, 1224, 1112]. Changed [249]. Changes [618, 1256, 1750, 1792, 112, 1768, 843, 775, 1059, 333, 191, 1680, 719, 588, 1014, 1443, 551, 455, 360, 1773, 178, 118, 1135, 71, 1839, 1186, 720, 546, 1790, 80, 264, 1757, 1099, 809, 471, 633, 1763, 243, 1625, 1364, 40, 1554]. Changing [687, 688, 1187, 1131, 1582, 1774, 205, 1337, 143, 833, 672, 1080, 1821, 349, 1151, 1490, 1782, 1003]. Channel [1330, 53, 944, 1523]. channelization [1198]. channels [1242]. Chaoborus [775, 1524]. char [895, 58, 495, 347, 150, 415, 1483, 1842, 679, 665, 1190, 1633, 811, 596, 1533, 1008, 62, 79]. character [1374]. characteristic [421]. Characteristics [858, 977, 808, 242, 1587, 1731, 20, 893, 1016, 39, 178, 118, 1287, 1324, 1568, 522, 408, 561, 394, 1501, 50, 126, 1428, 1283]. characterization [547, 465, 417, 1152]. characterize [1351, 1804]. characterized [1384]. Characterizing [1664, 1594, 1246]. charged [1559]. charr [428]. chasing [1501]. Chaumont [1540]. Checkley [686]. chemical [105, 1646, 1346, 1286]. chemically [731]. chemistry [115, 1806, 649, 728, 235, 189, 630, 1290, 618, 320, 1596, 1808, 1629, 98, 262]. chemoreception [1050]. chemosensory [790, 783, 234]. chemostat [996]. cherry [1214]. Chesapeake [378]. Chief [1581]. Chignik [600]. Chilko [1650]. Chilliwack [927]. chimaeras [1165]. China [1382, 1073]. Chinese [1437]. Chinook [26, 27, 925, 1520, 166, 195, 1736, 650, 140, 1345, 1626, 3, 1070, 1519, 200, 562, 1127, 770, 552, 1765, 53, 331, 1801, 492, 548, 778, 702, 870, 132, 943, 1808, 731, 1039, 1752, 230, 597, 577, 822, 310, 1674, 984, 1593, 1209, 1156, 1792, 1006, 592, 1052, 739, 1145, 1770, 1647, 25, 1608, 1740, 1555, 6, 156, 841]. Chionoecetes [1396, 878, 1664, 400, 703, 1228]. Chironomidae [303]. chlorine [1784]. chlorophyll [207, 919, 70]. choice [562, 1839, 533, 802, 729]. choices [1255]. Chondrichthyes [617]. Chondrostoma [1446, 1441]. Choosing [1310]. chromosome [3, 1070]. chronology [538]. Chrosomus [752]. Chrysophrys [1733, 466]. chrysophytes [1552]. chub [931, 1067]. chum [1337, 845, 735]. CI [1086]. CI-HCRs [1086]. circadian [640]. circulation [519, 277]. circuli [773]. cisco [1540, 4, 517, 1477, 896, 1600]. Cladoceran [1384, 112, 1622, 293]. cladocerans [1132]. Cladophora [1152]. Clair [1330]. clam [556, 1271]. clarity [1667]. clarkii [291, 42, 448, 430, 408, 41, 255]. clasper [617]. class [1128, 1182, 392]. classes [782, 1702, 953]. Classification [1392, 1346, 709, 127, 627, 1574, 426]. classifying [251]. clava [1219]. clear [1063]. clear-water [1063]. clearcutting [46]. Climate [1768, 1365, 913, 1319, 236, 314, 1187, 918, 1547, 1131, 1671, 16, 191, 1337, 54, 306, 253, 239, 553, 833, 298, 653, 17, 42, 1408, 787, 733, 1022, 1373, 753, 1676, 1099, 472, 433, 1714, 1782, 1428, 302, 261]. Climate-driven [1768]. Climate-induced [1365]. climate-sensitive [54]. Climatic [619, 738, 1649, 1688, 1283]. cline [188]. clines [753]. clogging [1334]. clonal [1591]. closed [289]. closely [603]. Closing [1795, 288, 1370]. closure [782, 1733, 1469, 632]. closures [1177, 1406, 988, 1407]. clubshell [1219]. clumped [1358]. Clupea [292, 11, 1685]. clupeaformis [1158, 1811, 826, 361, 1171]. cluster [465]. clustered [1473]. clustering [1327]. clusters [710, 465]. co [540, 1041, 1497, 1645, 1225, 1126]. co-constructed [1497]. co-existing [1041]. co-occurring [540, 1645]. Coarse [1718, 743]. Coast [1291, 1049, 606, 468, 407, 390, 577, 611, 439]. coast-wide [577]. Coastal [764, 972, 409, 906, 1236, 1331, 1561, 1213, 894, 47, 1585, 449, 658, 1502, 54, 1312, 618, 288, 653, 1491, 373, 349, 434, 275, 1416, 1619, 1635, 165, 1595, 70, 1196, 1259, 1208, 678, 262]. coastrange [871]. coastwide [975, 599]. cod [1420, 424, 1762, 1081, 1487, 1815, 1085, 1181, 667, 648, 588, 1249, 909, 812, 1459, 81, 141, 1080, 320, 1266, 61, 231, 406, 185, 1288, 1609, 1400, 693, 1746, 109, 1617, 734, 866, 1029, 1652, 1315, 1629, 383, 51, 834, 193, 1044, 1498, 1595, 254, 1009, 1196, 437, 366, 134, 399, 608]. coded [1362]. coded-wire [1362]. codend [1662]. codistributed [1095]. coefficients [276]. coevolution [1673]. coexistence [1116, 418]. Coherent [1542]. Coho [1749, 925, 882, 1362, 1653, 1836, 847, 870, 1022, 1485, 927, 1619, 1209, 439, 116, 722, 394, 1291, 50, 1046, 506, 802]. cohort [259, 245, 1752]. cohorts [197, 559]. Coincident [1083]. cold [74, 833, 966, 1198, 1282]. cold-water [74, 833, 966, 1198]. Coldwater [172]. Collaborative [1693, 1664, 1497]. collapse [1570, 1624, 275, 1588]. collapsed [307, 1652]. collapses [318]. collected [322]. colonial [1552]. colonies [441]. colonization [496, 1831]. colonizing [925]. colonorum [206]. colony [43]. colony-specific [43]. Colorado [152, 1023, 1263, 511, 1067, 487, 344, 1068]. colouration [208]. Columbia [26, 1694, 1594, 373, 182, 890, 435, 165, 6, 27, 1166, 113, 226, 1362, 1520, 1736, 65, 140, 1337, 1204, 770, 1316, 331, 944, 1454, 583, 1599, 991, 230, 310, 927, 1372, 1792, 739, 302, 261]. column [434]. combination [1374, 1822]. combinations [131, 959]. Combined [649, 720]. Combining [1027, 557, 1429, 1262, 1747, 1239, 1748]. comeback [1044]. Comment [176, 26, 72, 515, 686, 373, 657, 73, 177, 516, 374, 1139, 983, 27]. comments [1140]. commercial [789, 1181, 1775, 887, 1655, 1150, 1294, 1509, 396, 426, 387]. commercially [531, 681, 1484]. common [1173, 1346, 699, 1386, 399]. commons [160]. Communicating [1086]. communication [1691]. Communiqué [162, 549, 550, 1087, 1217]. communities [780, 1113, 956, 203, 719, 494, 1622, 1370, 395, 46, 451, 810, 968, 293, 1596, 1282, 881, 1107, 441, 1757, 1435, 529, 1351, 1126, 473, 715]. community [112, 1830, 138, 1360, 829, 133, 1751, 1193, 443, 874, 1118, 1648, 1198, 806, 323, 647, 1576, 1106, 220, 1776, 737, 817, 772, 1625, 1698, 110, 475, 1111, 1554]. community-based [1776]. comparability [192]. comparable [1480]. Comparative [1475, 1442, 1280, 729, 1832, 564, 1095, 711, 459]. compare [974, 789, 902]. Comparing [879, 1354, 446, 1120, 998, 1246, 1635, 1063]. Comparison [1540, 1005, 735, 247, 758, 490, 113, 226, 600, 844, 71, 230, 663, 1679, 592, 703, 1040, 276, 1255]. Comparisons [1419, 380, 1295, 82, 79]. Compensation [1082, 1767]. compensatory [258]. competition [1337, 259, 981, 150, 845, 1168, 1225, 596, 1509, 957]. Competitive [1538, 259, 1598]. competitors [981, 1151]. completion [1021]. Complex [1591, 88, 380, 689, 307, 547, 1643, 215, 895, 1713, 725, 1399, 1406, 1225, 1593, 1052, 1270]. complexities [1401, 1646]. complexity [882, 563, 1800, 130, 161, 1726, 933]. compliance [1204]. complicating [197]. component [119, 643]. components [1340, 141, 214, 1441]. composition [967, 112, 558, 1658, 8, 5, 73, 105, 1128, 1565, 72, 1027, 1193, 1147, 1346, 574, 806, 323, 647, 383, 152, 1355, 377, 768, 1472]. compositional [1670]. compound [1354]. compound-specific [1354]. compress [955]. compromise [1204]. computational [304]. computationally [1544]. concentration [1173, 105, 1708, 1839, 358, 1636, 1032]. concentrations [1161, 1585, 578, 343, 367, 1833, 477, 1798, 745, 1524]. conceptual [534, 1835]. concern [1322]. conclusions [1139]. concordance [874, 647]. Concordant [459]. Condition [1630, 1240, 1803, 1344, 1061, 1020, 1817, 1244, 1033, 621, 1160, 141, 1750, 15, 246, 434, 758, 760, 1328, 814, 763, 1709]. Condition-dependent [1630]. Conditional [717, 1682]. conditions [1639, 977, 102, 334, 1109, 1587, 1269, 1711, 1794, 352, 761, 1288, 1759, 147, 1151, 1490, 435, 1685, 739, 1814, 975, 1783]. conference [469]. Confidence [244, 1086]. confirmed [1425]. confirms [748]. Conflicting [332]. conflicts [1130]. confluentus [930, 1767, 336, 913, 418, 62]. confound [1278]. confounding [1660]. confront [1543]. congeneric [1735]. congrès [469]. connected [1713]. Connecticut [77]. connections [1022]. connectivity [882, 115, 519, 738, 1096, 1672, 474, 336, 877, 277, 831, 1069, 881, 1536, 954, 1610, 713, 1606, 425, 1629, 840, 1376, 1497, 206, 737, 98, 1452]. consecutive [1335]. consequence [1329, 1117]. Consequences [640, 892, 852, 1062, 1604, 171, 1074, 644, 97, 796, 485, 1197, 1179, 237, 827, 41, 1207, 937]. Conservation [1204, 272, 1793, 1034, 77, 612, 547, 544, 291, 893, 951, 931, 251, 1778, 1599, 1577, 1142, 1372, 752, 1332, 524, 1595, 1437, 1829, 1322, 1555]. conserve [220]. consideration [1324]. considerations [1835, 1404, 1412]. considering [739]. consistency [1541]. Consistent [1158, 94]. consolidation [1839]. conspecific [1545, 1436]. conspecifics [985]. constrain [327]. constraints [1122, 1148, 1219, 1212]. constructed [589, 1497, 769]. construction [353, 769]. consumer [363, 1212]. consumers [57, 1207]. consumption [1508, 176, 1345, 1586, 125, 1191, 1385, 7, 177, 1563, 1041, 815]. contact [1496, 976, 1442]. Contaminant [1061, 359]. contaminants [699]. contaminated [731]. contamination [1201, 411, 1478]. Contemporary [98, 446, 17, 1408, 647, 206, 85, 1511, 108]. content [1827, 1278, 1128, 211, 1432, 1614, 527, 1684, 1685]. Contents [180, 248, 403, 192]. contests [748]. context [771, 750]. continental [1076]. continous [131]. Continued [1652]. continuous [1546, 727]. continuously [941]. continuum [996, 1094]. contraction [1034]. Contrasting [347, 543, 1653, 629, 1292, 1709, 1472, 729]. contribute [1666]. contributing [1507]. Contribution [296, 1262, 559, 1228]. contributions [1378, 1254, 265, 1296]. control [209, 1166, 964, 1086, 1438, 28, 644, 1794, 951, 1021, 11, 1092, 1407, 1417, 19, 1163, 473, 1734, 1333]. controlled [440, 445, 1172, 676]. controlling [1067, 963, 274]. controls [797, 647]. conventional [928, 1635, 1544]. Convergent [1008]. cool [1379, 1020, 1537]. cool-water [1537]. cooling [1171, 173]. coordinates [417]. copepod [622]. copepods [1132]. coping [951]. copper [154, 1749, 1653, 1240]. coral [1406, 1106]. corals [1823]. coregonids [238]. Coregonus [1158, 84, 4, 517, 1477, 896, 1811, 826, 1600, 361, 1171]. cormorant [1065, 441]. cormorants [1819, 563, 972]. correct [197]. correcting [111]. correction [557]. correlate [428]. correlated [1369]. Correlates [579, 906, 1578, 777, 225, 984]. correlation [506]. Corresponding [1493]. Corridor [1629]. Corrigendum [256, 1101, 1561, 195, 8, 658, 584, 1070, 357, 273, 301, 500, 517, 1560, 1011, 1057, 1453, 814, 300, 1153, 716, 302, 570]. corroborating [1005]. cortisol [593, 237]. cost [1258, 28, 1842, 405]. cost-effective [1258]. costly [274]. Costs [492, 1349, 238, 1422, 1174, 783, 1097]. Cottus [871]. Could [1719, 746, 1076]. counter [1500]. counter-herding [1500]. countergradient [459]. Counterintuitive [1017]. counts [677, 1175]. coupled [740, 971]. Coupling [889, 260, 556, 785, 283]. covariates [1765, 664]. covariation [1519]. cover [1671, 633]. cowcod [724]. crab [1396, 878, 1664, 400, 1323, 1597, 703, 1228, 1307]. crabs [572, 502]. crayfish [1131, 125, 719, 644, 66, 22, 1798]. creation [1757, 1163]. credible [720]. Creek [291, 1240, 15]. creeks [1379, 1163]. crested [1819, 972]. critical [1699, 480, 1411, 106, 1482]. crop [876]. Cross [1206, 1644, 1767, 1207]. Cross-ecosystem [1206, 1207]. cross-fishery [1644]. cross-population [1767]. Crow [977]. crustacean [1596, 1500, 919]. crustaceans [196, 1785]. Cryptic [724, 1588]. cucumbers [681]. cue [1045]. cues [790, 1047, 234]. culling [1814]. culture [1048]. cultured [572, 626]. Cultus [65]. culvert [1642, 864]. culverts [877, 1066, 1320]. cumberlandensis [752]. cumulative [1311, 239, 169, 1537, 1092]. Current [1291, 1843, 1245, 825, 88, 1446, 1779, 391, 25, 102, 538, 560, 1022, 1714]. currents [523]. curve [438, 654, 1616, 747]. cusk [216, 1584]. CUSUM [1092]. cuticle [691]. cutthroat [1236, 1672, 624, 291, 42, 1469, 807, 448, 430, 408, 41, 255]. cuvier [82]. Cyanobacteria [1781, 1164]. cyanobacterial [879, 1814, 1366]. cyanobacterium [489]. CyanoHAB [861]. cycle [1128, 1021, 831, 284, 1661]. cycles [1020, 1250, 1063]. cyclic [238, 798]. cycling [249, 1743, 1705]. cygnus [16, 277]. cylindraceum [1158, 1499]. Cyprinella [146]. cyprinid [1317]. Cyprinus [1386].

D [457, 340]. D1 [1512]. dace [859, 752]. Daily [1756, 512, 1385, 407, 390]. dam [765, 1751, 770, 1700, 685, 1437, 1829]. damming [1282]. dams [1316, 867, 1528, 1235, 1786, 1255]. Danio [1482]. Danish [249]. dans [1387]. Danube [1446]. Daphnia [71, 513]. dark [1063]. dark-water [1063]. Data [1404, 271, 1420, 248, 339, 421, 319, 939, 1738, 1321, 710, 1503, 1716, 835, 824, 1340, 216, 819, 1392, 1546, 427, 1249, 322, 222, 1239, 1027, 979, 893, 1794, 292, 301, 280, 1227, 756, 311, 1237, 1788, 679, 1393, 328, 557, 1429, 1678, 1550, 837, 1754, 1746, 917, 663, 1515, 1234, 1352, 988, 1411, 1092, 282, 1747, 1635, 1139, 1514, 174, 459, 1772, 377, 1218, 131, 768, 1562, 1511, 315, 410, 714, 1395, 1544, 1670, 514, 1046, 276, 636, 1402, 357, 273]. data-limited [1321, 1794, 1092, 1747, 1514, 1772, 1562]. data-poor [1227, 1754, 1234, 1772, 714, 1046]. data-rich [1772]. database [1245, 1252]. dataset [578]. date [1074]. dating [512, 646, 1203, 1425, 773]. Davidson [983]. day [71, 94]. days [721, 830, 1692, 1650]. Daytime [790]. dead [1522, 557, 1132]. deathly [274]. d’eau [469]. debris [888, 135, 1565, 46]. Decadal [1763, 1680, 253, 414, 543]. decadal-scale [253]. decade [420]. decades [1643, 1256, 1469, 1555, 344]. decapod [196]. Decay [1565]. December [457]. decide [1177]. decision [1450, 1729]. decision-making [1729]. decisions [1529, 827]. Decline [606, 1768, 1543, 743, 1559, 1652, 1132, 398, 1588, 1117]. declined [1505]. declines [1170, 422, 745, 1745]. Declining [1353, 916, 31]. decomposers [905]. decomposition [299, 905, 1225]. Decoupling [1752, 1799]. decrease [99, 353, 458]. decreased [1564]. deep [646, 1136, 1697, 831, 919, 755]. deep-sea [1697, 831]. deep-water [646]. deepening [203]. DeepVision [677]. deepwater [512, 397]. defaults [532]. defense [1286]. deficiency [442, 941]. defining [92]. deglaciation [37]. Degradation [120, 420, 133, 221]. degree [609, 721, 830, 259, 895, 1692, 94]. degree-day [94]. degree-days [721, 830, 1692]. dehydrogenase [1186]. delay [900, 1055]. delay-difference [900, 1055]. Deleterious [1334]. Delta [1749, 1653, 1069, 566, 1059, 284, 1593, 1544]. delta-model [1544]. demand [1124]. demersal [1095, 588, 1350, 1459, 372, 391, 1623]. demographic [1536]. demographically [1096]. demographics [186]. demography [205, 1731]. demonstrate [1265]. dendritic [1248, 1831]. Denmark [359]. denning [1416]. densities [242, 1436, 1062, 815, 1159]. Density [1108, 503, 392, 596, 1732, 1034, 209, 1658, 1088, 35, 1358, 205, 258, 1337, 1427, 495, 1316, 845, 1318, 848, 852, 1252, 1590, 787, 190, 595, 1133, 741, 1597, 183, 284, 806, 896, 1107, 1725, 1197, 1504, 1674, 22, 660, 423, 1685, 363, 1276, 1618, 184, 1423, 1048, 86, 94, 716, 684, 1283, 412, 447, 309, 1056]. Density-dependent [1108, 503, 392, 596, 1358, 1337, 1427, 495, 845, 1318, 848, 741, 94, 716, 684]. Density-independent [1732]. dentatus [47, 1742]. depend [1288, 1760]. Dependence [1056, 1034, 1240, 190, 560, 284, 896, 1811, 1504, 423, 1231, 86, 707, 1453]. dependent [1108, 503, 558, 166, 195, 1358, 987, 335, 1337, 1427, 495, 639, 1630, 845, 1748, 1318, 1227, 848, 686, 1042, 741, 392, 1194, 1606, 1742, 1134, 596, 1094, 1100, 94, 716, 684, 313]. depends [1248, 1188]. depleted [1049, 1181, 1249, 1147, 724, 1082, 1489]. depletion [1230, 444]. deploying [1825]. deposition [298]. depredation [1623]. depressed [1034]. depth [558, 333, 417, 172, 391, 718, 1293]. depth-dependent [558]. depth-profiling [1293]. depth-related [391]. Deputy [1581]. derive [1826]. derived [10, 153, 469, 986, 1193, 1646, 1692, 515, 164, 286, 516, 1243, 1112]. Deriving [835, 144, 123]. describe [1692, 1532, 1416, 1511]. Describing [270, 704]. description [1272]. Design [1306, 1360, 638, 1654, 923, 278, 192, 1683, 1710]. designed [1157, 1333, 1442]. Designing [117, 1758]. designs [413]. desotoi [839]. despite [1161, 45, 1286]. detect [824, 234, 1134, 1818, 1580, 1112]. detected [1681, 1755]. Detecting [111, 215, 1028, 1005, 413, 1000, 1554]. Detection [586, 1836, 1546, 1457, 638, 627, 345, 63, 820, 949, 1695]. detects [1680]. determinants [741, 338]. determination [1058, 502, 617, 863]. determine [738, 1541, 666, 429, 1808, 1294, 1106, 904, 1423]. determined [1007, 1367, 791, 416, 467]. determines [1124, 1460, 44]. determining [1258, 647, 282, 533]. d’évaluation [1387]. develop [197]. developing [683]. Development [127, 1385, 1239, 1330, 965, 1695, 808, 337, 494, 11, 1286, 595, 690, 1446, 1759, 327, 480, 1568, 1335, 1488, 268, 1231, 1453, 25, 1073, 1171, 1740, 1769]. developmental [233]. device [1748, 1098, 1500]. devices [1007, 822]. diadromous [1224]. diagnostic [291]. diamond [1622, 1426]. Diapausing [45]. Diatom [398, 1258, 411, 360, 221, 473, 794]. diatom-based [1258, 360]. Diatom-inferred [398]. did [1646]. DIDSON [765]. Didymosphenia [501, 964, 1148, 771]. Diel [708, 775, 238, 1020, 1001, 1226, 1723, 896, 999, 694, 49, 1743, 1250, 110, 1063, 1755, 601]. diel-cycling [1743]. Diet [124, 453, 1658, 1819, 403, 125, 64, 154, 1263, 582, 860, 653, 342, 320, 1590, 1210, 287, 466, 1741, 1467, 1037, 1804, 340, 814, 763, 1091, 1834]. diet-induced [1091]. Dietary [535, 1059, 875, 1620, 1476]. diets [1354, 440, 1432, 1254, 1531, 1328, 341]. differ [1365, 50, 607]. difference [900, 791, 1055]. Differences [5, 73, 72, 1127, 352, 1091, 1805, 1420, 29, 189, 1462, 279, 1404, 754, 58, 552, 1765, 696, 38, 232, 1167, 145, 8]. Different [359, 1195, 1239, 659, 959, 1524, 24, 899, 616]. Differential [782, 1229, 1158, 701, 1539]. differentially [1512]. differentiate [849, 1253, 1526]. differentiation [895]. differing [1800]. differs [753]. difficult [674, 1824]. difficult-to-age [1824]. difficult-to-sample [674]. diffusion [1257]. diluviana [245]. dimension [266]. dimensional [402]. dimensions [497]. dimictic [629, 554]. dimorphism [1251]. dinoflagellates [409]. dioxide [393, 1267, 1199, 1715, 504, 1657]. diploid [897]. Diporeia [875]. Diptera [303]. Dipteran [603]. Direct [137, 164, 502, 356, 1428, 429, 513]. directed [1285]. direction [1090, 926]. Directional [616, 1596]. discard [1191, 711, 762, 1307]. discards [588, 989]. discharge [1088, 1446, 984, 561, 1171]. discontinuum [178, 118]. discovery [1526, 861]. discrete [541, 729]. discriminate [1303]. discrimination [1354, 465, 1263, 1027, 401, 697]. discus [40]. disease [368, 788]. diseases [1297, 689]. Disentangling [1169, 597, 725]. Dispersal [925, 1358, 971, 463, 608, 1088, 1096, 1024, 1095, 630, 481, 233, 1369, 1456, 504, 1495, 1446, 1326, 665, 1157, 1356, 68, 1368, 1437, 366, 483, 1441]. dispersed [1358]. disperser [1546]. dispersing [1223]. DISPLACE [749, 1101]. displacement [749, 1101]. displacing [1187]. display [1509]. displayed [95]. disproportional [1147]. disrupt [1723]. dissemination [567]. dissolved [1173, 235, 996, 452, 1590, 1833, 1032, 110, 1063, 1524, 1769]. Dissostichus [519]. distance [723, 1097, 969]. Distinct [1755, 58]. distributed [1741]. Distribution [20, 650, 333, 481, 452, 1309, 171, 669, 491, 661, 846, 372, 858, 1299, 1785, 55, 1532, 548, 923, 461, 1754, 303, 682, 1157, 413, 1816, 389, 471, 592, 862, 1647, 1300, 1437, 1418, 1291, 1688, 184, 315, 1390, 387, 103, 1634, 521, 570, 632, 729]. distributions [912, 1392, 666, 148, 792, 1440, 402, 394, 302, 261]. districts [384]. disturbance [1170, 100, 659, 610, 865, 1328, 243, 1575]. Disturbances [46, 135]. disturbed [1552]. ditropis [333]. divergence [895, 1573, 1025, 1533, 206, 718, 1292]. Divergent [1220, 577, 1673, 1823]. diverse [1643, 1262, 1057, 907, 1677, 1159]. diversification [892]. diversion [822, 805]. diversities [1566]. Diversity [1422, 130, 508, 112, 1570, 299, 138, 612, 895, 1751, 568, 555, 800, 753, 272, 522, 152, 1831, 1057, 907, 1145, 1511, 1698, 1296, 142, 1384, 1761, 161]. DLMs [1562]. DNA [1836, 3, 1070, 586, 1412, 104, 645, 1804, 1064, 85]. Do [305, 819, 1431, 1812, 11, 234, 731, 1610, 441, 1509, 106, 1250, 78, 899, 605, 577, 713, 1568, 988, 1598, 520, 324, 969, 50]. DOC [249]. document [1418]. documented [1266, 635]. Does [519, 250, 1343, 1733, 515, 286, 1813, 327, 290, 1466, 220, 1498, 1651, 313, 1323, 1597, 711, 316, 285, 1718, 516]. dogfish [669, 792, 767]. DOI [257]. DOIs [257]. Dolly [600, 1016, 976]. dolomieu [334, 1186]. dolphins [1623]. Domestic [32, 897]. domesticated [149, 985]. domestication [836, 1048]. dominance [35, 798, 326, 1366]. dominate [1132]. dopamine [1512]. doriferus [320]. Dorosoma [258]. Doryteuthis [1001]. dosage [1537]. dose [683]. Dosidicus [920]. double [1819, 972]. double-crested [1819, 972]. Dover [1788]. Down [288, 996, 597]. Downstream [1356, 1437, 1236, 1622, 1446, 546, 791]. downwind [780]. drain [1375]. drainage [1518, 717, 984]. draining [1653]. drains [1375]. Dramatic [411]. draw [969]. drawdown [388]. dream [1729]. dredge [1706, 1271]. Dreissena [970, 245, 530, 990, 163, 992]. dreissenid [875, 207]. dreissenids [403, 361]. drift [1587, 1129, 277, 496, 1708, 1325, 776, 1334, 1179, 1347, 1376, 1040, 1791, 1444]. drift-feeding [1347, 1040, 1791]. drift-foraging [1129]. drifting [1007, 1829]. drive [650, 1451, 130, 161, 605, 1839, 1783]. driven [1768, 1059, 381, 961, 743, 756, 720, 896, 1652, 22, 49]. driver [573, 1714, 1094]. Drivers [1841, 1671, 1741, 112, 1108, 1016, 619, 1304, 1078, 152]. drives [1182, 306, 414]. driving [57]. Drought [743, 859, 1213, 1225]. Drought-driven [743]. drum [965]. drummondhayi [646]. Dry [1491, 473, 1039]. dual [750]. due [1230, 1503, 381, 331, 1398, 321, 1652, 1364]. Dungeness [1323, 1307]. Dunlop [1140]. duration [1749, 842]. during [843, 1838, 1081, 1050, 650, 606, 1733, 613, 1202, 1556, 1020, 455, 1505, 889, 1456, 670, 1483, 1495, 997, 1689, 1813, 132, 546, 791, 597, 1799, 1650, 90, 972, 739, 142, 1474, 947, 103, 1444, 1632, 1341, 1718]. Dusky [1557]. dwelling [859, 1248, 347, 42, 363, 1476]. Dynamic [1059, 955, 1294, 1835, 749, 1101, 1716, 191, 499, 946, 532, 582, 708, 252, 574, 283, 1828, 1361, 1209, 1114]. Dynamics [1071, 1072, 1313, 219, 10, 153, 424, 1081, 821, 1136, 144, 123, 1115, 1061, 961, 666, 1542, 75, 662, 306, 215, 1178, 1713, 1481, 1075, 515, 1451, 245, 500, 450, 61, 270, 511, 1599, 744, 1399, 286, 516, 304, 1824, 732, 132, 634, 734, 1600, 1757, 1028, 1308, 435, 1747, 382, 934, 266, 1139, 704, 414, 1445, 1123, 692, 714, 514, 1625, 163, 1363, 1384, 1341, 309, 44]. dynamique [1071]. dystrophication [475].

Earlier [1676, 1730]. Early [1836, 482, 480, 1534, 1703, 841, 115, 1838, 1081, 650, 217, 1202, 534, 200, 1374, 1074, 455, 1505, 345, 523, 1495, 1750, 270, 943, 791, 597, 485, 1315, 1619, 68, 1649, 1740, 1341]. Earthquake [1029]. East [1029, 1382, 1073]. eastern [424, 421, 1585, 379, 1095, 400, 618, 1727, 1701, 301, 280, 311, 1237, 1262, 484, 1447, 1235, 303, 1376, 409, 1645, 1804, 703, 437, 1228, 262]. eats [1379]. eau [1206]. echo [1480, 1305]. echosounder [426]. eco [994]. eco-genetic [994]. Ecological [168, 702, 1357, 41, 1037, 1716, 1109, 135, 1344, 1800, 1713, 631, 178, 118, 67, 1324, 1821, 1754, 1485, 796, 1534, 801, 1580, 1554, 729]. ecologically [436, 1790, 1142]. écologie [1102]. ecology [1102, 333, 970, 819, 853, 1553, 143, 673, 1602, 787, 1186, 1326, 1560, 1492, 915, 1619, 1142, 1798]. ecology-based [853]. economic [1572, 1729, 667, 28, 951, 619, 732, 998]. economics [414, 397]. Ecopath [804]. ecoregion [1567]. ecoregion-level [1567]. Ecosim [804]. Ecosystem [469, 687, 1451, 560, 1574, 1714, 688, 1206, 102, 1841, 1819, 478, 1502, 1729, 34, 961, 759, 129, 306, 532, 1427, 1160, 119, 653, 1120, 1370, 1185, 296, 870, 1242, 720, 283, 1391, 1721, 1327, 1207, 414, 470, 809, 792, 767, 1531, 236, 707, 1569, 1580, 1843]. ecosystem-based [759, 129, 1120, 1721, 1531, 1569]. ecosystem-driven [961]. ecosystem-level [532]. écosystèmes [469, 1206]. ecosystems [420, 1557, 456, 465, 1104, 551, 1481, 574, 1828, 1141, 1243, 1823, 1789, 1259]. ecotype [1025]. ecotypes [1420, 1253]. ectoparasite [1178]. eddies [1440]. Editor [1581]. Editor-in-Chief [1581]. eDNA [1360]. edulis [120]. edwardsii [429]. eel [1582, 700, 853, 39, 1329, 1684, 1444]. eelgrass [611]. eels [855, 1605, 1201, 853, 263, 1534, 1281, 1787]. Effect [1815, 1784, 770, 372, 15, 439, 1840, 25, 975, 508, 16, 307, 452, 588, 659, 150, 207, 1662, 1406, 183, 175, 806, 711, 1167, 1407, 656, 641, 1052, 126, 1202]. effective [1258, 291, 568, 642, 1691, 769, 134, 1383]. Effectively [929]. Effectiveness [552, 1204, 1098, 927, 24, 1554]. Effects [469, 1236, 808, 203, 105, 836, 154, 905, 14, 948, 845, 1318, 1370, 269, 1699, 358, 680, 1434, 943, 246, 1107, 1636, 999, 1798, 561, 1305, 533, 340, 763, 184, 1403, 1555, 1769, 262, 10, 153, 1019, 256, 229, 1108, 649, 956, 556, 368, 1110, 1161, 1230, 299, 813, 1278, 1430, 824, 1181, 12, 135, 1115, 1378, 540, 233, 1360, 1541, 429, 1033, 743, 1471, 1342, 582, 1401, 208, 1199, 553, 568, 1295, 795, 225, 950, 848, 523, 17, 761, 39, 137, 1287, 701, 1648, 787, 451, 968, 1251, 1398, 1198, 751, 1797, 484, 121, 461, 733]. effects [851, 693, 598, 1558, 717, 597, 1176, 1617, 559, 1334, 742, 480, 371, 1015, 201, 1652, 1809, 1619, 423, 157, 1435, 1593, 41, 1539, 536, 1099, 472, 470, 1156, 838, 1276, 762, 1831, 1097, 433, 1631, 62, 1437, 730, 750, 1328, 1192, 815, 1048, 1740, 514, 1296, 326, 1428, 387, 346, 1668, 1208, 33, 173, 828, 1341, 1580, 1554, 814]. effets [469]. efficacy [659]. efficiencies [345]. Efficiency [793, 444, 1459, 701, 557, 663, 1706, 941]. Efficient [967, 370, 1544]. effluent [1622, 110]. Effort [345, 1835, 749, 1101, 1258, 1686, 1543, 835, 1109, 1239, 491, 215, 251, 1839, 921, 1463, 988, 384, 1351, 1300, 1589, 714, 716, 684, 276, 632, 729]. efforts [1796, 752, 108]. Egg [176, 7, 177, 202, 1614, 1749, 1666, 1127, 1456, 1568, 1167, 316, 285, 1156, 493, 1840, 1740, 947]. egg-to-fry [1156]. eggs [1687, 45, 523, 965, 1167, 587, 934]. eight [609]. elasmobranch [1278, 401]. elasmobranchs [997, 617]. electric [1157]. electrofishing [1680, 793, 1246, 949]. Electronic [181, 427]. element [1027, 260]. Elemental [289, 105, 627, 591, 555]. Elevated [454, 1225, 1478, 1199, 1126, 1657]. elevation [474, 1223, 237]. Eliminating [1535]. elimination [1784]. emamectin [386]. embayments [1585, 1366]. embryo [916, 1228]. embryological [1575]. embryonic [233, 1194]. embryos [808, 404, 1171]. emergence [1653, 1329, 297]. emergency [1019]. emerging [673, 1352, 689]. emigration [541]. emphasis [690]. Empirical [839, 633, 49, 1283, 558, 127, 1690, 401, 863, 462]. employs [1338]. enclosures [380]. encountered [1579]. end [1487, 1815, 1653]. end-of-season [1653]. endangered [859, 12, 948, 679, 1067, 539, 746, 106, 1219, 524, 1786, 884, 709, 467]. endocrine [546]. endurance [175]. enduring [420]. Energetic [1122, 238, 1174, 1803]. energetics [77, 346]. Energy [994, 132, 361, 335, 1253, 363, 520, 838, 1057, 907, 904, 1159, 1306]. engineering [761, 1458]. England [1146, 740, 1176, 1763, 1455, 1719]. Engraulis [465, 1313]. enhance [1018, 78]. enhanced [449, 658, 1627]. enhancement [1349, 568, 848, 1597, 742, 272]. enhances [621, 884]. Enhydra [681]. Enriched [670, 758, 535]. enrichment [480, 593, 1276, 755]. enrofloxacin [1280]. ensemble [1612]. entanglement [354]. Entiat [901]. entombment [404]. Entosphenus [1047, 1383]. entrainment [822, 173]. entrance [980]. entre [1206]. entry [613, 1220, 1662, 546, 1053]. entry-level [1053]. enumeration [1258]. Environment [293, 898, 909, 946, 582, 672, 1637, 1262, 1609, 943, 620, 526, 842, 1755]. Environmental [650, 233, 1620, 1148, 231, 428, 338, 566, 152, 649, 1258, 615, 821, 1155, 1272, 1273, 1836, 217, 1269, 719, 1014, 1427, 1047, 1594, 1456, 148, 1241, 761, 797, 422, 1412, 771, 741, 664, 349, 38, 303, 323, 1176, 1617, 1760, 861, 915, 1809, 963, 1078, 435, 645, 1685, 529, 1276, 767, 897, 1064, 1618, 1641, 722, 975, 1601, 585, 1112, 406]. Environmentally [844, 756, 720]. environments [1768, 853, 690, 541]. enzymatic [1281]. Eopsetta [538]. epidemiology [1787]. Epinephelus [646]. epizootic [1521]. epizootics [158]. epizootiology [1138]. equal [1041]. equation [8, 5, 73, 72]. equations [1338]. equivalence [838]. equivalent [436]. era [1352]. Erie [880, 1805, 250, 1303, 1720, 1443, 388, 523, 1431, 955, 1563, 861, 128, 1054, 528, 1466, 365, 295, 941, 1017]. Eriocheir [1280]. Eroding [1628]. erosion [626]. Erratum [257, 144, 161, 178, 406, 407, 448, 316, 59]. error [1486, 1616, 276]. errors [75]. erythraeum [476]. ESA [1526]. ESA-listed [1526]. Escanaba [1011, 958, 1631]. escape [1459, 1748, 1285]. escaped [20, 1468, 1695]. escapee [608]. escapement [438, 1661, 1046, 1651]. escaping [305]. Esox [848, 1292]. essential [1299, 1532]. establish [1774]. Establishing [1759, 1142, 336]. establishment [1836, 1223, 196, 245, 422, 801]. estimate [1658, 52, 12, 216, 1385, 1496, 1687, 1697, 240, 1336, 301, 280, 1429, 1810, 1254, 1234, 1661, 727, 174, 1175, 1647, 1383]. estimated [339, 545, 1249, 273, 222, 756, 321, 703, 678]. Estimates [1345, 1425, 1338, 1756, 671, 427, 129, 322, 948, 1193, 1005, 1252, 1001, 679, 1807, 328, 774, 595, 1813, 1406, 405, 356, 614, 839, 459, 862, 402, 1403, 1079, 1405, 1046, 1455]. Estimating [1736, 440, 1767, 1586, 1546, 1592, 1065, 1765, 1089, 867, 511, 744, 183, 63, 430, 1616, 645, 389, 949, 762, 377, 96, 276, 726, 967, 291, 312, 1737, 574, 1746, 1516, 2, 300, 1055, 1395, 514, 715, 448]. Estimation [248, 1316, 1706, 818, 779, 319, 974, 167, 136, 122, 505, 507, 531, 479, 1525, 674, 837, 1744, 1758, 91, 1702, 1562, 86]. estimator [370, 654, 917]. estimators [967, 974, 438]. estrogen [146]. estuaries [449, 658, 731]. Estuarine [460, 503, 1059, 168, 537, 582, 579, 1602, 132, 1190, 1758, 210, 1368, 1445, 1743]. Estuary [682, 1445, 1586, 288, 1606, 1510, 712, 206, 1427]. estuary-dependent [1606]. ethanol [1512]. etiologies [689]. Eubalaena [354]. Eumetopias [824]. Euphausia [103]. Eurasian [1439, 395, 66, 1460]. Europe [1108]. European [1768, 1201, 606, 64, 1329, 278, 1371, 1684, 1214, 498, 1444]. eutrophic [1781, 1355, 1126, 521, 570]. eutrophication [1348, 712, 70]. evaluate [144, 123, 375, 1135, 284, 1779, 410, 1442]. evaluated [1399, 1606]. Evaluating [453, 444, 1450, 581, 1137, 659, 1172, 1788, 1185, 1609, 1550, 807, 806, 1486, 210, 739, 1734, 715, 1627, 1146, 534, 1457, 332, 969, 1651]. Evaluation [544, 1013, 1751, 1324, 1398, 1477, 1359, 1684, 1390, 813, 1273, 1774, 1303, 1775, 1729, 1728, 1323, 582, 1795, 1777, 848, 1330, 810, 1828, 1780, 49, 1270, 1555, 1333, 1719, 801, 1695, 1709]. evaluations [1772]. EvEel [853]. event [1249]. events [499, 381, 622, 312, 1843]. Evidence [695, 386, 1483, 1461, 373, 725, 1042, 165, 784, 228, 166, 195, 29, 894, 1050, 1836, 404, 45, 380, 1274, 1624, 1412, 1367, 741, 991, 1026, 1414, 1486, 866, 915, 480, 1629, 1335, 1484, 365, 1770, 976, 748, 1283, 801, 1823]. evoke [1123]. Evolution [1448, 1373, 446, 1438, 143, 673, 338, 937]. Evolutionary [1162, 408, 853, 895, 1702, 1580]. exacerbate [1504]. exacerbates [385]. examination [1021, 185, 978]. examine [1655, 792]. Examining [824, 1808, 1308, 757, 713]. example [952, 1065, 292, 1318, 1324, 1445, 1584, 1403, 1383, 962]. exchange [986, 281, 45, 1211, 1167]. exclusion [1098]. excretion [1673, 919]. Executive [1581]. exert [1208]. exhibit [11, 718]. exhibits [1405]. existing [1041]. exit [1240]. exogenous [146]. exotic [305, 1836]. Expected [933, 1001, 1057, 907]. expedited [1825]. experience [866, 1618, 1693, 841]. experiences [856]. experiment [956, 1367, 1730, 1460, 157, 1467, 703]. Experimental [1303, 58, 1590, 966, 1814, 748, 1642, 876, 781, 905, 404, 1238, 386, 1471, 380, 848, 510, 1750, 164, 185, 1242, 463, 801, 1552]. experiments [45, 1172, 805, 366]. expert [801]. explain [1768, 1170, 659, 595]. explained [1337, 1443, 80]. explaining [323, 394]. explains [1539, 1417]. explicit [935, 839, 173]. exploitation [894, 1404, 266, 1011, 958, 1189, 513]. exploitative [150]. exploited [952, 446, 431, 232, 1484, 391, 486, 799, 1153, 1580]. Exploiting [612]. exploration [140]. Explorative [885]. explore [853, 181]. Exploring [75, 798, 771, 1406, 283, 1189, 1613, 1738, 667, 1329]. exposed [1837, 1790]. exposure [45, 659, 1199, 141, 968, 1822, 1810, 146, 656]. expression [141, 723]. extended [844]. extensions [512]. extent [1456, 807]. extinction [408]. extracted [849]. extraction [232]. Extreme [473, 1836, 1042, 1226, 1782]. extrinsic [649].

F1 [1434]. face [1547, 1638, 169, 1364]. facility [1776]. fact [926]. Factors [1067, 1513, 683, 1006, 816, 156, 649, 1354, 650, 847, 588, 1263, 1014, 1084, 1456, 1318, 148, 725, 916, 293, 1689, 690, 1821, 401, 1357, 428, 264, 758, 1598, 1476, 1618, 722, 437, 625, 1341]. Facultative [827]. FADs [1007]. Failure [326, 1580, 1166, 991, 146]. falciformis [786]. fall [770, 1647]. families [526]. family [864]. farm [373, 165]. farm-induced [373, 165]. farmed [20, 1801, 1468, 1297, 702, 1434, 109, 526, 608]. farms [823, 294, 374]. Faroe [337]. fat [1684]. fate [1604, 1660, 1607, 1243]. Fatty [279, 440, 64, 1172, 1432, 800]. fauna [305]. faunal [188]. Favorites [1191]. feasibility [1720]. Feasible [370]. Features [465, 1508, 609, 940]. February [469]. fecal [120]. fecundity [11, 916, 493, 1188, 1228]. fed [104]. federal [974, 936]. federally [752]. feed [305, 227]. feedback [1795, 1794, 1407]. Feeding [1492, 1259, 1768, 248, 440, 563, 1382, 1289, 177, 979, 1172, 708, 1080, 1347, 1798, 676, 1040, 1791, 40, 1056, 1560]. Female [1505, 1605, 1731, 400, 1817, 202, 1563, 385, 1335, 655, 828]. females [575]. fen [924]. fens [603]. ferruginea [935]. fertilization [1367, 183, 1474, 1458]. février [469]. fiction [926]. fidelity [1060, 601, 1061, 1609, 31, 1036]. field [1303, 125, 1193, 1690, 109, 1679, 1538, 40]. field-derived [1193]. filamentosus [512]. filtered [645]. filtering [1360]. fimbria [1274]. Fin [626, 432, 1232, 709, 535]. final [613, 1556]. find [1414]. finding [224]. Fine [710, 1518, 295, 1077, 1632, 1494, 808, 1756, 1620, 1016, 1573, 327, 382, 536, 1040, 1791, 1755, 678]. Fine-scale [710, 1518, 295, 1077, 1632, 1494, 1756, 1620, 1016, 1040, 1791, 1755, 678]. fine-scaled [1573]. fines [404]. fingerprints [289]. Finland [1348]. Finless [1382]. finned [1305]. fire [913, 1451, 57]. fire-prone [913, 1451]. First [1380, 1266, 476, 1081, 1801, 132, 1650]. first-generation [1801]. Fish [721, 830, 1020, 1360, 1545, 796, 1606, 1568, 1157, 907, 1123, 1375, 967, 1669, 453, 508, 906, 1113, 1665, 1108, 1034, 952, 859, 209, 503, 1806, 843, 749, 1101, 1331, 1561, 929, 1110, 558, 248, 1837, 1161, 1658, 1354, 601, 1583, 111, 912, 1268, 857, 1155, 1716, 205, 77, 223, 584, 1756, 1430, 898, 1109, 1712, 499, 52, 12, 630, 1290, 279, 1248, 1340, 281, 125, 578, 446, 1413, 544, 1681, 534, 671, 1223, 1553, 143, 154, 318, 465, 167, 1457, 877, 1265, 909, 946, 1007, 122, 1239, 1369, 981, 661, 133, 417, 208, 551, 668, 1261]. fish [1594, 455, 240, 766, 1512, 659, 1751, 1172, 631, 1592, 1696, 638, 889, 1065, 1701, 833, 531, 867, 1692, 443, 148, 852, 1773, 1252, 994, 1241, 589, 172, 1479, 137, 605, 1299, 1785, 504, 1042, 1812, 1602, 1412, 1690, 1590, 1648, 1285, 955, 1297, 968, 1737, 1532, 1689, 1222, 595, 1662, 966, 1465, 1759, 367, 304, 1813, 1282, 741, 121, 461, 294, 831, 785, 793, 881, 283, 890, 1493, 287, 1038, 1107, 63, 1558, 1790, 353, 663, 356, 921, 1560, 1492, 822, 425, 1347, 272, 1233, 1352, 1076, 441, 1356, 1435, 1078, 1298, 1028, 1260, 260, 1576, 1473, 1747]. fish [1724, 1246, 805, 152, 614, 330, 1484, 727, 1368, 922, 170, 1702, 90, 834, 611, 1106, 1551, 1823, 677, 1445, 391, 762, 1772, 1351, 1741, 1623, 463, 1332, 757, 1175, 206, 1440, 1703, 227, 1641, 1516, 1829, 340, 969, 426, 750, 737, 488, 817, 437, 1763, 1743, 96, 59, 21, 315, 565, 884, 1094, 772, 1188, 1783, 1625, 1698, 1843, 755, 1040, 1791, 1077, 142, 1651, 1283, 1683, 214, 33, 863, 1232, 241, 962, 709, 98, 1709, 899, 1247, 348, 715, 1580, 309, 636, 1452, 1224, 1112, 886, 44, 1057]. fish-guidance [822]. fishable [309]. fished [309]. fisher [1613, 1535, 1624, 1754, 1395]. Fisheries [257, 585, 1349, 1102, 36, 929, 370, 1049, 1686, 1487, 1507, 319, 1738, 892, 974, 813, 1422, 1205, 1774, 375, 856, 903, 478, 1086, 1392, 1535, 318, 427, 1728, 357, 271, 673, 844, 1323, 129, 1238, 332, 1104, 1777, 1794, 951, 756, 130, 161, 891, 1408, 1370, 774, 1118, 1285, 706, 740, 1550, 1424, 1571, 634, 1119, 1279, 1500, 1391, 713, 1234, 1577, 1463, 711, 1529, 232, 1092, 384, 1376, 247, 1509, 1514, 1772, 1271, 1275, 1714, 1531, 396, 377, 1009, 1677, 397, 1608, 1693, 410, 1390, 1442, 348, 1307, 1580]. fishermen [296, 989]. fishers [1509, 266, 678]. Fishery [1388, 1588, 1644, 1570, 1387, 209, 368, 1572, 421, 894, 1177, 1815, 705, 1543, 16, 835, 878, 824, 1775, 216, 667, 337, 54, 781, 1720, 1717, 588, 759, 1007, 429, 1323, 1238, 1459, 651, 887, 84, 619, 1010, 852, 1393, 1408, 1785, 1839, 706, 1615, 1824, 1406, 732, 1098, 1779, 998, 321, 1162, 1676, 965, 1606, 1675, 1724, 786, 834, 1018, 439, 267, 254, 730, 1395, 1200, 1403, 1651, 716, 684, 1284, 1554, 276]. fishery-independent [216]. fishery-induced [1554]. fishes [1173, 1213, 1671, 1278, 1083, 217, 1385, 23, 819, 211, 1697, 468, 743, 829, 590, 1138, 1456, 107, 372, 1346, 504, 1495, 1135, 500, 450, 1062, 1796, 574, 1537, 349, 1026, 1414, 1567, 527, 1183, 128, 820, 1308, 1635, 1798, 1154, 106, 1163, 341, 486, 799, 1153, 1667, 1296, 1668, 1710, 1125]. Fishing [339, 421, 337, 143, 169, 1589, 354, 989, 632, 906, 1051, 749, 1101, 1686, 1422, 1205, 835, 1638, 1181, 1733, 700, 429, 306, 532, 849, 491, 1204, 1089, 301, 280, 479, 97, 1370, 13, 955, 231, 406, 1397, 1090, 509, 1359, 1515, 921, 80, 711, 1529, 1294, 266, 220, 1514, 1275, 1789, 1300, 2, 300, 1100, 387, 1601, 1442, 462, 937, 729]. Fishing-gear [354]. Fishing-induced [143, 97, 937]. fishless [1236, 1295]. FISHSELECT [254]. Fishway [685]. fishways [1311, 1435, 1281]. fitness [1230, 225, 950, 485, 363, 156]. fitness-related [950]. five [344]. fixed [878, 63, 1247, 1710]. fixed-precision [1247, 1710]. fjord [1609, 890, 529]. flame [1054]. flatfish [266]. Flats [977]. flava [1199]. flavescens [1096, 105, 1443, 1431, 295, 1017]. flavicans [1524]. flavidus [188]. fleet [1835, 892, 844, 13, 1839, 634, 414, 1601]. Flemish [1655, 1294]. flesus [359]. flexibility [1091]. flexible [370]. FLEXSELECT [1500]. floating [758, 1589]. Flood [1163]. flooding [1451]. floodplain [1669, 1736, 1213, 829, 589, 1069, 1030, 927, 737, 1452]. floods [135]. Florida [887, 368, 148]. flounder [47, 379, 935, 1176, 1742, 828, 1761]. flounders [359]. flour [1832]. Flow [138, 1317, 1593, 1669, 26, 808, 590, 1016, 150, 1696, 1708, 858, 1226, 46, 269, 1038, 665, 30, 641, 473, 6, 27]. Flow-mediated [1593]. flowing [996, 1715]. flows [135, 1042, 566, 520]. fluctuating [1171]. fluctuation [999]. Fluctuations [22, 640, 269, 1210]. flume [1369, 805]. fluorescence [1193, 365, 1063]. fluorometry [880]. FluoroProbe [1193, 810]. fluvial [1042, 1465, 1441]. fluvial-dependent [1042]. fluviatilis [1460, 620]. flux [369, 583, 57, 520]. fluxes [275, 1179, 194]. FOBs [1589]. follow [603, 1231, 1453]. following [1570, 970, 1344, 1263, 1822, 37, 353, 1757, 1243, 1467, 1132, 1426, 1280]. fonctionnement [469]. fontanae [896]. Fontane [896]. fontinalis [564, 404, 803, 428, 1598, 490, 1328, 1012]. Food [644, 1035, 4, 968, 1119, 712, 525, 1122, 556, 1762, 1640, 159, 1385, 1553, 129, 532, 1621, 1481, 589, 1801, 510, 164, 573, 1563, 1211, 1151, 1054, 972, 1209, 980, 809, 340, 1426, 755, 1208, 818, 1056, 517]. footgear [1442]. foothills [1295, 371, 324]. Forage [437, 133, 417, 1512, 1560, 1492, 834, 1703, 817, 1077]. foragers [351]. Foraging [970, 701, 800, 1704, 333, 1385, 1129, 1553, 1711, 981, 347, 1277, 1062, 1186, 1253, 806, 776, 1151, 1509, 1040]. forced [1593]. forcing [253]. forecast [439, 933]. forecasted [1671]. Forecasting [1548, 55, 1355, 1146, 215, 252, 633]. forecasts [1760, 1390]. forefront [1412]. forensic [1583]. forest [99, 671, 1469, 46, 1210, 1347, 1641]. forested [138, 1621]. forests [1144]. Form [469, 93, 470]. formation [501, 1148, 1619]. forme [469]. formed [37]. forms [1338, 895]. formulations [8, 5, 73, 72]. forward [1796]. fouling [993]. foundations [1113, 1692]. four [1268, 1311, 1345, 1643, 168, 1456, 1469, 487, 1367, 1831, 1437]. Fourier [1027, 1701]. fractionation [47]. fractions [605]. Fragmentation [602, 1435]. fragmented [336]. frames [1729]. framework [1122, 675, 375, 1385, 534, 1687, 28, 531, 169, 1616, 1577]. frameworks [1252]. Francis [73]. Fraser [26, 6, 27, 239, 482, 435, 524, 497]. free [1762, 491]. free-ranging [1762]. French [1445, 1589, 70]. frequency [166, 195, 431, 1261, 1773, 1764, 917, 1562]. frequency-dependent [166, 195]. fresh [996, 1411, 1539, 1002, 975]. freshwater [556, 609, 1354, 1819, 1784, 898, 474, 159, 12, 578, 1013, 1014, 177, 1199, 770, 631, 1592, 846, 1842, 944, 451, 1796, 1465, 785, 1558, 522, 128, 106, 1243, 1209, 1741, 1703, 1100, 1292]. friendly [1233]. frog [1251, 31, 324]. frogs [683, 1130]. front [1487]. frontiers [1722]. fry [845, 848, 1488, 1156, 50]. fugitive [109]. Fukushima [873, 1478]. fullness [1564]. fully [445]. fulmars [43]. Fulmarus [43]. fulvescens [1506, 1640, 1731, 233, 362, 1330, 186, 1769]. function [368, 136, 317]. functional [1302, 1565, 100, 532, 383, 840, 152, 990, 236, 1426]. Functioning [469, 470]. functions [804, 1242]. Fundulus [133, 1743]. Fundy [1226, 916]. furnieri [1510]. further [741]. Fusconaia [1199]. future [1342, 619, 1408, 733, 278, 926, 933].

Gadget [1294]. Gadus [1420, 648, 909, 1459, 81, 1701, 141, 61, 231, 406, 1262, 693, 1746, 109, 866, 1029, 1652, 383, 51, 834, 1196, 437, 366, 134, 815, 399, 608]. gag [987]. gain [708]. gains [667]. galeata [513]. Galeocerdo [82]. gammarids [556]. gammarus [278, 1371]. Gape [904]. gaps [1245, 1457]. gas [1759, 1167, 707]. Gaspésie [771]. gauges [474]. Gaussian [1177, 1616]. gavage [1280]. gear [878, 1496, 1238, 1654, 1813, 461, 509, 663, 949, 2, 300, 354]. gear- [461]. gears [1350, 1239, 955]. geminata [501, 964, 1148, 771]. Gene [26, 6, 27, 860, 1016, 141, 665, 30, 399]. general [83, 1462]. generalist [481]. generalized [1059, 669, 792, 828]. Generation [255, 1801]. Generation-scale [255]. generational [1137]. generations [431, 170]. generic [1125]. genes [802]. Genetic [894, 198, 1735, 825, 155, 84, 1147, 950, 1499, 425, 742, 864, 1533, 976, 1270, 1284, 1570, 226, 1362, 1520, 738, 1507, 189, 612, 336, 562, 1137, 1540, 895, 1016, 92, 552, 188, 568, 994, 778, 1144, 702, 1573, 1536, 349, 971, 1329, 1279, 1682, 1025, 713, 272, 522, 260, 330, 498, 1831, 19, 295, 131, 1023, 134, 1734, 126, 1292, 602, 863, 1383]. genetic-based [552]. Genetically [227, 982, 1673]. genetics [181, 522, 186, 1009, 483, 98]. Genome [1573, 1549]. Genome-wide [1573, 1549]. Genomic [1595, 29, 1549, 1274]. genomics [1812]. genotype [412]. genotypes [399]. genotyping [778, 1279]. genotyping-by-sequencing [778]. genus [1735, 320]. geochemical [709]. geochemistry [1583, 922, 1196]. Geographic [205, 39, 915, 722, 1576, 902, 48]. geolocation [1400]. geomorphic [1587]. geophysical [963]. Georges [75, 1617, 1569]. Georgia [1594]. Georgian [1585, 618, 1366, 262]. geospatial [650, 1252]. geostatistics [417]. Geotria [1380]. gestion [1387]. Getting [1207]. ghost [1660]. gibbosus [970]. gigas [920]. gill [65]. gillnet [1604]. gillnets [679, 1247]. Gironde [1534]. give [983]. gizzard [258, 660]. glacial [1832, 833, 1499, 104, 1527, 1209]. glacial-fed [104]. glacialis [622, 43, 354]. glaciated [529]. glaciation [85]. Glacier [37]. glass [1281]. global [169, 410]. globally [1741]. Globicephala [1305]. glycolytic [1771]. goals [1651]. gobies [1325, 575]. gobiids [1800]. goby [403, 1803, 395, 55, 1766]. going [356]. goldfish [1289]. gonad [141]. good [780]. gorbuscha [571, 182, 175, 1357, 482, 434, 201, 435, 268, 1511]. Gorges [1437, 1829]. gouldi [1027]. governing [437]. GPS [1589]. GPS-monitored [1589]. gradient [454, 1244, 1169, 674, 67, 945, 603, 1368, 365, 70, 351, 1740]. gradients [303, 1534, 529, 1032, 941, 1829, 1283, 521, 570]. Grand [511, 1067, 1816, 1068]. grandis [159]. Grass [801]. gravel [986, 797]. gravel-bed [797]. gravels [316, 285]. gray [1538]. grayling [1377, 1704]. Great [646, 1579, 826, 972, 1574, 1734, 1366, 1065, 1439, 956, 1096, 167, 1415, 875, 1317, 1548, 586, 1406, 1029, 527, 1435, 442, 32, 801]. greater [1057, 907]. green [805, 1494]. Greenland [982, 580, 197, 124, 784, 1284]. Grey [1489, 1181]. grid [1748, 1662]. ground [74, 1456]. groundfish [1049, 1146, 253, 693, 1082, 1000, 815, 1455, 1719]. grounds [385, 1076, 266, 1036]. groundwater [1230, 1211, 603]. group [100]. grouper [646, 1793]. groupers [987]. groups [60, 1791, 359]. growing [1793, 667, 94]. Growth [1793, 1173, 615, 376, 582, 1746, 1743, 1195, 1203, 1302, 1338, 503, 1591, 728, 102, 35, 1358, 1421, 1116, 335, 721, 830, 606, 1475, 1337, 836, 1626, 64, 200, 1269, 1020, 154, 1374, 1415, 812, 1074, 429, 1244, 495, 621, 766, 893, 1692, 17, 1431, 218, 1483, 1479, 69, 1080, 1750, 1590, 378, 690, 1288, 367, 304, 1550, 1813, 751, 484, 132, 1563, 1030, 1485, 851, 806, 482, 1211, 353, 434, 1194, 1617, 1449, 87, 559, 1799, 1197, 1029, 1828, 237, 660, 682, 1490, 1078, 1488, 363, 1498, 459, 490, 1437, 1000, 1073, 1313, 2, 300, 184]. growth [773, 884, 326, 1091, 94, 818, 33, 1068, 1056]. growth-based [1313]. Guiana [1445]. guidance [822]. Guilty [1559]. Guinea [296]. Gulf [682, 86, 1464, 646, 855, 1605, 1396, 1394, 987, 1143, 1638, 306, 669, 67, 231, 406, 1779, 434, 1610, 1348, 1327, 834, 839, 1584, 1772, 712, 1538, 397, 366, 1688, 134, 815, 101, 1036]. guns [461]. Gwaii [1240]. Gyrodactylus [19].

H [1419]. Habitat [1508, 882, 1517, 1347, 1584, 925, 503, 1506, 930, 609, 1578, 912, 250, 242, 1738, 238, 1694, 12, 1496, 1004, 167, 1457, 1265, 171, 666, 1470, 743, 1800, 133, 1319, 380, 1751, 777, 846, 1765, 833, 1277, 252, 1168, 172, 42, 674, 530, 858, 1135, 1602, 269, 955, 451, 1330, 1532, 595, 1221, 1477, 246, 356, 1633, 1197, 1015, 1416, 927, 1298, 210, 827, 41, 1301, 106, 1209, 1598, 1156, 139, 520, 838, 641, 767, 471, 1645, 752, 463, 533, 62, 1789, 1516, 1423, 326, 1375, 506, 1726, 818, 344, 678, 1494, 978, 1718, 1761, 1125]. habitat- [1765]. Habitat-specific [1347, 451]. habitats [314, 1582, 168, 944, 234, 1699, 57, 1190, 1233, 1008, 1332, 1291, 317]. habits [1382, 40]. haddock [1033, 1459, 1748, 80, 254, 632]. Haida [1240]. hake [391]. half [1203]. half-century [1203]. Halibut [982, 580, 197, 1825, 784, 599, 1284, 1834]. halieutiques [257, 1102, 1387]. Haliotis [69, 40]. handle [804]. handling [208, 1822]. hanging [120]. hannai [40]. haplochromine [1559, 1117]. haplotypes [1279]. happy [405]. haptophyte [472]. Harbour [1322]. Hard [1255]. harengus [1685, 292, 11]. harmful [409]. Harvest [951, 99, 1805, 918, 1131, 499, 1343, 1462, 1086, 612, 1378, 1720, 1443, 1794, 1005, 515, 1324, 286, 516, 1760, 753, 22, 1092, 730, 350, 1333, 1455, 1580]. harvested [205, 531, 433, 714]. harvesting [1448, 667, 1150, 1093, 313, 44]. hastens [1045]. hatch [1074, 1194, 1811]. hatcheries [1349]. Hatchery [593, 1060, 29, 35, 223, 584, 572, 1378, 777, 1127, 885, 845, 552, 1607, 331, 568, 670, 352, 1253, 1526, 358, 1597, 546, 731, 791, 480, 1674, 883, 260, 1276, 1052, 637, 1003, 1501, 50, 1048, 1783, 1555, 607, 802, 156, 709, 873]. hatchery- [1127]. hatchery-cultured [572]. hatchery-produced [607]. hatchery-reared [223, 584, 885, 568, 670, 352, 1597, 731, 791, 480, 1276, 1003, 1501, 802, 873]. hatchery-supplemented [1052]. hatching [1653, 1687, 1167, 464, 44]. Hawaii [705]. Hawaii-based [705]. Hawaiian [1203, 1793, 1475]. Hawke [1370]. HCR [1092]. HCRs [1086]. head [1523]. headwater [138, 135, 1621, 803, 1469, 46, 1419, 24]. headwaters [1236]. health [1776, 788]. Hearing [1764]. heart [1553]. Heat [992, 1250]. Heat-induced [992]. Heavy [1564]. held [1071]. helping [989]. Hemimysis [1132].

Hemisphere [1083]. hemoglobin [399]. hemorrhagic [1138]. Henry [1103]. henshawi [408]. hepatopancreas [1280]. herbivorous [1136]. herding [1500]. Heritability [1702]. heritable [1264]. hermaphrodites [662]. herring [1180, 1728, 928, 1147, 1138, 292, 1075, 11, 158, 740, 664, 594, 1568, 1521, 1348, 1685, 1531, 975, 959, 978]. heteroclitus [133, 1743]. heterogeneity [393, 1712, 1269, 1594]. heteroscedasticity [995]. heterotrophic [243, 924]. hidden [878, 1400, 282]. Hierarchical [18, 1680, 136, 1525, 524, 444, 1381, 1767, 1340, 1160, 69, 1788, 1824, 663, 1706, 1742, 174, 433, 377, 1647, 799, 1153, 1189, 1046]. hierarchies [326]. hierarchy [1038]. higginsii [1657]. High [1261, 385, 1416, 477, 1473, 399, 1543, 835, 1638, 1223, 1711, 362, 479, 1637, 1226, 1190, 984, 1048]. high-effort [1543]. high-elevation [1223]. high-flow [1226]. High-frequency [1261]. high-latitude [1223]. High-resolution [1416, 399, 835]. higher [823, 364, 1436]. higher-order [1436]. highly [742, 1473, 677, 1126]. hind [646]. hinder [422]. hippoglossoides [580, 197, 784, 1284]. Hippoglossus [1825, 1834]. histocompatibility [412]. Historic [206]. Historical [424, 666, 222, 240, 1545, 653, 647, 1742, 1484, 1100, 98, 726, 273]. historically [541, 953]. histories [857, 1335, 656, 760, 96, 1709, 616]. history [1448, 512, 519, 1365, 1081, 821, 813, 205, 1802, 624, 217, 1378, 1202, 446, 819, 1643, 400, 1697, 1074, 1542, 322, 1016, 460, 1227, 1778, 869, 39, 1409, 1518, 807, 733, 1766, 38, 230, 577, 826, 1234, 723, 1260, 382, 827, 41, 957, 1831, 459, 1145, 490, 1123, 1649, 730, 1105, 718, 1562, 50, 410, 799, 1153, 126, 1283, 607, 1232, 1341, 462, 1627]. Hjort [1071, 1072]. hoc [939]. hoferi [158]. hold [1776]. holding [1301]. Homarus [695, 1666, 1505, 916, 1779, 693, 1194, 278, 1610, 1371, 1497, 689, 1688, 788]. Home [278, 1637]. Homing [613, 1556, 1414, 866, 1405]. Hommage [213]. homogeneity [155]. homogeneous [982]. honneur [1102]. honour [1102, 1104]. Hood [609, 1323]. hook [1654, 762]. hook-and-line [1654]. hooking [891]. Horizon [397]. horizontal [1266]. hormonal [843, 943]. host [481, 281, 1608]. Hot [1379, 781]. Hotspots [1339, 1163]. Hudson [1427, 817]. hull [993]. Human [494, 898, 1529, 840, 472, 48]. human-modified [840]. humic [701]. hump [1097]. hump-shaped [1097]. humpback [507, 1067]. humped [904]. Huron [1432, 1060, 1585, 1158, 889, 1160, 1461, 1563, 1139]. Husky [1518]. hybrid [1434, 526]. Hybridization [624, 256, 229, 777, 1545, 331, 807, 428, 965, 41, 1695, 985]. hybrids [1801]. hydraulic [1587, 1271]. hydraulics [1088]. Hydrodynamic [964, 488]. hydroelectric [765, 770, 353]. hydrogen [340]. Hydrogeomorphic [62, 930]. hydrographical [1534]. Hydrologic [980, 1705, 1452]. Hydrological [859, 1069]. hydrologie [469]. Hydrology [469, 733, 1183, 470, 85, 1428]. hydropeaking [776]. hydropower [1342]. hyperstability [318, 716, 684]. hypolimnetic [1833]. hyporheic [808, 986, 37]. Hyporthodus [1793]. Hyposmesus [284]. hypotheses [306, 1154]. hypothesis [122, 185]. Hypoxia [528, 250, 580, 1323, 1167, 1314, 1743]. hypoxia-tolerant [1314]. hypoxic [955, 890].

ice [147, 680, 633]. Iceland [648, 1521, 383]. Icelandic [1420, 124]. ichthyophoniasis [1521]. Ichthyophonus [1690, 158]. ichthyoplankton [1625]. iconic [1675]. ICP [1293, 1232]. Idaho [876]. ideal [491]. ideas [856]. Identification [1549, 67, 995, 355, 113, 1362, 1520, 1507, 1340, 1540, 685, 778, 1279, 1682, 965, 128, 1270, 616]. identified [778]. identifies [991]. identify [1465, 260, 959, 506]. Identifying [1644, 216, 1184, 432, 320]. identity [1541, 953]. ignoring [348]. II [1273, 830]. III [990]. illegal [1583]. Illinois [1155]. illusion [318]. illustrated [284]. image [139]. immature [330, 1770, 142]. Immersion [947]. immobility [997]. immune [802]. immune-relevant [802]. Impact [64, 580, 1074, 1637, 509, 1029, 865, 397, 1349, 1665, 857, 813, 1321, 700, 781, 1800, 1622, 1624, 451, 1282, 1235, 1162, 1141, 1829, 1455, 801, 937]. impacted [783, 1810]. Impacts [651, 395, 681, 378, 1348, 314, 780, 1201, 217, 478, 1147, 239, 329, 169, 1408, 1468, 284, 1093, 1789, 1786, 488, 1688, 1171, 1458]. impaired [783]. impairment [1307]. impeding [442]. imperiled [240, 931]. imperilled [1796]. implantation [997]. implausible [1535]. implementation [1829]. Implications [1034, 153, 649, 1827, 615, 421, 1672, 333, 77, 1290, 547, 20, 291, 666, 1047, 1713, 93, 893, 84, 638, 619, 239, 1089, 523, 42, 1409, 1152, 539, 1222, 732, 793, 1304, 246, 764, 278, 665, 954, 921, 1347, 1527, 1167, 1772, 524, 1714, 1032, 1631, 1440, 340, 1782, 1073, 565, 1489, 1555, 936, 447, 1769]. implicit [748]. importance [403, 993, 1248, 259, 177, 1168, 46, 1242, 623, 577, 1165, 289, 32, 341, 351, 145, 953]. Important [545, 1096, 1694, 1320, 681]. impossible [1078]. impounded [618]. impoundment [1260]. impoundments [1698]. improve [1487, 1430, 427, 322, 778, 1615, 862, 402]. Improved [1193, 837, 1744, 499, 1265, 1424, 1746, 1670]. improvement [221]. improves [1338, 1394, 200, 670, 679, 1501, 525, 1482]. Improving [1686, 1272, 1273, 1459, 774, 1285, 1691, 623, 1218, 1815, 296, 1567]. imputation [1503]. imputations [1359]. In-season [1725]. in-stream [299, 78]. inadequacy [391]. inbred [225]. incidence [1301]. incidental [1049, 1390]. Including [311, 1449, 614, 1085, 474, 253, 774, 634, 1529, 1224]. inclusion [1728]. incomes [1422]. incomplete [983, 96]. inconsistent [1026]. incorporate [1546]. Incorporating [1388, 191, 1269, 69, 118, 594, 1298, 497, 1387, 1507, 1729, 1392, 75, 1227, 1522, 891, 1813, 1176, 1577, 486, 309, 178]. Incorporation [1467, 1641, 962, 164, 121, 1742]. incorrect [1139]. increase [1349, 1257, 1590, 1730, 998, 1165, 811, 363, 316, 285]. Increased [1552, 1545, 1839, 1424, 290, 1670, 1364]. increases [1331, 1561, 26, 1161, 35, 337, 885, 643, 1133, 1090, 57, 1279, 6, 27, 1726]. Increasing [1504, 1792]. increasingly [492]. increment [512]. Incubation [1171, 1749]. independent [216, 1460, 614, 1732]. Index [179, 779, 1394, 1480, 1084, 1429]. Indexing [1507, 757]. Indian [198, 425, 786]. indicate [403, 875, 1160]. indicates [1752, 1292]. Indicator [779, 1830]. indicators [780, 1774, 1401, 443, 342, 706, 490, 1780, 1554]. indices [144, 123, 18, 1680]. indigenous [436]. indirect [429, 137, 164, 356]. indiscriminate [1119]. Individual [362, 43, 160, 569, 641, 760, 453, 749, 1101, 113, 1362, 1520, 1572, 248, 191, 217, 1169, 332, 1266, 283, 1779, 1197, 237, 711, 883, 827, 186, 788, 1040]. individual- [1169]. Individual-based [362, 453, 749, 1101, 1572, 191, 217, 332, 283, 1779, 186, 788]. individuals [1733, 1449, 1179, 340, 543]. induce [149]. induced [1365, 143, 386, 373, 97, 1523, 1167, 165, 170, 1200, 1091, 992, 1554, 937]. Induction [1012]. industrial [849, 71, 1552]. ineffective [1205]. infection [368, 1564, 201, 1150, 19]. Infections [1608, 65]. infectious [1297]. infer [878, 1681]. inference [705, 1682, 410, 467]. inferences [418]. inferred [1420, 938, 979, 971, 104, 398, 483, 98]. infestations [373, 182, 165]. Inflow [469, 474]. Influence [1832, 1088, 1642, 627, 629, 870, 1069, 349, 310, 723, 934, 1685, 767, 21, 1843, 163, 412, 802, 636, 1302, 609, 1838, 1660, 140, 986, 847, 1343, 1565, 1731, 1124, 1202, 1178, 1713, 1696, 515, 172, 1185, 1144, 1446, 286, 516, 807, 183, 1752, 713, 1150, 999, 660, 522, 435, 1167, 834, 980, 1645, 760, 730, 85, 1250, 1524, 985, 59]. influenced [1470, 1277, 1518, 1361]. Influences [1817, 1649, 920, 930, 600, 821, 393, 496, 380, 1751, 1620, 631, 736, 1608]. influencing [334, 1357, 1513, 683, 1476, 1006, 156]. influential [588, 1084]. influx [1076]. inform [1656, 129]. information [28, 766, 797, 1788, 1299, 1615, 63, 290, 1661, 1780, 486]. informing [1738]. infrared [1701]. infrastructure [1668]. ingest [1289]. ingredients [227]. ingress [460]. inhabiting [1831, 1533]. Initiation [1002]. injuries [310, 626]. injury [1837, 762]. inland [369, 1543, 1345, 1802, 856, 98]. Inlet [890, 1711]. inner [1226]. inorganic [1201]. input [1793, 130, 161, 1367]. inputs [888, 977, 573, 1670, 110]. Insecta [303]. insects [717]. inserted [1487]. inshore [1143, 1001, 1688]. Insights [1536, 642, 1382, 1129, 1711, 948, 431, 451, 1447, 800, 1063, 214, 605]. Instability [591]. installation [1545]. instars [775, 1524]. instrument [810]. insufficient [1026]. intake [1159, 1306]. intakes [173]. intégrant [1387]. Integrated [664, 1809, 131, 1660, 1245, 1546, 935, 1336, 552, 292, 672, 511, 1405, 636]. Integrating [1728, 1282, 1558, 1677, 749, 1101, 125, 544, 169, 1675, 306]. integration [1389, 1738, 1626]. intelligence [861]. intense [1770]. Intensified [1427]. intensity [906, 1666, 851]. intensively [1660, 803, 772]. Inter [854, 1381, 259, 694]. inter- [694]. Inter-annual [854]. inter-cohort [259]. inter-reach [1381]. interact [924]. Interacting [1541]. interaction [1640, 1743]. Interactions [334, 1423, 436, 379, 1430, 1337, 1182, 119, 1287, 1654, 1185, 185, 702, 870, 1098, 943, 66, 109, 485, 1529, 1157, 1134, 596, 1361, 528, 1598, 676, 1623, 1538, 1703, 1364, 678]. Interactive [1278, 1471, 1251, 1097, 1428]. interacts [1225]. Interannual [1669, 466, 1099, 1156, 1612, 1073, 1666]. interbreeding [149]. interception [1462]. intercohort [1168, 190]. intercompatibility [810]. interconnected [740]. Interface [1388, 1387]. interference [150]. interior [604, 302, 261]. intermittency [473]. intermittent [1491]. Internal [1411, 1366]. international [1205]. interplay [481, 1124, 665]. interpretation [1686, 1237, 657]. Interspecific [756, 845]. intersystem [1096]. intertissue [82]. interval [1086]. intervals [244]. interviews [1613]. intra [845, 854, 694]. intra- [845]. intra-annual [854]. intra-specific [694]. Intracohort [190]. Intrafragment [931]. intraspecific [1108]. intrinsic [952, 649, 1828, 1165]. Introduced [718, 508, 1471, 590, 1696, 761, 751, 22, 41, 1745, 1551, 236]. Introduction [1388, 1583, 1777, 409, 32]. Introgressive [331, 1545]. invaded [492]. invader [1439]. invasion [970, 1132, 817, 1117]. invasions [207, 163]. Invasive [1723, 1214, 575, 556, 675, 630, 1267, 1438, 644, 1045, 1436, 530, 504, 1648, 55, 451, 966, 66, 1766, 1493, 1053, 764, 1798, 1163, 1064, 1314, 274, 302, 261, 992, 416]. invasiveness [625]. invertebrate [1587, 496, 1708, 1179, 1347, 818, 1472]. invertebrates [279, 1309, 574, 371, 1745, 351]. investigate [1583, 1510]. Investigating [1572, 1711, 1654, 740, 828]. investigation [228]. investigations [1813]. investment [335]. involve [122]. involving [565, 1383]. ion [201]. Ireland [386, 114]. Iron [501, 388, 1781, 629, 1366]. iron-binding [1781]. Island [407, 390, 1190, 689, 103, 1259, 1325]. Islands [1203, 337]. isodar [491]. isodars [729]. isoforms [1539]. isolated [139]. Isolating [1038]. Isolation [193, 188, 969]. isolation-by-distance [969]. isoscapes [448, 430]. isotope [1354, 1278, 875, 849, 342, 401, 1419, 1510, 340, 947, 1232, 416, 40]. isotopes [403, 47, 563, 474, 551, 1737, 1560, 1492, 1467, 750, 79, 145, 535]. Isotopic [1263, 1447, 1382, 281, 1158, 555, 82, 750, 192, 1363]. Israel [283]. issue [469, 1071, 1777]. issues [257]. istiophorid [932]. Isurus [1393]. iteroparous [77].

jack [1073]. Jackson [516]. Jacobson [686]. Japan [1029, 1624, 219]. Japanese [742, 1313]. japonicus [1073, 1313, 219]. Jasper [298]. Jasus [429]. jenkinsi [240]. jeopardize [746]. Johan [1071, 1071, 1072]. John [464, 1632, 1634]. Jones [374]. jordani [538, 407, 390]. Journal [257, 257]. Juan [469]. jubatus [824]. July [172]. jumbo [920]. justice [1721]. Juvenile [1030, 1197, 805, 901, 1145, 1379, 650, 1767, 606, 105, 572, 117, 783, 154, 465, 666, 1257, 1220, 1542, 1564, 938, 621, 1751, 288, 1168, 950, 848, 1491, 352, 1801, 944, 373, 234, 787, 492, 190, 1288, 751, 680, 870, 943, 175, 392, 482, 822, 1504, 201, 1151, 927, 999, 383, 1593, 165, 374, 1732, 1209, 949, 520, 838, 569, 641, 490, 1196, 1291, 842, 1073, 184, 1423, 1322, 1055, 975, 953, 326, 1306, 346, 1632, 1634, 447]. juveniles [1358, 259, 1826, 1501].

Kajikia [330]. Kalama [1253]. Kansas [59, 21]. Kariba [1710]. kasmira [425]. Katsuwonus [198, 758]. kelts [181, 1335]. kernel [1658, 402]. key [1108, 631, 119, 1691, 1298, 587, 268]. keystone [1203]. kHz [426]. kidney [65]. kindtii [1314]. King [291, 421, 572, 484, 1597]. Kinneret [283]. kisutch [1749, 925, 1836, 847, 1619, 439, 116, 722, 394, 1291, 50, 506, 802]. Klamath [135]. Klickitat [331]. knowledge [1535, 296, 1754, 594, 1760, 1577, 1675]. known [1420, 758]. Kokanee [487, 1025, 1037]. krill [103].

LA-ICP-MS [1293]. labile [364]. laboratory [1263, 1485, 805, 1467, 1538]. laboratory-calibrated [1485]. Labrador [1370, 1533]. Lack [1204]. lactate [1186]. lactating [800]. ladder [1545]. lagoon [1619]. lagoons [70]. Lagrangian [1530]. lags [900]. Lahontan [1672, 408]. Lake [1506, 956, 65, 1585, 249, 1011, 958, 1631, 1060, 1640, 1050, 1136, 203, 258, 1109, 235, 1174, 1731, 233, 1158, 1553, 418, 362, 743, 1540, 754, 1319, 380, 1432, 388, 1461, 1614, 178, 118, 1750, 1518, 1590, 787, 1330, 1185, 1596, 966, 1465, 283, 1730, 15, 38, 1811, 1460, 826, 104, 264, 1600, 157, 384, 934, 1093, 361, 442, 1099, 1008, 186, 236, 541, 1229, 718, 79, 350, 1171, 243, 755, 1428, 475, 1063, 1755, 214, 1111, 317, 899, 1247, 1769, 1478, 1112, 228, 1805, 1060, 250, 1268, 1803, 1303, 381, 279, 819, 1158, 1720, 1443, 388, 360]. Lake [889, 1160, 523, 1431, 245, 55, 955, 1152, 1563, 1808, 283, 1559, 861, 1560, 1492, 919, 128, 1054, 528, 1466, 365, 1650, 972, 1139, 809, 295, 490, 1126, 265, 941, 1703, 1229, 49, 772, 1667, 1710, 832, 1017, 1117]. lake-wide [388]. Lakers [32]. Lakes [600, 780, 1108, 977, 1832, 1830, 1173, 615, 1131, 235, 1781, 1771, 719, 1817, 494, 411, 1622, 1319, 1433, 638, 940, 833, 207, 298, 629, 1295, 172, 137, 1461, 1753, 178, 118, 874, 422, 71, 293, 1737, 785, 1477, 1069, 1537, 303, 1107, 554, 1513, 1633, 1552, 371, 1833, 1527, 477, 264, 1298, 114, 865, 1215, 1301, 1551, 1628, 1057, 907, 324, 657, 1032, 633, 1426, 794, 1189, 707, 194, 398, 1428, 1063, 48, 1384, 521, 570, 1031, 1439, 1096, 1415, 875, 1432, 1548, 586, 1518, 1579, 880, 826, 527, 442, 1574, 32, 1734, 1366, 801]. Lamna [333, 1339]. Lampetra [355]. lamprey [1380, 675, 1311, 790, 1438, 28, 1047, 867, 1045, 1548, 643, 1021, 685, 591, 764, 1414, 1764, 704, 1417, 1804, 1229, 1734, 224, 1208, 1383, 616, 355]. lampreys [1316, 1053]. Lampsilis [1657]. land [112, 1206, 1110, 1671, 334, 454, 1751, 833, 1699, 323, 1215, 1291, 194, 816, 1224]. land-use [112, 323, 1291]. landings [410]. landlocked [1778, 1533]. Landscape [738, 393, 1716, 1527, 1831, 394, 1749, 334, 857, 1155, 913, 1016, 940, 1287, 1567, 840, 1298, 1598, 262]. landscape-level [1155]. Landscape-scale [1716, 1527, 1749]. landscapes [1291, 1782]. Large [855, 609, 746, 353, 134, 1669, 1793, 519, 29, 1557, 1059, 1136, 710, 1694, 835, 847, 604, 630, 1680, 64, 1129, 1643, 844, 431, 661, 196, 342, 1528, 1753, 1266, 1517, 269, 1185, 1821, 1030, 1513, 1606, 104, 1157, 1576, 614, 792, 767, 463, 1770, 543, 1625, 1270, 1159, 1474, 521, 570]. large- [844]. large-bodied [1753]. Large-scale [855, 609, 353, 519, 1557, 1694, 835, 1266, 1606, 1576]. largemouth [376, 1343, 582, 1433, 237, 1718]. Larger [1730]. larvae [695, 1074, 1540, 1369, 148, 523, 185, 246, 1313, 1444, 1441]. Larval [1096, 1368, 488, 173, 775, 1202, 233, 1374, 1033, 1666, 1594, 1751, 1316, 1431, 1299, 1325, 1465, 1446, 1406, 183, 806, 466, 682, 1376, 91, 90, 1445, 463, 1440, 1437, 1829, 730, 750, 1073, 1843, 1524]. laser [842, 1232]. lasting [1575]. late [1454]. late-run [1454]. Latent [1177]. lateral [1452]. Lates [95]. latitude [1331, 1561, 1223, 723]. Latitudinal [653, 484, 87, 1169, 145]. Laurentian [1579, 1523, 826, 972, 1574, 1734]. law [1247, 1710]. Lawrence [855, 872, 1344, 395, 489, 682, 855, 1605, 1396, 1638, 231, 406, 1610, 1534, 712, 366, 134]. layer [381, 919]. Layered [1033]. layers [820, 513]. lead [512, 154]. leads [1037]. Leaf [496, 299, 1251]. lean [754]. learned [1421, 1772]. leftovers [1191]. legacies [1291]. legacy [1023]. legged [31]. length [1122, 1338, 671, 1717, 1697, 75, 455, 1160, 1748, 1227, 1121, 1662, 664, 80, 220, 1514, 1645, 617, 768, 1562, 585]. length- [768]. length-at-age [80]. length-based [1121, 664, 220, 1514]. length-dependent [1748]. length-structured [1227]. lengths [677]. lengthy [656]. lentic [1412, 90]. Lepeophtheirus [281, 182, 175, 268, 1231, 1453]. Lepidopsetta [17]. Lepomis [325, 970, 241]. leptocephali [1326]. Leptodora [1314]. less [1433, 1424, 1314]. Lessons [1772, 420, 1421, 192]. Lethal [1657, 1684, 354]. Lethenteron [1804, 616]. leucas [537, 82]. leucomaenis [150]. Leucoraja [75]. level [1155, 743, 532, 668, 1062, 1397, 1424, 1053, 1567, 934, 827, 267, 1159, 1474, 317]. Levels [1253, 1081, 454, 364, 593, 1093, 794, 959]. levis [724]. lewisi [448, 430, 41]. lice [281, 386, 981, 182, 175, 201, 374, 1231, 1453]. lice-induced [386]. LIDAR [474, 1312]. Life [821, 1802, 217, 1378, 1643, 400, 1016, 1778, 1801, 38, 230, 1260, 1448, 512, 1203, 920, 115, 519, 1365, 1838, 1081, 857, 813, 205, 624, 1202, 446, 1697, 1074, 1542, 322, 455, 460, 1227, 869, 39, 1495, 1409, 1518, 270, 831, 733, 284, 1766, 485, 577, 826, 1234, 1315, 1661, 723, 1335, 68, 827, 41, 656, 957, 1831, 459, 1145, 490, 1123, 760, 1649, 730, 718, 1562, 50, 410, 799, 1153, 126, 1283, 607, 1232, 1709, 616, 1341, 462, 1627]. life-cycle [1661]. Life-history [1802, 1016, 1778, 1365, 1081, 813, 446, 1074, 322, 577, 826, 459, 1145, 730, 1562, 1283]. life-span [1203]. lifestyle [89]. lifetime [1357, 1188]. ligands [1781]. light [208, 1032, 941, 521, 570]. like [11]. likelihood [1310, 319, 2, 300]. Limanda [935]. Liming [199]. limit [564, 627, 869, 998, 1333]. limitation [876, 204, 904, 70]. limitations [231, 406]. Limited [1665, 1295, 371, 1426, 142, 1321, 1794, 1602, 1550, 1661, 1092, 1747, 1514, 1772, 1562, 1423, 1208, 937]. limiting [1056]. limits [1129, 1223, 966, 713, 128, 254]. limnological [977, 1692, 48]. Limnology [469, 1069]. Lindegren [686]. line [1392, 1654, 1406, 969, 632]. lineage [983]. lineages [331, 1511]. linear [669, 727, 533]. lines [1591, 120]. ling [1407]. linkage [1483]. linkages [1059, 1587, 1602, 1144]. linked [1781, 466, 1243]. Linking [1502, 719, 1312, 1456, 945, 1211, 883, 1724, 462, 652, 631, 1062, 1078, 1142, 827, 493]. lions [824, 800]. lipid [1827, 1278, 1128, 211, 1614, 358, 527, 1685]. lipids [145]. liquid [1289]. listed [1526, 752]. litter [299, 905, 1251]. little [1698]. littoral [1565, 380, 371, 755, 317]. live [154]. lived [512, 766]. liver [1084]. living [492, 963, 483]. Livingston [103]. loading [1411, 1296, 1366, 108]. loadings [865]. loads [854]. Lobster [687, 1497, 688, 695, 368, 16, 191, 1638, 277, 429, 1666, 1727, 1505, 916, 691, 690, 740, 1779, 998, 693, 1194, 278, 1610, 694, 1371, 1376, 697, 692, 788, 655, 678]. lobsters [1143, 696, 502, 689, 678]. Local [1666, 894, 857, 1109, 296, 1754, 63, 665, 441, 264, 289, 1484, 266, 1498, 1440, 1423, 1208]. Localization [911]. Location [1604, 362, 1045, 1839, 1515, 727, 1417, 410]. Location-specific [1604]. locations [1687, 174, 359, 1494]. locomotor [95]. loggerhead [887]. logging [1236, 100]. logistic [747]. logistic-based [747]. logs [758]. Long [508, 100, 537, 1727, 360, 141, 1599, 1596, 1198, 1221, 1833, 1076, 1757, 1490, 1625, 512, 1421, 120, 135, 1248, 1666, 766, 1199, 644, 479, 451, 1493, 353, 1600, 1260, 1093, 1099, 1575, 689]. long-lasting [1575]. long-lines [120]. long-lived [512, 766]. Long-term [508, 100, 537, 1727, 360, 141, 1599, 1596, 1198, 1221, 1833, 1076, 1757, 1490, 1625, 1421, 135, 1248, 1666, 1199, 644, 479, 451, 1493, 353, 1600, 1260, 1093, 1099]. longevity [646, 1425, 1793, 1475, 202]. longfin [1001]. longimanus [245, 1314, 48]. Longitudinal [829, 1829, 1452, 877, 178, 118, 1472]. longline [705, 1694, 835, 1496, 337, 13, 1509]. Lontra [125]. look [832]. loop [1795]. lose [908]. Loss [1570, 877, 1527, 1125]. losses [1265, 1666]. lotic [100, 1148, 164]. louse [1166, 1564, 268]. lousy [981]. Love [1274]. Low [405, 984, 1484, 674, 1042, 500, 450, 1053, 1764, 477, 1417, 1814, 1474]. Low-cost [405]. low-frequency [1764]. low-gradient [674]. low-level [1474]. low-nutrient [1814]. lower [713, 1792, 497, 1632, 1634, 1183, 682, 152]. lowland [871, 793, 68, 561, 1355]. lucius [1263, 848, 1292]. lupus [1416]. Lutjanus [117, 425, 1538]. lutrensis [146]. lutris [681]. Lyngbya [489]. Lynn [1180].

MacCall [686]. maccoyii [823, 854, 1200]. Mackenzie [84]. mackerel [742, 1073]. Macquaria [206]. Macrhybopsis [931]. macrocephalus [1029]. macrochirus [325, 241]. Macroecology [1648]. macroinvertebrate [1136, 138, 64, 46, 776, 623, 317, 24]. macroinvertebrates [1565, 100, 905, 14, 874, 1759]. macrophyte [323, 398]. macrorhynchus [1305]. macroscale [1813]. macrozooplankton [417]. magazine [1051]. magellanicus [1706, 1300]. magic [1258]. magister [1323, 1307]. magnify [1551]. magnitude [1213, 60, 1355]. Mahury [1445]. main [889, 1160]. Maine [1786, 1464, 1394, 1143, 67, 1779, 1327, 834, 1584, 1688, 688, 24]. mainstem [1379]. mainstream [717]. maintaining [1118]. Maintenance [229, 1375, 256]. major [20, 331, 1076, 1437, 412, 1588, 1501]. Makaira [1425]. makes [720]. making [1450, 1729]. mako [1393]. maladaptation [1676]. maladaptive [149]. male [1805, 166, 195, 1563, 617, 575, 841, 985]. male-biased [1805]. Mallotus [119, 682]. malma [1016, 1016, 976]. mammal [1345]. managed [892, 737, 772, 1580]. Management [153, 1388, 356, 1529, 1019, 1793, 1387, 209, 370, 1656, 1177, 1738, 1205, 1672, 789, 1774, 375, 856, 1775, 1729, 540, 54, 781, 1728, 28, 666, 759, 1323, 129, 1795, 1777, 1047, 1713, 93, 893, 942, 84, 638, 239, 1089, 252, 443, 1722, 1120, 643, 891, 1021, 1299, 169, 1408, 4, 517, 1152, 1185, 1465, 706, 1615, 1754, 1536, 1779, 1026, 764, 1725, 1617, 954, 965, 1577, 1675, 711, 1092, 1721, 1724, 1142, 1407, 1130, 414, 391, 439, 767, 1772, 19, 471, 739, 1714, 1531, 1595, 1789, 1009, 969, 1364, 219, 1284, 1719, 1569, 348, 1307]. Managing [1279, 929, 1794]. manganese [922]. mangrove [1445]. Manipulation [367, 25, 344]. Mantle [89]. mapping [1155, 1416, 1574, 1332]. maps [989]. marina [611]. Marine [1337, 1004, 979, 791, 89, 10, 153, 209, 918, 1331, 1561, 1838, 1578, 986, 1345, 856, 622, 217, 613, 117, 534, 1556, 200, 1309, 909, 429, 386, 322, 1178, 1512, 1147, 196, 342, 107, 515, 218, 1483, 1479, 1842, 567, 1121, 1240, 500, 450, 1370, 539, 1253, 690, 296, 1678, 286, 516, 401, 1026, 910, 482, 1752, 1391, 278, 866, 742, 711, 988, 1521, 1407, 1484, 529, 1243, 1209, 1018, 792, 767, 739, 1332, 1804, 1595, 1212, 1009, 969, 842, 341, 486, 799, 1153, 773]. marine-derived [10, 153, 986, 515, 286, 516, 1243]. marine-phase [1804]. marinus [790, 28, 1045, 1548, 643, 591, 764, 1764, 704, 1734, 224]. mark [376, 1249, 1522, 311, 1807, 483, 535]. marker [1303]. markers [226, 336, 3, 1070, 92, 1499, 383, 947]. Markov [878, 1400, 282]. marks [750]. marlin [1425, 76, 330]. marsh [474, 53, 349]. marshes [618, 133, 262]. Martin [516]. Maryland [263]. Marys [1021]. masks [318]. masou [87]. mass [1058, 671, 1697, 1673, 487, 992, 1056]. masu [845, 735, 87]. Match [185]. mate [562, 802]. material [1737, 980, 513]. Maternal [225, 485, 1631, 493, 813, 233, 950, 699, 1167, 730]. Maternally [750, 1646]. Mating [166, 467, 643, 412, 195]. matrices [417, 837]. matrix [551]. mats [964]. matter [996, 1610, 1568, 1419, 1063, 313]. matters [1613, 717, 864]. maturation [335, 1415, 141, 61, 87, 80, 1123]. mature [176, 7, 177]. maturity [1475, 836, 1443, 1505, 510, 197, 349, 1617, 1165, 1702, 617, 525, 94, 655, 828]. Maumee [614]. Maurolicus [1074]. mawsoni [519]. Maximizing [1118]. Maximum [319, 1165, 587, 542, 1129, 1661]. May [457, 1582, 1204]. MC [1232]. MC-ICP-MS [1232]. McAllister [457]. McClatchie [686]. McMurdo [473]. meadows [611]. mean [1239, 316, 285, 1651]. means [281]. measured [235, 941]. measurement [208]. measurements [818]. measures [810, 921, 396, 1437, 1111]. Measuring [1349, 1233, 1709, 952, 793]. Mechanics [1704]. Mechanisms [418, 699, 404, 679, 1356, 1313]. mechanistic [652, 891, 554, 716, 684]. mediate [1508, 1600]. mediated [695, 556, 583, 1593, 1100, 224]. Mediation [804, 521, 570]. medirostris [805, 1494]. Mediterranean [1487, 509, 1151, 820, 70, 476, 1364]. medium [844, 1235]. medium-scale [844]. medium-size [1235]. meet [856]. meeting [1729]. meets [1438]. Megaptera [507]. Mekong [1183]. Melanogrammus [1459, 80, 632]. melanostomus [403, 395, 55, 575]. Melicertus [421, 484]. melting [399]. meltwater [473]. membras [1685]. memory [1739]. Mendocino [188].

Menidia [1331, 1561, 189]. menu [1191]. Mercury [1331, 973, 1821, 1705, 1161, 454, 578, 343, 367, 121, 699, 124, 1798, 1561]. mesh [1487, 1748]. mesopelagic [820]. Mesoscale [1688, 1440]. meta [1350, 328, 1826, 587, 750, 78]. meta-analysis [1350, 328, 1826, 587, 750, 78]. metabarcoding [1360, 1804]. Metabolic [243, 1195, 640, 1201, 125, 860, 1149, 999, 1167, 170, 760, 1769]. metabolism [1841, 580, 876, 110, 1063, 1472]. metabolotypes [754]. Metacarcinus [1323]. metacommunity [1122, 1248]. metal [411, 1810]. metamorphosed [591]. Metapopulation [946, 1321, 336, 193]. metapopulations [713, 173]. meteorology [1428]. Method [638, 1658, 111, 1430, 1343, 127, 1717, 291, 1239, 208, 668, 890, 287, 438, 1254, 697, 842, 773, 1659, 467]. Method- [638]. methodology [1084]. methods [453, 1377, 144, 123, 474, 1374, 122, 1013, 1684, 365, 1562, 978]. methylation [29]. métier [1422]. metric [209, 94]. Metrics [413, 1686, 659, 631, 1592, 1759, 1477, 757, 326, 506, 1709]. Mexico [646, 987, 1772, 1538, 86, 1036]. MGMS [1239]. Michigan [1268, 1803, 279, 1432, 795, 55, 1330, 1808, 809, 772, 1667]. microbial [1225]. microcephalus [124]. microchemical [978]. microchemistry [1024, 1802, 1465, 128, 432, 247, 79, 748]. microcosm [956]. microcrustacean [1295]. microcystin [477]. microcystins [657]. Microcystis [861]. microhabitat [1704]. Micromesistius [54, 1184]. Microplastic [872]. Microplastics [1819]. Micropogonias [1510]. Micropterus [334, 376, 1186, 237]. microsatellite [648, 104, 498, 85, 616]. microsatellites [113]. microspectrometry [605]. microsporidian [795]. middle [1829]. midwater [1098]. migrants [1578]. migrating [1604, 15, 764, 63, 1636, 181, 1281, 614]. Migration [1761, 1464, 1639, 775, 1604, 1827, 855, 1605, 650, 238, 973, 1733, 1290, 613, 1556, 1004, 75, 88, 1047, 708, 1607, 1528, 670, 1001, 415, 1812, 1226, 635, 1133, 548, 358, 1573, 1447, 1493, 546, 1707, 896, 791, 735, 1676, 866, 1799, 723, 1356, 1635, 1154, 1650, 1301, 868, 1276, 1097, 739, 116, 1780, 1002, 49, 1783, 525, 953, 1046, 599, 1524]. migrations [1557, 656, 1618]. Migratory [1311, 539, 1420, 1838, 600, 47, 77, 612, 336, 547, 1220, 639, 1812, 1599, 1222, 871, 1609, 25, 1091, 224, 1530, 355]. mill [782]. mineral [605, 603]. mineralization [691]. minibicornis [65]. minimally [1552]. Minimum [186]. mining [1622, 1810, 1460, 1426, 1384]. mining-impacted [1810]. Minnesota [172]. minnow [508, 931]. minnows [380, 1814]. Miramichi [1586]. Misapplied [1139]. misdiagnosing [1401]. Mismatch [185, 1417]. mismatches [918]. missed [1175]. Missing [1026, 421, 1298, 377]. mission [1389]. Mississippi [432]. Missouri [432]. mitigate [1435]. mitigation [1623]. mix [1516]. Mixed [33, 226, 558, 1686, 1507, 1059, 1720, 951, 1785, 1210, 1279, 711, 426, 1627]. mixed-forest [1210]. mixed-stock [226, 1720, 951, 1279]. mixing [1409, 966]. mixture [1707, 131, 1516, 315]. mobile [601, 117, 1602]. mode [347]. Model [76, 705, 419, 900, 749, 1101, 1572, 558, 83, 1381, 939, 238, 1085, 1396, 1272, 1273, 789, 335, 1462, 581, 1680, 1385, 23, 1129, 961, 853, 1269, 427, 263, 576, 75, 129, 532, 332, 1369, 935, 798, 1336, 1708, 867, 1227, 1522, 312, 311, 1120, 1001, 674, 130, 161, 891, 270, 511, 557, 1330, 574, 740, 304, 1615, 1824, 1400, 720, 1826, 283, 1485, 664, 284, 1779, 806, 1725, 1176, 1486, 554, 1567, 1682, 1661, 1134, 282, 1294, 727, 1139, 704, 220, 174, 1584, 267, 433, 186, 533, 493, 1789, 1114, 1700, 96, 410, 799, 1153]. model [1544, 1670, 514, 772, 1189, 101, 1405, 1306, 1046, 1683, 1530, 937, 886, 1125]. model-based [410, 1670]. Modeling [986, 1155, 77, 159, 540, 1687, 669, 846, 1299, 1409, 691, 221, 1824, 1537, 851, 1816, 1130, 1786, 788, 464, 1444, 499, 1717, 418, 306, 1471, 1160, 672, 994, 686, 679, 401, 1810, 861, 713, 1809, 1584]. modelled [1076]. Modelling [1321, 987, 1369, 97, 1707, 520, 838, 1476, 453, 1377, 191, 217, 1340, 1191, 781, 1249, 1257, 531, 1398, 1529, 1612, 1402]. Models [10, 1585, 1310, 1113, 749, 1101, 369, 879, 1177, 1507, 319, 1059, 823, 878, 474, 903, 1729, 1681, 1535, 1546, 1309, 357, 271, 445, 669, 491, 505, 889, 1089, 672, 1252, 1, 1431, 69, 1788, 1807, 1690, 1532, 1397, 595, 1399, 634, 321, 1744, 663, 1194, 1706, 1449, 1828, 1655, 1742, 1298, 792, 1231, 1453, 471, 1175, 377, 862, 1647, 633, 1516, 315, 1055, 794, 506, 214, 33, 173, 1569, 933, 828, 1659, 309, 276, 729]. modern [1832, 230, 794]. modified [1030, 1616, 840, 361, 227, 426]. modifies [528]. modulate [1116, 860]. modulated [880]. modulates [236]. modulators [1512]. molecular [1382, 1702]. Moller [1717]. molluscs [89]. money [1651]. monitored [1660, 1589]. monitoring [339, 835, 235, 20, 1603, 251, 1759, 1221, 1515, 864, 694, 282, 727, 757, 1418, 1159, 1306]. month [1367]. moored [1480]. Moose [228]. Moran [1202]. morhua [1420, 648, 909, 1459, 81, 141, 61, 231, 406, 693, 1746, 109, 866, 1652, 383, 51, 834, 1196, 437, 366, 134, 399, 608]. Morone [1586, 1226, 378, 392, 459]. morph [841]. Morphological [325, 1382, 1790, 104, 1281, 718]. morphologically [58]. morphology [1346, 1099, 1097, 490, 1091, 1575, 841]. morphometry [697]. morphotypes [754, 1229]. Mortality [1837, 786, 900, 1264, 1081, 65, 1381, 1085, 1205, 1638, 427, 1415, 1249, 129, 386, 908, 1336, 1065, 1089, 479, 312, 1842, 891, 679, 97, 1222, 1397, 1090, 1550, 851, 385, 405, 917, 1486, 485, 559, 932, 1758, 1078, 839, 1514, 762, 1703, 1840, 543, 2, 300, 1200, 514, 1489, 1403, 1843, 1079, 1405, 992, 1307, 276, 462, 1718, 886]. mosaic [1209]. most [571]. mother [1646]. mothers [1665]. motivation [1320]. mottled [795]. Mountain [1023, 728, 1587, 905, 733, 1527, 1551, 1148]. Mountains [1806, 135, 1215, 298, 968]. mouths [1316]. Move [908, 1606, 599]. Movement [1024, 468, 1602, 1689, 1441, 728, 601, 987, 1712, 819, 427, 1257, 1470, 1369, 908, 1620, 288, 507, 1089, 867, 342, 1637, 511, 1398, 807, 1810, 693, 287, 864, 1157, 232, 1142, 569, 839, 1002, 1040, 1632, 1634, 1530, 255]. Movements [850, 1205, 1664, 1681, 445, 938, 979, 53, 960, 696, 1266, 764, 1610, 448, 430, 1197, 1510, 1466, 1423]. Moving [1796]. MS [1293, 1232]. MSE [1728, 1691]. MSY [1318]. mtDNA [983]. Mueller [1074]. muelleri [1074]. multi [1158, 1137, 1249, 1397, 1279, 1809, 315, 1363]. multi-area [1397]. multi-event [1249]. multi-generational [1137]. multi-SNP [1279]. multi-species [315]. multi-time [1158]. multi-timescale [1809]. multi-trophic [1363]. multibasin [1099]. Multibeam [765]. Multidecadal [590, 1224]. multidimensional [1040]. multielement [1510]. multifleet [276]. Multifractal [390, 407]. multifrequency [127]. multigear [1239]. multilake [451]. multilocus [1095]. multimodel [705]. Multiple [1543, 690, 782, 1354, 563, 1503, 474, 18, 1344, 149, 431, 1594, 360, 846, 1522, 373, 1022, 303, 1499, 9, 1747, 165, 561, 174, 1601, 1384, 985, 1566, 1224]. multiple-stressor [360]. multiproxy [228]. Multipurpose [1571]. multiregion [1405]. Multiscale [930, 417, 1664, 1698]. multisite [36]. Multispecies [86, 581, 1794, 1318, 1839, 1654, 1463, 1529, 1294, 220, 1275, 377, 254, 1114, 1395, 772, 101]. multistage [119, 284]. multistate [445, 674, 1810]. multistock [1788, 716, 684, 1402]. multistressor [1608]. multitrait [1488]. Multivariate [567, 1374, 67, 1655]. multiyear [636]. municipal [110]. muscle [1771, 1186, 358, 527, 1281, 1684, 214]. Muskoka [1093]. mussel [120, 1014, 1199, 846, 1219, 990, 817, 163]. mussels [970, 875, 1013, 1657, 992]. my [1646]. mycobacteriosis [378]. mykiss [609, 1264, 615, 29, 140, 273, 222, 288, 1607, 1454, 1812, 914, 1353, 328, 270, 269, 511, 1253, 751, 598, 310, 181, 723, 883, 1619, 157, 41, 1598, 897, 542, 760, 85, 1048, 126, 1159, 1306, 607, 447]. Myoxocephalus [1637, 1689]. Myriophyllum [66]. Mysis [245, 265]. Mytilus [120].

N [1516, 369, 249, 211, 551, 124, 1419, 1243, 1516, 755, 1834]. N-enrichment [755]. N-mix [1516]. N-mixture [1516]. namaycush [1060, 418, 1319, 1461, 1614, 1750, 1518, 38, 104, 442, 1229, 79, 350]. nase [1446, 1441]. Naso [1203]. nasus [1339, 84, 1446, 1441]. Natal [56, 1808, 555, 1636, 959, 1405, 142]. National [298, 1023, 350, 578, 1252]. native [305, 1131, 1640, 1136, 970, 624, 418, 825, 14, 1471, 590, 380, 1696, 345, 1295, 504, 4, 517, 1723, 733, 22, 41, 1163, 1703, 1314, 350, 1657, 816]. Natural [848, 9, 1081, 1841, 334, 1085, 223, 584, 1378, 819, 909, 590, 1696, 1708, 589, 858, 97, 1412, 1222, 1526, 1090, 1486, 1598, 839, 526, 737, 2, 300, 1783, 959, 1079, 1555, 1405, 709, 276]. natural-area [1378]. natural-occurring [334]. natural-origin [1526, 1555]. naturally [758]. nature [1194]. navigating [1577]. Navigation [936]. near [411, 1701, 869, 1262, 805]. near-bottom [1262]. nearly [308]. nearshore [381, 1637, 1190, 1698, 1667, 103, 1363]. nebulosus [1149]. need [1265]. needs [1404]. negatively [1173, 1674]. neglect [520]. neighboring [1033]. neighbour [417]. nekton [168]. nematode [1787]. Neogobius [403, 395, 55, 575]. Neophocaena [1382]. Nephrops [1285]. Nephrops-directed [1285]. nerka [113, 226, 65, 242, 308, 431, 239, 202, 1240, 487, 385, 15, 482, 1025, 458, 382, 1037, 541, 1840, 842]. nestling [1819]. nests [1380, 297]. Net [1501, 1159, 1349, 1265, 1026, 1334, 1306, 937]. Net-chasing [1501]. Network [1722, 1693, 99, 1644, 728, 1756, 1248, 1252, 1287, 717, 448, 430, 255]. networks [1430, 178, 118, 803, 59, 21]. neutral [92, 383]. Nevada [1451, 31]. Newfoundland [1816, 1838, 30, 1196, 608]. newly [925, 591]. next [1439]. niche [1158, 259, 1753, 861, 1560, 1492, 1645, 1363]. night [694]. nigricans [1425]. Nile [1559, 1117]. Ninilchik [181]. nitrate [388]. nitrogen [1354, 47, 449, 658, 1481, 137, 82, 477, 1467]. nitrogen-to-phosphorus [477]. No [1258, 29, 1018, 209, 235, 196, 1406, 480, 1745]. No-take [1018, 209, 196, 1406]. Node [1497]. noise [928, 1558]. noise-reduced [928]. noises [1790]. noisy [18]. nominal [1835]. Non [1136, 1131, 1640, 418, 1007, 345, 1295, 1414, 350]. non-homing [1414]. Non-native [1136, 1131, 1640, 418, 345, 1295, 350]. non-target [1007]. noncompliance [891]. nongame [23]. nonindigenous [436, 993, 32]. Nonlethal [1014, 1684, 1232]. nonlinear [319, 215]. nonnative [1703, 816]. nonphysical [504]. nonproportionality [216]. nonstationarity [1308]. Nonstationary [1797, 1824]. Nooksack [859]. nor [1839]. Nordborg [249]. Nordmøre [1662]. norm [1193]. norms [51]. north [207, 966, 945, 934, 314, 424, 76, 982, 217, 961, 1711, 468, 215, 81, 631, 1592, 1138, 292, 1456, 1324, 1330, 548, 720, 38, 1162, 1499, 458, 1376, 957, 1770, 1314, 116, 736, 1418, 1000, 976, 1625, 962, 1224]. north-temperate [207, 966]. Northeast [792, 767, 1081, 667, 1728, 1184, 945]. northeastern [600, 1519, 155]. Northern [1044, 1060, 1085, 1394, 545, 135, 1344, 1016, 43, 1065, 848, 17, 231, 406, 871, 1282, 434, 1799, 683, 1600, 1490, 1685, 1628, 1057, 907, 1538, 1426, 366, 1189, 1705, 1292, 1384, 1442, 1022, 1629, 1498]. northernmost [869]. Northwest [84, 663, 506, 522, 738, 894, 1205, 1735, 168, 1622, 1353, 1518, 1699, 850, 1082, 1674, 784]. northwestern [198, 1771]. Norvège [1071]. Norway [1071, 1842]. Norwegian [199, 158, 294, 1076]. Note [518, 698, 1043, 1216, 1410, 1611, 1820]. noticeable [1161]. Nototodarus [1027]. Notropis [1024]. Nouvelle [469]. Nova [1664, 579, 655]. novaeangliae [507]. Novel [1640, 761, 646, 1203, 1717, 571, 455, 1801, 1025, 1254, 965, 1037, 902, 863, 467]. November [688]. nuclear [873, 1478]. nucleotide [113, 1520, 291, 548, 778, 1511]. number [36, 1258]. numerical [140]. Numéro [1071]. numéros [257]. Nunavut [1190]. nursery [1832, 606, 1061, 682, 463, 1163, 541]. Nutrient [573, 880, 275, 919, 365, 1296, 1474, 108, 876, 494, 42, 164, 583, 477, 1757, 204, 1032, 941, 70, 1814, 163, 1208, 1627]. nutrient-enhanced [1627]. nutrient-limited [1208]. nutrient-rich [477]. Nutrients [924, 10, 153, 986, 515, 286, 516, 1243, 941, 1740]. nutrition [265]. Nutritional [1578, 11].

O [369, 56, 110]. obesus [343, 850]. objective [671, 1237]. objectives [332, 251, 1677]. objects [758, 1589]. obligate [1819]. obscurus [1557]. observation [1579, 1371, 276]. observational [1310]. observations [1380, 93, 1431, 1690, 1098]. observatory [890]. observe [890, 693]. observed [1170, 388]. observer [835, 1616]. obtain [671]. obtained [725]. occupancy [882, 1377, 167, 1470, 1316, 638, 674, 645, 901, 972, 949, 394, 1683]. occur [477]. Occurrence [1660, 887, 150, 1148, 46, 542, 1452]. occurring [334, 540, 1645]. Ocean [277, 1722, 831, 519, 102, 650, 823, 545, 1245, 1378, 1220, 45, 1353, 1797, 870, 132, 1022, 546, 597, 1368, 433, 592, 1647, 692, 142, 314, 973, 198, 1735, 1519, 343, 1138, 1393, 548, 1447, 945, 425, 786, 330, 957, 1645, 1770, 1418, 962]. ocean-harvested [433]. Ocean-scale [831]. oceanic [705, 869, 635, 1326, 1259]. Oceanographic [736, 1502, 67, 850]. Oceanography [469, 1742]. oceanography-dependent [1742]. oceans [600]. ocellata [75]. ocellatus [965]. October [1071]. octobre [1071]. Oddity [1387, 1388, 1389]. odontocetes [1509]. odor [274]. Odyssée [1387]. off [974, 465, 468, 417, 708, 944, 1370, 1009]. off-channel [944]. offs [1195, 1572, 1450, 375, 1415, 812, 332, 251, 1032, 1783]. offshore [1560, 1492, 809]. offsite [1005]. Offspring [536, 604, 1541, 171]. Ohio [614]. oil [780, 1289, 1348]. old [230, 1327, 1668, 1709, 977]. Oldman [454]. Olfactory [224]. Olfactory-mediated [224]. oligotrophic [364, 137, 1243, 1474]. Omic [1327]. Omic-style [1327]. once [1213]. Oncorhynchus [314, 1749, 10, 925, 176, 26, 113, 226, 166, 195, 609, 1264, 615, 29, 65, 242, 140, 1836, 847, 3, 1070, 200, 291, 576, 571, 7, 177, 273, 222, 308, 431, 1127, 288, 239, 1607, 53, 331, 251, 202, 1801, 42, 1240, 1454, 1812, 914, 1353, 328, 270, 269, 511, 182, 492, 1253, 548, 778, 487, 1210, 751, 132, 175, 1357, 385, 15, 482, 598, 230, 434, 597, 1025, 87, 448, 430, 822, 310, 201, 181, 723, 883, 1619, 157, 408, 413, 435, 458, 382, 41, 363, 1598, 1792, 439, 268, 1006, 897, 1770, 1037, 541, 542]. Oncorhynchus [760, 1840, 116, 722, 394, 1291, 842, 25, 85, 1511, 50, 1048, 1740, 126, 326, 1159, 1306, 6, 27, 506, 607, 802, 156, 841, 447, 255]. one [1528, 1274, 773]. only [63, 477]. onsite [1005]. Ontario [360, 1093, 819, 1771, 719, 1540, 411, 245, 71, 1152, 880, 1560, 1492, 1552, 1833, 610, 264, 745]. Ontogenetic [40, 1257, 455, 653, 1186, 1510, 792, 767, 1476, 1037, 1575]. ontogeny [534, 1798]. oocyte [304]. oocytes [197]. opacity [605]. open [823, 545, 1546, 45, 1689, 1718]. open-ocean [545, 45]. open-water [1689, 1718]. Opeongo [381]. operation [1750, 157]. operations [780, 770, 32]. opilio [1396, 878, 1664, 400, 703, 1228]. opportunistic [1051, 351]. opportunities [1049, 1421]. opportunity [28, 708]. Optimal [36, 1377, 891, 1760, 1351, 1704, 130, 161, 923, 1424, 1509, 520, 838, 1641, 1189, 313]. Optimization [595]. Optimizing [1678]. options [375, 384, 1734]. oral [1280]. Orconectes [451, 66]. order [1019, 1436, 1474]. ordination [1193]. Oregon [1316, 1545, 53, 439, 1145, 1291, 1307]. Organic [1201, 1173, 235, 996, 551, 605, 1590, 1737, 1419, 1032, 1063, 513]. organisms [875]. organochlorine [938, 699, 359]. orientalis [973]. orientation [1414]. oriented [418]. Origin [869, 1127, 1526, 1615, 1025, 432, 56, 959, 1555]. Origins [840, 416, 1808, 555, 9, 340, 142]. osmoregulation [1539]. osmoregulatory [268]. other [1138]. Otolith [1420, 320, 1465, 487, 1629, 128, 115, 1302, 649, 728, 1024, 1303, 1802, 189, 630, 1290, 671, 455, 627, 460, 1808, 1184, 555, 971, 287, 1799, 260, 210, 247, 56, 131, 1196, 79, 748, 1575, 1293, 98, 978]. otoliths [512, 1583, 105, 551, 1646, 1701, 1241, 605, 1346, 60, 121, 230, 947, 1232, 962, 416, 709, 535]. otter [125, 681]. outbreaks [1166]. outbred [225]. outcome [1825]. outcomes [1716]. outdoor [1267]. outmigration [1765, 1039, 1752, 722]. output [939, 130, 161]. outputs [1231, 1453]. outsized [1752]. ovarian [1731, 61]. ovary [11]. overcome [1579]. overcoming [1827]. overestimate [403]. overfishing [1333, 1719]. overlap [1686, 1735, 259, 875, 887, 1560, 1492, 1598]. Overlooked [732, 1766]. override [1598]. oversight [936]. overview [578, 1778]. Overwinter [1633, 559]. overwintering [1180, 383, 1769]. oxycline [417]. oxygen [89, 1230, 860, 1833, 163, 110, 1524, 1769]. oxyrinchus [839, 1632, 1634, 1393, 1632, 1634]. oxythermal [172, 1477, 1537]. oyster [713].

P [249]. Pacific [314, 168, 177, 1138, 548, 945, 957, 1770, 962, 314, 176, 600, 1180, 1379, 76, 973, 1519, 1711, 343, 576, 136, 1470, 155, 7, 177, 669, 1471, 215, 1047, 43, 1138, 251, 1, 686, 761, 1256, 1324, 1393, 685, 1353, 1699, 185, 1609, 560, 1797, 1447, 1304, 151, 654, 1029, 1015, 1674, 413, 330, 363, 611, 1825, 1009, 736, 1780, 1000, 326, 506, 599, 607, 1383, 1834, 355]. padus [1214]. Paenibacillus [442]. Pagrus [232, 1501]. pair [338]. paired [663, 1706, 1255]. paired-gear [663]. paired-release [1255]. Paiute [291]. paleolimnological [228, 1832, 411, 360]. paleolimnology [771]. PAM [880]. pan [622]. pan-Arctic [622]. Pandalus [1272, 1273, 419, 407, 390]. panel [1487, 1748, 1285]. panels [1695]. Panmixia [983, 1274, 784]. Panulirus [368, 16, 277]. paradigm [88]. paradox [1154]. Paralichthys [47, 1742]. Paralithodes [572]. Parallel [895, 38]. Parameter [1114, 532, 1813, 587]. parameterized [1431]. parameters [136, 322, 557, 1746, 1234, 1562, 2, 300, 96]. parametric [1659]. parasite [795, 373, 1150, 165, 1402]. parasites [481]. parasitic [854]. parasitism [1229]. Parastichopus [681]. parentage [1362, 1520, 1137, 1156, 637]. parentage-based [1362, 1520, 637]. parental [1033]. Park [1023, 350]. parks [298]. parr [640, 836, 64, 147, 1434, 87]. parrotfishes [1475]. Part [1272, 586, 1273]. Partial [896, 89, 1464, 764]. partially [1812, 341, 1091]. participation [1795]. particle [115]. particle-tracking [115]. particulate [513]. partitioning [1726]. Parvicapsula [65]. pass [1680]. passage [1311, 1413, 1642, 1066, 1320, 867, 1525, 685, 1226, 871, 1175, 1700, 863, 886]. passing [677]. passive [1660, 1546, 1064, 1418, 488, 1441]. past [63, 610, 1291, 1255]. Patagonia [1836]. patch [602]. patchy [912, 1621]. paternal [256, 229, 225]. pathogen [1690]. Paths [1495, 1601]. pathways [1469, 164, 1022, 1038, 1209]. Patinopecten [155]. pattern [418, 571, 366, 1292]. pattern-oriented [418]. Patterns [600, 60, 147, 1440, 70, 1644, 855, 1264, 248, 1049, 1838, 1180, 1815, 1716, 1083, 1024, 1109, 1337, 547, 1519, 1556, 1374, 343, 875, 1369, 347, 1620, 288, 889, 507, 1456, 1001, 914, 1353, 1517, 13, 1253, 1723, 1573, 1447, 1329, 287, 323, 407, 390, 647, 610, 1742, 384, 1635, 1154, 1488, 1327, 1093, 361, 1209, 1825, 641, 1741, 592, 524, 1032, 116, 730, 736, 1782, 1259, 953, 94, 1668, 1524, 108, 98, 1068, 255, 1441, 1017]. paucispinis [91]. PaV1 [368]. PBDEs [1054]. PCNMs [417]. Peace [566]. peak [1513]. peaked [747]. pealeii [1001]. pearlside [1074]. peat [513]. peatland [924]. Pecos [1024]. pecosensis [1024]. pedigree [1702, 483, 1383]. pedigrees [260]. pelagic [36, 556, 912, 1136, 1205, 710, 465, 364, 931, 653, 137, 1324, 13, 746, 706, 806, 90, 265, 772, 1390, 1364, 755, 715, 521, 570, 1031]. pelagic-broadcast [931]. pelagically [1687]. pelamis [198, 758]. pellets [120]. penalty [1615]. pengoi [245, 530]. Penobscot [1786]. Perca [1096, 105, 1443, 1431, 1460, 620, 295, 1017]. Perceived [1675]. perception [797]. perceptions [801]. perch [1096, 105, 1202, 1415, 1182, 1427, 1443, 860, 1431, 1465, 851, 1559, 1460, 620, 295, 206, 1017, 1117]. Percidae [240]. Percina [240]. Performance [526, 1019, 1195, 1146, 789, 564, 1311, 223, 584, 325, 581, 1413, 1323, 1066, 935, 552, 1431, 810, 1615, 509, 720, 201, 1828, 51, 247, 897, 1700, 1094, 933, 1482]. period [1613, 200, 554, 480]. periodic [400, 168]. periods [11, 1689]. periphyton [24]. permanence [138]. permanent [114]. permitting [1644]. Persistence [555, 1638, 338, 611, 1018]. personality [241]. perspective [250, 1638, 622, 1342, 411, 690, 628, 30, 232, 1164, 975, 313]. perspectives [996, 1352, 1215, 899]. perturbations [1344]. Peru [465, 417]. Peruvian [844]. pesticide [1438]. Peters [1102, 1103, 1104]. petrale [538]. Petromyzon [790, 28, 1045, 1548, 643, 591, 764, 1764, 704, 1734, 224]. pH [99, 1434, 1523]. PHABSIM [520]. Phalacrocorax [1819]. phase [1838, 613, 534, 1456, 132, 1804]. Phenology [1653, 167, 149, 1220, 1319, 945, 1373, 633, 1649]. phenotype [1541]. phenotypes [166, 195, 95, 1663, 985, 1769]. Phenotypic [1626, 297, 853, 1329]. pheromone [643, 764]. pheromones [575, 355]. Philippines [774]. phosphate [691]. Phosphorus [112, 501, 1585, 1170, 207, 629, 172, 137, 1367, 610, 477, 1411, 865, 745, 794]. phosphorus-poor [501]. photochemistry [941]. photographic [1706, 697]. phylogenetic [1592, 60]. Phylogeny [631]. Physical [785, 1233, 1156]. physicochemical [1821, 1219]. Physiological [1195, 843, 754, 1199, 546, 1626, 1020, 1013, 1014, 15, 297, 1229, 25, 1200, 1250, 326]. Physiology [897, 544, 1556, 237, 490]. Phytoplankton [556, 1784, 1193, 388, 874, 810, 1367, 880, 945, 365, 1355, 1126, 941, 521, 570]. pigment [911]. pike [1344, 848, 1292]. pikeminnow [1263]. pikeperch [1065]. pilchardus [498]. pilot [1305]. Pink [268, 1349, 571, 182, 175, 1357, 482, 434, 407, 390, 201, 435, 1407, 957, 1097, 1511]. pinniped [310, 1792, 902]. pinniped-caused [310]. pipes [822]. piscivore [1819, 573]. piscivores [1160, 4, 517, 416]. piscivorous [248, 742]. piscivory [889]. PIT [728]. pitfalls [939, 253]. placement [926]. Placopecten [1706, 1300]. plaice [606, 1162]. Plains [167, 1317, 1435]. planktivore [620, 528]. planktivorous [755]. planktivory [530]. plankton [956, 1116, 203, 1757, 1426]. planning [749, 1101, 567, 642, 1332, 1668]. plans [1678]. plant [1058, 494, 1225, 1745, 1478]. plants [227]. plastic [349]. plasticity [853, 725, 1329, 1041, 297, 1283]. platessa [606]. Platichthys [359]. playbacks [928]. plays [388, 268]. plebejus [421, 484]. Plenty [1753, 318]. pleuriticus [255]. Pleurobema [1219]. Pleuronectes [606]. plume [140, 1361, 90]. poaching [1321]. point [1308]. points [813, 136, 1401, 479, 1318, 1185, 628, 1617, 1300, 730, 486, 462]. policies [1021, 998]. Policy [356, 1372, 652, 384]. pollock [379, 619, 1701, 301, 280, 311, 672, 1262, 815, 387, 1601]. pollutant [656]. polluted [133]. pollution [872]. polymorpha [970, 990, 163, 992]. polymorphism [226, 1329, 1008, 1511]. polymorphisms [113, 1520, 291, 548, 778]. polyxystra [17]. pond [1481]. ponds [1267, 1367, 303, 1452]. pools [683, 901]. Poor [513, 501, 1227, 1754, 1234, 1772, 714, 1046, 924]. Pop [1770, 635]. Pop-up [1770, 635]. Population [1362, 1520, 1672, 189, 95, 662, 450, 1297, 635, 104, 660, 382, 91, 447, 10, 153, 1570, 1665, 1034, 1051, 749, 1101, 226, 601, 242, 821, 1096, 1421, 1136, 307, 1503, 191, 721, 830, 1767, 216, 1546, 1643, 277, 291, 666, 1169, 1542, 75, 88, 1401, 1178, 651, 770, 84, 1016, 507, 724, 568, 196, 1075, 1252, 245, 1042, 1518, 270, 1599, 1397, 1824, 751, 283, 1373, 555, 971, 284, 597, 1499, 559, 1197, 1652, 22, 30, 1165, 1134, 435, 1818, 827, 1685, 1702, 1139, 704, 472, 316, 285, 1018, 267, 1052, 524, 1145, 186, 295, 1618]. population [1786, 1511, 1423, 134, 1055, 514, 1079, 1270, 953, 126, 483, 1284, 348, 467, 44, 1627, 500]. population-level [1397, 827]. population-scale [1169]. Population-specific [95, 635]. populations [1187, 1768, 1349, 1639, 952, 26, 609, 1161, 1838, 600, 601, 1507, 1088, 1303, 223, 584, 281, 452, 3, 1070, 1519, 562, 877, 825, 1249, 149, 431, 58, 1673, 1433, 92, 553, 1089, 199, 251, 950, 218, 415, 1393, 1468, 1412, 744, 4, 517, 539, 741, 793, 38, 764, 1752, 275, 1179, 272, 1015, 522, 1490, 963, 260, 458, 1679, 1628, 1533, 324, 543, 79, 902, 1763, 714, 94, 6, 27, 1292, 607, 399, 344, 602, 98, 1580]. porbeagle [1339]. Pore [1219]. porpoises [1382]. Port [1717]. portfolio [307, 1052]. Portland [688]. pose [1205]. position [1248, 14, 342, 1750, 717, 904]. positional [1174]. positioning [59, 21]. positions [668, 1737]. positively [1100]. post [939, 909]. post-release [909]. postbomb [646]. posterior [912]. postexercise [201]. postglacial [481]. posthatch [1230]. postlogging [1469]. Postrelease [932, 35, 1133, 1840, 1200]. postsmolts [579, 539]. postspawned [639]. postzygotic [256, 229]. Potential [314, 877, 149, 568, 733, 66, 1376, 99, 1644, 780, 952, 918, 1024, 216, 452, 1438, 1771, 1720, 1342, 345, 1227, 848, 504, 1759, 1808, 1373, 1304, 246, 665, 290, 1082, 1504, 408, 1130, 1702, 1093, 32, 463, 1163, 1048, 1390, 1063, 1228, 818, 801]. potentially [409]. Potomac [263]. pouched [1380]. poutassou [54, 1184]. power [421, 337, 1279, 1818, 1247, 1710, 1478, 1112]. practices [1728, 849, 331, 471, 1786]. prairie [1841]. Prairies [1428]. prawn [421, 484]. pre [71, 1552]. pre-industrial [71, 1552]. Precambrian [1170]. Precise [1058]. precision [1515, 1670, 1247, 1710]. precocious [525]. preconditions [1366]. Predation [572, 783, 620, 1551, 1181, 1124, 836, 948, 708, 1065, 1765, 1436, 293, 681, 1242, 1107, 1652, 1315, 1568, 694, 1371, 883, 1294, 596, 1141, 1139, 1097, 1770, 1489, 1364]. predation-driven [1652]. predation-related [1765]. Predator [1600, 1361, 1364, 1365, 1180, 47, 379, 798, 380, 701, 1121, 451, 1185, 693, 1254, 1721, 529, 1123, 49, 815, 772, 1031]. predator-driven [49]. predators [710, 1345, 1182, 342, 1121, 746, 1259, 346]. Predatory [1436, 508, 1602, 1062, 1798]. predict [558, 1585, 819, 17, 1431, 1812, 1194, 471, 1631, 326, 1718, 886]. predicted [8, 5, 73, 171, 72]. Predicting [1131, 436, 1694, 942, 940, 553, 148, 1468, 489, 1754, 882, 211, 1515, 922]. Prediction [1800, 294, 912, 1708, 290, 1641]. predictions [1120, 891, 108, 220]. Predictive [1332, 879, 1390]. predictor [487]. predictors [48]. predicts [1260, 1159, 937]. Preference [1190, 1762]. preferences [696]. premature [65, 1154, 1301]. presence [375, 1343, 1496, 1781, 905, 1470, 891, 1226, 284, 1780, 48]. present [71, 361, 937]. present-day [71]. presentation [1237]. preserving [272]. prespawning [868]. pressure [89, 1777, 668, 1281, 1100]. pressures [1686]. pretagged [909]. prevalence [1791]. prevent [1157]. prevented [1719]. previous [1483, 22]. Prey [1129, 1365, 248, 1213, 1379, 563, 379, 824, 279, 1124, 125, 133, 798, 372, 701, 1121, 1062, 1185, 57, 1107, 1254, 466, 1600, 1361, 1685, 529, 1214, 520, 1123, 341, 772, 1364, 1031]. price [313]. prickly [871]. Primary [589, 1832, 34, 530, 573, 1032]. Prince [434]. principal [417]. Principles [1138]. prior [866, 1250, 525]. prioritization [685, 1218]. prioritizing [942]. priors [86]. Pristipomoides [512]. Probabilistic [1806, 379]. probabilities [638, 867, 63, 1744]. probability [1567, 949, 1403]. probable [76]. problem [544, 1707]. problems [1544]. procedures [370, 997, 1780]. process [75, 664, 1486]. Processes [1388, 1387, 1716, 1565, 1248, 1115, 534, 1169, 914, 1824, 1022, 634, 1691, 1705]. processing [1503]. processus [1387]. produce [604, 1225]. produced [849, 141, 607, 962]. Production [958, 900, 1832, 393, 12, 456, 34, 364, 129, 1666, 187, 589, 530, 1461, 183, 1030, 1211, 1347, 1828, 1183, 1600, 91, 1275, 1032, 1055, 899, 1011]. productive [34, 356, 114, 769, 1428]. Productivity [1582, 734, 957, 925, 375, 52, 1519, 770, 1590, 182, 1797, 945, 1357, 753, 1674, 984, 458, 1497, 1628, 1300, 1763, 1555]. profile [554]. profiles [64, 1432, 1607, 342, 359, 1472]. profiling [1293]. profundal [874]. program [1137, 552, 586, 642, 1006, 1776, 126, 1159, 636]. programs [1555]. progress [1372]. project [1668]. projected [1527]. Projecting [833]. projections [1146, 1627]. proliferation [771, 489]. prolonged [843]. promises [253]. promote [856, 4, 517]. prone [913, 1451]. properties [974]. proportional [1835]. proportionality [1338]. proportionality-based [1338]. proportions [514]. proportions-at-age [514]. proposed [122, 565]. Prosopium [728, 1158, 1499]. Prospects [19]. protect [1733, 822]. protected [508, 1343, 322, 866, 711, 1595]. protecting [24]. Protection [936, 1265, 1477]. protocol [1566]. protocols [377]. protogyny [1406]. provenience [999]. provide [1480]. provided [113, 1005]. provides [1299, 677]. providing [1772]. Provinces [235]. Prunus [1214]. pseudoharengus [975, 108]. Pseudophycis [320]. Pseudopleuronectes [828]. Ptychocheilus [1263]. Puerto [469]. Puget [200, 273, 222, 871, 731]. pulp [782]. pulse [880, 57, 598]. pulsed [499]. pumping [886]. pumpkinseed [970]. punctipennis [775]. punctuated [1047]. pupping [74]. purse [339, 1007, 786, 1589]. pusillus [320]. putative [1045, 355]. Pyganodon [159]. Pyrenean [825].

QFASA [1172]. quagga [163]. qualify [106]. qualitative [800, 1577, 1569]. quality [882, 556, 1213, 250, 1815, 1731, 496, 1127, 360, 42, 1788, 510, 1135, 916, 221, 1477, 1225, 272, 264, 493, 1428, 1228, 24, 636]. quality-improving [1815]. Quantification [1040, 965, 1334, 995]. quantified [110]. quantify [1412, 614]. Quantifying [918, 1421, 1712, 1178, 17, 1015, 1315, 1650, 769, 341, 1386, 1613, 1143, 889, 433, 377]. Quantitative [396, 1506, 1172, 800, 1207, 1402]. quantity [1319, 1127, 1477, 1825, 636]. Quantum [941]. Quaternary [577]. Que [1355]. Quebec [771, 794, 398]. quernus [1793]. question [88]. questions [720]. qui [469, 1071]. quick [1132]. Quinn [1739]. Quinte [530]. quo [1092]. quota [1086, 998]. quotas [711, 160].

races [577]. racetrack [1369]. radiata [1816]. radiation [683]. radio [1528]. radio-tagged [1528]. radioactive [1478]. radiocaesium [873]. radiocarbon [512, 646, 1203, 1425, 1475, 122, 538]. radium [512]. Rainbow [1470, 1598, 1195, 1591, 615, 624, 604, 1129, 914, 1663, 270, 269, 511, 1067, 751, 807, 598, 723, 157, 41, 897, 1703, 760, 85, 1755, 447, 1068]. rainforest [1312, 496]. raise [720]. Raman [605]. Rana [31]. ranched [386]. random [244, 671, 1369, 1679, 1641, 514]. Range [893, 1671, 1735, 810, 278, 1376, 902, 976]. Range-wide [893]. ranges [1637, 635]. ranging [1762]. Rapid [200, 1438, 1443, 810, 1526, 671, 889, 947]. rare [240, 1286]. raster [1369]. raster-based [1369]. rate [1302, 1827, 640, 1173, 823, 606, 1586, 125, 1020, 1673, 1807, 1286, 1149, 741, 1359, 405, 731, 917, 237, 999, 1165, 232, 786, 1167, 1635, 839, 760, 1405, 1159, 1306]. Rates [214, 607, 1613, 421, 1381, 705, 1421, 1085, 446, 1257, 1336, 1842, 679, 461, 401, 1197, 1742, 1758, 626, 1011, 958, 1623, 1002, 1073, 1189, 1843, 1455, 962, 1056]. rather [122]. ratio [209, 207, 1232, 416]. ratios [211, 1227, 1251, 477, 1510]. ray [432]. rays [1536, 1165, 1232, 709, 535]. RCRW [1369]. Re [1021, 151, 336, 801]. Re-assessment [151]. re-establishing [336]. re-evaluation [801]. Re-examination [1021]. reach [1381, 964, 867, 1067, 595, 1829, 1159]. reach-level [1159]. reach-specific [867]. reaches [1478]. reaction [928, 51]. real [420]. reality [1729, 1460]. realized [66]. realm [710]. realms [856]. reared [223, 584, 885, 568, 670, 352, 1597, 731, 791, 480, 1276, 1003, 1501, 802, 873]. Rearing [1264, 1591, 35, 670, 1133, 943, 546, 1276, 541, 760, 1048]. reassessing [769]. rebound [952]. Rebuilding [1547, 231, 406, 1498]. recall [1613]. recapture [376, 1249, 445, 766, 893, 1522, 311, 1807, 837, 1746, 626, 727, 96, 483]. receiver [63]. receivers [1681]. récentes [1387]. recently [37]. receptor [1512]. recession [1708]. recipient [1207]. reciprocal [1215, 194]. recolonization [9, 435]. recombinant [399]. recommendations [192]. Reconciling [1721]. reconditioned [1335]. reconnaissance [518, 698, 1043, 1216, 1410, 1611, 1820, 1261]. reconnection [1736, 737]. Reconsidering [547, 1078]. reconstruct [287]. Reconstructing [794, 83, 210, 1232]. reconstruction [1462, 831, 1702, 483, 1383, 467]. reconstructions [649, 1241]. record [273, 222, 476]. Recorded [1371]. records [1051, 240, 726]. recover [739]. recovered [817]. recoveries [979, 405]. recovering [544, 411]. recovery [782, 228, 235, 1181, 1170, 216, 100, 1249, 1522, 4, 517, 746, 1829, 437, 1588, 1627]. Recreational [906, 652, 1570, 1051, 929, 1543, 318, 332, 1204, 1005, 852, 130, 161, 891, 1550, 384, 170, 1200, 716, 684]. recreational-fisheries [130, 161]. recruit [576, 136, 686, 151, 587]. Recruitment [1071, 630, 1072, 1288, 741, 900, 1507, 1268, 813, 1083, 1143, 223, 584, 12, 1340, 1404, 1061, 534, 400, 1817, 136, 1033, 75, 1427, 1713, 93, 253, 942, 1433, 951, 1317, 119, 848, 1, 725, 511, 1615, 560, 785, 1022, 246, 991, 1176, 290, 995, 566, 289, 1078, 1308, 1294, 90, 442, 676, 1514, 1714, 1531, 1631, 493, 1703, 1641, 799, 1153, 975, 884, 953, 1386, 1683, 1017, 1341]. recruitments [321]. recrutement [1071]. red [987, 572, 117, 69, 1597, 146, 965, 1538, 1501]. redd [1616, 1383]. redds [923]. redfish [1655, 953]. redistribution [1498]. redistributions [1629]. reduce [1749, 250, 478, 117, 1286, 367, 1813, 1500, 989, 1442]. Reduced [35, 223, 1133, 928, 584]. reduces [1469, 1242, 385, 1214, 525]. reducing [1049, 1098, 1390]. reduction [258, 1163, 497]. Reductions [936, 918]. reef [205, 1457, 468, 148, 1406, 425, 1473, 1106, 762, 1772, 757, 1440]. reference [813, 1061, 136, 1401, 538, 1696, 479, 1318, 1237, 1185, 628, 1617, 490, 1300, 730, 486, 462, 262]. refilling [884]. reflect [1109]. reflex [1307]. refuge [1603]. refuges [1379]. refugia [859, 1277, 15, 1499]. regarding [720]. regenerated [34]. regeneration [1469]. regime [150, 1160, 1028, 972, 980]. regimes [590, 833, 858, 1699, 1563, 1304, 963, 1171, 985]. region [1549, 875, 75, 833, 1842, 1282, 1384, 801]. Regional [343, 647, 1196, 508, 1109, 1404, 188, 251, 1235, 1499, 264, 1818, 1668]. Regional-scale [647]. regions [879, 912, 373, 165]. regression [1592, 617]. regulate [977, 924]. regulated [269, 1038, 941]. regulating [264, 722]. regulation [588, 1427, 11, 1648, 741, 392, 561, 447]. regulations [130, 161, 1779, 1631, 1364]. regulators [1527]. rehabilitation [877].

Reinhardtius [580, 197, 784, 1284]. reintroduction [1450, 1137, 1721, 1006]. relate [1607]. related [1110, 1173, 1180, 1626, 1765, 1168, 950, 725, 1226, 1674, 391, 506, 1705]. relates [1667]. relating [598, 656]. relation [847, 1413, 1594, 846, 202, 1821, 147, 1810, 482, 1197, 1576, 626, 957]. relations [1806, 93, 676]. Relationship [1771, 1708, 1479, 1186, 434, 1302, 455, 921, 389, 1011, 958, 1097, 657, 904, 25, 1834]. Relationships [242, 1591, 1143, 1680, 1697, 306, 417, 631, 756, 530, 67, 567, 1121, 151, 1419, 995, 533, 707]. Relative [993, 1553, 259, 559, 48, 1341, 582, 172, 328, 46, 323, 663, 1552, 1292, 156, 309, 726]. release [376, 909, 1459, 1751, 329, 1724, 868, 1825, 1037, 1840, 1776, 575, 525, 1255]. released [120, 1036]. releases [386, 1674, 50]. relevance [881, 1294, 1284]. relevant [1790, 1142, 802]. reliably [52, 1385]. remain [45]. remarks [1105]. remediation [864]. remnant [884]. remote [1603]. removal [449, 658, 824, 435, 174, 514]. removals [1548]. renewed [599]. reorient [1575]. repatriation [1796]. repeatability [362]. replacement [1094]. replacing [1187]. replenishment [1665]. replication [595]. Reply [1140, 73, 516, 27, 177, 374]. reported [1085]. Reporting [1613, 1405]. represent [577]. reprioritizing [1437]. reproduction [1669, 102, 1050, 812, 362, 1317]. Reproductive [1222, 146, 109, 826, 1335, 335, 847, 325, 622, 1731, 143, 562, 81, 202, 328, 702, 1730, 1766, 1025, 290, 1231, 1453, 752, 1332, 1618, 1192, 1188, 1228, 802, 841, 985]. Republic [296]. required [226, 1351]. requirements [1826, 752]. requires [929]. rerio [1482]. research [1245, 1342, 928, 1001, 965, 247, 1497, 1693, 748]. researchers [1352, 899]. reserve [429, 278]. reserves [209, 918, 117, 910, 1018]. Reservoir [884, 393, 34, 573, 1062, 1757, 1458]. reservoirs [1161, 487, 353]. Residence [944, 1517, 597, 842]. residency [547, 53, 914, 142]. Resident [604, 1672, 814, 763]. resilience [1672, 54, 915, 243, 1601]. resilient [1433, 1663, 1375]. resistance [1438, 1042, 243]. resolution [835, 1416, 1473, 399]. resolved [1241]. Resolving [330]. Resonance [426]. resource [1206, 1124, 1158, 124, 1760, 1207, 1209, 541, 1212]. resources [1245, 54, 296, 1423]. respiration [1826]. respiratory [1167]. Response [737, 350, 344, 1448, 1639, 982, 376, 853, 95, 1020, 1014, 532, 380, 1592, 1621, 1436, 1548, 1750, 221, 1689, 943, 1373, 764, 1764, 264, 157, 1260, 983, 990, 897, 297, 1229, 488, 1743, 1100, 1250, 110, 1224, 44]. Responses [1267, 504, 776, 258, 499, 876, 1238, 590, 133, 411, 298, 196, 17, 1801, 349, 598, 1760, 811, 236, 1126, 1212, 1426, 1608, 1657, 475, 1384]. responsible [501]. ressources [1206]. restocked [64]. restoration [1672, 23, 1221, 927, 901, 926, 769, 108, 78]. restored [1841]. restores [199]. restriction [525]. result [290, 1376, 1139]. results [781, 129, 401, 1225, 366, 947]. resurgence [1152]. retardants [1054]. retention [823, 202, 68, 463, 1440]. Retrospective [1146, 991, 978, 1719]. return [1549, 308, 1399]. returning [1240, 1253]. returns [50]. Réunion [1325]. reveal [563, 710, 1756, 362, 548, 720, 1499, 289, 1770, 173]. revealed [1024, 1802, 630, 1432, 460, 591, 1190, 1825, 498]. revealing [1419]. reveals [198, 1147, 1602, 1822, 1221, 694, 1473, 1327, 1702, 868, 1625]. reversion [308]. review [1506, 914, 1558, 1411, 585]. revised [560]. revisited [560]. reward [1623]. rheoreaction [1369, 1441]. rheoreaction-based [1369]. Rhinichthys [859]. Rhizoprionodon [1735]. rich [404, 1622, 477, 1772]. richness [1751, 603, 317]. Ricker [493]. Rico [469, 469]. ridge [139]. riffle [859, 1248]. riffle-dwelling [859, 1248]. right [354]. Riley [1140]. ringens [465]. riparian [1451, 1469, 57, 106, 1243, 1328]. rise [1164]. Risk [1623, 375, 783, 1436, 1599, 489, 1424, 1026, 1416, 883, 408, 1141, 396, 354]. risk-assessment [375]. Risks [348, 1205, 825, 702]. risky [234]. River [1749, 26, 454, 1653, 263, 1427, 1342, 84, 1545, 944, 583, 1330, 1253, 1069, 230, 181, 927, 435, 901, 152, 90, 1093, 1209, 1445, 1145, 1786, 59, 21, 464, 6, 27, 302, 261, 1632, 1634, 156, 1494, 728, 1213, 1549, 1303, 996, 630, 1290, 613, 125, 64, 20, 1643, 171, 1470, 1603, 1316, 1147, 589, 1451, 178, 118, 1517, 1648, 269, 1282, 1030, 385, 1810, 1038, 731, 764, 1636, 448, 430, 522, 260, 1593, 432, 1361, 68, 561, 614, 1130, 1355, 980, 926, 739, 463, 526, 737, 1094, 1625, 1698, 959, 1428, 1046, 1474, 1588, 98, 899]. river [1472, 1478, 855, 1736, 140, 872, 77, 1344, 1586, 770, 1336, 239, 331, 670, 1021, 395, 511, 1067, 1599, 1446, 489, 991, 482, 1039, 310, 9, 1792, 739, 524, 1829, 637, 817, 497, 1068]. river-born [260]. river-floodplain [737]. River-plume [90]. river-spawning [1303]. Riverine [1731, 1582, 334, 1369, 1758, 1732, 1516, 884]. rivers [501, 452, 829, 199, 867, 1528, 797, 1287, 595, 793, 1826, 1235, 954, 984, 1183, 1356, 432, 1301, 110, 1711, 1454]. riverscape [931, 816, 522]. Rivière [1355]. road [1251]. Roanoke [1336]. Rob [1102, 1104]. Robert [1103]. robust [779, 1683]. robustness [479, 1399, 706, 1779, 1405]. rock [16, 277, 17]. Rockfish [1204, 1049, 102, 1694, 188, 139, 1714]. rockfishes [136]. Rockies [625]. Rocky [298, 1148, 968, 139, 1023]. Role [261, 859, 1131, 1050, 857, 853, 388, 1788, 283, 1329, 323, 866, 713, 610, 682, 264, 963, 68, 1139, 268, 714, 755, 1307, 1441, 302]. Roles [1143, 647, 1106]. room [1753]. rostrata [263, 39, 1444]. rostriformis [163]. round [403, 1803, 1158, 395, 55, 1766, 1499, 575]. routes [88]. routing [1593]. rufescens [69]. rule [1794, 1092]. rules [209, 1086, 951, 1333]. Run [1342, 1019, 1462, 1545, 1454, 1730, 1647, 1780]. Run-of-River [1342]. running [221, 1179]. runs [83, 363, 592]. Russia [522]. rusticus [451, 66]. rusty [66].

S [1339]. Saanich [890]. sablefish [908, 1274]. Sacramento [1039, 1494]. safety [1026]. sagax [1, 686, 560, 151, 1402]. saida [1266]. sailfin [219]. Saint [464, 1632, 1634]. salad [1058]. salar [256, 229, 640, 1230, 1838, 1358, 440, 1549, 545, 1586, 613, 452, 64, 259, 783, 171, 1257, 149, 386, 979, 93, 621, 639, 553, 199, 225, 312, 1528, 352, 218, 415, 869, 510, 234, 539, 635, 1133, 147, 680, 1434, 1826, 480, 593, 999, 522, 30, 811, 9, 260, 868, 536, 1598, 19, 526, 297, 1786, 184, 773, 1250, 525, 346, 1695, 467]. salaris [19]. salinities [349]. salinity [956, 582, 1178]. Salmo [256, 229, 640, 1230, 1838, 1578, 1088, 1381, 35, 1358, 440, 564, 1549, 545, 1586, 613, 452, 64, 259, 1004, 783, 171, 1257, 445, 149, 1244, 386, 979, 347, 93, 621, 639, 1630, 885, 553, 199, 225, 312, 1528, 352, 218, 415, 869, 1035, 510, 234, 190, 539, 635, 1133, 358, 147, 680, 1434, 1826, 642, 480, 593, 999, 522, 30, 423, 811, 9, 260, 922, 1488, 868, 536, 1598, 19, 1476, 526, 297, 1786, 184, 773, 1250, 525, 483, 346, 1695, 467]. salmoides [376, 237]. Salmon [299, 583, 1144, 374, 1372, 204, 1231, 314, 1749, 10, 153, 1019, 1768, 256, 229, 1349, 925, 1508, 882, 1639, 176, 1832, 1604, 1166, 26, 113, 226, 1362, 1520, 640, 166, 195, 1230, 1838, 1736, 738, 894, 65, 83, 242, 821, 1358, 440, 650, 140, 986, 1653, 333, 307, 789, 1549, 1836, 1345, 545, 847, 1462, 1775, 1586, 612, 1680, 613, 1337, 836, 1626, 452, 64, 3, 1070, 1519, 1556, 200, 876, 259, 783, 20, 1643, 1711, 562, 1137, 1717, 171, 666, 576, 1257, 1470, 149, 1603, 1169, 1220, 1542, 571, 386, 7, 177, 1471, 948]. salmon [308, 981, 1564, 431, 1727, 979, 798, 93, 621, 579, 639, 1620, 1127, 770, 845, 552, 239, 553, 1545, 1765, 53, 331, 199, 1168, 251, 225, 950, 202, 312, 1778, 1528, 670, 352, 761, 1801, 1525, 1256, 218, 415, 869, 944, 373, 510, 1240, 1135, 1468, 164, 234, 182, 492, 539, 635, 746, 1133, 548, 778, 702, 1822, 487, 923, 147, 680, 1797, 870, 1434, 132, 1826, 943, 175, 1808, 1022, 945, 1357, 1485, 555, 385, 15, 1304, 546, 482, 1707, 731, 1039, 1752, 1725, 1636, 654, 230, 434, 597, 735, 275, 577, 87, 822, 480, 593]. salmon [310, 1015, 1674, 201, 984, 927, 1661, 999, 1619, 522, 30, 413, 811, 9, 260, 1593, 435, 458, 626, 901, 1167, 165, 1732, 382, 1650, 1301, 1539, 868, 1243, 1209, 363, 536, 1598, 1156, 1214, 1276, 656, 569, 641, 1792, 957, 439, 268, 1097, 19, 1006, 592, 1052, 739, 1145, 1770, 236, 1647, 541, 526, 297, 1840, 1649, 116, 722, 736, 1780, 394, 1291, 1786, 1003, 1002, 842, 814, 763, 25, 1511, 1608, 184, 1423, 1192, 50, 1322, 1740, 773, 1250, 1270, 525, 326, 1555, 1046, 6, 27, 1651, 346, 506, 607, 802, 156, 1588, 933, 841]. salmon [1695, 985, 467, 1453]. salmon-derived [164]. Salmon-mediated [583]. salmonid [1365, 571, 150, 1479, 1135, 1812, 1699, 968, 327, 1221, 733, 1030, 1373, 1810, 1211, 438, 1799, 1154, 1516, 1100, 192, 947, 818, 412, 78]. salmonids [305, 1379, 223, 584, 1802, 1342, 347, 1491, 960, 787, 485, 1676, 827, 1130, 363, 316, 285, 520, 174, 351, 1458, 873]. Salmoniformes [60]. salmonis [281, 182, 175, 268, 1231, 1453]. Salt [956, 474, 133, 53, 1251, 1484]. saltwater [474]. Salvelinus [930, 1060, 564, 1767, 336, 913, 418, 404, 895, 58, 495, 347, 1319, 1016, 150, 415, 1461, 1614, 803, 1750, 1518, 428, 38, 665, 1190, 104, 811, 442, 1598, 1533, 490, 62, 1229, 1328, 79, 350, 976, 1012]. sample [226, 1258, 674, 768]. sampled [124, 1534, 1247]. samples [645, 498]. Sampling [1567, 1554, 967, 1377, 444, 1360, 1239, 1412, 1813, 923, 1406, 559, 413, 1473, 330, 174, 1351, 377, 1544, 359, 1247, 1710]. San [469]. sandbar [1619]. Sander [1570, 1805, 250, 1413, 1771, 1817, 1089, 523, 806, 566, 614, 911, 1467, 94]. sandercyanin [911]. sandfish [219]. sands [780]. sandwich [965]. Sandy [888]. Santa [1346]. sapidissima [1827, 77, 765, 92]. Sardina [498]. sardine [1116, 1, 686, 560, 151, 498, 1402]. Sardinops [1, 686, 560, 151, 1402].

Saskatchewan [1069]. satellite [635, 1770]. satisfaction [929]. saturation [1654]. sauger [1413]. sawsharks [1041]. saxatilis [1586, 1226, 378, 392, 459]. say [1646]. scalability [1309]. Scalable [1252]. scale [1749, 519, 855, 609, 1557, 1422, 710, 1694, 1716, 1784, 835, 1143, 1756, 964, 540, 1158, 1711, 844, 1169, 253, 1620, 1016, 1287, 1266, 1518, 1624, 1517, 367, 831, 353, 577, 1606, 425, 647, 927, 1527, 1576, 247, 414, 295, 533, 1789, 1040, 1791, 1077, 483, 506, 1755, 1386, 1632, 678, 1494, 255]. scaled [1573, 1552]. scales [112, 1109, 781, 661, 1594, 846, 797, 561, 382, 922, 536, 773, 1232, 709, 535]. Scaling [781]. scallop [1403]. scallops [155, 1706, 1300]. scan [1261, 509, 1571]. scans [460]. Scaphirhynchus [432, 1142]. scattering [820]. scavenging [1191, 629]. scenarios [643]. schedule [400]. school [465, 928]. Schooling [1482]. schools [478]. Sciaenops [965]. Science [687, 688, 1102, 544, 673, 1104, 1722, 1571, 1577, 257]. Sciences [469, 257, 1102]. scientific [1305]. scorpius [1637, 1689]. Scotia [1664, 579, 655]. Scotian [1237, 80, 437, 632]. Scotland [606, 1489]. Scottish [1549]. scour [62]. screens [805]. sculpin [795, 1637, 1689, 871]. sculpins [871, 814, 763]. Sea [981, 661, 1454, 182, 201, 1208, 1166, 675, 1311, 824, 281, 790, 1438, 28, 1697, 386, 1564, 867, 1045, 1548, 1479, 1080, 643, 1021, 13, 681, 591, 831, 175, 1053, 764, 1414, 800, 1764, 1706, 374, 1702, 704, 1417, 268, 1533, 1229, 1300, 1501, 1734, 773, 1403, 224, 424, 1177, 379, 1382, 1128, 217, 400, 588, 1459, 81, 619, 1065, 1701, 292, 119, 1456, 301, 280, 1318, 311, 672, 548, 1262, 720, 1746, 1162, 921, 441, 820, 1685, 1804, 703, 476, 1073, 1077, 219, 1601, 1228, 1292]. sea-accessible [1533]. Sea-to-sea [1454]. sea-winter [773]. seabed [1442]. Seabird [1531, 102]. seabirds [1191]. seafloor [1416]. seagrass [471]. seal [1181, 1098, 1674, 1489]. seals [320, 1322]. Séance [469]. seas [314, 1711]. seascape [1602, 1576, 1789]. seascape-scale [1789]. Season [1307, 1653, 1491, 732, 1725, 1718, 262]. Seasonal [775, 333, 252, 1080, 696, 1210, 1183, 682, 1645, 1618, 346, 1363, 1634, 1068, 558, 403, 1272, 1273, 1549, 1115, 1062, 1253, 484, 554, 114, 1466, 792, 911, 1647, 1126, 145, 1755, 497, 632, 288]. Seasonality [755]. Seasonally [1241, 1209]. seasons [1576]. seaward [1047, 739]. seawater [811, 1539]. seawater-acclimated [811]. Sebastes [1694, 136, 188, 724, 1655, 91, 139, 953]. Secondary [456, 1832, 187]. sedentary [1309]. Sediment [1242, 808, 120, 327, 598]. sediments [872]. seed [568]. seeking [593]. segmented [617]. segregation [895, 1753]. seine [1604, 1007, 786, 9]. seiners [339, 1589]. seismic [659, 461]. selected [583]. selecting [721, 830]. Selection [505, 166, 195, 1704, 705, 847, 574, 170, 767, 267, 747, 1516, 1048, 412, 1494, 616]. Selective [1281, 1275, 1448, 1508, 1264, 1638, 1343, 1813, 559, 220]. selectivity [1487, 1815, 1717, 1748, 1227, 679, 701, 1285, 1813, 1271, 254, 2, 300, 1659]. selects [1663]. seleniris [291]. Selenium [121, 343, 968]. Self [1092, 838].

Self-starting [1092]. self-thinning [838]. semi [946, 554, 1659]. semi-anadromous [946]. semi-mechanistic [554]. semi-parametric [1659]. Semidiscrete [499]. seminatural [352]. semiparametric [862]. semipelagic [557, 1429, 1262]. sensing [1603, 797]. sensitive [23, 1496, 54, 11, 1125]. Sensitivity [99, 1539, 443, 706, 1562]. sensory [797]. separate [1785]. Separating [900, 1340, 1289]. separation [254]. Sepioteuthis [289]. septicemia [1138]. sequencing [778]. Sequential [1710, 370, 662, 668, 505]. SERCA [1639]. series [215, 1084, 407, 390, 506]. service [1391]. session [469]. s’est [469, 1071]. sets [1747]. Setting [1300, 540, 1801]. settlement [695, 1666, 1473]. settling [120]. sevenfold [1161]. seventh [1474]. seventh-order [1474]. several [266]. severely [724]. severity [1541]. Sex [1805, 419, 1415, 1307, 205, 1475, 263, 1470, 1812, 1251, 1449, 723, 1618, 467]. Sex-based [1805, 1415]. sex-changing [205]. sex-specific [1449]. Sex-Structured [419, 263]. sexual [696, 1251, 412]. shad [1827, 77, 258, 765, 92, 971, 660, 1700]. shading [1621]. shale [1759]. shallow [977, 1781, 1732, 475]. shape [1420, 1027, 320, 1184, 707, 602]. shaped [1097]. shapes [1452]. shaping [1838]. share [1572, 1352]. shared [1297, 541]. shares [169, 1839]. sharing [856, 1788, 1352]. shark [705, 74, 1339, 537, 1324, 699, 915, 1259]. sharks [952, 1051, 1557, 1205, 333, 1095, 1735, 1238, 342, 1324, 1807, 82, 124, 1165, 786]. sharpnose [1735]. shed [823]. shelf [1128, 1076, 792, 767, 887, 1237, 80, 437, 632]. shell [788]. shellfisheries [713]. shelter [593]. shelter-seeking [593]. Shield [719, 1830, 1170, 494, 1319, 422, 293, 1833, 48]. shift [1427, 889, 1160, 972, 1037, 346]. shifting [1275]. Shifts [1116, 1671, 1481, 653, 1493, 694, 1183, 1028, 1476, 1123, 1698]. shiner [1024, 674, 146]. ship [409]. Shirreff [103]. Shoaling [1512]. shoals [315]. shore [1414]. shoreline [494]. Short [1575, 387, 1199, 405, 237, 1305]. short- [1199]. short-finned [1305]. Short-term [1575, 387, 405, 237]. shortcomings [173]. shortfin [1393]. shorthorn [1637, 1689]. shortnose [1464, 547, 464]. should [122, 1090]. shovelnose [893]. show [1041, 1126]. shows [1257]. shrimp [1177, 1394, 849, 1662, 407, 390, 68, 1442]. shrimps [502]. shy [405]. Si [249]. sibship [863, 467]. sibship-based [863]. side [1261, 509, 1571]. side-scan [1261, 509, 1571]. Sierra [1451, 31]. sierrae [31]. signal [1131]. signals [1288, 983, 1641]. signature [440, 1172]. signatures [279, 281, 1646, 627, 591, 555, 82, 1419, 709]. significance [328]. significant [112, 409]. significantly [324]. Siletz [1545]. siliquoidea [1657]. silky [786]. silver [855, 1605, 1267, 700, 291]. silverside [1331, 1561]. Simcoe [360]. similar [983]. Simple [438, 545, 544, 1227, 654, 842, 773]. simulate [1779]. simulated [420, 1837, 221, 1840, 715]. Simulating [123, 1408, 144]. Simulation [1405, 863, 974, 1273, 540, 825, 935, 1399, 706, 1234, 1009, 969]. simulation-based [706]. simulations [115, 789, 532]. simultaneously [2, 300, 748]. simus [1024]. since [1552]. sinensis [1280]. Single [548, 778, 1511, 113, 1520, 1549, 1680, 291, 169, 1397, 405, 315, 1255]. single- [1397, 315, 1255]. Single-nucleotide [548, 778, 1511]. single-pass [1680]. sinking [120]. siscowet [754]. sister [60]. Site [31, 1060, 601, 808, 563, 847, 1727, 1609, 697]. sites [36, 880, 1533]. situ [45, 1193, 810, 694]. six [1108, 1256, 1082, 103]. Size [1102, 1104, 1748, 1168, 1807, 82, 44, 1113, 920, 1108, 1302, 1830, 1110, 1258, 1264, 1487, 1815, 1396, 1653, 1116, 1272, 1273, 1430, 1109, 1638, 1115, 1124, 1129, 1553, 259, 1541, 291, 1169, 1220, 1443, 1673, 455, 1505, 568, 1227, 1451, 1001, 1256, 372, 1614, 1393, 1648, 71, 744, 1251, 1813, 837, 1235, 1730, 349, 998, 1107, 1744, 1811, 559, 713, 1197, 660, 883, 423, 1534, 1167, 1702, 536, 1106, 316, 285, 1271, 1231, 1453, 433, 186, 1531, 904, 1123, 254, 340, 1114, 1105, 1780, 351, 134, 486, 1188, 1100, 707, 655, 1111, 602, 1834, 447]. size [832, 313, 1112, 1125, 1307]. size-at-age [1534, 433, 1834]. size-at-hatch [1811]. size-at-maturity [1505, 349]. Size-based [1102, 1104, 82, 259, 1125]. size-dependent [1227, 1100, 313]. size-divergent [1673]. size-limit [998]. Size-related [1168, 1110]. size-selective [1264, 1638, 1813, 559]. Size-specific [1807, 660]. size-structured [1272, 1273, 1393, 744, 1744]. size-transition [837, 1744]. sizes [226, 768]. skate [75, 1816, 543]. skates [1536]. skill [1460, 439]. skip [1335]. skipjack [198, 758]. skipped [1630]. slackwater [561]. slow [1793]. slow-growing [1793]. Small [483, 912, 1422, 1587, 844, 1319, 931, 1481, 1624, 1517, 717, 915, 1347, 247, 543, 1364, 1125]. small- [844]. small-bodied [931, 915]. Small-scale [483, 1422, 1624, 1517]. smallmouth [334, 1344, 1186]. smelt [284, 1703]. Smith [1629, 193]. smolt [415, 1353, 358, 1276, 490, 50]. smoltification [571]. smolting [288, 1035, 25]. smolts [1586, 386, 948, 1607, 1528, 670, 1133, 546, 626, 116, 722, 1003, 1002, 525]. smooth [407, 390]. snail [451]. Snake [1454, 637]. snapper [512, 117, 466, 232, 1538]. snorkelling [1246]. snow [1396, 878, 1664, 400, 703, 1228]. SNP [1570, 1526, 1279, 1695]. SNPs [291, 778, 1526]. social [35, 1716, 544, 943, 660, 1721]. socially [130, 161]. Sockeye [308, 1639, 1832, 1604, 113, 226, 65, 242, 821, 1775, 1556, 1711, 1717, 1169, 1220, 1542, 981, 1564, 431, 798, 239, 202, 1240, 385, 15, 482, 1636, 458, 382, 1650, 957, 541, 1840, 1649, 842, 1627]. soft [71]. soft-water [71]. software [1233]. softwater [1622]. Solar [1001]. sole [1572, 538, 17, 1788]. solution [565]. solve [544]. somatic [1302, 1269, 1799]. some [1105]. Somniosus [124]. sonar [765, 1261, 509, 1571]. sorting [1748, 1825, 983]. Sound [200, 668, 1336, 731, 1629, 1461, 273, 222, 871, 434, 193, 689]. sounder [1480, 1305]. Sounds [461, 1764]. Source [977, 1714, 1507, 1766, 374]. Sources [192, 563, 630, 1654, 1465, 1183, 1419, 289, 340, 755]. south [1836, 719, 945, 1833, 610, 264, 472, 1402]. south-central [719, 1833, 610, 264, 472]. southcentral [1304, 976]. southeast [681, 204, 515, 286, 516]. southeastern [829, 507]. Southern [823, 1200, 1396, 1694, 1638, 1337, 555, 1176, 1610, 289, 1816, 745, 712, 542, 134, 794, 1843, 398, 1083, 579]. southernmost [1768]. Southwest [590]. southwestern [1339]. Space [1387, 1388, 1389, 1192, 1310, 1381, 1085, 1681, 576, 532, 1089, 867, 1522, 284, 1176, 1486, 1655, 1134, 1246, 1818, 402, 1000, 1040, 276]. space-use [1040]. SPAM [419]. span [1203, 920]. Spanish [742]. sparid [232]. Sparidae [1733, 466]. Spatial [1388, 403, 1587, 1519, 648, 171, 1257, 1432, 672, 874, 803, 1391, 527, 610, 1093, 266, 1417, 862, 1055, 1667, 655, 112, 1387, 1034, 749, 1101, 1389, 1658, 1180, 1686, 912, 1177, 144, 123, 1143, 1756, 1712, 1290, 1248, 1404, 23, 1681, 540, 961, 1392, 427, 666, 669, 491, 1401, 661, 417, 1027, 551, 651, 1594, 846, 1765, 252, 311, 1075, 1252, 1287, 567, 1299, 1408, 293, 1839, 190, 1222, 1397, 595, 1398, 1678, 1210, 1550, 923, 807, 1329, 1810, 910, 323, 1460, 866, 423, 384, 330, 382, 727, 922, 704, 767, 267, 592, 533, 1300, 1437]. spatial [59, 21, 1791, 1077, 506, 1787, 1726, 1386, 1363, 497, 828, 1068, 348, 729]. spatial-stream-network [1252]. spatiale [1387]. Spatially [678, 853, 215, 935, 1713, 1089, 1615, 1406, 1593, 1679, 1407, 839, 565, 173]. spatiaux [1387]. Spatio [728, 744, 536]. Spatio-temporal [728, 744, 536]. Spatiotemporal [1049, 1803, 1394, 887, 1393, 1445, 116, 1670, 424, 1656, 1578, 1396, 835, 903, 1733, 279, 1191, 1399, 1359, 1747, 1584, 1055, 1395]. spawn [197]. spawned [1687, 559]. spawner [183, 654, 363, 1006, 1383]. spawners [666]. Spawning [1060, 167, 486, 1805, 1380, 1506, 930, 808, 986, 1303, 1733, 1731, 1378, 648, 1687, 318, 445, 404, 1630, 931, 1456, 1227, 312, 1461, 725, 158, 304, 1615, 385, 1076, 413, 1335, 316, 285, 62, 1618, 1649, 184, 464, 1444, 802, 1036, 1494, 1017, 1761, 632]. Special [1071, 469, 1061, 1777, 1071]. spéciale [469]. specialist [1441]. specie [1376]. species [906, 1108, 929, 558, 1837, 1049, 1354, 1487, 299, 705, 1640, 1268, 1321, 993, 481, 279, 578, 117, 540, 534, 1438, 1309, 905, 1007, 306, 411, 766, 150, 631, 638, 531, 345, 1479, 530, 67, 567, 169, 1785, 1042, 320, 1517, 1648, 557, 1429, 4, 517, 1532, 1662, 1824, 461, 1754, 66, 890, 1606, 1568, 1416, 711, 338, 1134, 1041, 266, 656, 32, 1831, 1772, 1057, 907, 1741, 752, 1163, 533, 377, 862, 1064, 1418, 315, 1094, 1390, 1296, 953, 1364, 1091, 274, 302, 261, 607, 816, 98, 1530, 1709]. species-specific [638, 461, 1296]. specific [1604, 1354, 974, 667, 95, 43, 638, 867, 1237, 1807, 451, 635, 461, 401, 1449, 1347, 660, 694, 1647, 1296, 962]. specificity [443]. speckled [646, 931]. spectra [1116, 1109, 1193, 1107, 1111, 1112]. spectral [1193, 460, 810]. spectroscopy [1701]. spectrum [1113, 1830, 1115, 1114, 1105, 832]. speed [1591, 564, 1528, 1446, 707]. sperm [51]. spicatum [66]. spiders [1451]. spillovers [1644]. spills [1348]. spines [1232]. spinner [1058]. spiny [368, 429, 669, 792, 767]. split [1589]. sponges [1823]. sport [367]. spotted [150]. spp [314, 10, 176, 1694, 576, 875, 136, 7, 177, 251, 245, 530, 1655, 413, 363, 139, 953, 326, 607]. sprat [455, 1318, 1568]. Sprattus [455]. spread [32]. spring [1035, 158, 559, 603, 1301, 1006, 1145, 156]. spring- [559]. sprint [1827, 1591]. sprinting [564]. Squalus [669, 792]. square [1487, 1748]. square-mesh [1487]. squid [920, 1027, 1001]. Sr [416, 1712, 555, 210, 1232, 416]. Sr/ [416, 1712, 1232]. Sr/Ca [210]. SS [1092]. SS-CUSUM-HCR [1092]. SSN [1252]. ssp [859]. St [855, 1605, 1396, 1638, 1330, 231, 406, 1610, 682, 1534, 712, 366, 134]. St. [855, 872, 1344, 1021, 395, 489]. stability [26, 1712, 54, 431, 1424, 434, 954, 1076, 611, 1145, 1612, 134, 6, 27]. Stable [875, 551, 342, 1354, 403, 47, 563, 1278, 1382, 281, 849, 1737, 82, 1560, 1492, 1419, 1510, 340, 79, 145, 947, 1292, 40, 535]. stage [1838, 823, 1556, 662, 1824, 1568, 90, 641]. stage-structured [662]. stages [1365, 217, 270, 1315, 68, 617]. stakeholder [1728, 1795, 594]. standard [640, 1149, 999, 760, 1700]. standardization [1835, 1394, 878, 1239, 1359, 1463]. standardized [421, 208]. standardizing [705]. standing [876]. starting [1092]. starvation [843, 851]. State [1681, 639, 1818, 1310, 1381, 1085, 335, 54, 576, 1089, 867, 1522, 284, 1176, 1486, 1655, 1134, 1099, 1000, 276, 1345]. State-dependent [639, 335]. State-space [1681, 1818, 1310, 1381, 1085, 576, 1089, 867, 1522, 1176, 1486, 1655, 1134, 1000]. states [1245]. stations [63, 886]. Statistical [1175, 967, 1835, 974, 581, 357, 271, 75, 1084, 383, 1327, 533, 772, 1112]. statolith [1027, 591]. status [652, 76, 1401, 240, 251, 1778, 880, 1092, 365, 414, 297, 410]. status-quo [1092]. steelhead [1264, 29, 140, 273, 222, 288, 1607, 1545, 944, 1454, 914, 1353, 328, 1253, 1526, 923, 791, 1616, 1197, 1504, 310, 181, 723, 883, 1619, 1335, 1724, 901, 542, 760, 637, 1048, 1159, 1306, 1726, 607]. steep [188, 46]. steepness [143, 628]. stellate [1374]. stellatus [1374]. Steller [824]. stenolepis [1825, 1834]. stepwise [1234, 755]. stereo [139, 677]. stickleback [338]. stimulates [57]. stimulating [1459, 1285]. STM [554]. stochastic [1234]. Stock [1071, 1388, 1, 1072, 1610, 676, 652, 1310, 1387, 113, 226, 1362, 1520, 370, 111, 1507, 939, 1146, 1085, 1396, 1394, 1143, 499, 1181, 700, 1404, 961, 419, 1720, 576, 136, 357, 271, 129, 935, 1401, 1027, 651, 1713, 93, 253, 92, 1147, 951, 889, 568, 848, 1773, 686, 1001, 679, 1409, 1624, 270, 1185, 231, 406, 1222, 778, 1615, 1406, 1184, 628, 664, 594, 151, 910, 1279, 1682, 921, 742, 711, 128, 995, 1028, 1747, 587, 1407, 1348, 414, 1497, 1514, 1531, 493, 1647, 1218, 131, 969, 768, 714, 1670, 1489, 1403, 959, 1270, 497, 1455]. stock [1017, 636]. Stock-recruitment [1]. stock-specific [1647]. Stocked [324, 1605, 380, 1295, 428, 371, 260]. stocking [1358, 1340, 848, 852, 1597, 807, 642, 1786, 1023, 1627]. stocks [1071, 1387, 1146, 1272, 1273, 52, 198, 1626, 446, 419, 1415, 825, 1033, 1147, 951, 1065, 1227, 1725, 663, 1082, 1655, 1294, 1772, 19, 433, 969, 1489, 219]. stoichiometry [494]. Stokes [277]. stomach [248, 403, 1564, 1432, 192]. stomias [379]. stony [317]. storage [1420, 979, 363]. stored [1253]. straddling [54]. strains [1195, 897, 490]. Strait [1594]. Straits [148]. strategies [967, 1264, 1450, 789, 1542, 994, 39, 1324, 1408, 539, 826, 1679, 1040, 1791, 1601, 1719, 1255]. strategy [1774, 1775, 217, 1729, 1728, 1323, 1795, 1777, 1222, 943, 1529, 1407, 19, 412]. stratification [1350, 554]. stratified [244, 1679, 377]. stratified-random [244]. straying [607]. Stream [351, 99, 420, 1236, 299, 986, 857, 1672, 1756, 1587, 23, 167, 877, 905, 496, 187, 347, 240, 1696, 1491, 1252, 42, 674, 178, 118, 803, 1198, 1759, 1242, 733, 881, 1357, 1493, 37, 1567, 1334, 1197, 647, 1015, 984, 408, 963, 1298, 1260, 645, 901, 922, 745, 1745, 363, 1008, 1476, 739, 769, 542, 1212, 814, 763, 85, 59, 21, 483, 224, 1208, 78, 1472, 255]. stream-adjacent [1236]. stream-dwelling [347, 42, 363, 1476]. stream-finding [224]. stream-living [963, 483]. stream-resident [1672, 814, 763]. stream-size [351]. stream-type [739]. streamflow [1047, 1752]. streams [888, 1806, 1110, 1653, 1155, 138, 135, 819, 876, 877, 1244, 1621, 795, 1252, 515, 1469, 1135, 46, 583, 871, 1144, 286, 516, 1304, 1347, 610, 1419, 1598, 1214, 204, 473, 1328, 625, 344, 24]. streamwater [1312]. strength [392, 1052]. stress [544, 95, 1020, 1013, 1663, 1537, 349, 1276, 1200, 1250, 1575]. stress-resilient [1663]. stressor [360]. stressors [187, 1822, 943, 264, 1384, 1224]. strike [1837]. striped [76, 1586, 1336, 1226, 378, 392, 330, 459]. strong [953]. strongly [984, 941]. strontium [1303, 1290, 448, 430, 922, 416]. Structural [1569, 664, 587, 236]. Structure [529, 1108, 519, 558, 738, 138, 1735, 1404, 946, 88, 829, 1401, 1027, 551, 1751, 84, 1016, 724, 589, 1287, 67, 1518, 1118, 1648, 71, 1397, 1184, 1499, 1610, 104, 660, 522, 30, 330, 382, 1106, 1497, 809, 1823, 838, 391, 712, 524, 295, 493, 1123, 969, 1511, 59, 21, 959, 1667, 953, 483, 1284, 317, 602, 348, 1017, 467]. Structured [419, 1450, 319, 1272, 1273, 853, 263, 576, 662, 215, 889, 1089, 1227, 311, 672, 1393, 744, 1615, 664, 1744, 1176, 1486, 1679, 704, 493, 772, 1189, 101, 678, 1659]. structures [1800, 651, 1594, 901, 1163, 1306, 78]. structuring [46, 293, 1573, 302, 261, 1292]. studied [1484]. studies [1738, 440, 1727, 1690, 1538]. study [1832, 705, 1549, 721, 830, 376, 1095, 1174, 117, 540, 1720, 825, 445, 948, 360, 1701, 679, 685, 451, 713, 742, 1577, 114, 1407, 752, 463, 1009, 1023, 192, 1791, 599, 1341, 462, 1224]. studying [551, 899]. sturgeon [1464, 1506, 1640, 1731, 547, 233, 1374, 362, 893, 1330, 246, 991, 805, 1142, 839, 524, 186, 464, 497, 1632, 1634, 1494, 1769]. sturgeons [432]. stutchburyi [1150]. style [1327]. subalpine [755]. subarctic [977, 20, 1643, 553]. subdivision [1499]. subject [1406]. sublethal [1657, 992]. subpopulations [1061]. Subsampling [1679]. subsequent [1061, 1225, 1250]. subsidies [1206, 1215, 1207, 1551, 1214, 204, 1363]. subsidy [1213, 761, 1212]. subsistence [894, 84]. substantial [1702]. Substock [81]. substrate [246, 991, 1769]. substrates [404]. Subtropical [74, 1669, 1110, 1367, 243]. Suburbanization [1481]. subyearling [53]. success [166, 195, 1704, 675, 847, 1731, 1556, 562, 1817, 1433, 422, 328, 1133, 1597, 1221, 1730, 1053, 1752, 1417, 1006, 1192, 802, 816]. successful [1738, 1721]. successive [197]. sucker [782]. suckers [601]. suckleyi [669]. sufficient [1408]. suggests [188, 1097, 274]. suitability [1330, 1584, 471, 989]. suitable [912]. sum [1092]. Summer [1304, 541, 47, 606, 1545, 1483, 874, 1699, 1513, 559, 984, 1742, 1328]. summer-spawned [559]. summers [103]. summertime [1603]. sunfish [325, 1517, 241]. superba [103]. Superior [919, 490, 265, 1229, 49, 832]. superstorm [888]. supplementation [42, 126, 1555]. supplemented [1052]. supplementing [825]. supply [1059, 1124]. supply-driven [1059]. support [1059, 1573, 60, 1018, 1789]. supporting [1183, 1209, 784]. suppresses [1814]. surface [388, 1211, 1513, 762]. Surfing [1105]. surgical [997]. surplus [1828]. surprising [1484]. surrounding [940]. surveillance [586, 1064]. survey [319, 824, 216, 1480, 292, 1237, 679, 1654, 1139, 703, 315, 1079]. survey-based [292]. survey-comparable [1480]. surveyed [803]. surveys [444, 244, 1694, 659, 1005, 1252, 1429, 1654, 1288, 565, 1383]. Survival [1528, 1825, 50, 808, 35, 140, 1756, 1767, 700, 200, 1074, 1033, 948, 579, 1765, 950, 848, 523, 1491, 670, 1483, 1479, 1807, 803, 1454, 1353, 269, 539, 1286, 690, 147, 741, 870, 327, 1030, 385, 1810, 246, 731, 791, 1039, 1752, 434, 597, 1326, 466, 1029, 310, 883, 1165, 1593, 587, 727, 1650, 1156, 1792, 730, 1003, 1073, 1313, 1608, 1783, 678, 1769, 1255]. survivals [1825]. survivorship [1647]. Susitna [1209]. Suspended [598]. suspension [1289]. suspension-feeding [1289]. sustain [384]. sustainability [1119]. sustainable [856, 1010, 1677, 274]. sustained [1661]. swampy [674]. Swedish [1371, 1032]. swelling [947]. swim [325]. swimbladders [426]. Swimming [1413, 1827, 175, 735, 201, 805, 1482]. switch [1263, 1467]. switching [563, 238]. Sydney [1664]. sympatric [1158, 895, 415, 1517, 896, 1831, 1259]. Symposium [1071, 1072, 688, 1102, 1104, 687]. synchronization [1374]. synchronize [102]. Synchrony [1202, 218, 1365, 1088, 1268, 1386]. synergies [1294]. synthesis [673, 1690, 270]. synthesized [643]. System [1424, 339, 1059, 835, 630, 364, 1470, 360, 889, 1451, 1431, 367, 1515, 282, 596, 614, 677, 1831, 1003, 1625]. System-level [1424]. systemic [598]. systems [1572, 1116, 1148, 1579, 1534, 1758, 936, 1472].

table [844]. tables [168]. tactics [325, 1766]. tadpoles [1286]. tag [823, 427, 935, 1399, 837, 405, 1746, 1405]. tag-based [427]. tag-integrated [935, 1405]. tagged [908, 1528, 63]. tagging [1362, 1520, 728, 1336, 1635, 762, 637, 366, 599, 636]. tags [1420, 1362, 1660, 823, 979, 635, 181, 922, 1825, 1770, 959]. Taihu [1126]. taille [1102]. tailwater [1129]. take [209, 196, 1406, 1018]. Taking [92, 1324]. tank [593]. Target [1463, 318, 1007, 13, 1090, 1322, 1455]. Target-based [1463]. Targeting [478, 23, 1839, 266, 1395]. targets [540, 666, 241]. Tasmania [305]. taxon [623]. taxon-weighting [623]. taxonomic [1565, 1282, 145, 486]. Taylor [1247, 1710]. technical [1529]. technique [544, 241]. techniques [1382, 1558, 1233]. technological [337]. technology [509, 1571]. Telemetry [868, 855, 1664, 1174, 1586, 1720, 1007, 1603, 948, 668, 1579, 1609, 693, 1190, 1606, 1352, 1577, 1635, 1755, 678]. telemetry-based [948]. teleost [653, 170]. teleosts [124]. Temperate [1576, 1669, 879, 615, 1557, 205, 853, 1312, 905, 468, 1244, 496, 207, 196, 994, 1602, 1517, 966, 1210, 934, 463]. Temperature [381, 860, 560, 1563, 1141, 1488, 1094, 920, 1827, 640, 1591, 1838, 1088, 1784, 721, 830, 545, 105, 1020, 1603, 582, 1178, 1317, 1696, 172, 686, 510, 1226, 185, 1797, 733, 1235, 385, 15, 1304, 1493, 151, 1636, 1194, 554, 1513, 999, 22, 423, 811, 641, 1231, 1453, 297, 542, 760, 1649, 1782, 25, 1743, 1608, 815, 818, 828, 924, 985, 1769]. Temperature-dependent [1094, 686, 1194]. temperature-profile [554]. temperature-related [1226]. temperatures [1433, 1692, 1504, 1366, 992]. template [858, 963]. Temporal [1290, 13, 1325, 1617, 866, 384, 498, 1787, 112, 918, 26, 728, 1180, 1756, 1712, 1519, 1269, 719, 648, 1541, 1027, 551, 1696, 1765, 69, 744, 595, 1550, 793, 1810, 915, 527, 610, 30, 1473, 330, 536, 433, 592, 134, 6, 27, 655]. Temporally [1292]. temporary [1452]. tench [1439]. tendency [1573, 525]. tenu [469, 1071]. tenuis [391]. term [508, 1421, 135, 1248, 100, 537, 1666, 1727, 1199, 644, 360, 141, 479, 1599, 451, 1596, 1198, 1221, 1493, 405, 353, 1833, 237, 1076, 1600, 1757, 1490, 1260, 1093, 1099, 1625, 1575, 387]. terre [1206]. terrestrial [57, 1347, 1745, 1209, 980, 1551, 1214, 351, 513]. Territorial [1151, 1783]. territoriality [347]. Territories [1622, 1518, 84]. territory [447, 1189]. Terry [1739]. test [1784, 1309, 1717, 510]. Testing [115, 369, 1268, 935, 122, 966, 647, 128, 799, 1153, 1405]. tetraploids [1012]. Tetrapturus [76]. Texas [537]. Thailand [1613]. their [878, 1556, 940, 869, 60, 796, 1233, 834, 1831, 1011, 958, 340, 750, 1291, 575, 1306, 219]. theoretical [1113]. theory [1260, 1057, 907, 1105, 714, 799, 1153]. Theragra [379, 619, 301, 280, 311, 387]. Thermal [61, 1811, 51, 1212, 1782, 314, 1762, 1379, 1223, 95, 1020, 1603, 1244, 631, 1592, 1277, 17, 696, 1699, 1288, 966, 15, 1633, 1171, 1740, 1250]. thermally [695, 1653]. thermocline [203]. thermoregulation [1466]. Thermoregulatory [1762]. these [754]. thiamine [442]. thiaminolyticus [442]. thinning [838]. thorny [1816]. threat [354]. threatened [1421, 1137, 291, 883, 839, 752, 506]. Three [402, 1544, 1837, 974, 138, 1095, 853, 950, 1654, 170, 1328, 815, 1489, 1709, 1494, 1437, 1829]. Three-dimensional [402]. threefold [1161]. threespine [338]. thresher [699]. thresholds [532, 317]. throughout [600, 1138, 147]. Thunnus [823, 335, 973, 343, 88, 215, 854, 850, 56, 1200]. Thymallus [1377, 1704]. thynnus [335, 88, 56]. Tidal [53, 474, 944, 1163]. Tidally [1523, 1593]. tiger [82]. timber [515, 286, 516]. time [900, 1613, 375, 1729, 233, 1158, 781, 215, 208, 889, 1160, 1084, 221, 1725, 1194, 407, 390, 1828, 1593, 1246, 1818, 414, 506, 1659]. time-varying [375, 889, 1160, 1828, 1659]. times [944, 1552, 1231, 1453]. timescale [1809]. timid [241]. Timing [622, 415, 1019, 600, 1549, 362, 639, 460, 202, 1240, 1730, 15, 1676, 80, 1618, 1649, 116, 722, 1780, 464, 1046]. Tinca [1439]. tiré [1071]. Tissue [401, 1020, 1263]. Tissue-specific [401]. tissues [782, 47, 1467]. TN [657]. toad [1286]. together [1031]. tolerance [1484, 268]. tolerant [1314].

Tonic [997]. tool [456, 1720, 942, 965, 128, 210, 1531, 1390, 274]. tools [1374, 247, 1595]. toothfish [519]. top [1180, 1160, 597]. top-down [597]. topography [940]. Tornionjoki [670]. Total [1367, 1585, 1170, 671, 207, 172, 610, 614, 617, 514]. tows [1706]. toxic [1164, 472, 656]. TP [657]. trace [260]. tracers [938]. Trachurus [1073]. tracing [1243, 340]. track [281]. Tracking [1722, 1036, 115, 1712, 1678]. Trade [812, 251, 1032, 1783, 1195, 1572, 1450, 375, 1415, 332, 708]. trade-off [708]. Trade-offs [812, 251, 1032, 1783, 1195, 1572, 1450, 375, 1415, 332]. traditional [1338, 173]. traffic [1637, 409]. tragedy [160]. training [1501]. trait [1282, 623]. traits [205, 624, 1626, 446, 81, 950, 751, 1281, 1679, 1123, 462]. trajectories [370, 1596, 788]. trajectory [1575]. trans [973]. trans-Pacific [973]. Transatlantic [938, 399]. Transbaikalia [895]. Transboundary [1205]. transcription [860]. Transcriptome [1556, 1607, 1280]. Transfer [364, 968, 699, 1054, 707, 535]. Transferability [1309]. transferable [160]. transform [1701]. transformational [656]. transformative [1701]. Transgenerational [751, 1740]. transition [837, 1744]. Transmission [1690, 368]. transmitted [750]. transmitters [1174]. transmontanus [246, 991, 805, 524, 497]. transpacificus [284]. transparency [1771, 1527]. transponder [1660, 1546]. transport [993, 1708]. transported [1605]. trap [368, 1496, 643, 405, 1417]. trap-happy [405]. trap-shy [405]. trapezoidal [654]. trapped [1605]. trapping [675, 1045, 1053, 1417]. travail [469]. travel [1593]. trawl [1487, 1815, 244, 974, 781, 588, 1350, 1459, 1237, 1785, 557, 1429, 1839, 1098, 1500, 1825, 677, 254, 703, 565, 1442]. trawlable [244]. trawling [1370]. trawls [1262, 1623]. treated [386, 1135, 901]. treatment [1784, 249]. tree [1093, 533]. tree-based [533]. Trends [1344, 107, 1051, 1365, 235, 216, 537, 240, 527, 1833, 1600, 413, 1490, 1818, 1093, 543, 86, 1046, 1705]. trial [111]. tributaries [55, 717, 1458]. Tributary [1287, 1316]. Tribute [212]. Trichodesmium [476]. tridentata [355]. tridentatus [1047, 1383]. triggerfish [1538]. trip [974]. trip-specific [974]. triploid [897]. Trophic [119, 1054, 1209, 1106, 1354, 1640, 1059, 364, 14, 551, 187, 889, 342, 1451, 1750, 1737, 870, 401, 915, 1179, 1419, 1041, 365, 1099, 1645, 904, 1667, 1363, 521, 570]. trophy [332]. tropical [512, 1557, 144, 123, 1290, 1007, 1512, 1408, 1325, 1510, 786, 1445, 1645, 1589]. trouble [656]. troubled [488]. trout [1187, 1195, 256, 229, 1236, 930, 1591, 1264, 1060, 615, 1578, 808, 1088, 1381, 1450, 35, 1050, 1421, 1672, 564, 1767, 624, 604, 1174, 336, 1129, 1642, 1004, 913, 418, 825, 291, 1470, 445, 404, 1244, 332, 1471, 1066, 1320, 754, 1673, 347, 1319, 380, 1432, 1630, 777, 885, 1696, 1277, 1295, 42, 415, 1483, 1842, 1461, 1614, 1035, 803, 1469, 914, 1353, 1750, 1518, 328, 1663, 270, 269, 511, 1067, 1599, 190, 1286, 1723, 1526, 358, 1573, 1210, 751, 807, 642, 38, 791, 598, 776, 448, 430, 371, 310, 864, 1151, 104, 723, 1619, 423]. trout [157, 408, 963, 596, 626, 1724, 1818, 41, 1488, 442, 949, 1598, 838, 1476, 324, 897, 490, 1618, 542, 760, 1229, 718, 1023, 1328, 79, 350, 85, 483, 625, 1726, 1755, 1012, 302, 261, 344, 602, 447, 1068, 255]. trumps [293]. trutta [256, 229, 1578, 1088, 1381, 35, 564, 1004, 445, 1244, 347, 1630, 885, 653, 415, 1035, 190, 358, 642, 423, 922, 1488, 1476, 483]. tshawytscha [925, 26, 166, 195, 140, 3, 1070, 200, 1127, 53, 331, 1801, 492, 548, 778, 132, 230, 597, 822, 310, 1792, 1006, 1770, 25, 1740, 6, 27, 156, 841]. tsunami [1029]. tuna [339, 823, 335, 973, 198, 343, 1007, 88, 938, 252, 1409, 850, 1447, 786, 758, 56, 1589, 1200, 1364, 1036]. tunas [144, 123, 1645]. turbid [701, 793, 620]. turbidity [851, 980]. turbine [1837]. turbulent [620]. Turnover [47, 601, 1248, 1263, 401, 214]. turtle [13]. turtles [887]. twice [288]. two [1838, 8, 5, 73, 789, 823, 1802, 12, 217, 198, 1680, 318, 562, 72, 829, 1800, 58, 347, 298, 1528, 960, 1479, 803, 787, 968, 871, 1824, 1826, 66, 984, 1823, 730, 341, 773, 1259, 1555, 219, 1063, 344, 1472]. two-stage [823, 1824]. type [851, 1685, 739, 1188, 990]. types [1343, 1751, 1522, 1801, 1689, 796, 607]. typhoon [243]. typology [60].

ultraviolet [683]. uncertain [664]. uncertainties [1671]. Uncertainty [1455, 1503, 1086, 1392, 93, 1522, 1409, 664, 614, 587, 1139, 377, 1114, 1569]. Uncovering [1149]. under-reported [1085]. underestimate [520]. undergoing [712]. underlie [1466]. underlies [1096]. underlying [749, 1101, 1109, 1138]. underpin [1498]. underreported [111]. Understanding [1464, 1381, 1603, 1053, 1742, 254, 953, 929, 1430, 456, 1374, 28, 1075, 785, 1635, 1291, 716, 684, 599]. undertake [1557]. Underwater [1098, 1261, 565]. undisturbed [1821]. unequal [1403]. Uneven [888]. Unexpected [745]. unfed [50]. unfragmented [816]. unicornfish [1203]. unicornis [1203]. Unionidae [1657]. unique [1776]. unit [1656, 1109, 1239, 215, 1463, 716, 684]. universal [1231, 1453]. unknown [1495]. unmanaged [892]. unmonitored [1205, 1589]. unobserved [1525]. Unpalatable [1286]. unprotected [1286]. unravel [1747]. Unravelling [423]. unscreened [822]. unsexed [1449]. unstocked [1295]. untangles [1809]. untangling [1668]. until [233]. up-river [1636]. updated [898]. updating [1780]. uphold [1076]. Upland [579, 247]. upon [728, 120, 1281]. Upper [26, 27, 395, 1599, 991, 230, 1478, 6]. upriver [1604, 1301]. upriver-migrating [1604]. upstream [1642, 1316, 871, 735]. upstream-migration [735]. uptake [159, 1474, 947]. upwelling [1230, 1116, 381]. Urban [187, 1668, 877, 1214]. urea [1278]. urine [575]. Urophycis [391]. USA [228, 731, 1806, 956, 1155, 167, 1323, 537, 829, 944, 583, 37, 382, 472, 62, 941, 1291, 1786, 1023, 21, 302, 261, 24, 156, 59]. usable [1675]. usage [168]. Use [336, 429, 240, 1737, 1510, 112, 646, 503, 1110, 370, 1578, 334, 1379, 454, 939, 1422, 1712, 1158, 1004, 167, 1457, 1603, 1751, 1193, 833, 1277, 1168, 674, 567, 234, 1517, 269, 1699, 595, 1571, 246, 124, 323, 1633, 1294, 210, 1732, 41, 90, 1209, 641, 402, 1291, 1192, 1040, 48, 978, 1224, 1402]. used [8, 5, 73, 72, 1760, 1407, 768]. useful [209]. user [1233]. user-friendly [1233]. uses [1601]. Using [1051, 1583, 789, 1774, 878, 1496, 1223, 427, 1697, 129, 766, 1160, 280, 674, 1785, 1532, 1826, 1779, 693, 1194, 861, 1655, 1134, 471, 1647, 1418, 410, 1791, 1383, 115, 855, 248, 1058, 444, 1686, 1381, 1450, 319, 1059, 440, 1272, 1273, 1503, 1549, 1024, 1382, 545, 1664, 1174, 1586, 700, 117, 1535, 1546, 1137, 1720, 291, 1603, 75, 306, 1238, 332, 669, 417, 1172, 1592, 92, 1701, 1089, 627, 1692, 1252, 1241, 69, 605, 679, 67, 1393, 1807, 1412, 270, 511, 810, 635, 1526, 771, 1609, 837, 1537, 910, 124, 63, 1725, 438, 230, 1486, 1706]. using [1515, 448, 430, 1616, 1606, 383, 1308, 432, 645, 922, 949, 139, 762, 174, 459, 186, 617, 493, 254, 1641, 703, 426, 637, 842, 315, 1670, 794, 1403, 1189, 1040, 1306, 110, 828, 1834, 709, 978, 1710, 715, 467, 301]. Utah [956, 1458]. utility [824, 211, 1485, 757, 1009]. utilization [139, 402, 1761]. utilized [1067]. Utilizing [671, 1804, 1307].

validate [766]. validated [835]. Validation [122, 1794, 1400, 1293, 886, 440, 1475, 551, 538, 750, 637]. validations [122]. Valley [132, 592]. Valleys [473]. valuation [1391]. value [667, 1424, 634, 1188, 48, 937]. values [287, 340, 110]. Valuing [1773]. Vancouver [407, 390]. Varden [600, 1016, 976]. variabilité [1071]. Variability [1071, 1669, 1749, 728, 403, 1803, 381, 279, 217, 1666, 1432, 17, 1778, 803, 854, 1325, 1550, 741, 1022, 321, 232, 1156, 569, 641, 433, 1714, 192, 1452, 1072]. Variable [196, 787, 1085, 875, 1713, 147, 1514, 799, 1153, 1208]. variables [1110, 779, 1155, 1272, 1273, 846, 854, 158, 1576, 767, 48, 262]. variance [438, 654]. variances [664]. Variation [3, 1215, 1472, 1302, 1264, 738, 1381, 205, 1756, 1767, 1802, 325, 1290, 1626, 1191, 648, 1643, 1182, 171, 829, 895, 81, 1016, 951, 1696, 69, 1080, 1753, 1393, 1518, 1062, 870, 82, 1304, 1810, 1766, 230, 1617, 87, 915, 466, 181, 423, 1576, 498, 459, 911, 739, 490, 1647, 722, 1196, 1000, 85, 1055, 1395, 1667, 506, 1787, 655, 828, 1070]. variations [1578, 1222, 595, 1523, 922, 1126, 1073]. varied [1801]. vary [536]. varying [375, 889, 1160, 1436, 1828, 565, 933, 1659]. vector [409, 32]. vectors [993]. vegetable [1289]. vehicles [1261, 565]. velocity [707]. vendace [896]. VENUS [890]. verification [40]. vernal [683]. versa [1206]. versatile [1233]. versus [209, 1358, 474, 1340, 844, 164, 573, 1397, 595, 1149, 272, 647, 520, 1623, 533, 394, 737, 707, 755, 1441]. Vertical [1524, 775, 238, 1392, 1350, 979, 708, 148, 1001, 1785, 1226, 896, 49]. vessel [339, 421, 993, 835, 878, 928, 1637, 1515, 282]. vessel-characteristic [421]. vessels [301, 280]. veteran [1578]. via [256, 229, 1583, 1702, 1516]. viability [1321, 1840, 1012, 1442]. viable [45, 342, 186]. vice [1206]. Victoria [1190, 1559, 1117]. video [1496, 1233, 694, 757]. villosus [119, 682]. viral [1138]. virtual [1503]. visual [1540, 701, 1062, 620]. visualization [861]. vital [1742]. vitreus [1570, 1805, 250, 1771, 1817, 1089, 523, 806, 566, 614, 911, 1467, 94]. VMS [339, 835, 1392, 1515]. Volume [179, 180, 257]. volumes [1833]. Vulnerability [753, 906, 836, 913, 1007, 885, 955, 1460, 170, 1100, 1783]. vulpinus [699].

wadeable [1155]. wake [1721]. walk [1369]. Walleye [1433, 1570, 1805, 250, 379, 1303, 1771, 1720, 1817, 942, 619, 1701, 1089, 301, 280, 311, 672, 523, 1262, 1563, 806, 566, 1490, 614, 934, 1466, 1011, 958, 1628, 911, 1467, 1631, 815, 1189, 94, 387, 1601]. Wanted [1522]. Warm [1749, 615, 1671, 1020, 1182, 1094, 1366]. warm-temperate [615]. warm-water [1671, 1182, 1094]. warming [905, 1433, 492, 1649]. warning [1836]. was [1559]. Washington [1323, 731, 156, 1345, 200, 1316, 1253, 901, 62]. Wastewater [329]. Water [235, 264, 317, 89, 646, 1206, 977, 640, 1671, 1088, 74, 1784, 993, 474, 159, 1771, 1289, 1182, 618, 45, 1433, 1715, 1317, 360, 833, 141, 523, 701, 1062, 71, 221, 1596, 1689, 966, 1198, 121, 1537, 1235, 15, 764, 1211, 1636, 434, 1513, 822, 1504, 1527, 423, 645, 805, 1484, 934, 922, 1539, 1219, 471, 1714, 1649, 1782, 1002, 1171, 975, 1094, 1667, 1428, 1063, 1292, 173, 992, 24, 1718, 262]. water-diversion [822, 805]. waterbody [1430]. watermilfoil [66]. waters [369, 333, 1345, 856, 537, 1620, 388, 329, 1412, 221, 124, 1179, 1371, 1411, 1521, 56, 488, 354, 1036]. Watershed [927, 1093, 1660, 604, 288, 960, 1325, 1821, 1210, 1304, 954, 949]. Watershed-scale [927]. watersheds [1223, 1312, 674, 968]. wave [934]. wavelets [417]. Wawa [411]. Weak [795, 951, 1292]. Weakened [307]. Weakening [1052]. weather [12]. weathervane [155]. Weaver [15]. web [1640, 1553, 129, 532, 644, 1621, 1481, 589, 4, 517, 1054, 972, 809, 712, 755, 1208]. webs [164, 573, 1119, 980, 1426]. weight [991]. weight-of-evidence [991]. weighting [357, 271, 668, 623]. weirs [1525, 1175]. welfare [1276]. well [951, 1484, 397]. well-being [951]. well-studied [1484]. Wenatchee [156]. West [296, 1049, 606, 468, 407, 390, 949, 887, 1489]. west-central [949]. western [1806, 16, 277, 386, 778, 1746, 861, 425, 820, 458, 62, 1649, 1418, 1695, 16]. westslope [624, 807, 448, 430, 41]. wet [1058, 1039]. wetland [840, 865, 1351]. wetlands [1841, 589, 610]. whale [507, 1418]. whales [1305, 354, 1077, 726]. whaling [726]. whence [193]. Where [856, 969, 356]. which [1338]. while [1379, 1729, 1118]. white [782, 1427, 150, 1807, 246, 991, 805, 391, 524, 497]. white-spotted [150]. whitefish [728, 1158, 1540, 84, 1811, 1499, 826, 361, 1171]. whitetip [705]. whither [193]. whiting [54, 1184]. whole [258, 380, 1590, 367, 1730, 1460, 157, 214]. whole-lake [258, 380, 1590, 1460, 157, 214]. whole-system [367]. wide [1549, 893, 388, 586, 1573, 577, 1003]. widely [8, 5, 73, 72, 1823]. widespread [1736, 458]. widths [24]. Wild [1674, 1372, 1349, 1605, 1060, 29, 1174, 572, 1340, 20, 812, 149, 1564, 431, 661, 777, 1127, 845, 552, 1545, 352, 1801, 1614, 373, 1468, 234, 854, 1297, 1186, 1253, 358, 702, 1434, 1808, 109, 791, 272, 237, 883, 165, 374, 1702, 1598, 1276, 641, 19, 897, 526, 1003, 50, 773, 1783, 1791, 802, 156, 1588, 709, 1695, 608, 467]. wild- [1003]. wild-born [802]. wild-origin [1127]. wild-spawning [802]. wildfire [57]. Will [422]. Willamette [1145]. William [434]. williamsoni [728]. Wind [1031, 381, 934, 1843, 707, 194]. wind-driven [381]. Winter [510, 1226, 1541, 1244, 75, 1545, 1483, 147, 1069, 559, 1099, 773, 828, 1761]. wintering [43]. winters [1749]. Wintertime [102]. wire [1362]. Wisconsin [1817, 55, 1477, 1600, 1490, 1011, 958, 1628, 1631, 1189, 1063, 1705]. Wisconsinan [85]. within [721, 1248, 171, 1169, 1369, 331, 178, 118, 803, 1602, 574, 1090, 732, 717, 1744, 861, 1157, 1141, 1041, 1702, 611, 1792, 1533, 712, 541, 750, 59, 21, 1494, 1112]. within-population [721]. within-river [171]. within-season [732]. Withler [27]. without [1340, 1528]. wolffish [1416]. wollei [489]. women [774]. Wood [926, 847, 1251, 324]. woody [888, 1565, 743, 1306, 1718]. workflow [1670]. world [420, 492].

Y-chromosome [3, 1070]. Yamaska [1355]. Yangtze [1829]. Year [1182, 1213, 1653, 376, 1061, 1711, 132, 392, 230, 1782, 953, 1444, 1718]. Year-class [1182, 392]. years [1081, 235, 308, 197, 1039, 22, 769, 350]. yellow [1096, 1201, 105, 1202, 1415, 1182, 1443, 860, 1431, 1465, 851, 31, 1534, 295, 1017]. yellow-legged [31]. yellowfin [343]. Yellowstone [350, 350]. yellowtail [935, 188, 1176]. yield [1349, 479, 1424, 998, 1661]. yields [589, 1118]. York [228, 1806, 1061]. young [393, 1653, 1061, 1385, 994, 1330, 1356, 1718]. young-of-the-year [1061]. young-of-year [1653]. you’re [1274]. Yukon [977, 984].

Zealand [429, 800]. zebra [970, 990, 817, 163, 992]. zebrafish [1482]. zenith [1001]. Zn [60]. zone [196, 557, 972, 976, 521, 570, 1031]. zones [1502, 955, 106]. zooplanktivore [1314]. Zooplankton [780, 530, 145, 1830, 1136, 1128, 1502, 45, 1295, 1596, 806, 919, 1076, 157, 528]. Zostera [611].

References

Jacobson:1995:SRM

[1] Larry D. Jacobson and Alec D. MacCall. Stock-recruitment models for Pacific sardine (Sardinops sagax). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52(3):566–577, March 1995. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub. com/doi/full/10.1139/f95-057. See comments [151, 560] and reply [686].

Taylor:2005:NLS

[2] Nathan G. Taylor, Carl J. Walters, and Steven J. D. Martell. A new likelihood for simultaneously estimating von Bertalanffy growth parameters, gear selectivity, and natural and fishing mortality. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 62(1):215–223, January 2005. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f04-189. See corrigendum

[300].

Devlin:2005:VCD

[3] R. H. Devlin, L. Park, D. M. Sakhrani, J. D. Baker, A. R. Marshall, E. LaHood, S. E. Kolesar, M. R. Mayo, C. A. Biagi, and M. Uh. Variation of Y-chromosome DNA markers in Chinook salmon (Oncorhynchus tshawytscha) populations. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 62 (6):1386–1399, June 2005. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/

10.1139/f05-048. See corrigendum [1070].

Krueger:2005:FWA

[4] Damon M. Krueger and Thomas R. Hrabik. Food web alterations that promote native species: the recovery of cisco (Coregonus artedi) populations through management of native piscivores. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 62(10):2177–2188, October 2005. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f05-132. See corrigendum

[517].

Branch:2009:DPC

[5] Trevor A. Branch. Differences in predicted catch composition between two widely used catch equation formulations. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 66(1):126–132, January 2009. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/F08-196. See corrigendum

[8], comment [72], and reply [73].

Walter:2009:GFI

[6] Ryan P. Walter, Tutku Aykanat, David W. Kelly, J. Mark Shrimpton, and Daniel D. Heath. Gene flow increases temporal stability of Chinook salmon (Oncorhynchus tshawytscha) populations in the Upper Fraser River, British Columbia, Canada. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 66(2):167–176, February 2009. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/F08-201. See comment [26] and reply [27].

Garner:2009:ECM

[7] Shawn R. Garner, John W. Heath, and Bryan D. Neff. Egg consumption in mature Pacific salmon (Oncorhynchus spp.). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 66(9):1546–1553, September 2009. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:/

/cdnsciencepub.com/doi/full/10.1139/F09-103. See comment [176].

Branch:2009:CDP

[8] Trevor A. Branch. Corrigendum: Differences in predicted catch composition between two widely used catch equation formulations. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 66(9):1631, September 2009. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/F09-107. See [5].

Perrier:2010:NRS

[9] Charles Perrier, Guillaume Evanno, Jérôme Belliard, René Guyomard, and Jean-Luc Baglinière. Natural recolonization of the Seine River by Atlantic salmon (Salmo salar) of multiple origins. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 67(1):1–4, January 2010. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/F09-190.

Adkison:2010:MEM

[10] Milo D. Adkison. Models of the effects of marine-derived nutrients on salmon (Oncorhynchus spp.) population dynamics. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 67(1):5–15, January 2010. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/F09-161.

Kennedy:2010:DCB

[11] James Kennedy, Jon Egil Skjæraasen, Richard D. M. Nash, Anders Thorsen, Aril Slotte, Tom Hansen, and Olav S. Kjesbu. Do capital breeders like Atlantic herring (Clupea harengus) exhibit sensitive periods of nutritional control on ovary development and fecundity regulation? Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 67(1):16–27, January 2010. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f09-159.

Cooperman:2010:PEA

[12] Michael S. Cooperman, Douglas F. Markle, Mark Terwilliger, and David C. Simon. A production estimate approach to analyze habitat and weather effects on recruitment of two endangered freshwater fish. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 67(1):28–41, January 2010. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/F09-165.

Kot:2010:TPT

[13] Connie Y. Kot, Andre M. Boustany, and Patrick N. Halpin. Temporal patterns of target catch and sea turtle bycatch in the US Atlantic pelagic longline fishing fleet. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 67 (1):42–57, January 2010. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/F09-160.

Foster:2010:EBT

[14] Sophie E. Foster and W. Gary Sprules. Effects of Bythotrephes on the trophic position of native macroinvertebrates. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 67(1):58–69, January 2010. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/F09-164.

Mathes:2010:EWT

[15] M. Todd Mathes, Scott G. Hinch, Steven J. Cooke, Glenn T. Crossin, David A. Patterson, Andrew G. Lotto, and Anthony P. Farrell. Effect of water temperature, timing, physiological condition, and lake thermal refugia on migrating adult Weaver Creek sockeye salmon (Oncorhynchus nerka). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 67(1):70–84, January 2010. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/F09-158.

Caputi:2010:ECC

[16] Nick Caputi, Roy Melville-Smith, Simon de Lestang, Alan Pearce, and Ming Feng. The effect of climate change on the western rock lobster (Panulirus cygnus) fishery of Western Australia. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 67(1):85–96, January 2010. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/F09-167.

Hurst:2010:QTE

[17] Thomas P. Hurst, Alisa A. Abookire, and Brian Knoth. Quantifying thermal effects on contemporary growth variability to predict responses to climate change in northern rock sole (Lepidopsetta polyxystra). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 67(1):97–107, January 2010. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/F09-171.

Conn:2010:HAM

[18] Paul B. Conn. Hierarchical analysis of multiple noisy abundance indices. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 67(1):108–120, January 2010. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/F09-175.

Salte:2010:PGM

[19] Ragnar Salte, Hans Bernhard Bentsen, Thomas Moen, Smita Tripathy, Tor Andreas Bakke, Jørgen Ødegård, Stig Omholt, and Lars Petter Hansen. Prospects for a genetic management strategy to control Gyrodactylus salaris infection in wild Atlantic salmon (Salmo salar) stocks. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 67(1):121–129, January 2010. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/F09-168.

Erkinaro:2010:DBC

[20] Jaakko Erkinaro, Eero Niemelä, Juha-Pekka Vähä, Craig R. Primmer, Sturla Brørs, and Esa Hassinen. Distribution and biological characteristics of escaped farmed salmon in a major subarctic wild salmon river: implications for monitoring. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 67 (1):130–142, January 2010. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/F09-173.

Thornbrugh:2010:ISP

[21] Darren J. Thornbrugh and Keith B. Gido. Influence of spatial positioning within stream networks on fish assemblage structure in the Kansas River basin, USA. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 67(1):143–156, January 2010. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/

F09-169. See erratum [59].

Olsson:2010:FHN

[22] Karin Olsson, Wilhelm Granéli, Jörgen Ripa, and Per Nyström. Fluctuations in harvest of native and introduced crayfish are driven by temperature and population density in previous years. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 67(1):157–164, January 2010. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/F09-170.

Diebel:2010:SAM

[23] Matthew W. Diebel, Jeffrey T. Maxted, Olaf P. Jensen, and M. Jake Vander Zanden. A spatial autocorrelative model for targeting stream restoration to benefit sensitive nongame fishes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 67(1):165–176, January 2010. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/F09-156.

Wilkerson:2010:EDB

[24] Ethel Wilkerson, John M. Hagan, and Andrew A. Whitman. The effectiveness of different buffer widths for protecting water quality and macroinvertebrate and periphyton assemblages of headwater streams in Maine, USA. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 67(1):177–190, January 2010. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
F09-162.

Sykes:2010:ETC

[25] Gregory E. Sykes and J. Mark Shrimpton. Effect of temperature and current manipulation on smolting in Chinook salmon (Oncorhynchus tshawytscha): the relationship between migratory behaviour and physiological development. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 67(1):

191–201, January 2010. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/F09-166.

Beacham:2010:CGF

[26] Terry D. Beacham and Ruth E. Withler. Comment on “Gene flow increases temporal stability of Chinook salmon (Oncorhynchus tshawytscha) populations in the Upper Fraser River, British Columbia, Canada”. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 67(1):202–205, January 2010. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/F09-177. See

[6, 27].

Walter:2010:RCB

[27] Ryan P. Walter, J. Mark Shrimpton, and Daniel D. Heath. Reply to the comment by Beacham and Withler on “Gene flow increases temporal stability of Chinook salmon (Oncorhynchus tshawytscha) populations in the Upper Fraser River, British Columbia, Canada”. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 67(1):206–208, January 2010. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/F09-178. See [6, 26].

Fenichel:2010:OCI

[28] Eli P. Fenichel and Gretchen J. A. Hansen. The opportunity cost of information: an economic framework for understanding the balance between assessment and control in sea lamprey (Petromyzon marinus) management. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 67(1):209–216, January 2010. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/F09-163.

Blouin:2010:NEL

[29] Michael S. Blouin, Virginie Thuillier, Becky Cooper, Vindhya Amarasinghe, Laura Cluzel, Hitoshi Araki, and Christoph Grunau. No evidence for large differences in genomic methylation between wild and hatchery steelhead (Oncorhynchus mykiss). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 67(2):217–224, February 2010. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/F09-174.

Palstra:2010:TPP

[30] Friso P. Palstra and Daniel E. Ruzzante. A temporal perspective on population structure and gene flow in Atlantic salmon (Salmo salar) in Newfoundland, Canada. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 67 (2):225–242, February 2010. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/F09-176.

Matthews:2010:SFD

[31] Kathleen R. Matthews and Haiganoush K. Preisler. Site fidelity of the declining amphibian Rana sierrae (Sierra Nevada yellow-legged frog). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 67(2):243–255, February 2010. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/F09-172.

Rup:2010:DBO

[32] Michael P. Rup, Sarah A. Bailey, Chris J. Wiley, Mark S. Minton, A. Whitman Miller, Gregory M. Ruiz, and Hugh J. MacIsaac. Domestic ballast operations on the Great Lakes: potential importance of Lakers as a vector for introduction and spread of nonindigenous species. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 67(2):256–268, February 2010. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/F09-180.

Weisberg:2010:MEM

[33] Sanford Weisberg, George Spangler, and Laurie S. Richmond. Mixed effects models for fish growth. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 67 (2):269–277, February 2010. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/F09-181.

Domine:2010:NRP

[34] Leah M. Domine, Michael J. Vanni, and William H. Renwick. New and regenerated primary production in a productive reservoir ecosystem. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 67(2):278–287, February 2010. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/F09-182.

Brockmark:2010:RHR

[35] Sofia Brockmark and Jörgen I. Johnsson. Reduced hatchery rearing density increases social dominance, postrelease growth, and survival in brown trout (Salmo trutta). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 67(2): 288–295, February 2010. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/F09-185.

Auger:2010:ONS

[36] Pierre Auger, Christophe Lett, Ali Moussaoui, and Sylvain Pioch. Optimal number of sites in artificial pelagic multisite fisheries. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 67(2):296–303, February 2010. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/F09-188.

McDermott:2010:HAR

[37] Michael J. McDermott, Anne L. Robertson, Peter J. Shaw, and Alexander M. Milner. The hyporheic assemblage of a recently formed stream following deglaciation in Glacier Bay, Alaska, USA. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 67(2):304–313, February 2010. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/F09-179.

McDermid:2010:LHD

[38] Jenni L. McDermid, Brian J. Shuter, and Nigel P. Lester. Life history differences parallel environmental differences among North American lake trout (Salvelinus namaycush) populations. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 67(2):314–325, February 2010. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/F09-183.

Jessop:2010:GEA

[39] B. M. Jessop. Geographic effects on American eel (Anguilla rostrata) life history characteristics and strategies. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 67(2):326–346, February 2010. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/F09-189.

Won:2010:OCF

[40] Nam-Il Won, Tomohiko Kawamura, Hideki Takami, Tadakatsu Noro, Tatsuya Musashi, and Yoshiro Watanabe. Ontogenetic changes in the feeding habits of the abalone Haliotis discus hannai: field verification by stable isotope analyses. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 67 (2):347–356, February 2010. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/F09-187.

Rasmussen:2010:ECH

[41] Joseph B. Rasmussen, Michael D. Robinson, and Daniel D. Heath. Ecological consequences of hybridization between native westslope cutthroat (Oncorhynchus clarkii lewisi) and introduced rainbow (Oncorhynchus mykiss) trout: effects on life history and habitat use. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 67(2):357–370, February 2010. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/F09-191.

Jenkins:2010:BAH

[42] Amy R. Jenkins and Ernest R. Keeley. Bioenergetic assessment of habitat quality for stream-dwelling cutthroat trout (Oncorhynchus clarkii bouvieri) with implications for climate change and nutrient supplementation. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 67(2):371–385, February 2010. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/F09-193.

Hatch:2010:ICS

[43] Scott A. Hatch, Verena A. Gill, and Daniel M. Mulcahy. Individual and colony-specific wintering areas of Pacific northern fulmars (Fulmarus glacialis). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 67(2):386–400, February 2010. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/F09-184.

vanKooten:2010:SHD

[44] Tobias van Kooten, Jens Andersson, Pär Byström, Lennart Persson, and André M. de Roos. Size at hatching determines population dynamics and response to harvesting in cannibalistic fish. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 67(2):401–416, February 2010. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/F09-157.

Gray:2010:DZE

[45] Derek K. Gray and Hugh J. MacIsaac. Diapausing zooplankton eggs remain viable despite exposure to open-ocean ballast water exchange: evidence from in situ exposure experiments. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 67(2):417–426, February 2010. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/F09-192.

Kobayashi:2010:DSM

[46] Sohei Kobayashi, Takashi Gomi, Roy C. Sidle, and Yasuhiro Takemon. Disturbances structuring macroinvertebrate communities in steep headwater streams: relative importance of forest clearcutting and debris flow occurrence. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 67(2):427–444, February 2010. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
F09-186.

Buchheister:2010:TFC

[47] Andre Buchheister and Robert J. Latour. Turnover and fractionation of carbon and nitrogen stable isotopes in tissues of a migratory coastal predator, summer flounder (Paralichthys dentatus). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 67(3):445–461, March 2010. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/F09-196.

Weisz:2010:RVL

[48] Erika J. Weisz and Norman D. Yan. Relative value of limnological, geographic, and human use variables as predictors of the presence of Bythotrephes longimanus in Canadian Shield lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 67(3):462–472, March 2010. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/F09-197.

Stockwell:2010:EEP

[49] Jason D. Stockwell, Thomas R. Hrabik, Olaf P. Jensen, Daniel L. Yule, and Matthew Balge. Empirical evaluation of predator-driven diel vertical migration in Lake Superior. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 67 (3):473–485, March 2010. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/F09-198.

Theriault:2010:SLH

[50] Véronique Thériault, Gregory R. Moyer, and Michael A. Banks. Survival and life history characteristics among wild and hatchery Coho salmon (Oncorhynchus kisutch) returns: how do unfed fry differ from smolt releases? Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 67(3):486–497, March 2010. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/F09-200.

Purchase:2010:TRN

[51] Craig F. Purchase, Ian A. E. Butts, Alexandre Alonso-Fernández, and Edward A. Trippel. Thermal reaction norms in sperm performance of Atlantic cod (Gadus morhua). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 67 (3):498–510, March 2010. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/F10-001.

Conn:2010:WCW

[52] Paul B. Conn, Erik H. Williams, and Kyle W. Shertzer. When can we reliably estimate the productivity of fish stocks? Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 67(3):511–523, March 2010. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/F09-194.

Hering:2010:TMR

[53] David K. Hering, Daniel L. Bottom, Earl F. Prentice, Kim K. Jones, and Ian A. Fleming. Tidal movements and residency of subyearling Chinook salmon (Oncorhynchus tshawytscha) in an Oregon salt marsh channel. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 67(3):524–533, March 2010. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/F10-003.

Ekerhovd:2010:SRM

[54] Nils-Arne Ekerhovd. The stability and resilience of management agreements on climate-sensitive straddling fishery resources: the blue whiting (Micromesistius poutassou) coastal state agreement. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 67(3):534–552, March 2010. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/F09-195.

Kornis:2010:FDI

[55] Matthew S. Kornis and M. Jake Vander Zanden. Forecasting the distribution of the invasive round goby (Neogobius melanostomus) in Wisconsin tributaries to Lake Michigan. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 67 (3):553–562, March 2010. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/F10-002.

Schloesser:2010:NOA

[56] Ryan W. Schloesser, John D. Neilson, David H. Secor, and Jay R. Rooker. Natal origin of Atlantic bluefin tuna (Thunnus thynnus) from Canadian waters based on otolith δ ^{13}C and δ ^{18}O. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 67(3):563–569, March 2010. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/F10-005.

Malison:2010:FPW

[57] Rachel L. Malison and Colden V. Baxter. The fire pulse: wildfire stimulates flux of aquatic prey to terrestrial habitats driving increases in riparian consumers. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 67 (3):570–579, March 2010. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/F10-006.

Guenard:2010:EAB

[58] Guillaume Guénard, Daniel Boisclair, Ola Ugedal, Torbjørn Forseth, Bror Jonsson, and Ian A. Fleming. Experimental assessment of the bioenergetic and behavioural differences between two morphologically distinct populations of Arctic char (Salvelinus alpinus). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 67(3):580–595, March 2010. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/F10-004.

Thornbrugh:2010:EIS

[59] Darren J. Thornbrugh and Keith B. Gido. Erratum: Influence of spatial positioning within stream networks on fish assemblage structure in the Kansas River basin, USA. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 67(3):596, March 2010. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/F10-009. See [21].

Limburg:2010:PMZ

[60] Karin E. Limburg and Mikael Elfman. Patterns and magnitude of Zn : Ca in otoliths support the recent phylogenetic typology of Salmoniformes and their sister groups. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 67 (4):597–604, April 2010. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/F10-014.

Kjesbu:2010:TDO

[61] Olav Sigurd Kjesbu, David Righton, Maria Krüger-Johnsen, Anders Thorsen, Kathrine Michalsen, Merete Fonn, and Peter R. Witthames. Thermal dynamics of ovarian maturation in Atlantic cod (Gadus morhua). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 67(4):605–625, April 2010. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/F10-011.

Shellberg:2010:HEB

[62] Jeffrey G. Shellberg, Susan M. Bolton, and David R. Montgomery. Hydrogeomorphic effects on bedload scour in bull char (Salvelinus confluentus) spawning habitat, western Washington, USA. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 67(4):626–640, April 2010. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/F10-007.

Melnychuk:2010:EDP

[63] Michael C. Melnychuk and Carl J. Walters. Estimating detection probabilities of tagged fish migrating past fixed receiver stations using only local information. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 67(4):641–658, April 2010. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/F09-
199.

Descroix:2010:IMD

[64] Aurélie Descroix, Christian Desvilettes, Alexandre Bec, Patrick Martin, and Gilles Bourdier. Impact of macroinvertebrate diet on growth and fatty acid profiles of restocked 0+ Atlantic salmon (Salmo salar) parr from a large European river (the Allier). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 67(4):659–672, April 2010. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/F10-012.

Bradford:2010:PMI

[65] Michael J. Bradford, Jan Lovy, David A. Patterson, David J. Speare, William R. Bennett, Allen R. Stobbart, and Christine P. Tovey. Parvicapsula minibicornis infections in gill and kidney and the premature mortality of adult sockeye salmon (Oncorhynchus nerka) from Cultus Lake, British Columbia. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 67 (4):673–683, April 2010. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/F10-017.

Maezo:2010:PRI

[66] Maria José Maezo, Henri Fournier, and Beatrix E. Beisner. Potential and realized interactions between two aquatic invasive species: Eurasian watermilfoil (Myriophyllum spicatum) and rusty crayfish (Orconectes rusticus). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 67(4):684–700, April 2010. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/F10-016.

Jordaan:2010:IES

[67] Adrian Jordaan, Yong Chen, David W. Townsend, and Sally Sherman. Identification of ecological structure and species relationships along an oceanographic gradient in the Gulf of Maine using multivariate analysis with bootstrapping. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 67 (4):701–719, April 2010. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/F10-010.

Price:2010:RDR

[68] Amina E. Price and Paul Humphries. The role of dispersal and retention in the early life stages of shrimp in a lowland river. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 67(4):720–729, April 2010. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/F10-015.

Jiao:2010:ITV

[69] Yan Jiao, Laura Rogers-Bennett, Ian Taniguchi, John Butler, and Paul Crone. Incorporating temporal variation in the growth of red abalone (Haliotis rufescens) using hierarchical Bayesian growth models. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 67(4):730–742, April 2010. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/F10-019.

Souchu:2010:PNL

[70] Philippe Souchu, Béatrice Bec, Val H. Smith, Thierry Laugier, Annie Fiandrino, Laurent Benau, Valérie Orsoni, Yves Collos, and André Vaquer. Patterns in nutrient limitation and chlorophyll a along an anthropogenic eutrophication gradient in French Mediterranean coastal lagoons. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 67(4):743–753, April 2010. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/F10-018.

Korosi:2010:CPI

[71] Jennifer B. Korosi, Andrew M. Paterson, Anna M. DeSellas, and John P. Smol. A comparison of pre-industrial and present-day changes in Bosmina and Daphnia size structure from soft-water Ontario lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 67(4):754–762, April 2010. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/F10-013.

Francis:2010:CDP

[72] R. I. C. Chris Francis. Comment on “Differences in predicted catch composition between two widely used catch equation formulations”. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 67(4):763–765, April 2010. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/F10-023. See [5] and reply

[73].

Branch:2010:RCF

[73] Trevor A. Branch. Reply to the comment by Francis on “Differences in predicted catch composition between two widely used catch equation formulations”. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 67(4):766–768, April 2010. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/F10-

024. See [5, 72].

Campana:2010:SPG

[74] Steven E. Campana, Warren Joyce, and Mark Fowler. Subtropical pupping ground for a cold-water shark. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 67(5):769–773, May 2010. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/F10-020.

Frisk:2010:EPD

[75] M. G. Frisk, S. J. D. Martell, T. J. Miller, and K. Sosebee. Exploring the population dynamics of winter skate (Leucoraja ocellata) in the Georges Bank region using a statistical catch-at-age model incorporating length, migration, and recruitment process errors. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 67(5):774–792, May 2010. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/F10-008.

Brodziak:2010:MAP

[76] Jon Brodziak and Kevin Piner. Model averaging and probable status of North Pacific striped marlin, Tetrapturus audax. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 67(5):793–805, May 2010. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/F10-029.

Castro-Santos:2010:MME

[77] Theodore Castro-Santos and Benjamin H. Letcher. Modeling migratory energetics of Connecticut River American shad (Alosa sapidissima): implications for the conservation of an iteroparous anadromous fish. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 67(5):806–830, May 2010. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/F10-026.

Whiteway:2010:DSR

[78] Sarah L. Whiteway, Pascale M. Biron, André Zimmermann, Oscar Venter, and James W. A. Grant. Do in-stream restoration structures enhance salmonid abundance? A meta-analysis. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 67(5):831–841, May 2010. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/F10-021.

Swanson:2010:AAP

[79] Heidi K. Swanson, Karen A. Kidd, John A. Babaluk, Rick J. Wastle, Panseok P. Yang, Norman M. Halden, and James D. Reist. Anadromy in Arctic populations of lake trout (Salvelinus namaycush): otolith microchemistry, stable isotopes, and comparisons with Arctic char (Salvelinus alpinus). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 67(5):842–853, May 2010. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/F10-022.

Neuheimer:2010:CCL

[80] Anna B. Neuheimer and Christopher T. Taggart. Can changes in lengthat-age and maturation timing in Scotian Shelf haddock (Melanogrammus aeglefinus) be explained by fishing? Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 67(5):854–865, May 2010. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/F10-025.

Harrald:2010:SVR

[81] Marion Harrald, Peter J. Wright, and Francis C. Neat. Substock variation in reproductive traits in North Sea cod (Gadus morhua). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 67(5):866–876, May 2010. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/F10-030.

Matich:2010:SBV

[82] Philip Matich, Michael R. Heithaus, and Craig A. Layman. Size-based variation in intertissue comparisons of stable carbon and nitrogen isotopic signatures of bull sharks (Carcharhinus leucas) and tiger sharks (Galeocerdo cuvier). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 67 (5):877–885, May 2010. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/F10-037.

Branch:2010:GMR

[83] Trevor A. Branch and Ray Hilborn. A general model for reconstructing salmon runs. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 67(5):886–904, May 2010. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/F10-
032.

Harris:2010:GPS

[84] Les N. Harris and Eric B. Taylor. Genetic population structure of broad whitefish, Coregonus nasus, from the Mackenzie River, Northwest Territories: implications for subsistence fishery management. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 67(6):905–918, June 2010. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/F10-027.

Tamkee:2010:IWG

[85] P. Tamkee, E. Parkinson, and E. B. Taylor. The influence of wisconsinan glaciation and contemporary stream hydrology on microsatellite DNA variation in rainbow trout (Oncorhynchus mykiss). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 67(6):919–935, June 2010. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/F10-034.

Thorson:2010:MEB

[86] James T. Thorson and Jim Berkson. Multispecies estimation of Bayesian priors for catchability trends and density dependence in the US Gulf of Mexico. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 67(6):936–954, June 2010. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/F10-040.

Morita:2010:LVG

[87] Kentaro Morita and Toru Nagasawa. Latitudinal variation in the growth and maturation of masu salmon (Oncorhynchus masou) parr. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 67(6):955–965, June 2010. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/F10-028.

Galuardi:2010:CMR

[88] Benjamin Galuardi, François Royer, Walt Golet, John Logan, John Neilson, and Molly Lutcavage. Complex migration routes of Atlantic bluefin tuna (Thunnus thynnus) question current population structure paradigm. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 67(6):966–976, June 2010. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/F10-033.

Abele:2010:MCW

[89] Doris Abele, Melanie Kruppe, Eva E. R. Philipp, and Thomas Brey. Mantle cavity water oxygen partial pressure (pO_{2}) in marine molluscs aligns with lifestyle. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 67 (6):977–986, June 2010. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/F10-035.

Reichert:2010:RPU

[90] Julie M. Reichert, Brian J. Fryer, Kevin L. Pangle, Timothy B. Johnson, Jeff T. Tyson, Alison B. Drelich, and Stuart A. Ludsin. River-plume use during the pelagic larval stage benefits recruitment of a lentic fish. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 67(6):987–1004, June 2010. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/F10-036.

Ralston:2010:PEB

[91] Stephen Ralston and Bruce R. MacFarlane. Population estimation of bocaccio (Sebastes paucispinis) based on larval production. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 67(6):1005–1020, June 2010. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/F10-039.

Hasselman:2010:TSD

[92] Daniel J. Hasselman, Rod G. Bradford, and Paul Bentzen. Taking stock: defining populations of American shad (Alosa sapidissima) in Canada using neutral genetic markers. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 67 (6):1021–1039, June 2010. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/F10-031.

Gurney:2010:FUS

[93] William S. C. Gurney, Philip J. Bacon, Eddie McKenzie, Philip McGinnity, Julian Mclean, Gordon Smith, and Alan Youngson. Form and uncertainty in stock–recruitment relations: observations and implications for Atlantic salmon (Salmo salar) management. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 67(6):1040–1055, June 2010. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/F10-038.

Venturelli:2010:CPM

[94] Paul A. Venturelli, Nigel P. Lester, Terry R. Marshall, and Brian J. Shuter. Consistent patterns of maturity and density-dependent growth among populations of walleye (Sander vitreus): application of the growing degree-day metric. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 67 (7):1057–1067, July 2010. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/F10-041.

Edmunds:2010:PSL

[95] Richard C. Edmunds, Lynne van Herwerden, and Christopher J. Fulton. Population-specific locomotor phenotypes are displayed by barramundi, Lates calcarifer, in response to thermal stress. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 67(7):1068–1074, July 2010. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/F10-047.

Thompson:2010:EBM

[96] Bradley E. Thompson and Daniel B. Hayes. Estimating bioenergetics model parameters for fish with incomplete recapture histories. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 67(7):1075–1085, July 2010. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/F10-048.

Jorgensen:2010:MFI

[97] Christian Jørgensen and Øyvind Fiksen. Modelling fishing-induced adaptations and consequences for natural mortality. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 67(7):1086–1097, July 2010. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/F10-049.

Woods:2010:CHP

[98] Ryan J. Woods, Jed I. Macdonald, David A. Crook, Daniel J. Schmidt, and Jane M. Hughes. Contemporary and historical patterns of connectivity among populations of an inland river fish species inferred from genetics and otolith chemistry. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 67 (7):1098–1115, July 2010. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/F10-043.

Aagren:2010:SPB

[99] Anneli Ågren, Ishi Buffam, Kevin Bishop, and Hjalmar Laudon. Sensitivity of pH in a boreal stream network to a potential decrease in base cations caused by forest harvest. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 67(7):1116–1125, July 2010. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/F10-052.

Ely:2010:LTF

[100] Damon T. Ely and J. Bruce Wallace. Long-term functional group recovery of lotic macroinvertebrates from logging disturbance. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 67(7):1126–1134, July 2010. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/F10-045.

VanKirk:2010:MAS

[101] Kray F. Van Kirk, Terrance J. Quinn, and Jeremy S. Collie. A multispecies age-structured assessment model for the Gulf of Alaska. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 67(7):1135–1148, July 2010. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/F10-053.

Black:2010:WOC

[102] Bryan A. Black, Isaac D. Schroeder, William J. Sydeman, Steven J. Bograd, and Peter W. Lawson. Wintertime ocean conditions synchronize rockfish growth and seabird reproduction in the central California Current ecosystem. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 67(7):1149– 1158, July 2010. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
F10-055.

Warren:2010:ADA

[103] Joseph D. Warren and David A. Demer. Abundance and distribution of Antarctic krill (Euphausia superba) nearshore of Cape Shirreff, Livingston Island, Antarctica, during six austral summers between 2000 and 2007. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 67(7):1159–1170, July 2010. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/F10-042.

Northrup:2010:PSL

[104] Sara Northrup, Mark Connor, and Eric B. Taylor. Population structure of lake trout (Salvelinus namaycush) in a large glacial-fed lake inferred from microsatellite DNA and morphological analysis. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 67(7):1171–1186, July 2010. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/F10-054.

Collingsworth:2010:ETE

[105] Paris D. Collingsworth, Jason J. Van Tassell, John W. Olesik, and Elizabeth A. Marschall. Effects of temperature and elemental concentration on the chemical composition of juvenile yellow perch (Perca flavescens) otoliths. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 67(7):1187–1196, July 2010. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/F10-050.

Richardson:2010:DRZ

[106] John S. Richardson, Eric Taylor, Dolph Schluter, Mike Pearson, and Todd Hatfield. Do riparian zones qualify as critical habitat for endangered freshwater fishes? Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 67 (7):1197–1204, July 2010. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/F10-063.

Hutchings:2010:TAM

[107] Jeffrey A. Hutchings, Cóilín Minto, Daniel Ricard, Julia K. Baum, and Olaf P. Jensen. Trends in the abundance of marine fishes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 67(8):1205–1210, August 2010. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/F10-081.

West:2010:NLA

[108] Derek C. West, Annika W. Walters, Stephen Gephard, and David M. Post. Nutrient loading by anadromous alewife (Alosa pseudoharengus): contemporary patterns and predictions for restoration efforts. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 67(8):1211–1220, August 2010. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/F10-059.

Meager:2010:RIB

[109] Justin J. Meager, Jon Egil Skjæraasen, Anders Fernö, and Svein Løkkeborg. Reproductive interactions between fugitive farmed and wild Atlantic cod (Gadus morhua) in the field. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 67(8):1221–1231, August 2010. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/F10-066.

Wassenaar:2010:ACM

[110] L. I. Wassenaar, J. J. Venkiteswaran, S. L. Schiff, and G. Koehler. Aquatic community metabolism response to municipal effluent inputs in rivers quantified using diel δ ^{18}O values of dissolved oxygen. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 67(8):1232–1246, August 2010. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/F10-057.

Bousquet:2010:DCU

[111] Nicolas Bousquet, Noel Cadigan, Thierry Duchesne, and Louis-Paul Rivest. Detecting and correcting underreported catches in fish stock assessment: trial of a new method. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 67(8):1247–1261, August 2010. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/F10-051.

Albert:2010:PLU

[112] Marc Richard Albert, Guangjie Chen, Graham K. MacDonald, Jesse C. Vermaire, Elena M. Bennett, and Irene Gregory-Eaves. Phosphorus and land-use changes are significant drivers of cladoceran community composition and diversity: an analysis over spatial and temporal scales. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 67(8):1262–1273, August 2010. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/F10-058.

Beacham:2010:CSI

[113] Terry D. Beacham, B. McIntosh, and C. Wallace. A comparison of stock and individual identification for sockeye salmon (Oncorhynchus nerka) in British Columbia provided by microsatellites and single nucleotide polymorphisms. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 67(8): 1274–1290, August 2010. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/F10-061.

Pereira:2010:SLP

[114] Helder Cunha Pereira, Norman Allott, and Catherine Coxon. Are seasonal lakes as productive as permanent lakes? A case study from Ireland. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 67(8):1291–1302, August 2010. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/F10-062.

Ashford:2010:TEL

[115] Julian Ashford, Mario La Mesa, Bettina A. Fach, Christopher Jones, and Inigo Everson. Testing early life connectivity using otolith chemistry and particle-tracking simulations. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 67 (8):1303–1315, August 2010. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/F10-065.

Spence:2010:SPM

[116] Brian C. Spence and James D. Hall. Spatiotemporal patterns in migration timing of Coho salmon (Oncorhynchus kisutch) smolts in North America. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 67(8):1316–1334, August 2010. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
F10-060.

Diamond:2010:DMR

[117] Sandra L. Diamond, Kristin M. Kleisner, Daisy Englert Duursma, and Yuning Wang. Designing marine reserves to reduce bycatch of mobile species: a case study using juvenile red snapper (Lutjanus campechanus). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 67(8):1335–1349, August 2010. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/F10-044.

Jones:2010:ILW

[118] Nicholas E. Jones. Incorporating lakes within the river discontinuum: longitudinal changes in ecological characteristics in stream–lake networks. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 67(8):1350–1362, August 2010. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/F10-

069. See erratum [178].

Hjermann:2010:TIA

[119] Dag Øystein Hjermann, Bjarte Bogstad, Gjert Endre Dingsør, Harald Gjøsæter, Geir Ottersen, Anne Maria Eikeset, and Nils Christian Stenseth. Trophic interactions affecting a key ecosystem component: a multistage analysis of the recruitment of the Barents Sea capelin (Mallotus villosus). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 67(9):1363– 1375, September 2010. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/F10-064.

Carlsson:2010:DMM

[120] Marita Sundstein Carlsson, Ronnie Nøhr Glud, and Jens Kjerulf Petersen. Degradation of mussel (Mytilus edulis) fecal pellets released from hanging long-lines upon sinking and after settling at the sediment. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 67(9):1376–1387, September 2010. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/F10-067.

Lochet:2010:SIF

[121] Aude Lochet, Karin E. Limburg, Lars Rudstam, and Mario Montesdeoca. Selenium incorporation in fish otoliths: effects of selenium and mercury from the water. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 67(9):1388– 1397, September 2010. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/F10-070.

Francis:2010:VFA

[122] R. I. C. Chris Francis, Steven E. Campana, and Helen L. Neil. Validation of fish ageing methods should involve bias estimation rather than hypothesis testing: a proposed approach for bomb radiocarbon validations. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 67(9):1398–1408, September 2010. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/F10-068.

Carruthers:2010:SSD

[123] Thomas R. Carruthers, Murdoch K. McAllister, and Robert N. M. Ahrens. Simulating spatial dynamics to evaluate methods of deriving abundance indices for tropical tunas. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 67(9):1409–1427, September 2010. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/F10-056. See erratum [144].

McMeans:2010:DRU

[124] Bailey C. McMeans, Jörundur Svavarsson, Susan Dennard, and Aaron T. Fisk. Diet and resource use among Greenland sharks (Somniosus microcephalus) and teleosts sampled in Icelandic waters, using δ ^{13}C, δ 15N, and mercury. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 67(9):1428– 1438, September 2010. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/F10-072.

Dekar:2010:BAF

[125] Matthew P. Dekar, Daniel D. Magoulick, and Jeff Beringer. Bioenergetics assessment of fish and crayfish consumption by river otter (Lontra canadensis): integrating prey availability, diet, and field metabolic rate. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 67(9):1439–1448, September 2010. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/F10-074.

VanDoornik:2010:ESP

[126] Donald M. Van Doornik, Barry A. Berejikian, Lance A. Campbell, and Eric C. Volk. The effect of a supplementation program on the genetic and life history characteristics of an Oncorhynchus mykiss population. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 67(9):1449–1458, September 2010. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/F10-073.

DeRobertis:2010:DAE

[127] Alex De Robertis, Denise R. McKelvey, and Patrick H. Ressler. Development and application of an empirical multifrequency method for backscatter classification. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 67 (9):1459–1474, September 2010. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/

doi/full/10.1139/F10-075.

Pangle:2010:OMS

[128] Kevin L. Pangle, Stuart A. Ludsin, and Brian J. Fryer. Otolith microchemistry as a stock identification tool for freshwater fishes: testing its limits in Lake Erie. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 67(9): 1475–1489, September 2010. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/F10-076.

Gaichas:2010:UFW

[129] Sarah K. Gaichas, Kerim Y. Aydin, and Robert C. Francis. Using food web model results to inform stock assessment estimates of mortality and production for ecosystem-based fisheries management. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 67(9):1490–1506, September 2010. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/F10-071.

Johnston:2010:DCA

[130] Fiona D. Johnston, Robert Arlinghaus, and Ulf Dieckmann. Diversity and complexity of angler behaviour drive socially optimal input and output regulations in a bioeconomic recreational-fisheries model. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 67(9):1507–1531, September 2010. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/F10-046. See erra-

tum [161].

Smith:2010:ISM

[131] Stephen J. Smith and Steven E. Campana. Integrated stock mixture analysis for continous and categorical data, with application to genetic– otolith combinations. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 67(10): 1533–1548, October 2010. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/F10-078.

MacFarlane:2010:EDG

[132] R. Bruce MacFarlane. Energy dynamics and growth of Chinook salmon (Oncorhynchus tshawytscha) from the Central Valley of California during the estuarine phase and first ocean year. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 67(10):1549–1565, October 2010. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/F10-080.

Goto:2010:BRB

[133] Daisuke Goto and William G. Wallace. Bioenergetic responses of a benthic forage fish (Fundulus heteroclitus) to habitat degradation and altered prey community in polluted salt marshes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 67(10):1566–1584, October 2010. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/F10-082.

Therkildsen:2010:LEP

[134] Nina Overgaard Therkildsen, Einar Eg Nielsen, Douglas P. Swain, and Jes Søe Pedersen. Large effective population size and temporal genetic stability in Atlantic cod (Gadus morhua) in the southern Gulf of St. Lawrence. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 67(10):1585–1595, October 2010. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
F10-084.

Cover:2010:CDH

[135] Matthew R. Cover, Juan A. de la Fuente, and Vincent H. Resh. Catastrophic disturbances in headwater streams: the long-term ecological effects of debris flows and debris floods in the Klamath Mountains, northern California. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 67(10):1596– 1610, October 2010. CODEN CJFSDX. ISSN 0706-652X (print), 1205-

7533 (electronic). URL https://cdnsciencepub.com/doi/full/10.
1139/F10-079.

Forrest:2010:HBE

[136] Robyn E. Forrest, Murdoch K. McAllister, Martin W. Dorn, Steven J. D. Martell, and Richard D. Stanley. Hierarchical Bayesian estimation of recruitment parameters and reference points for Pacific rockfishes (Sebastes spp.) under alternative assumptions about the stock–recruit function. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 67(10):1611–1634, October 2010. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/F10-077.

Johnson:2010:DIE

[137] Cody R. Johnson, Chris Luecke, Stephen C. Whalen, and Mary Anne Evans. Direct and indirect effects of fish on pelagic nitrogen and phosphorus availability in oligotrophic Arctic Alaskan lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 67(10):1635–1648, October 2010. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/F10-085.

Clarke:2010:FPA

[138] Amber Clarke, Ralph Mac Nally, Nick Bond, and P. S. Lake. Flow permanence affects aquatic macroinvertebrate diversity and community structure in three headwater streams in a forested catchment. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 67(10):1649–1657, October 2010. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/F10-087.

Rooper:2010:AHU

[139] Christopher N. Rooper, Gerald R. Hoff, and Alex De Robertis. Assessing habitat utilization and rockfish (Sebastes spp.) biomass on an isolated rocky ridge using acoustics and stereo image analysis. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 67(10):1658–1670, October 2010. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/F10-088.

Burla:2010:ICR

[140] Michela Burla, António M. Baptista, Edmundo Casillas, John G. Williams, and Douglas M. Marsh. The influence of the Columbia River plume on the survival of steelhead (Oncorhynchus mykiss) and Chinook salmon (Oncorhynchus tshawytscha): a numerical exploration. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 67(10):1671–1684, October 2010. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/F10-083.

Holth:2010:LTE

[141] Tor Fredrik Holth, Anders Thorsen, Pål A. Olsvik, and Ketil Hylland. Long-term exposure of Atlantic cod (Gadus morhua) to components of produced water: condition, gonad maturation, and gene expression. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 67(10):1685–1698, October 2010. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/F10-089.

Walther:2010:LDN

[142] Benjamin D. Walther and Simon R. Thorrold. Limited diversity in natal origins of immature anadromous fish during ocean residency. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 67(10):1699–1707, October 2010. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/F10-086.

Enberg:2010:FIE

[143] Katja Enberg, Christian Jørgensen, and Marc Mangel. Fishing-induced evolution and changing reproductive ecology of fish: the evolution of steepness. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 67(10):1708–1719, October 2010. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
F10-090.

Carruthers:2010:ESS

[144] Thomas R. Carruthers, Murdoch K. McAllister, and Robert N. M. Ahrens. Erratum: Simulating spatial dynamics to evaluate methods of deriving abundance indices for tropical tunas. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 67(10):1720, October 2010. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/F10-111. See [123].

Syvaranta:2010:ZLS

[145] Jari Syväranta and Milla Rautio. Zooplankton, lipids and stable isotopes: importance of seasonal, latitudinal, and taxonomic differences. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 67(11):1721–1729, November 2010. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/F10-091.

McGree:2010:RFR

[146] Michelle M. McGree, Dana L. Winkelman, Nicole K. M. Vieira, and Alan M. Vajda. Reproductive failure of the red shiner (Cyprinella lutrensis) after exposure to an exogenous estrogen. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 67(11):1730–1743, November 2010. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/F10-092.

Linnansaari:2010:PAS

[147] Tommi Linnansaari and Richard A. Cunjak. Patterns in apparent survival of Atlantic salmon (Salmo salar) parr in relation to variable ice conditions throughout winter. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 67 (11):1744–1754, November 2010. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/

doi/full/10.1139/F10-093.

Huebert:2010:PVD

[148] Klaus B. Huebert, Su Sponaugle, and Robert K. Cowen. Predicting the vertical distributions of reef fish larvae in the Straits of Florida from environmental factors. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 67(11):

1755–1767, November 2010. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/F10-116.

Fraser:2010:PDW

[149] Dylan J. Fraser, Cóilín Minto, Anna M. Calvert, James D. Eddington, and Jeffrey A. Hutchings. Potential for domesticated–wild interbreeding to induce maladaptive phenology across multiple populations of wild Atlantic salmon (Salmo salar). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 67(11):1768–1775, November 2010. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/F10-094.

Hasegawa:2010:EFR

[150] Koh Hasegawa and Shoichiro Yamamoto. The effect of flow regime on the occurrence of interference and exploitative competition in a salmonid species, white-spotted char (Salvelinus leucomaenis). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 67(11):1776–1781, November 2010. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/F10-100.

McClatchie:2010:RAS

[151] Sam McClatchie, Ralf Goericke, Guillermo Auad, and Kevin Hill. Reassessment of the stock–recruit and temperature–recruit relationships for Pacific sardine (Sardinops sagax). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 67(11):1782–1790, November 2010. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/F10-101. See [1] and comment [560].

Pool:2010:EDF

[152] Thomas K. Pool, Julian D. Olden, Joanna B. Whittier, and Craig P. Paukert. Environmental drivers of fish functional diversity and composition in the Lower Colorado River Basin. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 67(11):1791–1807, November 2010. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/F10-095.

Adkison:2010:MIE

[153] Milo D. Adkison. Management implications of the effects of marinederived nutrients on salmon population dynamics. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 67(11):1808–1815, November 2010. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/F10-097.

Erickson:2010:ECC

[154] Russell J. Erickson, David R. Mount, Terry L. Highland, J. Russell Hockett, Edward N. Leonard, Vincent R. Mattson, Timothy D. Dawson, and Kevin G. Lott. Effects of copper, cadmium, lead, and arsenic in a live diet on juvenile fish growth. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 67(11): 1816–1826, November 2010. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/F10-098.

Gaffney:2010:GHW

[155] Patrick M. Gaffney, Carita M. Pascal, Jeffery Barnhart, W. Stewart Grant, and James E. Seeb. Genetic homogeneity of weathervane scallops (Patinopecten caurinus) in the northeastern Pacific. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 67(11):1827–1839, November 2010. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/F10-096.

Williamson:2010:FIR

[156] Kevin S. Williamson, Andrew R. Murdoch, Todd N. Pearsons, Eric J. Ward, and Michael J. Ford. Factors influencing the relative fitness of hatchery and wild spring Chinook salmon (Oncorhynchus tshawytscha) in the Wenatchee River, Washington, USA. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 67(11):1840–1851, November 2010. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/F10-099.

Paterson:2010:RZW

[157] Michael J. Paterson, Cheryl L. Podemski, Wilhelmina J. Findlay, David L. Findlay, and Alex G. Salki. The response of zooplankton in a whole-lake experiment on the effects of a cage aquaculture operation for rainbow trout (Oncorhynchus mykiss). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 67(11):1852–1861, November 2010. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/F10-106.

Kramer-Schadt:2010:AVA

[158] Stephanie Kramer-Schadt, Jens Christian Holst, and Dankert Skagen. Analysis of variables associated with the Ichthyophonus hoferi epizootics in Norwegian spring spawning herring, 1992–2008. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 67(11):1862–1873, November 2010. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/F10-102.

Cooper:2010:MCU

[159] Sophie Cooper, Landis Hare, and Peter G. C. Campbell. Modeling cadmium uptake from water and food by the freshwater bivalve Pyganodon grandis. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 67(11):1874–1888, November 2010. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
F10-103.

Parslow:2010:ITQ

[160] John Parslow. Individual transferable quotas and the “tragedy of the commons”. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 67(11):1889–1896, November 2010. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
F10-104.

Johnston:2010:EDC

[161] Fiona D. Johnston, Robert Arlinghaus, and Ulf Dieckmann. Erratum: Diversity and complexity of angler behaviour drive socially optimal input and output regulations in a bioeconomic recreational-fisheries model. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 67(11):1897–1898, November 2010. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/F10-113. See

[130].

Anonymous:2010:AC

[162] Anonymous. Announcement / Communiqué. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 67(12), December 2010. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f10-902.

Turner:2010:IZD

[163] Caroline B. Turner. Influence of zebra (Dreissena polymorpha) and quagga (Dreissena rostriformis) mussel invasions on benthic nutrient and oxygen dynamics. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 67(12): 1899–1908, December 2010. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/F10-107.

Kiernan:2010:DVI

[164] Joseph D. Kiernan, Brett N. Harvey, and Michael L. Johnson. Direct versus indirect pathways of salmon-derived nutrient incorporation in experimental lotic food webs. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 67(12): 1909–1924, December 2010. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/F10-110.

Price:2010:EFI

[165] M. H. H. Price, A. Morton, and J. D. Reynolds. Evidence of farminduced parasite infestations on wild juvenile salmon in multiple regions of coastal British Columbia, Canada. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 67(12):1925–1932, December 2010. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/F10-105. See comment [373].

Berejikian:2010:MSA

[166] Barry A. Berejikian, Donald M. Van Doornik, Rob C. Endicott, Timothy L. Hoffnagle, Eugene P. Tezak, Megan E. Moore, and Jeffery Atkins. Mating success of alternative male phenotypes and evidence for frequency-dependent selection in Chinook salmon, Oncorhynchus tshawytscha. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 67(12):1933– 1941, December 2010. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/

10.1139/F10-112. See corrigendum [195].

Falke:2010:SPH

[167] Jeffrey A. Falke, Kurt D. Fausch, Kevin R. Bestgen, and Larissa L. Bailey. Spawning phenology and habitat use in a Great Plains, USA, stream fish assemblage: an occupancy estimation approach. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 67(12):1942–1956, December 2010. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/F10-109.

Ferraro:2010:EPT

[168] Steven P. Ferraro and Faith A. Cole. Ecological periodic tables for nekton usage of four US Pacific Northwest estuarine habitats. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 67(12):1957–1967, December 2010. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/F10-114.

Kaplan:2010:FCS

[169] Isaac C. Kaplan, Phillip S. Levin, Merrick Burden, and Elizabeth A. Fulton. Fishing catch shares in the face of global change: a framework for integrating cumulative impacts and single species management. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 67(12):1968–1982, December 2010. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/F10-118.

Redpath:2010:MBB

[170] Tara D. Redpath, Steven J. Cooke, Cory D. Suski, Robert Arlinghaus, Patrice Couture, David H. Wahl, and David P. Philipp. The metabolic and biochemical basis of vulnerability to recreational angling after three generations of angling-induced selection in a teleost fish. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 67(12):1983–1992, December 2010. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/F10-120.

Finstad:2010:SDA

[171] Anders Gravbrøt Finstad, Sigurd Einum, Leif Magnus Sættem, and Bjart Are Hellen. Spatial distribution of Atlantic salmon (Salmo salar) breeders: among- and within-river variation and predicted consequences for offspring habitat availability. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 67(12):1993–2001, December 2010. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/F10-122.

Jacobson:2010:CFO

[172] Peter C. Jacobson, Heinz G. Stefan, and Donald L. Pereira. Coldwater fish oxythermal habitat in Minnesota lakes: influence of total phosphorus, July air temperature, and relative depth. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 67(12):2002–2013, December 2010. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/F10-115.

White:2010:LEC

[173] J. Wilson White, Kerry J. Nickols, Linden Clarke, and John L. Largier. Larval entrainment in cooling water intakes: spatially explicit models reveal effects on benthic metapopulations and shortcomings of traditional assessments. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 67(12):2014– 2031, December 2010. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/F10-108.

Ruiz:2010:HME

[174] Philippe Ruiz and Christophe Laplanche. A hierarchical model to estimate the abundance and biomass of salmonids by using removal sampling and biometric data from multiple locations. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 67(12):2032–2044, December 2010. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/F10-123.

Mages:2010:ESL

[175] Paul A. Mages and Lawrence M. Dill. The effect of sea lice (Lepeophtheirus salmonis) on juvenile pink salmon (Oncorhynchus gorbuscha) swimming endurance. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 67(12): 2045–2051, December 2010. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/F10-121.

Armstrong:2010:CEC

[176] Jonathan B. Armstrong. Comment on “Egg consumption in mature Pacific salmon (Oncorhynchus spp.)”. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 67(12):2052–2054, December 2010. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/F10-127. See [7].

Garner:2010:IFF

[177] Shawn R. Garner, John W. Heath, and Bryan D. Neff. The importance of freshwater feeding in mature Pacific salmon: a reply to the comment by Armstrong on “Egg consumption in mature Pacific salmon (Oncorhynchus spp.)”. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 67(12): 2055–2057, December 2010. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/F10-128.

Jones:2010:EIL

[178] Nicholas E. Jones. Erratum: Incorporating lakes within the river discontinuum: longitudinal changes in ecological characteristics in stream–lake networks. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 67(12):2058, December 2010. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/F10-142. See

[118].

Anonymous:2010:AIV

[179] Anonymous. Author index: Volume 67, 2010. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 67(12), December 2010. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f10-903.

Anonymous:2010:CV

[180] Anonymous. Contents: Volume 67, 2010. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 67(12), December 2010. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f10-904.

Nielsen:2011:ETG

[181] Jennifer L. Nielsen, Sara M. Turner, and Christian E. Zimmerman. Electronic tags and genetics explore variation in migrating steelhead kelts (Oncorhynchus mykiss), Ninilchik River, Alaska. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 68(1):1–16, January 2011. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/F10-124.

Krkosek:2011:SLL

[182] Martin Krkošek and Ray Hilborn. Sea lice (Lepeophtheirus salmonis) infestations and the productivity of pink salmon (Oncorhynchus gorbuscha) in the Broughton Archipelago, British Columbia, Canada. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 68(1):17–29, January 2011. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/F10-137.

Lundquist:2011:ELP

[183] Carolyn J. Lundquist and Louis W. Botsford. Estimating larval production of a broadcast spawner: the influence of density, aggregation, and the fertilization Allee effect. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 68 (1):30–42, January 2011. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/F10-125.

Teichert:2011:ESD

[184] Maxim A. K. Teichert, Anders Foldvik, Torbjørn Forseth, Ola Ugedal, Sigurd Einum, Anders G. Finstad, Richard D. Hedger, and Edwige Bellier. Effects of spawning distribution on juvenile Atlantic salmon (Salmo salar) density and growth. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 68 (1):43–50, January 2011. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/F10-141.

Laurel:2011:EET

[185] Benjamin J. Laurel, Thomas P. Hurst, and Lorenzo Ciannelli. An experimental examination of temperature interactions in the Match–Mismatch Hypothesis for Pacific cod larvae. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 68(1):51–61, January 2011. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/F10-130.

Schueller:2011:MVP

[186] Amy M. Schueller and Daniel B. Hayes. Minimum viable population size for lake sturgeon (Acipenser fulvescens) using an individual-based model of demographics and genetics. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 68(1):62–73, January 2011. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/F10-129.

Gucker:2011:USA

[187] Björn Gücker, Mario Brauns, Angelo G. Solimini, Maren Voss, Norbert Walz, and Martin T. Pusch. Urban stressors alter the trophic basis of secondary production in an agricultural stream. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 68(1):74–88, January 2011. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/F10-126.

Hess:2011:SGC

[188] Jon E. Hess, Russell D. Vetter, and Paul Moran. A steep genetic cline in yellowtail rockfish, Sebastes flavidus, suggests regional isolation across the Cape Mendocino faunal break. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 68(1):89–104, January 2011. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/F10-131.

Clarke:2011:PDO

[189] Lora M. Clarke, Simon R. Thorrold, and David O. Conover. Population differences in otolith chemistry have a genetic basis in Menidia menidia. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 68(1):105–114, January 2011. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/F10-147.

Kvingedal:2011:IIS

[190] Eli Kvingedal and Sigurd Einum. Intracohort and intercohort spatial density dependence in juvenile brown trout (Salmo trutta). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 68(1):115–121, January 2011. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/F10-132.

Chang:2011:ICC

[191] Yi-Jay Chang, Chi-Lu Sun, Yong Chen, Yuying Zhang, and Su-Zan Yeh. Incorporating climate changes into population dynamic modelling: an individual-based modelling approach for lobster. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 68(1):122–136, January 2011. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/F10-148.

Vinson:2011:SVC

[192] Mark R. Vinson and Phaedra Budy. Sources of variability and comparability between salmonid stomach contents and isotopic analyses: study design lessons and recommendations. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 68(1):137–151, January 2011. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/F10-117.

Rose:2011:IMW

[193] George A. Rose, R. John Nelson, and Luiz G. S. Mello. Isolation or metapopulation: whence and whither the Smith Sound cod? Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 68(1):152–169, January 2011. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/F10-135.

VanderZanden:2011:BWR

[194] M. Jake Vander Zanden and Claudio Gratton. Blowin’ in the wind: reciprocal airborne carbon fluxes between lakes and land. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 68(1):170–182, January 2011. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/F10-157.

Berejikian:2011:CMS

[195] Barry A. Berejikian, Donald M. Van Doornik, Rob C. Endicott, Timothy L. Hoffnagle, Eugene P. Tezak, Megan E. Moore, and Jeffery Atkins. Corrigendum: Mating success of alternative male phenotypes and evidence for frequency-dependent selection in Chinook salmon, Oncorhynchus tshawytscha. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 68 (1):183, January 2011. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/F10-155. See [166].

Hoskin:2011:VPR

[196] M. G. Hoskin, R. A. Coleman, E. von Carlshausen, and C. M. Davis. Variable population responses by large decapod crustaceans to the establishment of a temperate marine no-take zone. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 68(2):185–200, February 2011. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/F10-143.

Kennedy:2011:GHR

[197] James Kennedy, Agnes C. Gundersen, Åge S. Høines, and Olav S. Kjesbu. Greenland halibut (Reinhardtius hippoglossoides) spawn annually but successive cohorts of oocytes develop over 2 years, complicating correct assessment of maturity. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 68(2): 201–209, February 2011. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/F10-149.

Dammannagoda:2011:GAR

[198] Sudath T. Dammannagoda, David A. Hurwood, and Peter B. Mather. Genetic analysis reveals two stocks of skipjack tuna (Katsuwonus pelamis) in the northwestern Indian Ocean. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 68(2):210–223, February 2011. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://

cdnsciencepub.com/doi/full/10.1139/F10-136.

Hesthagen:2011:LRA

[199] Trygve Hesthagen, Bjørn M. Larsen, and Peder Fiske. Liming restores Atlantic salmon (Salmo salar) populations in acidified Norwegian rivers. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 68(2):224–231, February 2011. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/F10-133.

Duffy:2011:RGE

[200] Elisabeth J. Duffy and David A. Beauchamp. Rapid growth in the early marine period improves the marine survival of Chinook salmon (Oncorhynchus tshawytscha) in Puget Sound, Washington. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 68(2):232–240, February 2011. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/F10-144.

Nendick:2011:SLI

[201] L. Nendick, M. Sackville, S. Tang, C. J. Brauner, and A. P. Farrell. Sea lice infection of juvenile pink salmon (Oncorhynchus gorbuscha): effects on swimming performance and postexercise ion balance. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 68(2):241–249, February 2011. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/F10-150.

Hruska:2011:ERR

[202] Kimberly A. Hruska, Scott G. Hinch, David A. Patterson, and Michael C. Healey. Egg retention in relation to arrival timing and reproductive longevity in female sockeye salmon (Oncorhynchus nerka). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 68(2):250–259, February 2011. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/F10-153.

Cantin:2011:ETD

[203] Ariane Cantin, Beatrix E. Beisner, John M. Gunn, Yves T. Prairie, and Jennifer G. Winter. Effects of thermocline deepening on lake plankton communities. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 68(2):260–276, February 2011. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
F10-138.

Ruegg:2011:SSA

[204] Janine Rüegg, Scott D. Tiegs, Dominic T. Chaloner, Peter S. Levi, Jennifer L. Tank, and Gary A. Lamberti. Salmon subsidies alleviate nutrient limitation of benthic biofilms in southeast Alaska streams. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 68(2):277–287, February 2011. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/F10-145.

Caselle:2011:GVD

[205] Jennifer E. Caselle, Scott L. Hamilton, Donna M. Schroeder, Milton S. Love, Julie D. Standish, Jorge A. Rosales-Casián, and Oscar Sosa-Nishizaki. Geographic variation in density, demography, and life history traits of a harvested, sex-changing, temperate reef fish. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 68(2):288–303, February 2011. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/F10-140.

Shaddick:2011:HDC

[206] Kim Shaddick, Dean M. Gilligan, Christopher P. Burridge, Dean R. Jerry, Kiet Truong, and Luciano B. Beheregaray. Historic divergence with contemporary connectivity in a catadromous fish, the estuary perch (Macquaria colonorum). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 68(2): 304–318, February 2011. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/F10-139.

Higgins:2011:EDI

[207] S. N. Higgins, M. J. Vander Zanden, L. N. Joppa, and Y. Vadeboncoeur. The effect of dreissenid invasions on chlorophyll and the chlorophyll : total phosphorus ratio in north-temperate lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 68(2):319–329, February 2011. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/F10-134.

Gray:2011:EHT

[208] Suzanne M. Gray, Francine L. Hart, Maude E. M. Tremblay, Thomas J. Lisney, and Craig W. Hawryshyn. The effects of handling time, ambient light, and anaesthetic method, on the standardized measurement of fish colouration. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 68(2): 330–342, February 2011. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/F10-151.

Babcock:2011:HUR

[209] Elizabeth A. Babcock and Alec D. MacCall. How useful is the ratio of fish density outside versus inside no-take marine reserves as a metric for fishery management control rules? Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 68(2):343–359, February 2011. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/F10-146.

Phillis:2011:EOS

[210] Corey C. Phillis, David J. Ostrach, B. Lynn Ingram, and Peter K. Weber. Evaluating otolith Sr/Ca as a tool for reconstructing estuarine habitat use. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 68(2):360–373, February 2011. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/F10-152.

Fagan:2011:AUC

[211] Kelly-Anne Fagan, Marten A. Koops, Michael T. Arts, and Michael Power. Assessing the utility of C:N ratios for predicting lipid content in fishes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 68(2):374–385, February 2011. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/F10-119.

Anonymous:2011:T

[212] Anonymous. Tribute. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 68(3), March 2011. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f10-163.

Anonymous:2011:H

[213] Anonymous. Hommage. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 68(3), March 2011. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f2011-900.

Weidel:2011:RCC

[214] Brian C. Weidel, Stephen R. Carpenter, James F. Kitchell, and M. Jake Vander Zanden. Rates and components of carbon turnover in fish muscle: insights from bioenergetics models and a whole-lake ^{13}C addition.

Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 68(3):387–399, March 2011. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/F10-158.

Glaser:2011:DFC

[215] Sarah M. Glaser, Hao Ye, Mark Maunder, Alec MacCall, Michael Fogarty, and George Sugihara. Detecting and forecasting complex nonlinear dynamics in spatially structured catch-per-unit-effort time series for North Pacific albacore (Thunnus alalunga). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 68(3):400–412, March 2011. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/F10-160.

Davies:2011:INF

[216] Trevor D. Davies and Ian D. Jonsen. Identifying nonproportionality of fishery-independent survey data to estimate population trends and assess recovery potential for cusk (Brosme brosme). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 68(3):413–425, March 2011. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/F10-165.

Daewel:2011:LHS

[217] Ute Daewel, Myron A. Peck, and Corinna Schrum. Life history strategy and impacts of environmental variability on early life stages of two marine fishes in the North Sea: an individual-based modelling approach. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 68(3):426–443, March 2011. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/F10-164.

Jensen:2011:SMG

[218] Arne Johan Jensen, Peder Fiske, Lars Petter Hansen, Bjørn Ove Johnsen, Kjell Arne Mork, and Tor Fredrik Næsje. Synchrony in marine growth among Atlantic salmon (Salmo salar) populations. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 68(3):444–457, March 2011. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/F10-156.

Watanabe:2011:DTS

[219] Kyuji Watanabe, Kazumi Sakuramoto, Hideki Sugiyama, and Naoki Suzuki. Dynamics of two sailfin sandfish (Arctoscopus japonicus) stocks in the Sea of Japan, and their management. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 68(3):458–468, March 2011. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/F10-161.

Rochet:2011:DSF

[220] Marie-Joëlle Rochet, Jeremy S. Collie, Simon Jennings, and Stephen J. Hall. Does selective fishing conserve community biodiversity? Predictions from a length-based multispecies model. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 68(3):469–486, March 2011. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/F10-159.

Lacoursiere:2011:MRT

[221] Sophie Lacoursière, Isabelle Lavoie, Marco A. Rodríguez, and Stéphane Campeau. Modeling the response time of diatom assemblages to simulated water quality improvement and degradation in running waters. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 68(3):487–497, March 2011. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/F10-162.

Gayeski:2011:HAP

[222] N. Gayeski, B. McMillan, and P. Trotter. Historical abundance of Puget Sound steelhead, Oncorhynchus mykiss, estimated from catch record data. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 68(3):498–510, March 2011. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/F10-166. See

corrigendum [273].

Chilcote:2011:RRP

[223] M. W. Chilcote, K. W. Goodson, and M. R. Falcy. Reduced recruitment performance in natural populations of anadromous salmonids associated with hatchery-reared fish. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 68 (3):511–522, March 2011. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/

10.1139/F10-168. See corrigendum [584].

Vrieze:2011:OMS

[224] Lance A. Vrieze, Roger A. Bergstedt, and Peter W. Sorensen. Olfactorymediated stream-finding behavior of migratory adult sea lamprey (Petromyzon marinus). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 68 (3):523–533, March 2011. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/F10-169.

Houde:2011:MPE

[225] Aimee Lee S. Houde, Dylan J. Fraser, Patrick O’Reilly, and Jeffrey A. Hutchings. Maternal and paternal effects on fitness correlates in outbred and inbred Atlantic salmon (Salmo salar). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 68(3):534–549, March 2011. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f2011-001.

Beacham:2011:CPG

[226] Terry D. Beacham, B. McIntosh, and C. G. Wallace. A comparison of polymorphism of genetic markers and population sample sizes required for mixed-stock analysis of sockeye salmon (Oncorhynchus nerka) in British Columbia. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 68 (3):550–562, March 2011. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/F10-167.

Sissener:2011:GMP

[227] Nini Hedberg Sissener, Monica Sanden, Åshild Krogdahl, Anne-Marie Bakke, Lene Elisabeth Johannessen, and Gro-Ingunn Hemre. Genetically modified plants as fish feed ingredients. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 68(3):563–574, March 2011. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/F10-154.

Arseneau:2011:REB

[228] Kristina M. A. Arseneau, Charles T. Driscoll, Lindsay M. Brager, Karen A. Ross, and Brian F. Cumming. Recent evidence of biological recovery from acidification in the Adirondacks (New York, USA): a multiproxy paleolimnological investigation of Big Moose Lake. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 68(4):575–592, April 2011. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f2011-003.

Alvarez:2011:MAH

[229] David Álvarez and Eva Garcia-Vazquez. Maintenance of asymmetric hybridization between Atlantic salmon (Salmo salar) and brown trout (Salmo trutta) via postzygotic barriers and paternal effects. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 68(4):593–602, April 2011. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:/

/cdnsciencepub.com/doi/full/10.1139/f2011-004. See corrigendum

[256].

Miller:2011:LHV

[230] Jessica A. Miller, Virginia L. Butler, Charles A. Simenstad, David H. Backus, and Adam J. R. Kent. Life history variation in upper Columbia River Chinook salmon (Oncorhynchus tshawytscha): a comparison using modern and ≈ 500-year-old archaeological otoliths. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 68(4):603–617, April 2011. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f2011-002.

Lambert:2011:EFL

[231] Yvan Lambert. Environmental and fishing limitations to the rebuilding of the northern Gulf of St. Lawrence cod stock (Gadus morhua). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 68(4):618–631, April 2011. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f2011-007. See er-

ratum [406].

Parsons:2011:FPB

[232] Darren M. Parsons, Mark A. Morrison, Jeremy R. McKenzie, Bruce W. Hartill, Richard Bian, and R. I. C. Chris Francis. A fisheries perspective of behavioural variability: differences in movement behaviour and extraction rate of an exploited sparid, snapper (Pagrus auratus). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 68(4):632–642, April 2011. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f2011-005.

Duong:2011:EME

[233] Thuy Yen Duong, Kim T. Scribner, James A. Crossman, Patrick S. Forsythe, and Edward A. Baker. Environmental and maternal effects on embryonic and larval developmental time until dispersal of lake sturgeon (Acipenser fulvescens). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 68 (4):643–654, April 2011. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f2011-008.

Kim:2011:DJA

[234] Jae-Woo Kim, James W. A. Grant, and Grant E. Brown. Do juvenile Atlantic salmon (Salmo salar) use chemosensory cues to detect and avoid risky habitats in the wild? Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 68 (4):655–662, April 2011. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f2011-011.

Clair:2011:WCD

[235] Thomas A. Clair, Ian F. Dennis, and Robert Vet. Water chemistry and dissolved organic carbon trends in lakes from Canada’s Atlantic provinces: no recovery from acidification measured after 25 years of lake monitoring. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 68(4):663–674, April 2011. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f2011-013.

Selbie:2011:CCM

[236] Daniel T. Selbie, Jon N. Sweetman, Peter Etherton, Kim D. Hyatt, D. Paul Rankin, Bruce P. Finney, and John P. Smol. Climate change modulates structural and functional lake ecosystem responses to introduced anadromous salmon. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 68 (4):675–692, April 2011. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f2011-006.

OConnor:2011:CST

[237] Constance M. O’Connor, Kathleen M. Gilmour, Robert Arlinghaus, Shuichi Matsumura, Cory D. Suski, David P. Philipp, and Steven J.

Cooke. The consequences of short-term cortisol elevation on individual physiology and growth rate in wild largemouth bass (Micropterus salmoides). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 68(4):693–705, April 2011. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f2011-009.

Busch:2011:ECB

[238] Susan Busch, Brett M. Johnson, and Thomas Mehner. Energetic costs and benefits of cyclic habitat switching: a bioenergetics model analysis of diel vertical migration in coregonids. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 68(4):706–717, April 2011. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f2011-012.

Healey:2011:CIC

[239] Michael Healey. The cumulative impacts of climate change on Fraser River sockeye salmon (Oncorhynchus nerka) and implications for management. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 68(4):718–737, April 2011. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f2011-010.

Hagler:2011:URH

[240] Megan M. Hagler, Mary C. Freeman, Seth J. Wenger, Byron J. Freeman, Patrick L. Rakes, and J. R. Shute. Use of recent and historical records to estimate status and trends of a rare and imperiled stream fish, Percina jenkinsi (Percidae). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 68 (5):739–748, May 2011. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f2011-014.

Wilson:2011:CTF

[241] Alexander D. M. Wilson, Thomas R. Binder, Keegan P. McGrath, Steven J. Cooke, and Jean-Guy J. Godin. Capture technique and fish personality: angling targets timid bluegill sunfish, Lepomis macrochirus. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 68(5):749–757, May 2011. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f2011-019.

Braun:2011:RBH

[242] Douglas C. Braun and John D. Reynolds. Relationships between habitat characteristics and breeding population densities in sockeye salmon (Oncorhynchus nerka). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 68 (5):758–767, May 2011. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f2011-015.

Tsai:2011:MCR

[243] Jeng-Wei Tsai, Timothy K. Kratz, Paul C. Hanson, Nobuaki Kimura, Wen-Cheng Liu, Fang-Pan Lin, Hsiu-Mei Chou, Jiunn-Tzong Wu, and Chih-Yu Chiu. Metabolic changes and the resistance and resilience of a subtropical heterotrophic lake to typhoon disturbance. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 68(5):768–780, May 2011. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f2011-024.

Cadigan:2011:CIT

[244] Noel G. Cadigan. Confidence intervals for trawlable abundance from stratified-random bottom trawl surveys. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 68(5):781–794, May 2011. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f2011-026.

Johannsson:2011:MDP

[245] Ora E. Johannsson, Kelly L. Bowen, Kristen T. Holeck, and Maureen G. Walsh. Mysis diluviana population and cohort dynamics in Lake Ontario before and after the establishment of Dreissena spp., Cercopagis pengoi, and Bythotrephes longimanus. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 68(5):795–811, May 2011. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f2011-028.

McAdam:2011:ESC

[246] Steven O. McAdam. Effects of substrate condition on habitat use and survival by white sturgeon (Acipenser transmontanus) larvae and potential implications for recruitment. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 68 (5):812–822, May 2011. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f2011-021.

Ramsay:2011:CPS

[247] Alice L. Ramsay, Nigel J. Milner, Roger N. Hughes, and Ian D. Mc-Carthy. Comparison of the performance of scale and otolith microchemistry as fisheries research tools in a small upland catchment. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 68(5):823–833, May 2011. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f2011-027.

Berg:2011:EFP

[248] Casper Willestofte Berg and Axel Temming. Estimation of feeding patterns for piscivorous fish using individual prey data from stomach contents. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 68(5):834–841, May 2011. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f2011-020.

Egemose:2011:CCP

[249] Sara Egemose, Inmaculada de Vicente, Kasper Reitzel, Mogens R. Flindt, Frede Ø. Andersen, Torben L. Lauridsen, Martin Søndergaard, Erik Jeppesen, and Henning S. Jensen. Changed cycling of P, N, Si, and DOC in Danish Lake Nordborg after aluminum treatment. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 68(5):842–856, May 2011. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f2011-016.

Brandt:2011:DHR

[250] Stephen B. Brandt, Marco Costantini, Sarah Kolesar, Stuart A. Ludsin, Doran M. Mason, Christopher M. Rae, and Hongyan Zhang. Does hypoxia reduce habitat quality for Lake Erie walleye (Sander vitreus)? A bioenergetics perspective. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 68 (5):857–879, May 2011. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f2011-018.

Holt:2011:TOB

[251] Kendra R. Holt, Randall M. Peterman, and Sean P. Cox. Trade-offs between monitoring objectives and monitoring effort when classifying regional conservation status of Pacific salmon (Oncorhynchus spp.) populations. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 68(5):880–897, May 2011. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f2011-022.

Hobday:2011:SFT

[252] Alistair J. Hobday, Jason R. Hartog, Claire M. Spillman, and Oscar Alves. Seasonal forecasting of tuna habitat for dynamic spatial management. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 68(5):898–911, May 2011. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f2011-031.

Haltuch:2011:PPI

[253] Melissa A. Haltuch and André E. Punt. The promises and pitfalls of including decadal-scale climate forcing of recruitment in groundfish stock assessment. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 68(5):912–926, May 2011. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f2011-030.

Sistiaga:2011:ULC

[254] Manu Sistiaga, Bent Herrmann, Kåre N. Nielsen, and Roger B. Larsen. Understanding limits to cod and haddock separation using size selectivity in a multispecies trawl fishery: an application of FISHSELECT. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 68(5):927, May 2011. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f2011-017.

Young:2011:GSM

[255] Michael K. Young. Generation-scale movement patterns of cutthroat trout (Oncorhynchus clarkii pleuriticus) in a stream network. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 68(5):941–951, May 2011. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f2011-023.

Alvarez:2011:CMA

[256] David Álvarez and Eva Garcia-Vazquez. Corrigendum: Maintenance of asymmetric hybridization between Atlantic salmon (Salmo salar) and brown trout (Salmo trutta) via postzygotic barriers and paternal effects.

Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 68(5):952, May 2011. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f2011-063. See [229].

Anonymous:2011:EDV

[257] Anonymous. Erratum: DOIs in volume 68, issues 3 and 4, of the Canadian Journal of Fisheries and Aquatic Science / erratum : les DOI du volume 68, numéros 3 et 4, du Journal canadien des sciences halieutiques et aquatiques. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 68(5): 953, May 2011. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/

f2011-068. From the erratum: “Inadvertently, papers were assigned DOIs identical to ones assigned to previously published papers. The DOIs on the web versions have been changed so that the papers could be deposited with our aggregators. However, the print versions have the incorrect DOIs and readers should refer to the table below for their correct DOIs.”.

Catalano:2011:WLD

[258] Matthew J. Catalano and Micheal S. Allen. A whole-lake density reduction to assess compensatory responses of gizzard shad Dorosoma cepedianum. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 68(6):955–968, June 2011. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f2011-036.

Einum:2011:RIS

[259] Sigurd Einum and Eli Kvingedal. Relative importance of size-based competitive ability and degree of niche overlap in inter-cohort competition of Atlantic salmon (Salmo salar) juveniles. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 68(6):969–976, June 2011. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f2011-042.

Perrier:2011:CGO

[260] Charles Perrier, Françoise Daverat, Guillaume Evanno, Christophe Pécheyran, Jean-Luc Bagliniere, and Jean-Marc Roussel. Coupling genetic and otolith trace element analyses to identify river-born fish with hatchery pedigrees in stocked Atlantic salmon (Salmo salar) populations. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 68(6):977–987, June 2011. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f2011-040.

Wenger:2011:RCI

[261] Seth J. Wenger, Daniel J. Isaak, Jason B. Dunham, Kurt D. Fausch, Charles H. Luce, Helen M. Neville, Bruce E. Rieman, Michael K. Young, David E. Nagel, Dona L. Horan, and Gwynne L. Chandler. Role of climate and invasive species in structuring trout distributions in the interior Columbia River Basin, USA. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 68 (6):988–1008, June 2011. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/

10.1139/f2011-034. See corrigendum [302].

deCatanzaro:2011:ELV

[262] Rachel deCatanzaro and Patricia Chow-Fraser. Effects of landscape variables and season on reference water chemistry of coastal marshes in eastern Georgian Bay. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 68 (6):1009–1023, June 2011. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f2011-035.

Fenske:2011:ASS

[263] Kari H. Fenske, Michael J. Wilberg, David H. Secor, and Mary C. Fabrizio. An age- and sex-structured assessment model for American eels (Anguilla rostrata) in the Potomac River, Maryland. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 68(6):1024–1037, June 2011. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f2011-038.

Palmer:2011:WQC

[264] Michelle E. Palmer, Norman D. Yan, Andrew M. Paterson, and Robert E. Girard. Water quality changes in south-central Ontario lakes and the role of local factors in regulating lake response to regional stressors. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 68(6):1038–1050, June 2011. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f2011-041.

Sierszen:2011:BPC

[265] Michael E. Sierszen, John R. Kelly, Timothy D. Corry, Jill V. Scharold, and Peder M. Yurista. Benthic and pelagic contributions to Mysis nutrition across Lake Superior. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 68 (6):1051–1063, June 2011. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f2011-033.

Rijnsdorp:2011:SDE

[266] Adriaan D. Rijnsdorp, Jan Jaap Poos, and Floor J. Quirijns. Spatial dimension and exploitation dynamics of local fishing grounds by fishers targeting several flatfish species. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 68 (6):1064–1076, June 2011. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f2011-032.

Sampson:2011:SMF

[267] David B. Sampson and Robert D. Scott. A spatial model for fishery age-selection at the population level. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 68(6):1077–1086, June 2011. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f2011-044.

Sackville:2011:PSO

[268] M. Sackville, S. Tang, L. Nendick, A. P. Farrell, and C. J. Brauner. Pink salmon (Oncorhynchus gorbuscha) osmoregulatory development plays a key role in sea louse (Lepeophtheirus salmonis) tolerance. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 68(6):1087–1096, June 2011. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f2011-037.

Korman:2011:EFF

[269] Josh Korman, Carl Walters, S. J. D. Martell, W. E. Pine, and Andrew Dutterer. Effects of flow fluctuations on habitat use and survival of age-0 rainbow trout (Oncorhynchus mykiss) in a large, regulated river. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 68(6):1097–1109, June 2011. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f2011-045.

Korman:2011:DPD

[270] Josh Korman, S. J. D. Martell, and Carl Walters. Describing population dynamics for early life stages of rainbow trout (Oncorhynchus mykiss) using a stock synthesis model. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 68 (6):1110–1123, June 2011. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f2011-046.

Francis:2011:DWS

[271] R. I. C. Chris Francis. Data weighting in statistical fisheries stock assessment models. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 68(6):1124– 1138, June 2011. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/

f2011-025. See corrigendum [357].

Neff:2011:CEW

[272] Bryan D. Neff, Shawn R. Garner, and Trevor E. Pitcher. Conservation and enhancement of wild fish populations: preserving genetic quality versus genetic diversity. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 68 (6):1139–1154, June 2011. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f2011-029.

Gayeski:2011:CHA

[273] N. Gayeski, B. McMillan, and P. Trotter. Corrigendum: Historical abundance of Puget Sound steelhead, Oncorhynchus mykiss, estimated from catch record data. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 68(6):1155, June 2011. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/

f2011-043. See [222].

Wagner:2011:DOS

[274] C. Michael Wagner, Eric M. Stroud, and Trevor D. Meckley. A deathly odor suggests a new sustainable tool for controlling a costly invasive species. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 68(7):1157–1160, July 2011. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f2011-072.

Moore:2011:NFR

[275] Jonathan W. Moore, Sean A. Hayes, Walter Duffy, Sean Gallagher, Cyril J. Michel, and David Wright. Nutrient fluxes and the recent collapse of coastal California salmon populations. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 68(7):1161–1170, July 2011. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f2011-054.

Zhou:2011:EMC

[276] Shijie Zhou, André E. Punt, Roy Deng, and Janet Bishop. Estimating multifleet catchability coefficients and natural mortality from fishery catch and effort data: comparison of Bayesian state–space and observation error models. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 68 (7):1171–1181, July 2011. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f2011-052.

Feng:2011:OCS

[277] Ming Feng, Nick Caputi, James Penn, Dirk Slawinski, Simon de Lestang, Evan Weller, and Alan Pearce. Ocean circulation, Stokes drift, and connectivity of western rock lobster (Panulirus cygnus) population. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 68(7):1182–1196, July 2011. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f2011-065.

Moland:2011:HRE

[278] Even Moland, Esben Moland Olsen, Kristian Andvord, Jan Atle Knutsen, and Nils Chr. Stenseth. Home range of European lobster (Homarus gammarus) in a marine reserve: implications for future reserve design. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 68(7):1197–1210, July 2011. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f2011-053.

Czesny:2011:FAS

[279] Sergiusz J. Czesny, Jacques Rinchard, S. Dale Hanson, John M. Dettmers, and Konrad Dabrowski. Fatty acid signatures of Lake Michigan prey fish and invertebrates: among-species differences and spatiotemporal variability. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 68(7):1211– 1230, July 2011. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f2011-048.

Honkalehto:2011:UAD

[280] Taina Honkalehto, Patrick H. Ressler, Richard H. Towler, and Christopher D. Wilson. Using acoustic data from fishing vessels to estimate walleye pollock (Theragra chalcogramma) abundance in the eastern Bering Sea. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 68(7):1231–1242, July 2011. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/

f2011-050. See corrigendum [301].

Dean:2011:ASI

[281] Susan Dean, Claudio DiBacco, and Robert Scott McKinley. Assessment of stable isotopic signatures as a means to track the exchange of sea lice (Lepeophtheirus salmonis) between host fish populations. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 68(7):1243–1251, July 2011. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f2011-039.

Peel:2011:HMM

[282] David Peel and Norman M. Good. A hidden Markov model approach for determining vessel activity from vessel monitoring system data. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 68(7):1252–1264, July 2011. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f2011-055.

Makler-Pick:2011:ERF

[283] Vardit Makler-Pick, Gideon Gal, James Shapiro, and Matthew R. Hipsey. Exploring the role of fish in a lake ecosystem (Lake Kinneret, Israel) by coupling an individual-based fish population model to a dynamic ecosystem model. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 68(7):1265– 1284, July 2011. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f2011-051.

Maunder:2011:SSM

[284] Mark N. Maunder and Richard B. Deriso. A state–space multistage life cycle model to evaluate population impacts in the presence of density dependence: illustrated with application to delta smelt (Hyposmesus transpacificus). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 68(7):1285– 1306, July 2011. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f2011-071.

Rollinson:2011:WDE

[285] Njal Rollinson and Jeffrey A. Hutchings. Why does egg size of salmonids increase with the mean size of population spawning gravels? Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 68(8):1307–1315, August 2011. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f2011-069. See er-

ratum [118].

Levi:2011:DTH

[286] Peter S. Levi, Jennifer L. Tank, Scott D. Tiegs, Janine Rüegg, Dominic T. Chaloner, and Gary A. Lamberti. Does timber harvest influence the dynamics of marine-derived nutrients in Southeast Alaska streams? Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 68(8):1316–1329, August 2011. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/

f2011-067. See comment [515] and reply [516].

McMahon:2011:NMR

[287] Kelton W. McMahon, Marilyn L. Fogel, Beverly J. Johnson, Leah A. Houghton, and Simon R. Thorrold. A new method to reconstruct fish diet and movement patterns from δ ^{13}C values in otolith amino acids. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 68(8):1330–1340, August 2011. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f2011-070.

Hayes:2011:STS

[288] Sean A. Hayes, Morgan H. Bond, Chad V. Hanson, Andrew W. Jones, Arnold J. Ammann, Jeffrey A. Harding, Alison L. Collins, Jeffrey Perez, and R. Bruce MacFarlane. Down, up, down and “smolting” twice? Seasonal movement patterns by juvenile steelhead (Oncorhynchus mykiss) in a coastal watershed with a bar closing estuary. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 68(8):1341–1350, August 2011. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f2011-062.

Pecl:2011:EFS

[289] Gretta T. Pecl, Sean R. Tracey, Leonid Danyushevsky, Simon Wotherspoon, and Natalie A. Moltschaniwskyj. Elemental fingerprints of southern calamary (Sepioteuthis australis) reveal local recruitment sources and allow assessment of the importance of closed areas. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 68(8):1351–1360, August 2011. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f2011-059.

Morgan:2011:DII

[290] M. Joanne Morgan, Alfonso Perez-Rodriguez, and Fran Saborido-Rey. Does increased information about reproductive potential result in better prediction of recruitment? Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 68 (8):1361–1368, August 2011. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f2011-049.

Finger:2011:AME

[291] Amanda J. Finger, Eric C. Anderson, Molly R. Stephens, and Bernard P. May. Application of a method for estimating effective population size and admixture using diagnostic single nucleotide polymorphisms (SNPs): implications for conservation of threatened Paiute cutthroat trout (Oncorhynchus clarkii seleniris) in Silver King Creek, California. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 68(8):1369–1386, August 2011. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f2011-075.

Hillary:2011:BIS

[292] Richard Hillary. Bayesian integrated survey-based assessments: an example applied to North Sea herring (Clupea harengus) survey data. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 68(8):1387–1407, August 2011. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f2011-047.

Kurek:2011:ETP

[293] Joshua Kurek, Russ C. Weeber, and John P. Smol. Environment trumps predation and spatial factors in structuring cladoceran communities from Boreal Shield lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 68(8): 1408–1419, August 2011. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f2011-081.

Loland:2011:PBN

[294] Anders Løland, Magne Aldrin, Gunnhildur Högnadóttir Steinbakk, Ragnar Bang Huseby, and Jon Arne Grøttum. Prediction of biomass in Norwegian fish farms. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 68(8): 1420–1434, August 2011. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f2011-078.

Sepulveda-Villet:2011:FSP

[295] Osvaldo J. Sepulveda-Villet and Carol A. Stepien. Fine-scale population genetic structure of the yellow perch Perca flavescens in Lake Erie. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 68(8):1435–1453, August 2011. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f2011-077.

LeFur:2011:CLF

[296] Jean Le Fur, Athanase Guilavogui, and Antoine Teitelbaum. Contribution of local fishermen to improving knowledge of the marine ecosystem and resources in the Republic of Guinea, West Africa. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 68(8):1454–1469, August 2011. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f2011-061.

Skoglund:2011:PPP

[297] Helge Skoglund, Sigurd Einum, Torbjørn Forseth, and Bjørn Torgeir Barlaup. Phenotypic plasticity in physiological status at emergence from nests as a response to temperature in Atlantic salmon (Salmo salar). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 68(8):1470–1479, August 2011. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f2011-056.

Hobbs:2011:BRT

[298] William O. Hobbs, Rolf D. Vinebrooke, and Alexander P. Wolfe. Biogeochemical responses of two alpine lakes to climate change and atmospheric deposition, Jasper and Banff National parks, Canadian Rocky Mountains. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 68(8):1480–1494, August 2011. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f2011-058.

Bretherton:2011:SCA

[299] Welles D. Bretherton, John S. Kominoski, Dylan G. Fischer, and Carri J. LeRoy. Salmon carcasses alter leaf litter species diversity effects on instream decomposition. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 68(8): 1495–1506, August 2011. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f2011-082.

Taylor:2011:CNL

[300] Nathan G. Taylor, Carl J. Walters, and Steven J. D. Martell. Corrigendum: A new likelihood for simultaneously estimating von Bertalanffy growth parameters, gear selectivity, and natural and fishing mortality. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 68(8):1507, August 2011. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f2011-100. See [2].

Honkalehto:2011:CUA

[301] Taina Honkalehto, Patrick H. Ressler, Richard H. Towler, and Christopher D. Wilson. Corrigendum: Using acoustic data from fishing vessels to estimate walleye pollock (Theragra chalcogramma) abundance in the eastern Bering Sea. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 68(8): 1508–1509, August 2011. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f2011-101. See [280].

Wenger:2011:CRC

[302] Seth J. Wenger, Daniel J. Isaak, Jason B. Dunham, Kurt D. Fausch, Charles H. Luce, Helen M. Neville, Bruce E. Rieman, Michael K. Young, David E. Nagel, Dona L. Horan, and Gwynne L. Chandler. Corrigendum: Role of climate and invasive species in structuring trout distributions in the interior Columbia River Basin, USA. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 68(8):1510, August 2011. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f2011-102. See [261].

Medeiros:2011:DCI

[303] Andrew S. Medeiros and Roberto Quinlan. The distribution of the Chironomidae (Insecta: Diptera) along multiple environmental gradients in lakes and ponds of the eastern Canadian Arctic. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 68(9):1511–1527, September 2011. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f2011-076.

Li:2011:CMA

[304] Zhenhong Li, Daniel L. Villeneuve, Kathleen M. Jensen, Gerald T. Ankley, and Karen H. Watanabe. A computational model for asynchronous oocyte growth dynamics in a batch-spawning fish. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 68(9):1528–1538, September 2011. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f2011-066.

Abrantes:2011:DES

[305] Kátya Gisela Abrantes, Jeremy Martin Lyle, Peter D. Nichols, and Jayson Mark Semmens. Do exotic salmonids feed on native fauna after escaping from aquaculture cages in Tasmania, Australia? Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 68(9):1539–1551, September 2011. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f2011-057.

Gaichas:2011:WDD

[306] Sarah K. Gaichas, Kerim Y. Aydin, and Robert C. Francis. What drives dynamics in the Gulf of Alaska? Integrating hypotheses of species, fishing, and climate relationships using ecosystem modeling. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 68(9):1553–1578, September 2011. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f2011-080.

Carlson:2011:WPE

[307] Stephanie Marie Carlson and William Hallowell Satterthwaite. Weakened portfolio effect in a collapsed salmon population complex. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 68(9):1579–1589, September 2011. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f2011-084.

Godbout:2011:SSO

[308] L. Godbout, C. C. Wood, R. E. Withler, S. Latham, R. J. Nelson, L. Wetzel, R. Barnett-Johnson, M. J. Grove, A. K. Schmitt, and K. D. McKeegan. Sockeye salmon (Oncorhynchus nerka) return after an absence of nearly 90 years: a case of reversion to anadromy. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 68(9):1590–1602, September 2011. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f2011-089.

Zhou:2011:BFB

[309] Shijie Zhou, André E. Punt, Roy Deng, Marco Kienzle, and Wayne Rochester. Bayesian fishable biomass dynamics models incorporating fished area and relative fish density. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 68(9):1603–1614, September 2011. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f2011-083.

Naughton:2011:IPC

[310] George P. Naughton, Matthew L. Keefer, Tami S. Clabough, Michael A. Jepson, Steven R. Lee, Christopher A. Peery, and Christopher C. Caudill. Influence of pinniped-caused injuries on the survival of adult Chinook salmon (Oncorhynchus tshawytscha) and steelhead trout (Oncorhynchus mykiss) in the Columbia River basin. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 68(9):1615–1624, September 2011. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f2011-064.

Hulson:2011:IMR

[311] Peter-John F. Hulson, Sara E. Miller, James N. Ianelli, and Terrance J. Quinn. Including mark–recapture data into a spatial age-structured model: walleye pollock (Theragra chalcogramma) in the eastern Bering Sea. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 68(9):1625–1634, September 2011. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f2011-060.

Hubley:2011:MEM

[312] P. Bradford Hubley and A. Jamie F. Gibson. A model for estimating mortality of Atlantic salmon, Salmo salar, between spawning events. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 68(9):1635–1650, September 2011. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f2011-074.

Zimmermann:2011:DSM

[313] Fabian Zimmermann, Mikko Heino, and Stein Ivar Steinshamn. Does size matter? A bioeconomic perspective on optimal harvesting when price is size-dependent. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 68(9):1651– 1659, September 2011. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f2011-093.

Abdul-Aziz:2011:PCC

[314] Omar I. Abdul-Aziz, Nathan J. Mantua, and Katherine W. Myers. Potential climate change impacts on thermal habitats of Pacific salmon (Oncorhynchus spp.) in the North Pacific Ocean and adjacent seas. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 68(9):1660–1680, September 2011. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f2011-079.

Thorson:2011:AFS

[315] James T. Thorson, Ian J. Stewart, and André E. Punt. Accounting for fish shoals in single- and multi-species survey data using mixture distribution models. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 68(9):1681– 1693, September 2011. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f2011-086.

Rollinson:2011:EWD

[316] Njal Rollinson and Jeffrey A. Hutchings. Erratum: Why does egg size of salmonids increase with the mean size of population spawning gravels? Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 68(9):1694, September 2011. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/

f2011-121. See [285].

White:2011:WLT

[317] Michael S. White, Marguerite A. Xenopoulos, Robert A. Metcalfe, and Keith M. Somers. Water level thresholds of benthic macroinvertebrate richness, structure, and function of boreal lake stony littoral habitats. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 68(10):1695–1704, October 2011. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f2011-094.

Erisman:2011:IPH

[318] Brad E. Erisman, Larry G. Allen, Jeremy T. Claisse, Daniel J. Pondella, Eric F. Miller, and Jason H. Murray. The illusion of plenty: hyperstability masks collapses in two recreational fisheries that target fish spawning aggregations. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 68(10): 1705–1716, October 2011. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f2011-090.

Brinch:2011:MLE

[319] Christian N. Brinch, Anne Maria Eikeset, and Nils Chr. Stenseth. Maximum likelihood estimation in nonlinear structured fisheries models using survey and catch-at-age data. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 68(10):1717–1731, October 2011. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f2011-085.

Kemp:2011:OCM

[320] Jodie Kemp, Stephen E. Swearer, Gregory P. Jenkins, and Simon Robertson. Otolith chemistry is more accurate than otolith shape in identifying cod species (genus Pseudophycis) in the diet of Australian fur seals (Arctocephalus pusillus doriferus). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 68(10):1732–1743, October 2011. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f2011-088.

Methot:2011:ABD

[321] Richard D. Methot, Jr. and Ian G. Taylor. Adjusting for bias due to variability of estimated recruitments in fishery assessment models. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 68(10):1744–1760, October 2011. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f2011-092.

Garrison:2011:CDC

[322] Thomas M. Garrison, Owen S. Hamel, and André E. Punt. Can data collected from marine protected areas improve estimates of life-history parameters? Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 68(10):1761–1777, October 2011. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f2011-073.

Mikulyuk:2011:RRE

[323] Alison Mikulyuk, Sapna Sharma, Scott Van Egeren, Eric Erdmann, Michelle E. Nault, and Jennifer Hauxwell. The relative role of environmental, spatial, and land-use patterns in explaining aquatic macrophyte community composition. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 68(10): 1778–1789, October 2011. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f2011-095.

Schank:2011:STD

[324] C. M. M. Schank, C. A. Paszkowski, W. M. Tonn, and G. J. Scrimgeour. Stocked trout do not significantly affect wood frog populations in boreal foothills lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 68(10): 1790–1801, October 2011. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f2011-097.

Colborne:2011:MSP

[325] S. F. Colborne, M. C. Bellemare, P. R. Peres-Neto, and B. D. Neff. Morphological and swim performance variation among reproductive tactics of bluegill sunfish (Lepomis macrochirus). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 68(10):1802–1810, October 2011. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f2011-091.

VanLeeuwen:2011:FPM

[326] Travis E. Van Leeuwen, Jordan S. Rosenfeld, and Jeffrey G. Richards. Failure of physiological metrics to predict dominance in juvenile Pacific salmon (Oncorhynchus spp.): habitat effects on the allometry of growth in dominance hierarchies. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 68(10):1811–1818, October 2011. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f2011-099.

Louhi:2011:DFS

[327] P. Louhi, M. Ovaska, A. Mäki-Petäys, J. Erkinaro, and T. Muotka. Does fine sediment constrain salmonid alevin development and survival? Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 68(10):1819–1826, October 2011. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f2011-106.

Kitada:2011:BSR

[328] Shuichi Kitada, Hirohisa Kishino, and Katsuyuki Hamasaki. Bias and significance of relative reproductive success estimates based on steelhead trout (Oncorhynchus mykiss) data: a Bayesian meta-analysis. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 68(10):1827–1835, October 2011. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f2011-087.

Holeton:2011:WRI

[329] Claire Holeton, Patricia A. Chambers, and Laura Grace. Wastewater release and its impacts on Canadian waters. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 68(10):1836–1859, October 2011. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f2011-096.

Purcell:2011:RGS

[330] Catherine M. Purcell and Suzanne Edmands. Resolving the genetic structure of striped marlin, Kajikia audax, in the Pacific Ocean through spatial and temporal sampling of adult and immature fish. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 68(11):1861–1875, November 2011. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f2011-104.

Hess:2011:IHA

[331] Jon E. Hess, Andrew P. Matala, Joseph S. Zendt, Chris R. Frederiksen, Bill Sharp, and Shawn R. Narum. Introgressive hybridization among major Columbia River Chinook salmon (Oncorhynchus tshawytscha) lineages within the Klickitat River due to hatchery practices. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 68(11):1876–1891, November 2011. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f2011-107.

Garcia-Asorey:2011:COT

[332] Martín Ignacio García-Asorey, Gabriela Escati-Peñaloza, Ana María Parma, and Miguel Alberto Pascual. Conflicting objectives in trophy trout recreational fisheries: evaluating trade-offs using an individualbased model. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 68(11):1892– 1904, November 2011. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f2011-108.

Carlisle:2011:SCD

[333] Aaron B. Carlisle, Christopher R. Perle, Kenneth J. Goldman, and Barbara A. Block. Seasonal changes in depth distribution of salmon sharks (Lamna ditropis) in Alaskan waters: implications for foraging ecology. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 68(11):1905–1921, November 2011. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f2011-105.

Brewer:2011:IBN

[334] Shannon K. Brewer and Charles F. Rabeni. Interactions between naturaloccurring landscape conditions and land use influencing the abundance of riverine smallmouth bass, Micropterus dolomieu. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 68(11):1922–1933, November 2011. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f2011-110.

Chapman:2011:ABT

[335] Erik W. Chapman, Christian Jørgensen, and Molly E. Lutcavage. Atlantic bluefin tuna (Thunnus thynnus): a state-dependent energy allocation model for growth, maturation, and reproductive investment. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 68(11):1934–1951, November 2011. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f2011-109.

DeHaan:2011:UGM

[336] Patrick W. DeHaan, Shana R. Bernall, Joseph M. DosSantos, Lawrence L. Lockard, and William R. Ardren. Use of genetic markers to aid in re-establishing migratory connectivity in a fragmented metapopulation of bull trout (Salvelinus confluentus). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 68(11):1952–1969, November 2011. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f2011-098.

Eigaard:2011:FPI

[337] Ole Ritzau Eigaard, Bjarti Thomsen, Holger Hovgaard, Anders Nielsen, and Adriaan D. Rijnsdorp. Fishing power increases from technological development in the Faroe Islands longline fishery. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 68(11):1970–1982, November 2011. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f2011-103.

Ormond:2011:EDT

[338] Chad I. Ormond, Jordan S. Rosenfeld, and Eric B. Taylor. Environmental determinants of threespine stickleback species pair evolution and persistence. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 68(11):1983–1997, November 2011. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f2011-113.

Bez:2011:FAT

[339] Nicolas Bez, Emily Walker, Daniel Gaertner, Jacques Rivoirard, and Philippe Gaspar. Fishing activity of tuna purse seiners estimated from vessel monitoring system (VMS) data. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 68(11):1998–2010, November 2011. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f2011-114.

Soto:2011:ESD

[340] David X. Soto, Leonard I. Wassenaar, Keith A. Hobson, and Jordi Catalan. Effects of size and diet on stable hydrogen isotope values (δ d) in fish: implications for tracing origins of individuals and their food sources. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 68(11):2011–2019, November 2011. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f2011-112.

Swanson:2011:QIM

[341] Heidi K. Swanson, Karen A. Kidd, and James D. Reist. Quantifying importance of marine prey in the diets of two partially anadromous fishes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 68(11):2020–2028, November 2011. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f2011-111.

Hussey:2011:SIP

[342] Nigel E. Hussey, Sheldon F. J. Dudley, Ian D. McCarthy, Geremy Cliff, and Aaron T. Fisk. Stable isotope profiles of large marine predators: viable indicators of trophic position, diet, and movement in sharks? Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 68(12):2029–2045, December 2011. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f2011-115.

Ferriss:2011:RPM

[343] Bridget E. Ferriss and Timothy E. Essington. Regional patterns in mercury and selenium concentrations of yellowfin tuna (Thunnus albacares) and bigeye tuna (Thunnus obesus) in the Pacific Ocean. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 68(12):2046–2056, December 2011. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f2011-120.

White:2011:RTP

[344] Shannon L. White, Charles Gowan, Kurt D. Fausch, Josh G. Harris, and W. Carl Saunders. Response of trout populations in five Colorado streams two decades after habitat manipulation. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 68(12):2057–2063, December 2011. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f2011-125.

Hoffman:2011:EPE

[345] Joel C. Hoffman, John R. Kelly, Anett S. Trebitz, Greg S. Peterson, and Corlis W. West. Effort and potential efficiencies for aquatic non-native species early detection. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 68(12): 2064–2079, December 2011. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f2011-117.

Ward:2011:SSE

[346] Darren M. Ward, Keith H. Nislow, and Carol L. Folt. Seasonal shift in the effects of predators on juvenile Atlantic salmon (Salmo salar) energetics. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 68(12):2080–2089, December 2011. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f2011-123.

Gunnarsson:2011:CPT

[347] Gudmundur Smári Gunnarsson and Stefán Óli Steingrímsson. Contrasting patterns of territoriality and foraging mode in two stream-dwelling salmonids, Arctic char (Salvelinus alpinus) and brown trout (Salmo trutta). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 68(12):2090–2100, December 2011. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f2011-127.

Ying:2011:RIF

[348] Yiping Ying, Yong Chen, Longshan Lin, and Tianxian Gao. Risks of ignoring fish population spatial structure in fisheries management. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 68(12):2101–2120, December 2011. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f2011-116.

Martin:2011:IES

[349] Shannon B. Martin and Paul L. Leberg. Influence of environmental stress on age- and size-at-maturity: genetic and plastic responses of coastal marsh fishes to changing salinities. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 68(12):2121–2131, December 2011. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f2011-119.

Syslo:2011:RNN

[350] John M. Syslo, Christopher S. Guy, Patricia E. Bigelow, Philip D. Doepke, Brian D. Ertel, and Todd M. Koel. Response of non-native lake trout (Salvelinus namaycush) to 15 years of harvest in Yellowstone Lake, Yellowstone National Park. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 68(12):2132–2145, December 2011. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f2011-122.

Syrjanen:2011:SSO

[351] J. Syrjänen, K. Korsu, P. Louhi, R. Paavola, and T. Muotka. Stream salmonids as opportunistic foragers: the importance of terrestrial invertebrates along a stream-size gradient. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 68(12):2146–2156, December 2011. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f2011-118.

Jackson:2011:DAB

[352] Christopher D. Jackson and Grant E. Brown. Differences in antipredator behaviour between wild and hatchery-reared juvenile Atlantic salmon (Salmo salar) under seminatural conditions. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 68(12):2157–2166, December 2011. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f2011-129.

Milbrink:2011:LSL

[353] Göran Milbrink, Tobias Vrede, Lars J. Tranvik, and Emil Rydin. Largescale and long-term decrease in fish growth following the construction of hydroelectric reservoirs. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 68(12): 2167–2173, December 2011. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f2011-131.

Vanderlaan:2011:FGT

[354] Angelia S. M. Vanderlaan, R. Kent Smedbol, and Christopher T. Taggart. Fishing-gear threat to right whales (Eubalaena glacialis) in Canadian waters and the risk of lethal entanglement. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 68(12):2174–2193, December 2011. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f2011-124.

Yun:2011:IPM

[355] Sang-Seon Yun, Andrew J. Wildbill, Michael J. Siefkes, Mary L. Moser, Andrew H. Dittman, Stephen C. Corbett, Weiming Li, and David A. Close. Identification of putative migratory pheromones from Pacific lamprey (Lampetra tridentata). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 68 (12):2194–2203, December 2011. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/

doi/full/10.1139/f2011-140.

Minns:2011:DIE

[356] Charles K. Minns, Robert G. Randall, Karen E. Smokorowski, Keith D. Clarke, Antonio Vélez-Espino, Robert S. Gregory, Simon Courtenay, and Patrice LeBlanc. Direct and indirect estimates of the productive capacity of fish habitat under Canada’s policy for the management of fish habitat: where have we been, where are we now, and where are we going? Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 68(12):2204–2227, December 2011. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f2011-130.

Francis:2011:CDW

[357] R. I. C. Chris Francis. Corrigendum: Data weighting in statistical fisheries stock assessment models. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 68(12):2228, December 2011. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f2011-165. See [271].

Larsson:2012:EML

[358] Stefan Larsson, Ignacio Serrano, and Lars-Ove Eriksson. Effects of muscle lipid concentration on wild and hatchery brown trout (Salmo trutta) smolt migration. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 69(1):1–12, January 2012. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f2011-128.

Vorkamp:2012:DOC

[359] Katrin Vorkamp, Tore Christian Svendsen, Bent Rønsholdt, and Martin Mørk Larsen. Different organochlorine contaminant profiles in groups of flounders (Platichthys flesus) from sampling locations around Denmark. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 69(1):13–23, January 2012. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f2011-141.

Hawryshyn:2012:LTW

[360] Jessica Hawryshyn, Kathleen M. Rühland, Roberto Quinlan, and John P. Smol. Long-term water quality changes in a multiple-stressor system: a diatom-based paleolimnological study of Lake Simcoe (Ontario, Canada). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 69(1):24–40, January 2012. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f2011-134.

Rennie:2012:EAA

[361] Michael D. Rennie, Timothy B. Johnson, and W. Gary Sprules. Energy acquisition and allocation patterns of lake whitefish (Coregonus clupeaformis) are modified when dreissenids are present. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 69(1):41–59, January 2012. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f2011-126.

Forsythe:2012:IBA

[362] Patrick S. Forsythe, James A. Crossman, Nora M. Bello, Edward A. Baker, and Kim T. Scribner. Individual-based analyses reveal high repeatability in timing and location of reproduction in lake sturgeon (Acipenser fulvescens). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 69 (1):60–72, January 2012. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f2011-132.

Rinella:2012:PSO

[363] Daniel J. Rinella, Mark S. Wipfli, Craig A. Stricker, Ron A. Heintz, and Matthew J. Rinella. Pacific salmon (Oncorhynchus spp.) runs and consumer fitness: growth and energy storage in stream-dwelling salmonids increase with salmon spawner density. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 69(1):73–84, January 2012. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f2011-133.

Faithfull:2012:TBP

[364] Carolyn Faithfull, Magnus Huss, Tobias Vrede, Jan Karlsson, and Ann-Kristin Bergström. Transfer of bacterial production based on labile carbon to higher trophic levels in an oligotrophic pelagic system. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 69(1):85–93, January 2012. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f2011-142.

Rattan:2012:NSP

[365] Kim J. Rattan, William D. Taylor, and Ralph E. H. Smith. Nutrient status of phytoplankton across a trophic gradient in Lake Erie: evidence from new fluorescence methods. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 69(1):94–111, January 2012. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f2011-135.

Tamdrari:2012:DPB

[366] Hacène Tamdrari, Martin Castonguay, Jean-Claude Brêthes, Peter S. Galbraith, and Daniel E. Duplisea. The dispersal pattern and behaviour of Atlantic cod (Gadus morhua) in the northern Gulf of St. Lawrence: results from tagging experiments. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 69(1):112–121, January 2012. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f2011-137.

Lepak:2012:MGR

[367] Jesse M. Lepak, Kristoph-Dietrich Kinzli, Eric R. Fetherman, William M. Pate, Adam G. Hansen, Eric I. Gardunio, C. Nathan Cathcart, William L. Stacy, Zachary E. Underwood, Mandi M. Brandt, Christopher A. Myrick, and Brett M. Johnson. Manipulation of growth to reduce mercury concentrations in sport fish on a whole-system scale. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 69(1):122–135, January 2012. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f2011-136.

Behringer:2012:PIF

[368] Donald C. Behringer, Mark J. Butler, Jessica Moss, and Jeffrey D. Shields. PaV1 infection in the Florida spiny lobster (Panulirus argus) fishery and its effects on trap function and disease transmission. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 69(1):136–144, January 2012. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f2011-146.

Baulch:2012:TMA

[369] Helen M. Baulch, Sherry L. Schiff, Roxane Maranger, and Peter J. Dillon. Testing models of aquatic N_{2}O flux for inland waters. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 69(1):145–160, January 2012. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f2011-144.

Bentley:2012:FST

[370] Nokome Bentley and Adam D. Langley. Feasible stock trajectories: a flexible and efficient sequential estimator for use in fisheries management procedures. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 69(1):161–177, January 2012. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f2011-143.

Nasmith:2012:LES

[371] Leslie E. Nasmith, William M. Tonn, Cynthia A. Paszkowski, and Garry J. Scrimgeour. Limited effects of stocked trout on littoral invertebrates in boreal foothills lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 69(1):178–190, January 2012. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f2011-145.

Johnson:2012:EPA

[372] Andrew F. Johnson, Maria Valls, Joan Moranta, Stuart R. Jenkins, Jan G. Hiddink, and Hilmar Hinz. Effect of prey abundance and size on the distribution of demersal fishes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 69(1):191–200, January 2012. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f2011-138.

Jones:2012:CEF

[373] Simon R. M. Jones and Richard J. Beamish. Comment on “Evidence of farm-induced parasite infestations on wild juvenile salmon in multiple regions of coastal British Columbia, Canada”. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 69(1):201–203, January 2012. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f2011-159. See [165].

Price:2012:SFS

[374] M. H. H. Price and J. D. Reynolds. Salmon farms as a source of sea lice on juvenile wild salmon; reply to the comment by Jones and Beamish. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 69(1):204–207, January 2012. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f2011-164.

Collie:2012:FRA

[375] Jeremy S. Collie, Randall M. Peterman, and Brett M. Zuehlke. A fisheries risk-assessment framework to evaluate trade-offs among management options in the presence of time-varying productivity. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 69(2):209–223, February 2012. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f2011-148.

Cline:2012:GRL

[376] Timothy J. Cline, Brian C. Weidel, James F. Kitchell, and James R. Hodgson. Growth response of largemouth bass (Micropterus salmoides) to catch-and-release angling: a 27-year mark–recapture study. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 69(2):224–230, February 2012. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f2011-150.

Shelton:2012:ESC

[377] Andrew O. Shelton, E. J. Dick, Donald E. Pearson, Stephen Ralston, and Marc Mangel. Estimating species composition and quantifying uncertainty in multispecies fisheries: hierarchical Bayesian models for stratified sampling protocols with missing data. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 69(2):231–246, February 2012. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f2011-152.

Latour:2012:IMG

[378] Robert J. Latour, David T. Gauthier, James Gartland, Christopher F. Bonzek, Kathleen A. McNamee, and Wolfgang K. Vogelbein. Impacts of mycobacteriosis on the growth of striped bass (Morone saxatilis) in Chesapeake Bay. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 69(2):247–258, February 2012. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f2011-158.

Chen:2012:PCA

[379] Kun Chen, Kung-Sik Chan, Kevin M. Bailey, Kerim Aydin, and Lorenzo Ciannelli. A probabilistic cellular automata approach for predator–prey interactions of arrowtooth flounder (Atheresthes stomias) and walleye pollock (Theragra chalcogramma) in the eastern Bering Sea. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 69(2):259–272, February 2012. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f2011-160.

Hanisch:2012:CLH

[380] J. R. Hanisch, W. M. Tonn, C. A. Paszkowski, and G. J. Scrimgeour. Complex littoral habitat influences the response of native minnows to stocked trout: evidence from whole-lake comparisons and experimental predator enclosures. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 69(2): 273–281, February 2012. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f2011-162.

Coman:2012:TVN

[381] Melissa Anne Coman and Mathew Graeme Wells. Temperature variability in the nearshore benthic boundary layer of Lake Opeongo is due to wind-driven upwelling events. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 69 (2):282–296, February 2012. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f2011-167.

Quinn:2012:PDA

[382] Thomas P. Quinn, Harry B. Rich, Dido Gosse, and Nicolas Schtickzelle. Population dynamics and asynchrony at fine spatial scales: a case history of sockeye salmon (Oncorhynchus nerka) population structure in Alaska, USA. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 69(2):297–306, February 2012. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f2011-147.

Pampoulie:2012:CAO

[383] Christophe Pampoulie, Anna Kristín Daníelsdóttir, Vilhjálmur Thorsteinsson, Einar Hjörleifsson, Gudrún Marteinsdóttir, and Daniel E. Ruzzante. The composition of adult overwintering and juvenile aggregations of Atlantic cod (Gadus morhua) around Iceland using neutral and functional markers: a statistical challenge. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 69(2):307–320, February 2012. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f2011-151.

Post:2012:TSP

[384] John. R. Post and Eric A. Parkinson. Temporal and spatial patterns of angler effort across lake districts and policy options to sustain recreational fisheries. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 69(2):321–329, February 2012. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f2011-163.

Martins:2012:HRT

[385] Eduardo G. Martins, Scott G. Hinch, David A. Patterson, Merran J. Hague, Steven J. Cooke, Kristina M. Miller, David Robichaud, Karl K. English, and Anthony P. Farrell. High river temperature reduces survival of sockeye salmon (Oncorhynchus nerka) approaching spawning grounds and exacerbates female mortality. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 69(2):330–342, February 2012. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f2011-154.

Gargan:2012:ESL

[386] P. G. Gargan, G. Forde, N. Hazon, D. J. F. Russell, and C. D. Todd. Evidence for sea lice-induced marine mortality of Atlantic salmon (Salmo salar) in western Ireland from experimental releases of ranched smolts treated with emamectin benzoate. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 69(2):343–353, February 2012. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f2011-155.

Walline:2012:STE

[387] Paul D. Walline, Christopher D. Wilson, Anne B. Hollowed, and Sarah C. Stienessen. Short-term effects of commercial fishing on the distribution and abundance of walleye pollock (Theragra chalcogramma). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 69(2):354–368, February 2012. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f2011-166.

Havens:2012:IPR

[388] Sonya M. Havens, Christel S. Hassler, Rebecca L. North, Stephanie J. Guildford, Greg Silsbe, Steven W. Wilhelm, and Michael R. Twiss. Iron plays a role in nitrate drawdown by phytoplankton in Lake Erie surface waters as observed in lake-wide assessments. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 69(2):369–381, February 2012. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f2011-157.

Rindorf:2012:ERB

[389] Anna Rindorf and Peter Lewy. Estimating the relationship between abundance and distribution. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 69 (2):382–397, February 2012. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f2011-153.

Montes:2012:MPD

[390] Rodrigo M. Montes, R. Ian Perry, Evgeny A. Pakhomov, Andrew M. Edwards, and James A. Boutillier. Multifractal patterns in the daily catch time series of smooth pink shrimp (Pandalus jordani) from the west coast of Vancouver Island, Canada. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 69 (2):398–413, February 2012. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f2011-156. See erratum [407].

Roy:2012:BDE

[391] Denis Roy, Thomas R. Hurlbut, and Daniel E. Ruzzante. Biocomplexity in a demersal exploited fish, white hake (Urophycis tenuis): depth-related structure and inadequacy of current management approaches. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 69(3):415–429, March 2012. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f2011-178.

Martino:2012:DDR

[392] E. J. Martino and E. D. Houde. Density-dependent regulation of yearclass strength in age-0 juvenile striped bass (Morone saxatilis). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 69(3):430–446, March 2012. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f2011-149.

Brothers:2012:LHI

[393] Soren M. Brothers, Paul A. del Giorgio, Cristian R. Teodoru, and Yves T. Prairie. Landscape heterogeneity influences carbon dioxide production in a young boreal reservoir. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 69 (3):447–456, March 2012. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f2011-174.

Steel:2012:LCC

[394] E. A. Steel, D. W. Jensen, K. M. Burnett, K. Christiansen, J. C. Firman, B. E. Feist, K. J. Anlauf, and D. P. Larsen. Landscape characteristics and Coho salmon (Oncorhynchus kisutch) distributions: explaining abundance versus occupancy. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 69 (3):457–468, March 2012. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f2011-161.

Kipp:2012:IER

[395] Rebekah Kipp and Anthony Ricciardi. Impacts of the Eurasian round goby (Neogobius melanostomus) on benthic communities in the upper St. Lawrence River. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 69 (3):469–486, March 2012. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f2011-139.

Sethi:2012:QRM

[396] Suresh Andrew Sethi, Michael Dalton, and Ray Hilborn. Quantitative risk measures applied to Alaskan commercial fisheries. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 69(3):487–498, March 2012. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f2011-170.

Sumaila:2012:IDH

[397] U. Rashid Sumaila, Andrés M. Cisneros-Montemayor, Andrew Dyck, Ling Huang, William Cheung, Jennifer Jacquet, Kristin Kleisner, Vicky Lam, Ashley McCrea-Strub, Wilf Swartz, Reg Watson, Dirk Zeller, and Daniel Pauly. Impact of the Deepwater Horizon well blowout on the economics of US Gulf fisheries. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 69 (3):499–510, March 2012. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f2011-171.

Vermaire:2012:DID

[398] Jesse C. Vermaire, Yves T. Prairie, and Irene Gregory-Eaves. Diatominferred decline of macrophyte abundance in lakes of southern Quebec, Canada. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 69(3):511–524, March 2012. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f2011-169.

Wetten:2012:HRM

[399] Ola Frang Wetten, Robert C. Wilson, and Øivind Andersen. Highresolution melting analysis of common and recombinant genotypes of the Atlantic cod (Gadus morhua) hemoglobin β 1 gene in transatlantic populations. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 69(3):525–531, March 2012. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f2011-176.

Ernst:2012:LHS

[400] Billy Ernst, David A. Armstrong, Julián Burgos, and J. M. (Lobo) Orensanz. Life history schedule and periodic recruitment of female snow crab (Chionoecetes opilio) in the eastern Bering Sea. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 69(3):532–550, March 2012. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f2011-173.

Malpica-Cruz:2012:TSI

[401] Luis Malpica-Cruz, Sharon Z. Herzka, Oscar Sosa-Nishizaki, and Juan Pablo Lazo. Tissue-specific isotope trophic discrimination factors and turnover rates in a marine elasmobranch: empirical and modeling results. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 69(3):551–564, March 2012. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f2011-172.

Simpfendorfer:2012:TDK

[402] Colin A. Simpfendorfer, Esben M. Olsen, Michelle R. Heupel, and Even Moland. Three-dimensional kernel utilization distributions improve estimates of space use in aquatic animals. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 69(3):565–572, March 2012. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f2011-179.

Brush:2012:SSV

[403] Jaclyn M. Brush, Aaron T. Fisk, Nigel E. Hussey, and Timothy B. Johnson. Spatial and seasonal variability in the diet of round goby (Neogobius melanostomus): stable isotopes indicate that stomach contents overestimate the importance of dreissenids. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 69(3):573–586, March 2012. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f2012-001.

Franssen:2012:AEM

[404] Jan Franssen, Catherine Blais, Michel Lapointe, Francis Bérubé, Normand Bergeron, and Pierre Magnan. Asphyxiation and entombment mechanisms in fines rich spawning substrates: experimental evidence with brook trout (Salvelinus fontinalis) embryos. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 69(3):587–599, March 2012. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f2011-168.

McGarvey:2012:LCE

[405] Richard McGarvey and Janet M. Matthews. Low-cost estimates of mortality rate from single tag recoveries: addressing short-term trap-happy and trap-shy bias. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 69(3):600–611, March 2012. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f2011-175.

Lambert:2012:EEF

[406] Yvan Lambert. Erratum: Environmental and fishing limitations to the rebuilding of the northern Gulf of St. Lawrence cod stock (Gadus morhua). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 69(3):612, March 2012. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/

f2012-014. See [231].

Montes:2012:EMP

[407] Rodrigo M. Montes, R. Ian Perry, Evgeny A. Pakhomov, Andrew M. Edwards, and James A. Boutillier. Erratum: Multifractal patterns in the daily catch time series of smooth pink shrimp (Pandalus jordani) from the west coast of Vancouver Island, Canada. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 69(3):613, March 2012. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f2012-015. See [390].

Peacock:2012:EPE

[408] Mary M. Peacock and Ned A. Dochtermann. Evolutionary potential but not extinction risk of Lahontan cutthroat trout (Oncorhynchus clarkii henshawi) is associated with stream characteristics. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 69(4):615–626, April 2012. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f2012-006.

Roy:2012:CST

[409] Suzanne Roy, Marie Parenteau, Oscar Casas-Monroy, and André Rochon. Coastal ship traffic: a significant introduction vector for potentially harmful dinoflagellates in eastern Canada. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 69(4):627–644, April 2012. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f2012-008.

Thorson:2012:UMB

[410] James T. Thorson, Trevor A. Branch, and Olaf P. Jensen. Using modelbased inference to evaluate global fisheries status from landings, location, and life history data. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 69 (4):645–655, April 2012. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f2012-016.

Greenaway:2012:DDS

[411] Christine M. Greenaway, Andrew M. Paterson, (Bill) W. Keller, and John P. Smol. Dramatic diatom species assemblage responses in lakes recovering from acidification and metal contamination near wawa, Ontario, Canada: a paleolimnological perspective. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 69(4):656–669, April 2012. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f2012-003.

Weir:2012:IDM

[412] Laura K. Weir, Jeffrey A. Hutchings, and Daniel D. Heath. Influence of density and major histocompatibility genotype on sexual selection in a salmonid alternative mating strategy. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 69(4):670–680, April 2012. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f2012-007.

Peacock:2012:MSD

[413] Stephanie J. Peacock and Carrie A. Holt. Metrics and sampling designs for detecting trends in the distribution of spawning Pacific salmon (Oncorhynchus spp.). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 69 (4):681–694, April 2012. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f2012-004.

Rochet:2012:CMS

[414] Marie-Joëlle Rochet, Fabienne Daurès, and Verena M. Trenkel. Capacity management, not stock status or economics, drives fleet dynamics in the Bay of Biscay ecosystem on a decadal time scale. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 69(4):695–710, April 2012. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f2012-002.

Jensen:2012:TSM

[415] Arne J. Jensen, Bengt Finstad, Peder Fiske, Nils Arne Hvidsten, Audun H. Rikardsen, and Laila Saksgård. Timing of smolt migration in sympatric populations of Atlantic salmon (Salmo salar), brown trout (Salmo trutta), and Arctic char (Salvelinus alpinus). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 69(4):711–723, April 2012. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f2012-005.

Wolff:2012:OIP

[416] Brian A. Wolff, Brett M. Johnson, André R. Breton, Patrick J. Martinez, and Dana L. Winkelman. Origins of invasive piscivores determined from the strontium isotope ratio (^{87}Sr/^{86}Sr) of otoliths. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 69(4):724–739, April 2012. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f2012-009.

Grados:2012:MCS

[417] Daniel Grados, Ronan Fablet, Michael Ballón, Nicolas Bez, Ramiro Castillo, Ainhoa Lezama-Ochoa, and Arnaud Bertrand. Multiscale characterization of spatial relationships among oxycline depth, macrozooplankton, and forage fish off Peru using geostatistics, principal coordinates of neighbour matrices (PCNMs), and wavelets. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 69(4):740–754, April 2012. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f2012-017.

Ferguson:2012:MCB

[418] Jake M. Ferguson, Mark L. Taper, Christopher S. Guy, and John M. Syslo. Mechanisms of coexistence between native bull trout (Salvelinus confluentus) and non-native lake trout (Salvelinus namaycush): inferences from pattern-oriented modeling. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 69(4):755–769, April 2012. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f2011-177.

Drouineau:2012:SSS

[419] Hilaire Drouineau, Louise Savard, Mathieu Desgagnés, and Daniel Duplisea. SPAM (sex-structured Pandalus assessment model): a stock assessment model for Pandalus stocks. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 69(4):770–783, April 2012. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f2012-011.

Bailey:2012:BSE

[420] Robert C. Bailey, Garry Scrimgeour, David Coté, Dan Kehler, Simon Linke, and Yong Cao. Bioassessment of stream ecosystems enduring a decade of simulated degradation: lessons for the real world. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 69(4):784–796, April 2012. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f2012-010.

Braccini:2012:FPS

[421] J. M. Braccini, M. F. O’Neill, A. B. Campbell, G. M. Leigh, and A. J. Courtney. Fishing power and standardized catch rates: implications of missing vessel-characteristic data from the Australian eastern king prawn (Melicertus plebejus) fishery. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 69 (5):797–809, May 2012. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f2012-023.

Kim:2012:WEC

[422] Natalie Kim, Bjørn Walseng, and Norman David Yan. Will environmental calcium declines hinder Bythotrephes establishment success in Canadian Shield lakes? Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 69 (5):810–820, May 2012. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f2012-024.

Parra:2012:UEW

[423] Irene Parra, Ana Almodóvar, Daniel Ayllón, Graciela G. Nicola, and Benigno Elvira. Unravelling the effects of water temperature and density dependence on the spatial variation of brown trout (Salmo trutta) body size. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 69(5):821–832, May 2012. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f2012-025.

Bartolino:2012:HSD

[424] Valerio Bartolino, Massimiliano Cardinale, Henrik Svedäng, Hans W. Linderholm, Michele Casini, and Anders Grimwall. Historical spatiotemporal dynamics of eastern North Sea cod. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 69(5):833–841, May 2012. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f2012-028.

Muths:2012:GCR

[425] D. Muths, G. Gouws, M. Mwale, E. Tessier, and J. Bourjea. Genetic connectivity of the reef fish Lutjanus kasmira at the scale of the western Indian Ocean. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 69(5):842–853, May 2012. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f2012-012.

Stanton:2012:RCM

[426] Timothy K. Stanton, Cynthia J. Sellers, and J. Michael Jech. Resonance classification of mixed assemblages of fish with swimbladders using a modified commercial broadband acoustic echosounder at 1–6 kHz. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 69(5):854–868, May 2012. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f2012-013.

Eveson:2012:UET

[427] J. Paige Eveson, Marinelle Basson, and Alistair J. Hobday. Using electronic tag data to improve mortality and movement estimates in a tagbased spatial fisheries assessment model. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 69(5):869–883, May 2012. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f2012-026.

Marie:2012:EFC

[428] Amandine D. Marie, Louis Bernatchez, and Dany Garant. Environmental factors correlate with hybridization in stocked brook charr (Salvelinus fontinalis). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 69(5):884–893, May 2012. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f2012-027.

Freeman:2012:UMR

[429] Debbie J. Freeman, Paul A. Breen, and Alison B. MacDiarmid. Use of a marine reserve to determine the direct and indirect effects of fishing on growth in a New Zealand fishery for the spiny lobster Jasus edwardsii. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 69(5):894–905, May 2012. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f2012-032.

Muhlfeld:2012:EWC

[430] Clint C. Muhlfeld, Simon R. Thorrold, Thomas E. McMahon, and Brian Marotz. Estimating westslope cutthroat trout (Oncorhynchus clarkii lewisi) movements in a river network using strontium isoscapes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 69(5):906–915, May 2012. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f2012-033. See er-

ratum [448].

Gomez-Uchida:2012:AFS

[431] Daniel Gomez-Uchida, James E. Seeb, Christopher Habicht, and Lisa W. Seeb. Allele frequency stability in large, wild exploited populations over multiple generations: insights from Alaska sockeye salmon (Oncorhynchus nerka). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 69 (5):916–929, May 2012. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f2012-029.

Phelps:2012:IRO

[432] Quinton E. Phelps, Gregory W. Whitledge, Sara J. Tripp, Kurt T. Smith, James E. Garvey, David P. Herzog, David E. Ostendorf, Joseph W. Ridings, Jason W. Crites, Robert A. Hrabik, Wyatt J. Doyle, and Tracy D. Hill. Identifying river of origin for age-0 Scaphirhynchus sturgeons in the Missouri and Mississippi rivers using fin ray microchemistry. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 69(5):930–941, May 2012. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f2012-038.

Satterthwaite:2012:BHM

[433] William Hallowell Satterthwaite, Michael Steven Mohr, Michael Robert O’Farrell, and Brian Kenneth Wells. A Bayesian hierarchical model of size-at-age in ocean-harvested stocks — quantifying effects of climate and temporal variability. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 69 (5):942–954, May 2012. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f2012-036.

Miller:2012:RWC

[434] Sara E. Miller, Milo Adkison, and Lewis Haldorson. Relationship of water column stability to the growth, condition, and survival of pink salmon (Oncorhynchus gorbuscha) in the northern coastal Gulf of Alaska and Prince William Sound. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 69 (5):955–969, May 2012. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f2012-031.

Pess:2012:IPD

[435] G. R. Pess, R. Hilborn, K. Kloehn, and T. P. Quinn. The influence of population dynamics and environmental conditions on pink salmon (Oncorhynchus gorbuscha) recolonization after barrier removal in the Fraser River, British Columbia, Canada. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 69(5):970–982, May 2012. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f2012-030.

Breen:2012:PIB

[436] Erin Breen and Anna Metaxas. Predicting the interactions between “ecologically equivalent” indigenous and nonindigenous brachyurans. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 69(5):983–995, May 2012. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f2012-037.

Swain:2012:FFF

[437] Douglas P. Swain and Robert K. Mohn. Forage fish and the factors governing recovery of Atlantic cod (Gadus morhua) on the eastern Scotian Shelf. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 69(6):997–1001, June 2012. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f2012-045.

Millar:2012:SES

[438] Russell B. Millar, Sam McKechnie, and Chris E. Jordan. Simple estimators of salmonid escapement and its variance using a new areaunder-the-curve method. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 69 (6):1002–1015, June 2012. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f2012-034.

Rupp:2012:EFS

[439] David E. Rupp, Thomas C. Wainwright, and Peter W. Lawson. Effect of forecast skill on management of the Oregon coast Coho salmon (Oncorhynchus kisutch) fishery. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 69 (6):1016–1032, June 2012. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f2012-040.

Budge:2012:EDA

[440] Suzanne M. Budge, Sarah N. Penney, and Santosh P. Lall. Estimating diets of Atlantic salmon (Salmo salar) using fatty acid signature analyses; validation with controlled feeding studies. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 69(6):1033–1046, June 2012. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f2012-039.

Ostman:2012:DCC

[441] Örjan Östman, Mikaela Bergenius, Maria Kristina Boström, and Sven-Gunnar Lunneryd. Do cormorant colonies affect local fish communities in the Baltic Sea? Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 69(6): 1047–1055, June 2012. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f2012-042.

Richter:2012:PTC

[442] Catherine A. Richter, Allison N. Evans, Maureen K. Wright-Osment, James L. Zajicek, Scott A. Heppell, Stephen C. Riley, Charles C. Krueger, and Donald E. Tillitt. Paenibacillus thiaminolyticus is not the cause of thiamine deficiency impeding lake trout (Salvelinus namaycush) recruitment in the Great Lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 69(6):1056–1064, June 2012. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f2012-043.

Houle:2012:ASS

[443] Jennifer E. Houle, Keith D. Farnsworth, Axel G. Rossberg, and David G. Reid. Assessing the sensitivity and specificity of fish community indicators to management action. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 69 (6):1065–1079, June 2012. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f2012-044.

Bohrmann:2012:ESE

[444] Thomas F. Bohrmann and Mary C. Christman. Evaluating sampling efficiency in depletion surveys using hierarchical Bayes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 69(6):1080–1090, June 2012. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f2012-035.

Frank:2012:ABT

[445] Béatrice M. Frank, Olivier Gimenez, and Philippe V. Baret. Assessing brown trout (Salmo trutta) spawning movements with multistate capture–recapture models: a case study in a fully controlled Belgian brook. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 69(6):1091–1104, June 2012. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f2012-041.

Devine:2012:CRC

[446] Jennifer A. Devine, Peter J. Wright, Heidi E. Pardoe, and Mikko Heino. Comparing rates of contemporary evolution in life-history traits for exploited fish stocks. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 69(6):1105– 1120, June 2012. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f2012-047.

Wood:2012:PDT

[447] J. L. A. Wood, J. W. A. Grant, and M. H. Belanger. Population density and territory size in juvenile rainbow trout, Oncorhynchus mykiss: implications for population regulation. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 69(6):1121–1128, June 2012. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f2012-048.

Muhlfeld:2012:EEW

[448] Clint C. Muhlfeld, Simon R. Thorrold, Thomas E. McMahon, and Brian Marotz. Erratum: Estimating westslope cutthroat trout (Oncorhynchus clarkii lewisi) movements in a river network using strontium isoscapes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 69(6):1129– 1130, June 2012. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/

f2012-049. See [430].

Carmichael:2012:BEN

[449] Ruth H. Carmichael, William Walton, and Heidi Clark. Bivalve-enhanced nitrogen removal from coastal estuaries. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 69(7):1131–1149, July 2012. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f2012-057. See corrigendum [658].

Keith:2012:PDM

[450] David M. Keith and Jeffrey A. Hutchings. Population dynamics of marine fishes at low abundance. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 69 (7):1150–1163, July 2012. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/

10.1139/f2012-055. See corrigendum [500].

Kreps:2012:IIP

[451] Timothy A. Kreps, Ashley K. Baldridge, and David M. Lodge. The impact of an invasive predator (Orconectes rusticus) on freshwater snail communities: insights on habitat-specific effects from a multilake longterm study. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 69(7):1164– 1173, July 2012. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f2012-052.

Dennis:2012:DDA

[452] Ian F. Dennis and Thomas A. Clair. The distribution of dissolved aluminum in Atlantic salmon (Salmo salar) rivers of Atlantic Canada and its potential effect on aquatic populations. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 69(7):1174–1183, July 2012. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f2012-053.

Ahlbeck:2012:EFD

[453] Ida Ahlbeck, Sture Hansson, and Olle Hjerne. Evaluating fish diet analysis methods by individual-based modelling. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 69(7):1184–1201, July 2012. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f2012-051.

Brinkmann:2012:EML

[454] Lars Brinkmann and Joseph B. Rasmussen. Elevated mercury levels in biota along an agricultural land use gradient in the Oldman River basin, Alberta. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 69(7):1202–1213, July 2012. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f2012-056.

Gunther:2012:NLB

[455] Claudia C. Günther, Axel Temming, Hannes Baumann, Bastian Huwer, Christian Möllmann, Catriona Clemmesen, and Jens-Peter Herrmann.

A novel length back-calculation approach accounting for ontogenetic changes in the fish length–otolith size relationship during the early life of sprat (Sprattus sprattus). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 69 (7):1214–1229, July 2012. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f2012-054.

Dolbeth:2012:SPT

[456] M. Dolbeth, M. Cusson, R. Sousa, and M. A. Pardal. Secondary production as a tool for better understanding of aquatic ecosystems. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 69(7):1230–1253, July 2012. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f2012-050.

Anonymous:2012:CDM

[457] Anonymous. Carey D. McAllister (3 May 1931–17 December 2011). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 69(8), August 2012. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f2012-067.

Peterman:2012:WDP

[458] Randall M. Peterman and Brigitte Dorner. A widespread decrease in productivity of sockeye salmon (Oncorhynchus nerka) populations in western North America. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 69(8): 1255–1260, August 2012. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f2012-063.

Rypel:2012:CEC

[459] Andrew L. Rypel. Concordant estimates of countergradient growth variation in striped bass (Morone saxatilis) using comparative life-history data. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 69(8):1261–1265, August 2012. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f2012-069.

Hoover:2012:EIT

[460] Reneé R. Hoover, Cynthia M. Jones, and Chester E. Grosch. Estuarine ingress timing as revealed by spectral analysis of otolith life history scans.

Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 69(8):1266–1277, August 2012. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f2012-058.

Lokkeborg:2012:SSA

[461] Svein Løkkeborg, Egil Ona, Aud Vold, and Are Salthaug. Sounds from seismic air guns: gear- and species-specific effects on catch rates and fish distribution. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 69(8): 1278–1291, August 2012. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f2012-059.

Zhou:2012:LFM

[462] Shijie Zhou, Shaowu Yin, James T. Thorson, Anthony D. M. Smith, and Michael Fuller. Linking fishing mortality reference points to life history traits: an empirical study. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 69 (8):1292–1301, August 2012. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f2012-060.

Schludermann:2012:DRL

[463] Elisabeth Schludermann, Michael Tritthart, Paul Humphries, and Hubert Keckeis. Dispersal and retention of larval fish in a potential nursery habitat of a large temperate river: an experimental study. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 69(8):1302–1315, August 2012. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f2012-061.

Usvyatsov:2012:MTS

[464] Sima Usvyatsov, Jeffrey Picka, Ryan Scott Hardy, Travis Dawson Shepherd, James Watmough, and Matthew Kenneth Litvak. Modeling the timing of spawning and hatching of shortnose sturgeon, Acipenser brevirostrum, in the Saint John River, New Brunswick, Canada. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 69(8):1316–1328, August 2012. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f2012-062.

Fablet:2012:BFF

[465] Ronan Fablet, Paul Gay, Salvador Peraltilla, Cecilia Peña, Ramiro Castillo, and Arnaud Bertrand. Bags-of-features for fish school cluster characterization in pelagic ecosystems: application to the discrimination of juvenile and adult anchovy (Engraulis ringens) clusters off Peru. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 69(8):1329–1339, August 2012. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f2012-064.

Murphy:2012:IVL

[466] Hannah M. Murphy, Gregory P. Jenkins, Paul A. Hamer, and Stephen E. Swearer. Interannual variation in larval survival of snapper (Chrysophrys auratus, Sparidae) is linked to diet breadth and prey availability. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 69(8):1340–1351, August 2012. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f2012-066.

deMestral:2012:MSE

[467] Louise G. de Mestral, Christophe M. Herbinger, and Patrick T. O’Reilly. Mating structure of an endangered population of wild Atlantic salmon (Salmo salar) as determined using sibship reconstruction and a novel method of sex inference. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 69 (8):1352–1361, August 2012. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f2012-065.

Freiwald:2012:MAT

[468] Jan Freiwald. Movement of adult temperate reef fishes off the west coast of North America. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 69(8): 1362–1374, August 2012. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f2012-068.

Anonymous:2012:NHI

[469] Anonymous. A new hydrology: Inflow effects on ecosystem form and functioning. A special session derived from the February 2011 Association for the Sciences of Limnology and Oceanography (ASLO) Aquatic Sciences conference in San Juan, Puerto Rico / Nouvelle hydrologie : effets des apports d’eau sur la forme et le fonctionnement des écosystèmes.

Séance de travail spéciale issue du congrès de l’Association for the Sciences of Limnology and Oceanography (ASLO) qui s’est tenu en février 2011 à San Juan (Puerto Rico). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 69(8), August 2012. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f2012-075.

Roelke:2012:NHE

[470] Daniel L. Roelke, Sofie Spatharis, and Simon M. Mitrovic. A new hydrology: effects on ecosystem form and functioning. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 69(8):1377–1379, August 2012. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f2012-085.

Santos:2012:UHS

[471] Rolando O. Santos and Diego Lirman. Using habitat suitability models to predict changes in seagrass distribution caused by water management practices. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 69(8):1380–1388, August 2012. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f2012-018.

Roelke:2012:AHP

[472] Daniel L. Roelke, Bryan W. Brooks, James P. Grover, George M. Gable, Leslie Schwierzke-Wade, and Natalie C. Hewitt. Anticipated human population and climate change effects on algal blooms of a toxic haptophyte in the south-central USA. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 69 (8):1389–1404, August 2012. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f2012-019.

Stanish:2012:ESF

[473] Lee F. Stanish, Tyler J. Kohler, Rhea M. M. Esposito, Breana L. Simmons, Uffe N. Nielsen, Diana H. Wall, Diana R. Nemergut, and Diane M. McKnight. Extreme streams: flow intermittency as a control on diatom communities in meltwater streams in the McMurdo Dry Valleys, Antarctica. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 69(8):1405–1419, August 2012. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f2012-022.

Colon-Rivera:2012:SMC

[474] Ricardo J. Colón-Rivera, Rusty A. Feagin, Jason B. West, and Kevin M. Yeager. Salt marsh connectivity and freshwater versus saltwater inflow: multiple methods including tidal gauges, water isotopes, and LI-DAR elevation models. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 69(8): 1420–1432, August 2012. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f2012-046.

Waters:2012:ACR

[475] Matthew N. Waters, Michael F. Piehler, Joseph M. Smoak, and Thomas S. Bianchi. Algal community responses to shallow lake dystrophication. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 69(8):1433–1443, August 2012. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f2012-021.

Spatharis:2012:FRT

[476] Sofie Spatharis, Nikolaos Skliris, Alexandra Meziti, and Konstantinos Ar. Kormas. First record of a Trichodesmium erythraeum bloom in the Mediterranean Sea. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 69(8): 1444–1455, August 2012. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f2012-020.

Orihel:2012:HMC

[477] Diane M. Orihel, David F. Bird, Michael Brylinsky, Huirong Chen, Derek B. Donald, Dorothy Y. Huang, Alessandra Giani, David Kinniburgh, Hedy Kling, Brian G. Kotak, Peter R. Leavitt, Charlene C. Nielsen, Sharon Reedyk, Rebecca C. Rooney, Sue B. Watson, Ron W. Zurawell, and Rolf D. Vinebrooke. High microcystin concentrations occur only at low nitrogen-to-phosphorus ratios in nutrient-rich Canadian lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 69(9):1457–1462, September 2012. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/

f2012-088. See comment [657].

Dagorn:2012:TBS

[478] Laurent Dagorn, John D. Filmalter, Fabien Forget, Monin Justin Amandè, Martin A. Hall, Peter Williams, Hilario Murua, Javier Ariz, Pierre Chavance, and Nicolas Bez. Targeting bigger schools can reduce ecosystem impacts of fisheries. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 69(9):1463–1467, September 2012. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f2012-089.

Horbowy:2012:ERA

[479] Jan Horbowy and Anna Luzeńczyk. The estimation and robustness of F_{MSY} and alternative fishing mortality reference points associated with high long-term yield. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 69(9): 1468–1480, September 2012. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f2012-070.

Naslund:2012:EEE

[480] Joacim Näslund, Kim Aarestrup, Søren T. Thomassen, and Jörgen I. Johnsson. Early enrichment effects on brain development in hatcheryreared Atlantic salmon (Salmo salar): no evidence for a critical period. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 69(9):1481–1490, September 2012. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f2012-074.

Cyr:2012:ADG

[481] Frédéric Cyr, Pedro Peres-Neto, and Bernard Angers. Allopatric distribution of generalist parasites: interplay between postglacial dispersal and host species. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 69(9):1491– 1498, September 2012. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f2012-076.

McKinnell:2012:EMG

[482] Skip McKinnell and Maxine Reichardt. Early marine growth of juvenile Fraser River sockeye salmon (Oncorhynchus nerka) in relation to juvenile pink (Oncorhynchus gorbuscha) and sockeye salmon abundance. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 69(9):1499–1512, September 2012. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f2012-078.

Vollestad:2012:SSD

[483] Leif Asbjørn Vøllestad, Dimitar Serbezov, Arthur Bass, Louis Bernatchez, Esben Moland Olsen, and Annette Taugbøl. Small-scale dispersal and population structure in stream-living brown trout (Salmo trutta) inferred by mark–recapture, pedigree reconstruction, and population genetics. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 69(9):1513–1524, September 2012. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f2012-073.

Lloyd-Jones:2012:LSE

[484] Luke R. Lloyd-Jones, You-Gan Wang, Anthony J. Courtney, Andrew J. Prosser, and Steven S. Montgomery. Latitudinal and seasonal effects on growth of the Australian eastern king prawn (Melicertus plebejus). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 69(9):1525–1538, September 2012. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f2012-072.

Mogensen:2012:MFC

[485] Stephanie Mogensen and Jeffrey A. Hutchings. Maternal fitness consequences of interactions among agents of mortality in early life of salmonids. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 69(9):1539–1555, September 2012. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f2012-071.

Thorson:2012:SBR

[486] James T. Thorson, Jason M. Cope, Trevor A. Branch, and Olaf P. Jensen. Spawning biomass reference points for exploited marine fishes, incorporating taxonomic and body size information. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 69(9):1556–1568, September 2012. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f2012-077.

Lepak:2012:OMP

[487] Jesse M. Lepak, C. Nathan Cathcart, and Mevin B. Hooten. Otolith mass as a predictor of age in Kokanee salmon (Oncorhynchus nerka) from four Colorado reservoirs. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 69(10): 1569–1575, October 2012. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f2012-081.

Stoll:2012:LFT

[488] Stefan Stoll and Peter Beeck. Larval fish in troubled waters — is the behavioural response of larval fish to hydrodynamic impacts active or passive? Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 69(10):1576–1584, October 2012. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f2012-086.

Levesque:2012:PRP

[489] David Lévesque, Antonella Cattaneo, Christiane Hudon, and Pierre Gagnon. Predicting the risk of proliferation of the benthic cyanobacterium Lyngbya wollei in the St. Lawrence River. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 69(10):1585–1595, October 2012. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f2012-087.

Serfas:2012:CGP

[490] Carla A. V. Serfas, Anna Varian, Rachel Holman, Lindsey M. Watch, Jesse Karner, and Jill B. K. Leonard. Comparison of growth physiology, morphology, and smolt indicators in juvenile Lake Superior brook trout (Salvelinus fontinalis) strains in reference to life history variation. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 69(10):1596–1609, October 2012. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f2012-083.

Gillis:2012:AAI

[491] D. M. Gillis and A. van der Lee. Advancing the application of the ideal free distribution to spatial models of fishing effort: the isodar approach. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 69(10):1610–1620, October 2012. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f2012-091.

Kuehne:2012:CLJ

[492] Lauren M. Kuehne, Julian D. Olden, and Jeffrey J. Duda. Costs of living for juvenile Chinook salmon (Oncorhynchus tshawytscha) in an increasingly warming and invaded world. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 69(10):1621–1630, October 2012. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f2012-094.

Shelton:2012:MAF

[493] Andrew O. Shelton, Stephan B. Munch, David Keith, and Marc Mangel. Maternal age, fecundity, egg quality, and recruitment: linking stock structure to recruitment using an age-structured Ricker model. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 69(10):1631–1641, October 2012. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f2012-082.

Frost:2012:HSD

[494] Paul C. Frost and Andrea L. Hicks. Human shoreline development and the nutrient stoichiometry of aquatic plant communities in Canadian Shield lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 69(10):1642– 1650, October 2012. CODEN CJFSDX. ISSN 0706-652X (print), 1205-

7533 (electronic). URL https://cdnsciencepub.com/doi/full/10.
1139/f2012-080.

Guenard:2012:BDD

[495] Guillaume Guénard, Daniel Boisclair, Ola Ugedal, Torbjørn Forseth, Ian A. Fleming, and Bror Jonsson. The bioenergetics of densitydependent growth in Arctic char (Salvelinus alpinus). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 69(10):1651–1662, October 2012. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f2012-093.

Garcia:2012:LQI

[496] Liliana García, John S. Richardson, and Isabel Pardo. Leaf quality influences invertebrate colonization and drift in a temperate rainforest stream. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 69(10):1663–1673, October 2012. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f2012-090.

Whitlock:2012:ISS

[497] Rebecca E. Whitlock and Murdoch K. McAllister. Incorporating spatial and seasonal dimensions in a stock reduction analysis for lower Fraser River white sturgeon (Acipenser transmontanus). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 69(10):1674–1697, October 2012. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f2012-079.

Ruggeri:2012:TGV

[498] Paolo Ruggeri, Andrea Splendiani, Sara Bonanomi, Enrico Arneri, Nando Cingolani, Alberto Santojanni, Andrea Belardinelli, Massimo Giovannotti, and Vincenzo Caputo. Temporal genetic variation as revealed by a microsatellite analysis of European sardine (Sardina pilchardus) archived samples. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 69(10):1698– 1709, October 2012. CODEN CJFSDX. ISSN 0706-652X (print), 1205-

7533 (electronic). URL https://cdnsciencepub.com/doi/full/10.
1139/f2012-092.

Colvin:2012:SBD

[499] Michael E. Colvin, Clay L. Pierce, and Timothy W. Stewart. Semidiscrete biomass dynamic modeling: an improved approach for assessing fish stock responses to pulsed harvest events. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 69(10):1710–1721, October 2012. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f2012-084.

Keith:2012:CPD

[500] David M. Keith and Jeffrey A. Hutchings. Corrigendum: Population dynamics of marine fishes at low abundance. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 69(10):1722, October 2012. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f2012-100. See [450].

Bothwell:2012:IRD

[501] Max L. Bothwell, Cathy Kilroy, Brad W. Taylor, Eric T. Ellison, Daniel A. James, Carole-Anne Gillis, Kevin D. Bladon, and Uldis Silins. Iron is not responsible for Didymosphenia geminata bloom formation in phosphorus-poor rivers. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 69(11):

1723–1727, November 2012. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f2012-112.

Kilada:2012:DDA

[502] Raouf Kilada, Bernard Sainte-Marie, Rémy Rochette, Neill Davis, Caroline Vanier, and Steven Campana. Direct determination of age in shrimps, crabs, and lobsters. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 69 (11):1728–1733, November 2012. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/

doi/full/10.1139/cjfas-2012-0254.

Bacheler:2012:DDH

[503] Nathan M. Bacheler, Jeffrey A. Buckel, and Lee M. Paramore. Densitydependent habitat use and growth of an estuarine fish. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 69(11):1734–1747, November 2012. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f2012-098.

Kates:2012:RNI

[504] Dan Kates, Clark Dennis, Matthew R. Noatch, and Cory D. Suski. Responses of native and invasive fishes to carbon dioxide: potential for a nonphysical barrier to fish dispersal. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 69(11):1748–1759, November 2012. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f2012-102.

Gudmundsson:2012:SES

[505] Gudmundur Gudmundsson and Thorvaldur Gunnlaugsson. Selection and estimation of sequential catch-at-age models. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 69(11):1760–1772, November 2012. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f2012-095.

Ward:2012:ATS

[506] Eric J. Ward, George R. Pess, Kara Anlauf-Dunn, and Chris E. Jordan. Applying time series models with spatial correlation to identify the scale of variation in habitat metrics related to threatened Coho salmon (Oncorhynchus kisutch) in the Pacific northwest. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 69(11):1773–1782, November 2012. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f2012-096.

Hendrix:2012:BEH

[507] A. N. Hendrix, J. Straley, C. M. Gabriele, and S. M. Gende. Bayesian estimation of humpback whale (Megaptera novaeangliae) population abundance and movement patterns in southeastern Alaska. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 69(11):1783–1797, November 2012. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f2012-101.

Aiken:2012:LTA

[508] James K. Aiken, C. Scott Findlay, and François Chapleau. Long-term assessment of the effect of introduced predatory fish on minnow diversity in a regional protected area. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 69(11):1798–1805, November 2012. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f2012-105.

Lucchetti:2012:IPM

[509] Alessandro Lucchetti and Antonello Sala. Impact and performance of Mediterranean fishing gear by side-scan sonar technology. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 69(11):1806–1816, November 2012. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f2012-107.

Jonsson:2012:WTF

[510] Bror Jonsson, Anders G. Finstad, and Nina Jonsson. Winter temperature and food quality affect age at maturity: an experimental test with Atlantic salmon (Salmo salar). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 69(11):1817–1826, November 2012. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f2012-108.

Korman:2012:ERD

[511] Josh Korman, Steven J. D. Martell, Carl J. Walters, Andrew S. Makinster, Lewis G. Coggins, Michael D. Yard, and William R. Persons. Estimating recruitment dynamics and movement of rainbow trout (Oncorhynchus mykiss) in the Colorado River in Grand Canyon using an integrated assessment model. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 69 (11):1827–1849, November 2012. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/

doi/full/10.1139/f2012-097.

Andrews:2012:LLL

[512] Allen H. Andrews, Edward E. DeMartini, Jon Brodziak, Ryan S. Nichols, and Robert L. Humphreys. A long-lived life history for a tropical, deepwater snapper (Pristipomoides filamentosus): bomb radiocarbon and lead–radium dating as extensions of daily increment analyses in otoliths. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 69(11):1850–1869, November 2012. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f2012-109.

Wenzel:2012:PDE

[513] Anja Wenzel, Ann-Kristin Bergström, Mats Jansson, and Tobias Vrede. Poor direct exploitation of terrestrial particulate organic material from peat layers by Daphnia galeata. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 69(11):1870–1880, November 2012. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f2012-110.

Trenkel:2012:REP

[514] Verena M. Trenkel, Mark V. Bravington, and Pascal Lorance. A random effects population dynamics model based on proportions-at-age and removal data for estimating total mortality. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 69(11):1881–1893, November 2012. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f2012-103.

Jackson:2012:CDT

[515] C. Rhett Jackson and Douglas J. Martin. Comment on “Does timber harvest influence the dynamics of marine-derived nutrients in Southeast Alaska streams?”. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 69(11): 1894–1897, November 2012. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/

10.1139/f2012-104. See [286] and reply [516].

Levi:2012:RCJ

[516] Peter S. Levi, Jennifer L. Tank, Scott D. Tiegs, Janine Rüegg, Dominic T. Chaloner, and Gary A. Lamberti. Reply to comment by Jackson and Martin on “Does timber harvest influence the dynamics of marine-derived nutrients in Southeast Alaska streams?”. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 69(11):1898–1901, November 2012. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f2012-106. See [286, 515].

Krueger:2012:CFW

[517] D. M. Krueger and T. R. Hrabik. Corrigendum: Food web alterations that promote native species: the recovery of cisco (Coregonus artedi) populations through management of native piscivores. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 69(11):1902, November 2012. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2012-0462. See [4].

Anonymous:2012:NAN

[518] Anonymous. Note of appreciation/note de reconnaissance. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 69(12), December 2012. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f2012-122.

Ashford:2012:DLS

[519] Julian Ashford, Michael Dinniman, Cassandra Brooks, Allen H. Andrews, Eileen Hofmann, Gregor Cailliet, Christopher Jones, and Nakul Ramanna. Does large-scale ocean circulation structure life history connectivity in Antarctic toothfish (Dissostichus mawsoni)? Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 69(12):1903–1919, December 2012. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f2012-111.

Rosenfeld:2012:MAH

[520] Jordan S. Rosenfeld and Ron Ptolemy. Modelling available habitat versus available energy flux: do PHABSIM applications that neglect prey abundance underestimate optimal flows for juvenile salmonids? Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 69(12):1920–1934, December 2012.

CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f2012-115.

Zhang:2012:DPA

[521] Min Zhang, Yang Yu, Zhen Yang, Xiaoli Shi, and Fanxiang Kong. The distribution of phytoplankton along trophic gradients and its mediation by available light in the pelagic zone of large eutrophic lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 69(12):1935–1946, December 2012. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f2012-113. See cor-

rigendum [570].

Ozerov:2012:RGR

[522] Mikhail Yu. Ozerov, Alexey E. Veselov, Jaakko Lumme, and Craig R. Primmer. “Riverscape” genetics: river characteristics influence the genetic structure and diversity of anadromous and freshwater Atlantic salmon (Salmo salar) populations in northwest Russia. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 69(12):1947–1958, December 2012. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f2012-114.

Humphrey:2012:EWC

[523] Sarah Humphrey, Yingming Zhao, and Dennis Higgs. The effects of water currents on walleye (Sander vitreus) eggs and larvae and implications for the early survival of walleye in Lake Erie. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 69(12):1959–1967, December 2012. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f2012-116.

Schreier:2012:HPP

[524] Andrea Drauch Schreier, Brian Mahardja, and Bernie May. Hierarchical patterns of population structure in the endangered Fraser River white sturgeon (Acipenser transmontanus) and implications for conservation. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 69(12):1968–1980, December 2012. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f2012-120.

Vainikka:2012:FRP

[525] Anssi Vainikka, Riina Huusko, Pekka Hyvärinen, Pekka K. Korhonen, Tapio Laaksonen, Juha Koskela, Jouni Vielma, Heikki Hirvonen, and Matti Salminen. Food restriction prior to release reduces precocious maturity and improves migration tendency of Atlantic salmon (Salmo salar) smolts. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 69(12):1981–1993, December 2012. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f2012-119.

Skaala:2012:PFH

[526] Øystein Skaala, Kevin A. Glover, Bjørn T. Barlaup, Terje Svåsand, Francois Besnier, Michael M. Hansen, and Reidar Borgstrøm. Performance of farmed, hybrid, and wild Atlantic salmon (Salmo salar) families in a natural river environment. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 69(12): 1994–2006, December 2012. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f2012-118.

Neff:2012:STT

[527] Margaret R. Neff, Satyendra P. Bhavsar, and Janice Xin Yan Chin. Spatial and temporal trends of muscle lipid content in Great Lakes fishes: 1970s–2008. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 69(12):2007–2017, December 2012. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f2012-121.

Pothoven:2012:HMP

[528] Steven A. Pothoven, Henry A. Vanderploeg, Tomas O. Höök, and Stuart A. Ludsin. Hypoxia modifies planktivore–zooplankton interactions in Lake Erie. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 69(12):2018– 2028, December 2012. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2012-0144.

Renner:2012:SMP

[529] Martin Renner, Mayumi L. Arimitsu, and John F. Piatt. Structure of marine predator and prey communities along environmental gradients in a glaciated fjord. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 69(12): 2029–2045, December 2012. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f2012-117.

Johannsson:2012:ZPB

[530] Ora E. Johannsson and Kelly L. Bowen. Zooplankton production in the Bay of Quinte 1975–2008: relationships with primary production, habitat, planktivory, and aquatic invasive species (Dreissena spp. and Cercopagis pengoi). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 69(12): 2046–2063, December 2012. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2012-0162.

Hirst:2012:BMF

[531] David Hirst, Geir Storvik, Hanne Rognebakke, Magne Aldrin, Sondre Aanes, and Jon Helge Vølstad. A Bayesian modelling framework for the estimation of catch-at-age of commercially harvested fish species. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 69(12):2064–2076, December 2012. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2012-0075.

Gaichas:2012:BDF

[532] Sarah K. Gaichas, Garrett Odell, Kerim Y. Aydin, and Robert C. Francis. Beyond the defaults: functional response parameter space and ecosystem-level fishing thresholds in dynamic food web model simulations. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 69(12):2077–2094, December 2012. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f2012-099.

Sharma:2012:ESS

[533] Sapna Sharma, Pierre Legendre, Daniel Boisclair, and Stéphane Gauthier. Effects of spatial scale and choice of statistical model (linear versus tree-based) on determining species–habitat relationships. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 69(12):2095–2111, December 2012. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2011-0505.

Doyle:2012:NCF

[534] Miriam J. Doyle and Kathryn L. Mier. A new conceptual framework for evaluating the early ontogeny phase of recruitment processes among marine fish species. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 69(12): 2112–2129, December 2012. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2012-0171.

Woodcock:2013:DTE

[535] Skye H. Woodcock, Casey A. Grieshaber, and Benjamin D. Walther. Dietary transfer of enriched stable isotopes to mark otoliths, fin rays, and scales. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 70(1):1–4, January 2013. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
cjfas-2012-0389.

Robertsen:2013:OSE

[536] Grethe Robertsen, Helge Skoglund, and Sigurd Einum. Offspring size effects vary over fine spatio-temporal scales in Atlantic salmon (Salmo salar). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 70(1):5–12, January 2013. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
cjfas-2012-0152.

Froeschke:2013:LTT

[537] John T. Froeschke, Bridgette F. Froeschke, and Charlotte M. Stinson. Long-term trends of bull shark (Carcharhinus leucas) in estuarine waters of Texas, USA. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 70 (1):13–21, January 2013. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2012-0037.

Haltuch:2013:CCB

[538] Melissa A. Haltuch, Owen S. Hamel, Kevin R. Piner, Patrick McDonald, Craig R. Kastelle, and John C. Field. A California Current bomb radiocarbon reference chronology and petrale sole (Eopsetta jordani) age validation. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 70(1):22–31, January 2013. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
cjfas-2011-0504.

Lacroix:2013:MSA

[539] Gilles L. Lacroix. Migratory strategies of Atlantic salmon (Salmo salar) postsmolts and implications for marine survival of endangered populations. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 70(1):32–48, January 2013. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
cjfas-2012-0270.

Dougherty:2013:MCO

[540] Dawn T. Dougherty, Ray Hilborn, André E. Punt, and Ian J. Stewart. Modeling co-occurring species: a simulation study on the effects of spatial scale for setting management targets. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 70(1):49–56, January 2013. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/cjfas-2012-0042.

Simmons:2013:SER

[541] Ryan K. Simmons, Thomas P. Quinn, Lisa W. Seeb, Daniel E. Schindler, and Ray Hilborn. Summer emigration and resource acquisition within a shared nursery lake by sockeye salmon (Oncorhynchus nerka) from historically discrete rearing environments. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 70(1):57–63, January 2013. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/cjfas-2012-0159.

Sloat:2013:MST

[542] Matthew R. Sloat and Ann-Marie K. Osterback. Maximum stream temperature and the occurrence, abundance, and behavior of steelhead trout (Oncorhynchus mykiss) in a southern California stream. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 70(1):64–73, January 2013. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2012-0228.

Swain:2013:CDT

[543] Douglas P. Swain, Ian D. Jonsen, James E. Simon, and Trevor D. Davies. Contrasting decadal trends in mortality between large and small individuals in skate populations in Atlantic Canada. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 70(1):74–89, January 2013. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/cjfas-2012-0179.

Donaldson:2013:EST

[544] M. R. Donaldson, G. D. Raby, V. N. Nguyen, S. G. Hinch, D. A. Patterson, A. P. Farrell, M. A. Rudd, L. A. Thompson, C. M. O’Connor, A. H. Colotelo, S. H. McConnachie, K. V. Cook, D. Robichaud, K. K. English, and S. J. Cooke. Evaluation of a simple technique for recovering fish from capture stress: integrating physiology, biotelemetry, and social science to solve a conservation problem. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 70 (1):90–100, January 2013. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2012-0218.

Chittenden:2013:IOO

[545] Cedar M. Chittenden, Per Fauchald, and Audun H. Rikardsen. Important open-ocean areas for northern Atlantic salmon (Salmo salar) — as estimated using a simple ambient-temperature approach. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 70(1):101–104, January 2013. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2012-0215.

McCormick:2013:PEC

[546] Stephen D. McCormick, Timothy F. Sheehan, Björn Thrandur Björnsson, Christine Lipsky, John F. Kocik, Amy M. Regish, and Michael F. O’Dea. Physiological and endocrine changes in Atlantic salmon smolts during hatchery rearing, downstream migration, and ocean entry. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 70(1):105–118, January 2013. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
cjfas-2012-0151.

Dionne:2013:RRC

[547] Phillip E. Dionne, Gayle B. Zydlewski, Michael T. Kinnison, Joseph Zydlewski, and Gail S. Wippelhauser. Reconsidering residency: characterization and conservation implications of complex migratory patterns of shortnose sturgeon (Acispenser brevirostrum). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 70(1):119–127, January 2013. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2012-0196.

Larson:2013:SNP

[548] Wesley A. Larson, Fred M. Utter, Katherine W. Myers, William D. Templin, James E. Seeb, Charles M. Guthrie III, Alexander V. Bugaev, and Lisa W. Seeb. Single-nucleotide polymorphisms reveal distribution and migration of Chinook salmon (Oncorhynchus tshawytscha) in the Bering Sea and North Pacific Ocean. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 70 (1):128–141, January 2013. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2012-0233.

Anonymous:2013:ACa

[549] Anonymous. Announcement / Communiqué. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 70(2), February 2013. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/cjfas-2013-0049.

Anonymous:2013:ACb

[550] Anonymous. Announcement / Communiqué. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 70(2), February 2013. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/cjfas-2013-0040.

Gronkjaer:2013:SCI

[551] Peter Grønkjær, Jens B. Pedersen, Torben T. Ankjærø, Henrik Kjeldsen, Jan Heinemeier, Petur Steingrund, Jens M. Nielsen, and Jens T. Christensen. Stable N and C isotopes in the organic matrix of fish otoliths: validation of a new approach for studying spatial and temporal changes in the trophic structure of aquatic ecosystems. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 70(2):143–146, February 2013. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2012-0386.

Hayes:2013:EIH

[552] Michael C. Hayes, Reginald R. Reisenbichler, Stephen P. Rubin, Deanne C. Drake, Karl D. Stenberg, and Sewall F. Young. Effectiveness of an integrated hatchery program: can genetic-based performance differences between hatchery and wild Chinook salmon be avoided? Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 70(2):147–158, February 2013. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2012-0138.

Hedger:2013:PCC

[553] Richard D. Hedger, Line E. Sundt-Hansen, Torbjørn Forseth, Ola Ugedal, Ola H. Diserud, Ånund S. Kvambekk, and Anders G. Finstad. Predicting climate change effects on subarctic–Arctic populations of Atlantic salmon (Salmo salar). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 70 (2):159–168, February 2013. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2012-0205.

Minns:2013:SMS

[554] Charles K. Minns and Brian J. Shuter. A semi-mechanistic seasonal temperature-profile model (STM) for the period of stratification in dimictic lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 70(2):169–181, February 2013. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
cjfas-2012-0253.

Martin:2013:PSA

[555] Jean Martin, Gilles Bareille, Sylvain Berail, Christophe Pécheyran, François Gueraud, Frédéric Lange, Françoise Daverat, Noëlle Bru, Eddy Beall, David Barracou, and Olivier Donard. Persistence of a southern Atlantic salmon population: diversity of natal origins from otolith elemental and Sr isotopic signatures. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 70 (2):182–197, February 2013. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2012-0284.

Basen:2013:PFQ

[556] Timo Basen, Rene Gergs, Karl-Otto Rothhaupt, and Dominik Martin-Creuzburg. Phytoplankton food quality effects on gammarids: benthic– pelagic coupling mediated by an invasive freshwater clam. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 70(2):198–207, February 2013. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2012-0188.

Kotwicki:2013:CBT

[557] Stan Kotwicki, Alex De Robertis, James N. Ianelli, André E. Punt, and John K. Horne. Combining bottom trawl and acoustic data to model acoustic dead zone correction and bottom trawl efficiency parameters for semipelagic species. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 70(2): 208–219, February 2013. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2012-0321.

Benoit:2013:EMS

[558] Hugues P. Benoît. An empirical model of seasonal depth-dependent fish assemblage structure to predict the species composition of mixed catches. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 70(2):220–232, February 2013. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2012-0166.

Morley:2013:RCS

[559] James W. Morley, Jeffrey A. Buckel, and Thomas E. Lankford. Relative contribution of spring- and summer-spawned bluefish cohorts to the adult population: effects of size-selective winter mortality, overwinter growth, and sampling bias. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 70 (2):233–244, February 2013. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2012-0332.

Lindegren:2013:TDP

[560] Martin Lindegren and David M. Checkley. Temperature dependence of Pacific sardine (Sardinops sagax) recruitment in the California Current ecosystem revisited and revised. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 70(2):245–252, February 2013. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/cjfas-2012-0211. See [1, 686].

Price:2013:EDR

[561] Amina E. Price, Paul Humphries, Ben Gawne, and Martin C. Thoms. Effects of discharge regulation on slackwater characteristics at multiple scales in a lowland river. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 70 (2):253–262, February 2013. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2012-0164.

Evans:2013:BGA

[562] Melissa L. Evans, Bryan D. Neff, and Daniel D. Heath. Behavioural and genetic analyses of mate choice and reproductive success in two Chinook salmon populations. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 70(2): 263–270, February 2013. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2012-0415.

Bugajski:2013:CCS

[563] Aleksandra Bugajski, Matthew W. Reudink, Jennifer L. Doucette, Samantha E. Franks, Björn Wissel, and Christopher M. Somers. The complexity of cormorants: stable isotopes reveal multiple prey sources and feeding site switching. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 70 (2):271–279, February 2013. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2012-0237.

Castro-Santos:2013:BSL

[564] Theodore Castro-Santos, Francisco Javier Sanz-Ronda, and Jorge Ruiz-Legazpi. Breaking the speed limit — comparative sprinting performance of brook trout (Salvelinus fontinalis) and brown trout (Salmo trutta). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 70(2):280–293, February 2013. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2012-0186.

Thorson:2013:ISV

[565] James T. Thorson, M. Elizabeth Clarke, Ian J. Stewart, and André E. Punt. The implications of spatially varying catchability on bottom trawl surveys of fish abundance: a proposed solution involving underwater vehicles. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 70(2):294–306, February 2013. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
cjfas-2012-0330.

Paul:2013:EFR

[566] Andrew J. Paul. Environmental flows and recruitment of walleye (Sander vitreus) in the peace–athabasca delta. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 70(2):307–315, February 2013. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/cjfas-2012-0279.

Jordaan:2013:MDS

[567] Adrian Jordaan, Michael G. Frisk, Lewis S. Incze, Nicholas H. Wolff, Lindsay Hamlin, and Yong Chen. Multivariate dissemination of species relationships for use in marine spatial planning. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 70(2):316–329, February 2013. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2011-0516.

Hold:2013:PES

[568] Natalie Hold, Lee G. Murray, Michel J. Kaiser, Hilmar Hinz, Andrew R. Beaumont, and Martin I. Taylor. Potential effects of stock enhancement with hatchery-reared seed on genetic diversity and effective population size. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 70(2):330–338, February 2013. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
cjfas-2012-0378.

Roy:2013:IVM

[569] Mathieu L. Roy, André G. Roy, James W. A. Grant, and Normand E. Bergeron. Individual variability in the movement behaviour of juvenile Atlantic salmon. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 70(2):

339–347, February 2013. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2012-0234.

Zhang:2013:CDP

[570] Min Zhang, Yang Yu, Zhen Yang, Xiaoli Shi, and Fanxiang Kong. Corrigendum: The distribution of phytoplankton along trophic gradients and its mediation by available light in the pelagic zone of large eutrophic lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 70(2):348, February 2013. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
cjfas-2013-0018. See [521].

Gallagher:2013:NPS

[571] Z. S. Gallagher, J. S. Bystriansky, A. P. Farrell, and C. J. Brauner. A novel pattern of smoltification in the most anadromous salmonid: pink salmon (Oncorhynchus gorbuscha). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 70(3):349–357, March 2013. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/cjfas-2012-0390.

Daly:2013:PHC

[572] Benjamin Daly, Ginny L. Eckert, and Timothy D. White. Predation of hatchery-cultured juvenile red king crabs (Paralithodes camtschaticus) in the wild. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 70(3):358–366, March 2013. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
cjfas-2012-0377.

Kinter:2013:NIV

[573] Bryan T. Kinter and Stuart A. Ludsin. Nutrient inputs versus piscivore biomass as the primary driver of reservoir food webs. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 70(3):367–380, March 2013. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2012-0369.

Lauretta:2013:CMS

[574] Matthew Vincent Lauretta, Edward Vincent Camp, William Earl Pine, and Thomas Kerry Frazer. Catchability model selection for estimating the composition of fishes and invertebrates within dynamic aquatic ecosystems. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 70(3):381–392, March 2013. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
cjfas-2012-0319.

Tierney:2013:IMR

[575] Keith B. Tierney, Matthew Kereliuk, Yogesh Kumar Katare, Alexander P. Scott, Stephen J. Loeb, and Barbara Zielinski. Invasive male round gobies (Neogobius melanostomus) release pheromones in their urine to attract females. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 70(3):393–400, March 2013. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
cjfas-2012-0246.

Fleischman:2013:ASS

[576] Steven J. Fleischman, Matthew J. Catalano, Robert A. Clark, and David R. Bernard. An age-structured state-space stock–recruit model for Pacific salmon (Oncorhynchus spp.). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 70(3):401–414, March 2013. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/cjfas-2012-0112.

Moran:2013:DLH

[577] Paul Moran, David J. Teel, Michael A. Banks, Terry D. Beacham, M. Renee Bellinger, Scott M. Blankenship, John R. Candy, John Carlos Garza, Jon E. Hess, Shawn R. Narum, Lisa W. Seeb, William D. Templin, Colin G. Wallace, and Christian T. Smith. Divergent life-history races do not represent Chinook salmon coast-wide: the importance of scale in quaternary biogeography. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 70 (3):415–435, March 2013. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2012-0135.

Depew:2013:OMC

[578] David C. Depew, Neil M. Burgess, M. Robin Anderson, Randy Baker, Satyendra P. Bhavsar, R. A. (Drew) Bodaly, Chris S. Eckley, Marlene S. Evans, Nikolaus Gantner, Jennifer A. Graydon, Kevin Jacobs, Jason E. LeBlanc, Vincent L. St. Louis, and Linda M. Campbell. An overview of mercury concentrations in freshwater fish species: a national fish mercury dataset for Canada. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 70 (3):436–451, March 2013. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2012-0338.

Halfyard:2013:CES

[579] Edmund A. Halfyard, A. Jamie F. Gibson, Michael J. W. Stokesbury, Daniel E. Ruzzante, and Frederick G. Whoriskey. Correlates of estuarine survival of Atlantic salmon postsmolts from the Southern Upland, Nova Scotia, Canada. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 70 (3):452–460, March 2013. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2012-0287.

Dupont-Prinet:2013:IHM

[580] Aurélie Dupont-Prinet, Marie Vagner, Denis Chabot, and Céline Audet. Impact of hypoxia on the metabolism of Greenland halibut (Reinhardtius hippoglossoides). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 70 (3):461–469, March 2013. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2012-0327.

Curti:2013:EPM

[581] Kiersten L. Curti, Jeremy S. Collie, Christopher M. Legault, and Jason S. Link. Evaluating the performance of a multispecies statistical catch-at-age model. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 70 (3):470–484, March 2013. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2012-0229.

Glover:2013:GLB

[582] David C. Glover, Dennis R. DeVries, and Russell A. Wright. Growth of largemouth bass in a dynamic estuarine environment: an evaluation of the relative effects of salinity, diet, and temperature. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 70(3):485–501, March 2013. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2012-0295.

Kohler:2013:SMN

[583] André E. Kohler, Paul C. Kusnierz, Timothy Copeland, David A. Venditti, Lytle Denny, Josh Gable, Bert A. Lewis, Ryan Kinzer, Bruce Barnett, and Mark S. Wipfli. Salmon-mediated nutrient flux in selected streams of the Columbia River basin, USA. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 70(3):502–512, March 2013. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/cjfas-2012-0347.

Chilcote:2013:CRR

[584] M. W. Chilcote, K. W. Goodson, and M. R. Falcy. Corrigendum: Reduced recruitment performance in natural populations of anadromous salmonids associated with hatchery-reared fish. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 70(3):513–515, March 2013. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:/

/cdnsciencepub.com/doi/full/10.1139/cjfas-2012-0542. See [223].

deKerckhove:2013:LER

[585] Derrick Tupper de Kerckhove, Charles Kenneth Minns, and Brian John Shuter. The length of environmental review in Canada under the Fisheries Act. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 70(4):517–521, April 2013. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
cjfas-2012-0411.

Jerde:2013:DAC

[586] Christopher L. Jerde, W. Lindsay Chadderton, Andrew R. Mahon, Mark A. Renshaw, Joel Corush, Michelle L. Budny, Sagar Mysorekar, and David M. Lodge. Detection of Asian carp DNA as part of a Great Lakes basin-wide surveillance program. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 70(4):522–526, April 2013. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/cjfas-2012-0478.

Pulkkinen:2013:MSE

[587] Henni Pulkkinen and Samu Mäntyniemi. Maximum survival of eggs as the key parameter of stock–recruit meta-analysis: accounting for parameter and structural uncertainty. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 70(4):527–533, April 2013. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/cjfas-2012-0268.

Feekings:2013:ERC

[588] Jordan Feekings, Peter Lewy, and Niels Madsen. The effect of regulation changes and influential factors on Atlantic cod discards in the Baltic Sea demersal trawl fishery. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 70 (4):534–542, April 2013. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2012-0273.

Jackson:2013:PPF

[589] Andrew T. Jackson, Alphonse Adite, Katherine A. Roach, and Kirk O. Winemiller. Primary production, food web structure, and fish yields in constructed and natural wetlands in the floodplain of an African river. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 70(4):543–553, April 2013. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2012-0403.

Gido:2013:MRN

[590] Keith B. Gido, David L. Propst, Julian D. Olden, and Kevin R. Bestgen. Multidecadal responses of native and introduced fishes to natural and altered flow regimes in the American southwest. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 70(4):554–564, April 2013. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/cjfas-2012-0441.

Lochet:2013:ISE

[591] Aude Lochet, J. Ellen Marsden, Brian J. Fryer, and Stuart A. Ludsin. Instability of statolith elemental signatures revealed in newly metamorphosed sea lamprey (Petromyzon marinus). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 70(4):565–573, April 2013. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/cjfas-2012-0410.

Satterthwaite:2013:CTP

[592] William H. Satterthwaite, Michael S. Mohr, Michael R. O’Farrell, and Brian K. Wells. A comparison of temporal patterns in the ocean spatial distribution of California’s Central Valley Chinook salmon runs. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 70(4):574–584, April 2013. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2012-0395.

Naslund:2013:HTE

[593] Joacim Näslund, Malin Rosengren, Diego Del Villar, Lars Gansel, Johnny R. Norrgård, Lo Persson, John James Winkowski, and Eli Kvingedal. Hatchery tank enrichment affects cortisol levels and shelterseeking in Atlantic salmon (Salmo salar). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 70(4):585–590, April 2013. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/cjfas-2012-0302.

Mantyniemi:2013:ISK

[594] Samu Mäntyniemi, Päivi Haapasaari, Sakari Kuikka, Raimo Parmanne, Maiju Lehtiniemi, and Joni Kaitaranta. Incorporating stakeholders’ knowledge to stock assessment: Central Baltic herring. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 70(4):591–599, April 2013. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2012-0316.

Lanthier:2013:OTV

[595] Gabriel Lanthier, Daniel Boisclair, Guillaume Bourque, Pierre Legendre, Michel Lapointe, and Bernard Angers. Optimization of temporal versus spatial replication in the development of habitat use models to explain among-reach variations of fish density estimates in rivers. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 70(4):600–609, April 2013. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2012-0216.

Persson:2013:DDI

[596] Lennart Persson, Per Arne Amundsen, André M. De Roos, Rune Knudsen, Raul Primicerio, and Anders Klemetsen. Density-dependent interactions in an Arctic char–brown trout system: competition, predation, or both? Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 70(4):610– 616, April 2013. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
cjfas-2012-0175.

Miller:2013:DBT

[597] Jessica A. Miller, David J. Teel, Antonio Baptista, and Cheryl A. Morgan. Disentangling bottom-up and top-down effects on survival during early ocean residence in a population of Chinook salmon (Oncorhynchus tshawytscha). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 70(4):617–629, April 2013. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
cjfas-2012-0354.

Michel:2013:SSP

[598] Christian Michel, Heike Schmidt-Posthaus, and Patricia Burkhardt-Holm. Suspended sediment pulse effects in rainbow trout (Oncorhynchus mykiss) — relating apical and systemic responses. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 70(4):630–641, April 2013. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2012-0376.

Webster:2013:PHM

[599] Raymond A. Webster, William G. Clark, Bruce M. Leaman, and Joan E. Forsberg. Pacific halibut on the move: a renewed understanding of adult migration from a coastwide tagging study. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 70(4):642–653, April 2013. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/cjfas-2012-0371.

Bond:2013:PID

[600] Morgan H. Bond and Thomas P. Quinn. Patterns and influences on Dolly Varden migratory timing in the Chignik Lakes, Alaska, and comparison of populations throughout the northeastern Pacific and Arctic oceans. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 70(5):655–665, May 2013. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2012-0416.

Booth:2013:HMF

[601] Michael T. Booth, Nelson G. Hairston, and Alexander S. Flecker. How mobile are fish populations? Diel movement, population turnover, and site fidelity in suckers. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 70 (5):666–677, May 2013. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2012-0334.

Whiteley:2013:FPS

[602] Andrew R. Whiteley, Jason A. Coombs, Mark Hudy, Zachary Robinson, Amanda R. Colton, Keith H. Nislow, and Benjamin H. Letcher. Fragmentation and patch size shape genetic structure of brook trout populations. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 70(5):678–688, May 2013. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2012-0493.

Omelkova:2013:DAS

[603] Markéta Omelková, Vít Syrovátka, Vendula Křoupalová, Vanda Rádková, Jindřiška Bojková, Michal Horsák, Marie Zhai, and Jan Helešic. Dipteran assemblages of spring fens closely follow the gradient of groundwater mineral richness. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 70(5):689–700, May 2013. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/cjfas-2013-0026.

Courter:2013:RRT

[604] Ian I. Courter, David B. Child, James A. Hobbs, Thomas M. Garrison, Justin J. G. Glessner, and Shadia Duery. Resident rainbow trout produce anadromous offspring in a large interior watershed. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 70(5):701–710, May 2013. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2012-0457.

Jolivet:2013:HDO

[605] Aurélie Jolivet, Jean-François Bardeau, Ronan Fablet, Yves-Marie Paulet, and Hélène de Pontual. How do the organic and mineral fractions drive the opacity of fish otoliths? Insights using Raman microspectrometry. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 70(5):711–719, May 2013.

CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2012-0298.

Ciotti:2013:DGR

[606] Benjamin J. Ciotti, Timothy E. Targett, and Michael T. Burrows. Decline in growth rate of juvenile European plaice (Pleuronectes platessa) during summer at nursery beaches along the west coast of Scotland. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 70(5):720–734, May 2013. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2012-0331.

Westley:2013:RSH

[607] Peter A. H. Westley, Thomas P. Quinn, and Andrew H. Dittman. Rates of straying by hatchery-produced Pacific salmon (Oncorhynchus spp.) and steelhead (Oncorhynchus mykiss) differ among species, life history types, and populations. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 70 (5):735–746, May 2013. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2012-0536.

Zimmermann:2013:DWE

[608] Emily W. Zimmermann, Craig F. Purchase, Ian A. Fleming, and John Brattey. Dispersal of wild and escapee farmed Atlantic cod (Gadus morhua) in Newfoundland. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 70 (5):747–755, May 2013. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2012-0428.

Berejikian:2013:LSF

[609] Barry A. Berejikian, Lance A. Campbell, and Megan E. Moore. Largescale freshwater habitat features influence the degree of anadromy in eight Hood Canal Oncorhynchus mykiss populations. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 70(5):756–765, May 2013. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2012-0491.

OBrien:2013:STP

[610] Heather D. O’Brien, M. Catherine Eimers, Shaun A. Watmough, and Nora J. Casson. Spatial and temporal patterns in total phosphorus in south-central Ontario streams: the role of wetlands and past disturbance. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 70(5):766–774, May 2013. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2012-0474.

Robinson:2013:PSF

[611] Clifford L. K. Robinson and Jennifer Yakimishyn. The persistence and stability of fish assemblages within eelgrass meadows (Zostera marina) on the Pacific coast of Canada. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 70 (5):775–784, May 2013. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2012-0339.

Dann:2013:EGD

[612] Tyler H. Dann, Christopher Habicht, Timothy T. Baker, and James E. Seeb. Exploiting genetic diversity to balance conservation and harvest of migratory salmon. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 70 (5):785–793, May 2013. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2012-0449.

Davidsen:2013:HBA

[613] Jan Grimsrud Davidsen, Audun Håvard Rikardsen, Eva Bonsak Thorstad, Elina Halttunen, Hiromichi Mitamura, Kim Præbel, Jofrid Skardhamar, and Tor Fredrik Næsje. Homing behaviour of Atlantic salmon (Salmo salar) during final phase of marine migration and river entry. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 70(5):794–802, May 2013. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2012-0352.

Pritt:2013:IIE

[614] Jeremy J. Pritt, Mark R. DuFour, Christine M. Mayer, Patrick M. Kocovsky, Jeffrey T. Tyson, Eric J. Weimer, and Christopher S. Vandergoot. Including independent estimates and uncertainty to quantify total abundance of fish migrating in a large river system: walleye (Sander vitreus) in the Maumee River, Ohio. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 70 (5):803–814, May 2013. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2012-0484.

Blair:2013:GRT

[615] Jennifer M. Blair, Ilia Ostrovsky, Brendan J. Hicks, Robert J. Pitkethley, and Paul Scholes. Growth of rainbow trout (Oncorhynchus mykiss) in warm-temperate lakes: implications for environmental change. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 70(5):815–823, May 2013. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2012-0409.

Yamazaki:2013:DSA

[616] Yuji Yamazaki and Terumi Nagai. Directional selection against different life histories in the Arctic lamprey (Lethenteron camtschaticum): identification by microsatellite analysis. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 70(6):825–829, June 2013. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/cjfas-2013-0090.

Segura:2013:DMS

[617] A. M. Segura, A. C. Milessi, R. Vögler, F. Galván-Magaña, and V. Muggeo. The determination of maturity stages in male elasmobranchs (Chondrichthyes) using a segmented regression of clasper length on total length. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 70(6):830– 833, June 2013. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
cjfas-2012-0545.

Fracz:2013:CWC

[618] Amanda Fracz and Patricia Chow-Fraser. Changes in water chemistry associated with beaver-impounded coastal marshes of eastern Georgian Bay. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 70(6):834–840, June 2013. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
cjfas-2012-0431.

Haynie:2013:CED

[619] Alan C. Haynie and Lisa Pfeiffer. Climatic and economic drivers of the Bering Sea walleye pollock (Theragra chalcogramma) fishery: implications for the future. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 70 (6):841–853, June 2013. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2012-0265.

Pekcan-Hekim:2013:PVP

[620] Zeynep Pekcan-Hekim, Laura Joensuu, and Jukka Horppila. Predation by a visual planktivore perch (Perca fluviatilis) in a turbulent and turbid environment. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 70(6):854– 859, June 2013. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
cjfas-2012-0466.

Guyette:2013:CAA

[621] Margaret Q. Guyette, Cynthia S. Loftin, and Joseph Zydlewski. Carcass analog addition enhances juvenile Atlantic salmon (Salmo salar) growth and condition. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 70 (6):860–870, June 2013. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2012-0496.

Daase:2013:TRE

[622] Malin Daase, Stig Falk-Petersen, Øystein Varpe, Gerald Darnis, Janne E. Søreide, Anette Wold, Eva Leu, Jørgen Berge, Benoit Philippe, and Louis Fortier. Timing of reproductive events in the marine copepod Calanus glacialis: a pan-Arctic perspective. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 70(6):871–884, June 2013. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/cjfas-2012-0401.

Monaghan:2013:IAA

[623] Kieran A. Monaghan and Amadeu M. V. M. Soares. Improving the accuracy of aquatic macroinvertebrate trait assignment — the importance of taxon-weighting. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 70 (6):885–894, June 2013. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2012-0402.

Corsi:2013:HRT

[624] Matthew P. Corsi, Lisa A. Eby, and Craig A. Barfoot. Hybridization with rainbow trout alters life history traits of native westslope cutthroat trout. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 70(6):895–904, June 2013. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2012-0312.

Warnock:2013:ABF

[625] Will G. Warnock and Joseph B. Rasmussen. Abiotic and biotic factors associated with brook trout invasiveness into bull trout streams of the Canadian Rockies. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 70 (6):905–914, June 2013. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2012-0387.

Petersson:2013:FEI

[626] Erik Petersson, Lars Karlsson, Bjarne Ragnarsson, Marcus Bryntesson, Anders Berglund, Stefan Stridsman, and Sara Jonsson. Fin erosion and injuries in relation to adult recapture rates in cultured smolts of Atlantic salmon and brown trout. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 70 (6):915–921, June 2013. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2012-0247.

Higgins:2013:ILD

[627] Ruth M. Higgins, Bret S. Danilowicz, Deirdre Brophy, Audrey J. Geffen, and Ted McGowan. Influence of the limit of detection on classification using otolith elemental signatures. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 70(6):922–929, June 2013. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/cjfas-2012-0464.

Mangel:2013:PSR

[628] Marc Mangel, Alec D. MacCall, Jon Brodziak, E. J. Dick, Robyn E. Forrest, Roxanna Pourzand, and Stephen Ralston. A perspective on steepness, reference points, and stock assessment. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 70(6):930–940, June 2013. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/cjfas-2012-0372.

Hoffman:2013:IPS

[629] Adam R. Hoffman, David E. Armstrong, and Richard C. Lathrop. Influence of phosphorus scavenging by iron in contrasting dimictic lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 70(7):941–952, July 2013. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2012-0391.

Crook:2013:RSD

[630] David A. Crook, Jed I. Macdonald, Dale G. McNeil, Dean M. Gilligan, Martin Asmus, Roland Maas, and Jon Woodhead. Recruitment sources and dispersal of an invasive fish in a large river system as revealed by otolith chemistry analysis. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 70 (7):953–963, July 2013. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2012-0504.

Hasnain:2013:PIR

[631] Sarah S. Hasnain, Brian J. Shuter, and Charles K. Minns. Phylogeny influences the relationships linking key ecological thermal metrics for North American freshwater fish species. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 70(7):964–972, July 2013. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/cjfas-2012-0217.

vanderLee:2013:FLC

[632] A. van der Lee, D. M. Gillis, P. Comeau, and P. Hurley. Fishing the line: catch and effort distribution around the seasonal haddock (Melanogrammus aeglefinus) spawning closure on the Scotian Shelf. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 70(7):973–981, July 2013. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2012-0341.

Shuter:2013:EMF

[633] B. J. Shuter, C. K. Minns, and S. R. Fung. Empirical models for forecasting changes in the phenology of ice cover for Canadian lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 70(7):982–991, July 2013. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2012-0437.

Marchal:2013:WAV

[634] Paul Marchal, José A. A. De Oliveira, Pascal Lorance, Loïc Baulier, and Lionel Pawlowski. What is the added value of including fleet dynamics processes in fisheries models? Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 70 (7):992–1010, July 2013. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2012-0326.

Lacroix:2013:PSR

[635] Gilles L. Lacroix. Population-specific ranges of oceanic migration for adult Atlantic salmon (Salmo salar) documented using pop-up satellite archival tags. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 70(7):1011– 1030, July 2013. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
cjfas-2013-0038.

Ziegler:2013:IDQ

[636] Philippe Eric Ziegler. Influence of data quality and quantity from a multiyear tagging program on an integrated fish stock assessment. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 70(7):1031–1045, July 2013. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2012-0413.

Steele:2013:VPB

[637] Craig A. Steele, Eric C. Anderson, Michael W. Ackerman, Maureen A. Hess, Nathan R. Campbell, Shawn R. Narum, and Matthew R. Campbell. A validation of parentage-based tagging using hatchery steelhead in the Snake River basin. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 70 (7):1046–1054, July 2013. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2012-0451.

Haynes:2013:MSS

[638] Trevor B. Haynes, Amanda E. Rosenberger, Mark S. Lindberg, Matthew Whitman, and Joel A. Schmutz. Method- and species-specific detection probabilities of fish occupancy in Arctic lakes: implications for design and management. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 70 (7):1055–1062, July 2013. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2012-0527.

Halttunen:2013:SDM

[639] Elina Halttunen, Jenny Lovisa Alexandra Jensen, Tor Fredrik Næsje, Jan Grimsrud Davidsen, Eva Bonsak Thorstad, Cedar Marget Chittenden, Sandra Hamel, Raul Primicerio, and Audun Håvard Rikardsen. State-dependent migratory timing of postspawned Atlantic salmon (Salmo salar). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 70(7):1063– 1071, July 2013. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
cjfas-2012-0525.

Beauregard:2013:CCF

[640] David Beauregard, Eva Enders, and Daniel Boisclair. Consequences of circadian fluctuations in water temperature on the standard metabolic rate of Atlantic salmon parr (Salmo salar). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 70(7):1072–1081, July 2013. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://

cdnsciencepub.com/doi/full/10.1139/cjfas-2012-0342.

Roy:2013:IVW

[641] Mathieu L. Roy, André G. Roy, James W. A. Grant, and Normand E. Bergeron. Individual variability of wild juvenile Atlantic salmon activity patterns: effect of flow stage, temperature, and habitat use. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 70(7):1082–1091, July 2013. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2012-0274.

Marco-Rius:2013:IPE

[642] Francisco Marco-Rius, Graciela Sotelo, Pablo Caballero, and Paloma Morán. Insights for planning an effective stocking program in anadromous brown trout (Salmo trutta). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 70(7):1092–1100, July 2013. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/cjfas-2013-0084.

Johnson:2013:SMP

[643] Nicholas S. Johnson, Michael J. Siefkes, C. Michael Wagner, Heather Dawson, Huiyong Wang, Todd Steeves, Michael Twohey, and Weiming Li. A synthesized mating pheromone component increases adult sea lamprey (Petromyzon marinus) trap capture in management scenarios. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 70(7):1101–1108, July 2013. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2013-0080.

Hansen:2013:FWC

[644] Gretchen J. A. Hansen, Catherine L. Hein, Brian M. Roth, M. Jake Vander Zanden, Jereme W. Gaeta, Alexander W. Latzka, and Stephen R. Carpenter. Food web consequences of long-term invasive crayfish control. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 70(7):1109–1122, July 2013. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2012-0460.

Pilliod:2013:EOA

[645] David S. Pilliod, Caren S. Goldberg, Robert S. Arkle, and Lisette P. Waits. Estimating occupancy and abundance of stream amphibians using environmental DNA from filtered water samples. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 70(8):1123–1130, August 2013. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2013-0047.

Andrews:2013:GLS

[646] Allen H. Andrews, Beverly K. Barnett, Robert J. Allman, Ryan P. Moyer, and Hannah D. Trowbridge. Great longevity of speckled hind (Epinephelus drummondhayi), a deep-water grouper, with novel use of postbomb radiocarbon dating in the Gulf of Mexico. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 70(8):1131–1140, August 2013. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2012-0537.

Neff:2013:RSP

[647] Margaret R. Neff and Donald A. Jackson. Regional-scale patterns in community concordance: testing the roles of historical biogeography versus contemporary abiotic controls in determining stream community composition. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 70(8):1141–1150, August 2013. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
cjfas-2012-0497.

Eiriksson:2013:STM

[648] Gudni Magnús Eiríksson and Einar Árnason. Spatial and temporal microsatellite variation in spawning Atlantic cod, Gadus morhua, around Iceland. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 70(8):1151–1158, August 2013. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
cjfas-2012-0494.

Barnes:2013:CEE

[649] Thomas C. Barnes and Bronwyn M. Gillanders. Combined effects of extrinsic and intrinsic factors on otolith chemistry: implications for environmental reconstructions. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 70 (8):1159–1166, August 2013. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2012-0442.

Burke:2013:EGF

[650] Brian J. Burke, Martin C. Liermann, David J. Teel, and James J. Anderson. Environmental and geospatial factors drive juvenile Chinook salmon distribution during early ocean migration. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 70(8):1167–1177, August 2013. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2012-0505.

Guan:2013:IPF

[651] Wenjiang Guan, Jie Cao, Yong Chen, and Matthew Cieri. Impacts of population and fishery spatial structures on fishery stock assessment. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 70(8):1178–1189, August 2013.

CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2012-0364.

Abbott:2013:AAM

[652] Joshua K. Abbott and Eli P. Fenichel. Anticipating adaptation: a mechanistic approach for linking policy and stock status to recreational angler behavior. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 70(8):1190–1208, August 2013. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
cjfas-2012-0517.

Hughes:2013:LOH

[653] Julian M. Hughes, John Stewart, Jeremy M. Lyle, Jaime McAllister, Jerom R. Stocks, and Iain M. Suthers. Latitudinal, ontogenetic, and historical shifts in the diet of a carnivorous teleost, Arripis trutta, in a coastal pelagic ecosystem altered by climate change. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 70(8):1209–1230, August 2013. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2013-0083.

Millar:2013:SVE

[654] Russell B. Millar and Chris E. Jordan. A simple variance estimator for the trapezoidal area-under-the-curve estimator of the spawner abundance of Pacific salmon. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 70(8): 1231–1239, August 2013. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2012-0499.

Watson:2013:STV

[655] Fiona L. Watson, Robert J. Miller, and Samuel Alan Stewart. Spatial and temporal variation in size at maturity for female American lobster in Nova Scotia. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 70(8): 1240–1251, August 2013. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2012-0480.

Ross:2013:TSR

[656] P. S. Ross, C. J. Kennedy, L. K. Shelley, K. B. Tierney, D. A. Patterson, W. L. Fairchild, and R. W. Macdonald. The trouble with salmon: relating pollutant exposure to toxic effect in species with transformational life histories and lengthy migrations. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 70(8):1252–1264, August 2013. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/cjfas-2012-0540.

Scott:2013:CAI

[657] J. Thad Scott, Mark J. McCarthy, Timothy G. Otten, Morgan M. Steffen, Bryant C. Baker, Erin M. Grantz, Steven W. Wilhelm, and Hans W. Paerl. Comment: An alternative interpretation of the relationship between TN:TP and microcystins in Canadian lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 70(8):1265–1268, August 2013. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:/

/cdnsciencepub.com/doi/full/10.1139/cjfas-2012-0490. See [477].

Carmichael:2013:CBE

[658] Ruth H. Carmichael, William Walton, and Heidi Clark. Corrigendum: Bivalve-enhanced nitrogen removal from coastal estuaries. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 70(8):1269, August 2013. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:/

/cdnsciencepub.com/doi/full/10.1139/cjfas-2013-0190. See [449].

Handegard:2013:EES

[659] Nils Olav Handegard, Tron Vedul Tronstad, and Jens Martin Hovem. Evaluating the effect of seismic surveys on fish — the efficacy of different exposure metrics to explain disturbance. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 70(9):1271–1277, September 2013. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2012-0465.

Oplinger:2013:PSS

[660] Randall W. Oplinger, Matthew J. Diana, and David H. Wahl. Population social structure and gizzard shad density influence the size-specific growth of bluegill. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 70(9):1278– 1288, September 2013. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2013-0055.

Goodbrand:2013:SCA

[661] L. Goodbrand, M. V. Abrahams, and G. A. Rose. Sea cage aquaculture affects distribution of wild fish at large spatial scales. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 70(9):1289–1295, September 2013. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2012-0317.

Fujiwara:2013:PDS

[662] Masami Fujiwara and Can Zhou. Population dynamics of stagestructured sequential hermaphrodites. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 70(9):1296–1305, September 2013. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2012-0520.

Miller:2013:CHM

[663] Timothy J. Miller. A comparison of hierarchical models for relative catch efficiency based on paired-gear data for US Northwest Atlantic fish stocks. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 70(9):1306–1316, September 2013. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
cjfas-2013-0136.

Mantyniemi:2013:IAS

[664] Samu Mäntyniemi, Laura Uusitalo, Heikki Peltonen, Päivi Haapasaari, and Sakari Kuikka. Integrated, age-structured, length-based stock assessment model with uncertain process variances, structural uncertainty, and environmental covariates: case of Central Baltic herring. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 70(9):1317–1326, September 2013. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2012-0315.

Moore:2013:IBD

[665] Jean-Sébastien Moore, Les N. Harris, Ross F. Tallman, and Eric B. Taylor. The interplay between dispersal and gene flow in anadromous Arctic char (Salvelinus alpinus): implications for potential for local adaptation. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 70(9):1327–1338, September 2013. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
cjfas-2013-0138.

Finstad:2013:HAS

[666] Anders Gravbrøt Finstad, Leif Magnus Sættem, and Sigurd Einum. Historical abundance and spatial distributions of spawners determine juvenile habitat accessibility in salmon: implications for population dynamics and management targets. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 70(9):1339–1345, September 2013. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/cjfas-2012-0455.

Diekert:2013:GVA

[667] Florian K. Diekert. The growing value of age: exploring economic gains from age-specific harvesting in the Northeast Arctic cod fishery. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 70(9):1346–1358, September 2013. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2012-0471.

Grothues:2013:SPL

[668] Thomas Marcellin Grothues and Walker Christian Davis. Sound pressure level weighting of the center of activity method to approximate sequential fish positions from acoustic telemetry. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 70(9):1359–1371, September 2013. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/cjfas-2013-0056.

Gasper:2013:MSD

[669] Jason R. Gasper and Gordon H. Kruse. Modeling of the spatial distribution of Pacific spiny dogfish (Squalus suckleyi) in the Gulf of Alaska using generalized additive and generalized linear models. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 70(9):1372–1385, September 2013. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2012-0535.

Hyvarinen:2013:ERI

[670] Pekka Hyvärinen and Petra Rodewald. Enriched rearing improves survival of hatchery-reared Atlantic salmon smolts during migration in the River Tornionjoki. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 70(9):1386– 1395, September 2013. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2013-0147.

Dub:2013:URF

[671] Joshua D. Dub, Rebecca A. Redman, David H. Wahl, and Sergiusz J. Czesny. Utilizing random forest analysis with otolith mass and total fish length to obtain rapid and objective estimates of fish age. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 70(9):1396–1401, September 2013. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2013-0169.

Hulson:2013:SMB

[672] Peter-John F. Hulson, Terrance J. Quinn, Dana H. Hanselman, and James N. Ianelli. Spatial modeling of Bering Sea walleye pollock with integrated age-structured assessment models in a changing environment. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 70(9):1402–1416, September 2013. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
cjfas-2013-0020.

Fraser:2013:ESE

[673] Dylan J. Fraser. The emerging synthesis of evolution with ecology in fisheries science. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 70(9):1417– 1428, September 2013. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2013-0171.

Jensen:2013:UMO

[674] Timothy Jensen and Jason C. Vokoun. Using multistate occupancy estimation to model habitat use in difficult-to-sample watersheds: bridle shiner in a low-gradient swampy stream. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 70(10):1429–1437, October 2013. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2013-0011.

Bravener:2013:BFT

[675] Gale A. Bravener and Robert L. McLaughlin. A behavioural framework for trapping success and its application to invasive sea lamprey. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 70(10):1438–1446, October 2013. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2012-0473.

Rossberg:2013:SRR

[676] Axel G. Rossberg, Jennifer E. Houle, and Kieran Hyder. Stock– recruitment relations controlled by feeding interactions alone. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 70(10):1447–1455, October 2013. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2012-0531.

Rosen:2013:DSC

[677] Shale Rosen, Terje Jörgensen, Darren Hammersland-White, and Jens Christian Holst. DeepVision: a stereo camera system provides highly accurate counts and lengths of fish passing inside a trawl. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 70(10):1456–1467, October 2013. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2013-0124.

Wiig:2013:SSI

[678] Jørgen Ree Wiig, Even Moland, Thrond O. Haugen, and Esben Moland Olsen. Spatially structured interactions between lobsters and lobster fishers in a coastal habitat: fine-scale behaviour and survival estimated from acoustic telemetry. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 70(10): 1468–1476, October 2013. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2013-0209.

Jonsson:2013:ACM

[679] Tomas Jonsson, Malin Setzer, John G. Pope, and Alfred Sandström. Addressing catch mechanisms in gillnets improves modeling of selectivity and estimates of mortality rates: a case study using survey data on an endangered stock of Arctic char. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 70(10):1477–1487, October 2013. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/cjfas-2012-0472.

Linnansaari:2013:EIB

[680] Tommi Linnansaari and Richard A. Cunjak. Effects of ice on behaviour of juvenile Atlantic salmon (Salmo salar). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 70(10):1488–1497, October 2013. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2012-0236.

Larson:2013:ISO

[681] Sean D. Larson, Zachary N. Hoyt, Ginny L. Eckert, and Verena A. Gill. Impacts of sea otter (Enhydra lutris) predation on commercially important sea cucumbers (Parastichopus californicus) in southeast Alaska. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 70(10):1498–1507, October 2013. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2013-0025.

Ouellet:2013:SDA

[682] Patrick Ouellet, Alice Olga Victoria Bui, Diane Lavoie, Joël Chassé, Nicolas Lambert, Nadia Ménard, and Pascal Sirois. Seasonal distribution, abundance, and growth of larval capelin (Mallotus villosus) and the role of the Lower Estuary (Gulf of St. Lawrence, Canada) as a nursery area. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 70(10):1508–1530, October 2013. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2013-0227.

Olker:2013:FIU

[683] Jennifer H. Olker, Lucinda B. Johnson, Richard P. Axler, and Catherine M. Johnson. Factors influencing ultraviolet radiation dose to developing frogs in northern vernal pools. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 70(10):1531–1541, October 2013. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/cjfas-2013-0137.

Ward:2013:MUH

[684] Hillary G. M. Ward, Paul J. Askey, and John R. Post. A mechanistic understanding of hyperstability in catch per unit effort and density-dependent catchability in a multistock recreational fishery.

Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 70(10):1542–1550, October 2013. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/

cjfas-2013-0264. See corrigendum [716].

Keefer:2013:FPB

[685] Matthew L. Keefer, Christopher C. Caudill, Tami S. Clabough, Michael A. Jepson, Eric L. Johnson, Christopher A. Peery, Megan D. Higgs, and Mary L. Moser. Fishway passage bottleneck identification and prioritization: a case study of Pacific lamprey at Bonneville Dam. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 70(10):1551–1565, October 2013. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2013-0164.

Jacobson:2013:CTD

[686] Larry D. Jacobson and Sam McClatchie. Comment on temperaturedependent stock–recruit modeling for Pacific sardine (Sardinops sagax) in Jacobson and MacCall (1995), McClatchie et al. (2010), and lindegren and checkley (2013). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 70(10): 1566–1569, October 2013. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2013-0128. See [1, 151, 560].

Anonymous:2013:ALC

[687] Anonymous. The American lobster in a changing ecosystem: a US– Canada Science Symposium. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 70(11), November 2013. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2013-0485.

Wahle:2013:ALC

[688] Richard A. Wahle, Andrea Battison, Louis Bernatchez, Stephanie Boudreau, Kathy Castro, Jonathan H. Grabowski, Spencer J. Greenwood, Carla Guenther, Rémy Rochette, and Jim Wilson. The American lobster in a changing ecosystem: a US–Canada Science Symposium, 27– 30 November 2012, Portland, Maine. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 70(11):1571–1575, November 2013. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/cjfas-2013-0465.

Shields:2013:CEE

[689] Jeffrey D. Shields. Complex etiologies of emerging diseases in lobsters (Homarus americanus) from Long Island Sound. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 70(11):1576–1587, November 2013. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2013-0050.

Laufer:2013:MFM

[690] Hans Laufer, Ming Chen, Bryan Baclaski, James M. Bobbitt, James D. Stuart, Yuegang Zuo, and Molly W. Jacobs. Multiple factors in marine environments affecting lobster survival, development, and growth, with emphasis on alkylphenols: a perspective. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 70(11):1588–1600, November 2013. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2013-0148.

Kunkel:2013:MCP

[691] Joseph G. Kunkel. Modeling the calcium and phosphate mineralization of American lobster cuticle. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 70(11): 1601–1611, November 2013. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2013-0034.

Steneck:2013:ALD

[692] Robert S. Steneck and Richard A. Wahle. American lobster dynamics in a brave new ocean. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 70(11): 1612–1624, November 2013. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2013-0094.

McMahan:2013:UAT

[693] Marissa D. McMahan, Damian C. Brady, Diane F. Cowan, Jonathan H. Grabowski, and Graham D. Sherwood. Using acoustic telemetry to observe the effects of a groundfish predator (Atlantic cod, Gadus morhua) on movement of the American lobster (Homarus americanus). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 70(11):1625–1634, November 2013. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2013-0065.

Oppenheim:2013:CNS

[694] Noah G. Oppenheim and Richard A. Wahle. Cannibals by night? In situ video monitoring reveals diel shifts in inter- and intra-specific predation on the American lobster. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 70(11): 1635–1640, November 2013. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2013-0099.

Annis:2013:ETM

[695] Eric R. Annis, Carl J. Wilson, Robert Russell, and Philip O. Yund. Evidence for thermally mediated settlement in lobster larvae (Homarus americanus). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 70(11):1641– 1649, November 2013. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2013-0060.

Jury:2013:SSD

[696] Steven H. Jury and Winsor H. Watson. Seasonal and sexual differences in the thermal preferences and movements of American lobsters. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 70(11):1650–1657, November 2013. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2013-0061.

Rycroft:2013:PML

[697] Nathan Rycroft, Kristin Radcliffe, and Jelle Atema. A photographic method for lobster morphometry aimed at site discrimination. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 70(11):1658–1665, November 2013. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2013-0103.

Anonymous:2013:NAN

[698] Anonymous. Note of appreciation / note de reconnaissance. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 70(12), December 2013. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2013-0518.

Lyons:2013:MMT

[699] Kady Lyons and Christopher G. Lowe. Mechanisms of maternal transfer of organochlorine contaminants and mercury in the common thresher shark (Alopias vulpinus). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 70(12): 1667–1672, December 2013. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2013-0222.

Dekker:2013:AFI

[700] Willem Dekker and Niklas B. Sjöberg. Assessment of the fishing impact on the silver eel stock in the Baltic using survival analysis. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 70(12):1673–1684, December 2013. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2013-0250.

Jonsson:2013:FEP

[701] Mikael Jönsson, Lynn Ranåker, P. Anders Nilsson, and Christer Brönmark. Foraging efficiency and prey selectivity in a visual predator: differential effects of turbid and humic water. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 70(12):1685–1690, December 2013. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2013-0150.

Lehnert:2013:EGR

[702] Sarah J. Lehnert, John W. Heath, and Daniel D. Heath. Ecological and genetic risks arising from reproductive interactions between wild and farmed Chinook salmon. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 70(12): 1691–1698, December 2013. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2013-0181.

Somerton:2013:CSC

[703] David A. Somerton, Kenneth L. Weinberg, and Scott E. Goodman. Catchability of snow crab (Chionoecetes opilio) by the eastern Bering Sea bottom trawl survey estimated using a catch comparison experiment. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 70(12):1699–1708, December 2013. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
cjfas-2013-0100.

Robinson:2013:SAS

[704] Jason M. Robinson, Michael J. Wilberg, Jean V. Adams, and Michael L. Jones. A spatial age-structured model for describing sea lamprey (Petromyzon marinus) population dynamics. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 70(12):1709–1722, December 2013. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2012-0375.

Brodziak:2013:MSM

[705] Jon Brodziak and William A. Walsh. Model selection and multimodel inference for standardizing catch rates of bycatch species: a case study of oceanic whitetip shark in the Hawaii-based longline fishery. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 70(12):1723–1740, December 2013. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2013-0111.

Lehuta:2013:SBA

[706] Sigrid Lehuta, Stéphanie Mahévas, Pascal Le Floc’h, and Pierre Petitgas. A simulation-based approach to assess sensitivity and robustness of fisheries management indicators for the pelagic fishery in the Bay of Biscay. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 70(12):1741–1756, December 2013. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
cjfas-2013-0066.

Vachon:2013:ESS

[707] Dominic Vachon and Yves T. Prairie. The ecosystem size and shape dependence of gas transfer velocity versus wind speed relationships in lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 70(12):1757–1764, December 2013. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
cjfas-2013-0241.

Harrison:2013:DVM

[708] P. M. Harrison, L. F. G. Gutowsky, E. G. Martins, D. A. Patterson, A. Leake, S. J. Cooke, and M. Power. Diel vertical migration of adult burbot: a dynamic trade-off among feeding opportunity, predation avoidance, and bioenergetic gain. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 70 (12):1765–1774, December 2013. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/

doi/full/10.1139/cjfas-2013-0183.

Wolff:2013:CHW

[709] Brian A. Wolff, Brett M. Johnson, and Chad M. Landress. Classification of hatchery and wild fish using natural geochemical signatures in otoliths, fin rays, and scales of an endangered catostomid. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 70(12):1775–1784, December 2013. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2013-0116.

Capello:2013:FSC

[710] M. Capello, P. Bach, and E. Romanov. Fine-scale catch data reveal clusters of large predators in the pelagic realm. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 70(12):1785–1791, December 2013. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2013-0149.

Ono:2013:HDS

[711] Kotaro Ono, Daniel S. Holland, and Ray Hilborn. How does species association affect mixed stock fisheries management? A comparative analysis of the effect of marine protected areas, discard bans, and individual fishing quotas. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 70(12): 1792–1804, December 2013. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2013-0046.

Schein:2013:FWS

[712] Allison Schein, Simon C. Courtenay, Karen A. Kidd, K. Alexander Campbell, and Michael R. van den Heuvel. Food web structure within an estuary of the southern Gulf of St. Lawrence undergoing eutrophication. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 70(12):1805–1812, December 2013. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
cjfas-2013-0251.

Munroe:2013:HDS

[713] Daphne M. Munroe, Eileen E. Hofmann, Eric N. Powell, and John M. Klinck. How do shellfisheries influence genetic connectivity in metapopulations? A modeling study examining the role of lower size limits in oyster fisheries. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 70(12):1813– 1828, December 2013. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2013-0089.

Thorson:2013:NRE

[714] James T. Thorson, Cóilín Minto, Carolina V. Minte-Vera, Kristin M. Kleisner, and Catherine Longo. A new role for effort dynamics in the theory of harvested populations and data-poor stock assessment. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 70(12):1829–1844, December 2013. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2013-0280.

Yule:2013:EAA

[715] Daniel L. Yule, Jean V. Adams, David M. Warner, Thomas R. Hrabik, Patrick M. Kocovsky, Brian C. Weidel, Lars G. Rudstam, and Patrick J. Sullivan. Evaluating analytical approaches for estimating pelagic fish biomass using simulated fish communities. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 70(12):1845–1857, December 2013. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2013-0072.

Ward:2013:CMU

[716] Hillary G. M. Ward, Paul J. Askey, and John R. Post. Corrigendum: A mechanistic understanding of hyperstability in catch per unit effort and density-dependent catchability in a multistock recreational fishery. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 70(12):1858, December 2013. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
cjfas-2013-0547. See [684].

Milesi:2014:CEA

[717] Silvia Vendruscolo Milesi and Adriano Sanches Melo. Conditional effects of aquatic insects of small tributaries on mainstream assemblages: position within drainage network matters. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 71(1):1–9, January 2014. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/cjfas-2013-0092.

Stafford:2014:ILT

[718] Craig P. Stafford, Megan V. McPhee, Lisa A. Eby, and Fred W. Allendorf. Introduced lake trout exhibit life history and morphological divergence with depth. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 71 (1):10–20, January 2014. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2013-0115.

Edwards:2014:LTC

[719] Brie A. Edwards, Donald A. Jackson, and Keith M. Somers. Linking temporal changes in crayfish communities to environmental changes in boreal shield lakes in south-central Ontario. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 71(1):21–30, January 2014. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/cjfas-2013-0232.

Mackinson:2014:CAR

[720] Steven Mackinson. Combined analyses reveal environmentally driven changes in the North Sea ecosystem and raise questions regarding what makes an ecosystem model’s performance credible? Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 71(1):31–46, January 2014. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2013-0173.

Chezik:2014:FGDa

[721] Kyle A. Chezik, Nigel P. Lester, and Paul A. Venturelli. Fish growth and degree-days i: selecting a base temperature for a within-population study. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 71(1):47–55, January 2014. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2013-0295.

Spence:2014:GVE

[722] Brian C. Spence and E. J. Dick. Geographic variation in environmental factors regulating outmigration timing of Coho salmon (Oncorhynchus kisutch) smolts. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 71(1):56–69, January 2014. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
cjfas-2012-0479.

Ohms:2014:ISM

[723] Haley A. Ohms, Matthew R. Sloat, Gordon H. Reeves, Chris E. Jordan, and Jason B. Dunham. Influence of sex, migration distance, and latitude on life history expression in steelhead and rainbow trout (Oncorhynchus mykiss). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 71 (1):70–80, January 2014. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2013-0274.

Hess:2014:CPS

[724] Jon E. Hess, Paul Chittaro, Anna Elz, Elizabeth A. Gilbert-Horvath, Victor Simon, and John Carlos Garza. Cryptic population structure in the severely depleted cowcod, Sebastes levis. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 71(1):81–92, January 2014. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/cjfas-2012-0510.

Kaemingk:2014:EBS

[725] Mark A. Kaemingk, Kristopher J. Stahr, Jeffrey C. Jolley, Richard S. Holland, and David W. Willis. Evidence for bluegill spawning plasticity obtained by disentangling complex factors related to recruitment. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 71(1):93–105, January 2014. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2013-0282.

delaMare:2014:ERA

[726] William K. de la Mare. Estimating relative abundance of whales from historical Antarctic whaling records. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 71(1):106–119, January 2014. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/cjfas-2013-0016.

Raabe:2014:SCR

[727] Joshua K. Raabe, Beth Gardner, and Joseph E. Hightower. A spatial capture–recapture model to estimate fish survival and location from linear continuous monitoring arrays. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 71(1):120–130, January 2014. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/cjfas-2013-0198.

Benjamin:2014:STV

[728] Joseph R. Benjamin, Lisa A. Wetzel, Kyle D. Martens, Kimberly Larsen, and Patrick J. Connolly. Spatio-temporal variability in movement, age, and growth of mountain whitefish (Prosopium williamsoni) in a river network based upon PIT tagging and otolith chemistry. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 71(1):131–140, January 2014. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2013-0279.

vanderLee:2014:CAS

[729] A. van der Lee, D. M. Gillis, and P. Comeau. Comparative analysis of the spatial distribution of fishing effort contrasting ecological isodars and discrete choice models. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 71 (1):141–150, January 2014. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2012-0511.

Spencer:2014:IME

[730] Paul D. Spencer, Sarah B. M. Kraak, and Edward A. Trippel. The influence of maternal effects in larval survival on fishery harvest reference points for two life-history patterns. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 71(1):151–161, January 2014. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/cjfas-2013-0253.

Meador:2014:DCC

[731] James P. Meador. Do chemically contaminated river estuaries in Puget Sound (Washington, USA) affect the survival rate of hatchery-reared Chinook salmon? Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 71(1):162–180, January 2014. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
cjfas-2013-0130.

Liu:2014:OBE

[732] Xiaozi Liu and Mikko Heino. Overlooked biological and economic implications of within-season fishery dynamics. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 71(2):181–188, February 2014. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2013-0029.

MacDonald:2014:PFC

[733] Ryan J. MacDonald, Sarah Boon, James M. Byrne, Mike D. Robinson, and Joseph B. Rasmussen. Potential future climate effects on mountain hydrology, stream temperature, and native salmonid life history. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 71(2):189–202, February 2014. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2013-0221.

Minto:2014:PDA

[734] Cóilín Minto, Joanna Mills Flemming, Gregory Lee Britten, and Boris Worm. Productivity dynamics of Atlantic cod. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 71(2):203–216, February 2014. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2013-0161.

Miyoshi:2014:CSA

[735] Koji Miyoshi, Kazufumi Hayashida, Taku Sakashita, Makoto Fujii, Hisaya Nii, Katsuya Nakao, and Hiroshi Ueda. Comparison of the swimming ability and upstream-migration behavior between chum salmon and masu salmon. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 71(2):217–225, February 2014. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
cjfas-2013-0480.

Stachura:2014:OIP

[736] Megan M. Stachura, Nathan J. Mantua, and Mark D. Scheuerell. Oceanographic influences on patterns in North Pacific salmon abundance. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 71(2):226–235, February 2014. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2013-0367.

Stoffels:2014:RFF

[737] R. J. Stoffels, K. R. Clarke, R. A. Rehwinkel, and B. J. McCarthy. Response of a floodplain fish community to river-floodplain connectivity: natural versus managed reconnection. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 71(2):236–245, February 2014. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/cjfas-2013-0042.

Bradbury:2014:LSC

[738] Ian R. Bradbury, Lorraine C. Hamilton, Martha J. Robertson, Chuck E. Bourgeois, Atef Mansour, and J. Brian Dempson. Landscape structure and climatic variation determine Atlantic salmon genetic connectivity in the Northwest Atlantic. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 71 (2):246–258, February 2014. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2013-0240.

Schaller:2014:ERM

[739] Howard A. Schaller, Charles E. Petrosky, and Eric S. Tinus. Evaluating river management during seaward migration to recover Columbia River stream-type Chinook salmon considering the variation in marine conditions. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 71(2):259–271, February 2014. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
cjfas-2013-0226.

Lehuta:2014:IIF

[740] Sigrid Lehuta, Daniel S. Holland, and Andrew J. Pershing. Investigating interconnected fisheries: a coupled model of the lobster and herring fisheries in New England. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 71 (2):272–289, February 2014. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2013-0185.

Lobon-Cervia:2014:RSR

[741] Javier Lobón-Cerviá. Recruitment and survival rate variability in fish populations: density-dependent regulation or further evidence of environmental determinants? Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 71 (2):290–300, February 2014. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2013-0320.

Nakajima:2014:GEM

[742] Kaori Nakajima, Shuichi Kitada, Yoko Habara, Shoko Sano, Emi Yokoyama, Takuma Sugaya, Akio Iwamoto, Hirohisa Kishino, and Katsuyuki Hamasaki. Genetic effects of marine stock enhancement: a case study based on the highly piscivorous Japanese Spanish mackerel. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 71(2):301–314, February 2014. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2013-0418.

Gaeta:2014:DDL

[743] Jereme W. Gaeta, Greg G. Sass, and Stephen R. Carpenter. Droughtdriven lake level decline: effects on coarse woody habitat and fishes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 71(2):315–325, February 2014. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2013-0451.

Kristensen:2014:EST

[744] Kasper Kristensen, Uffe Høgsbro Thygesen, Ken Haste Andersen, and Jan E. Beyer. Estimating spatio-temporal dynamics of size-structured populations. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 71(2):326–336, February 2014. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
cjfas-2013-0151.

Raney:2014:UDS

[745] Shanel M. Raney and M. Catherine Eimers. Unexpected declines in stream phosphorus concentrations across southern Ontario. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 71(3):337–342, March 2014. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2013-0300.

Lacroix:2014:LPP

[746] Gilles L. Lacroix. Large pelagic predators could jeopardize the recovery of endangered Atlantic salmon. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 71 (3):343–350, March 2014. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2013-0458.

Smith:2014:PLB

[747] James A. Smith and Matthew D. Taylor. A peaked logistic-based selection curve. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 71(3):351–355, March 2014. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
cjfas-2013-0401.

Veinott:2014:EES

[748] Geoff Veinott, Peter A. H. Westley, Craig F. Purchase, and Lucas Warner. Experimental evidence simultaneously confirms and contests assumptions implicit to otolith microchemistry research. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 71(3):356–365, March 2014. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2013-0224.

Bastardie:2014:DDI

[749] Francois Bastardie, J. Rasmus Nielsen, and Tanja Miethe. DISPLACE: a dynamic, individual-based model for spatial fishing planning and effort displacement — integrating underlying fish population models. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 71(3):366–386, March 2014. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2013-0126. See cor-

rigendum [1101].

Starrs:2014:MTI

[750] Danswell Starrs, Jacqueline T. Davis, Jodie Schlaefer, Brendan C. Ebner, Stephen M. Eggins, and Christopher J. Fulton. Maternally transmitted isotopes and their effects on larval fish: a validation of dual isotopic marks within a meta-analysis context. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 71(3):387–397, March 2014. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/cjfas-2013-0416.

Liberoff:2014:TEA

[751] Ana Laura Liberoff, Jessica Adele Miller, Carla Marcela Riva-Rossi, Fernando José Hidalgo, Marilyn Louise Fogel, and Miguel Alberto Pascual. Transgenerational effects of anadromy on juvenile growth traits in an introduced population of rainbow trout (Oncorhynchus mykiss). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 71(3):398–407, March 2014. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2013-0466.

Scherer:2014:ARR

[752] Avery E. Scherer and Nicholas Santangelo. Assessment of reproductive requirements in habitat conservation efforts: a case study on blackside dace (Chrosomus cumberlandensis), a federally listed threatened species. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 71(3):408–415, March 2014. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2013-0404.

Mogensen:2014:VHA

[753] Stephanie Mogensen, John R. Post, and Michael G. Sullivan. Vulnerability to harvest by anglers differs across climate, productivity, and diversity clines. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 71(3):416–426, March 2014. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
cjfas-2013-0336.

Goetz:2014:PDB

[754] Frederick Goetz, Andrew Jasonowicz, Ronald Johnson, Peggy Biga, Greg Fischer, and Shawn Sitar. Physiological differences between lean and siscowet lake trout morphotypes: Are these metabolotypes? Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 71(3):427–435, March 2014. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2013-0463.

Visconti:2014:SLV

[755] Anna Visconti, Pietro Volta, Amedeo Fadda, Antonio Di Guardo, and Marina Manca. Seasonality, littoral versus pelagic carbon sources, and stepwise ^{15}N-enrichment of pelagic food web in a deep subalpine lake: the role of planktivorous fish. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 71 (3):436–446, March 2014. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2013-0178.

Hosack:2014:IRE

[756] Geoffrey R. Hosack, Gareth W. Peters, and Stuart A. Ludsin. Interspecific relationships and environmentally driven catchabilities estimated from fisheries data. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 71 (3):447–463, March 2014. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2013-0236.

Schobernd:2014:EUA

[757] Zeb H. Schobernd, Nathan M. Bacheler, and Paul B. Conn. Examining the utility of alternative video monitoring metrics for indexing reef fish abundance. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 71(3):464–471, March 2014. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
cjfas-2013-0086.

Robert:2014:CCF

[758] Marianne Robert, Laurent Dagorn, Nathalie Bodin, Fabrice Pernet, Eve-Julie Arsenault-Pernet, and Jean Louis Deneubourg. Comparison of condition factors of skipjack tuna (Katsuwonus pelamis) associated or not with floating objects in an area known to be naturally enriched with logs. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 71(3):472–478, March 2014. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2013-0389.

Fogarty:2014:AEB

[759] Michael J. Fogarty. The art of ecosystem-based fishery management. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 71(3):479–490, March 2014.

CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2013-0203.

Sloat:2014:ICS

[760] Matthew R. Sloat and Gordon H. Reeves. Individual condition, standard metabolic rate, and rearing temperature influence steelhead and rainbow trout (Oncorhynchus mykiss) life histories. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 71(4):491–501, April 2014. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/cjfas-2013-0366.

Janetski:2014:NEC

[761] David J. Janetski, Dominic T. Chaloner, Ashley H. Moerke, Peter S. Levi, and Gary A. Lamberti. Novel environmental conditions alter subsidy and engineering effects by introduced Pacific salmon. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 71(4):502–513, April 2014. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2013-0292.

Rudershausen:2014:ERF

[762] P. J. Rudershausen, J. A. Buckel, and J. E. Hightower. Estimating reef fish discard mortality using surface and bottom tagging: effects of hook injury and barotrauma. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 71 (4):514–520, April 2014. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2013-0337.

Swain:2014:ESD

[763] Noel R. Swain, Morgan D. Hocking, Jennifer N. Harding, and John D. Reynolds. Effects of salmon on the diet and condition of streamresident sculpins. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 71(4):521– 532, April 2014. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/

cjfas-2013-0159. See corrigendum [814].

Meckley:2014:CMM

[764] Trevor D. Meckley, C. Michael Wagner, and Eliezer Gurarie. Coastal movements of migrating sea lamprey (Petromyzon marinus) in response to a partial pheromone added to river water: implications for management of invasive populations. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 71 (4):533–544, April 2014. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2013-0487.

Grote:2014:MSD

[765] Ann B. Grote, Michael M. Bailey, Joseph D. Zydlewski, and Joseph E. Hightower. Multibeam sonar (DIDSON) assessment of American shad (Alosa sapidissima) approaching a hydroelectric dam. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 71(4):545–558, April 2014. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2013-0308.

Hamel:2014:UMR

[766] Martin J. Hamel, Jeff D. Koch, Kirk D. Steffensen, Mark A. Pegg, Jeremy J. Hammen, and Mathew L. Rugg. Using mark–recapture information to validate and assess age and growth of long-lived fish species. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 71(4):559–566, April 2014. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2013-0393.

Sagarese:2014:IES

[767] Skyler R. Sagarese, Michael G. Frisk, Thomas J. Miller, Kathy A. Sosebee, John A. Musick, and Paul J. Rago. Influence of environmental, spatial, and ontogenetic variables on habitat selection and management of spiny dogfish in the northeast (US) shelf large marine ecosystem. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 71(4):567–580, April 2014. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2013-0259.

Stewart:2014:BSS

[768] Ian J. Stewart and Owen S. Hamel. Bootstrapping of sample sizes for length- or age-composition data used in stock assessments. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 71(4):581–588, April 2014. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2013-0289.

Scrimgeour:2014:QER

[769] Garry J. Scrimgeour, William M. Tonn, and Nicholas E. Jones. Quantifying effective restoration: reassessing the productive capacity of a constructed stream 14 years after construction. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 71(4):589–601, April 2014. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/cjfas-2013-0354.

Harnish:2014:EHD

[770] Ryan A. Harnish, Rishi Sharma, Geoffrey A. McMichael, Russell B. Langshaw, and Todd N. Pearsons. Effect of hydroelectric dam operations on the freshwater productivity of a Columbia River fall Chinook salmon population. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 71 (4):602–615, April 2014. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2013-0276.

Lavery:2014:EEC

[771] J. M. Lavery, J. Kurek, K. M. Rühland, C. A. Gillis, M. F. J. Pisaric, and J. P. Smol. Exploring the environmental context of recent Didymosphenia geminata proliferation in Gaspésie, Quebec, using paleolimnology. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 71(4):616–626, April 2014. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2013-0442.

Tsehaye:2014:MSA

[772] Iyob Tsehaye, Michael L. Jones, James R. Bence, Travis O. Brenden, Charles P. Madenjian, and David M. Warner. A multispecies statistical age-structured model to assess predator–prey balance: application to an intensively managed Lake Michigan pelagic fish community. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 71(4):627–644, April 2014. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2013-0313.

Todd:2014:SMD

[773] Christopher D. Todd, Bryce D. M. Whyte, Julian C. MacLean, Crawford W. Revie, Michael E. Lonergan, and Nora N. Hanson. A simple method of dating marine growth circuli on scales of wild one seawinter and two sea-winter Atlantic salmon (Salmo salar). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 71(5):645–655, May 2014. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2013-0359.

Kleiber:2014:IFE

[774] Danika Kleiber, Leila M. Harris, and Amanda C. J. Vincent. Improving fisheries estimates by including women’s catch in the Central Philippines. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 71(5):656–664, May 2014. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2013-0177.

Barth:2014:SCD

[775] Lauren Emily Barth, W. Gary Sprules, Mathew Wells, and Melissa Coman. Seasonal changes in the diel vertical migration of Chaoborus punctipennis larval instars. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 71 (5):665–674, May 2014. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2013-0440.

Miller:2014:RMD

[776] Scott W. Miller and Sarah Judson. Responses of macroinvertebrate drift, benthic assemblages, and trout foraging to hydropeaking. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 71(5):675–687, May 2014. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2013-0562.

Harbicht:2014:AHC

[777] Andrew B. Harbicht, Mohammed Alshamlih, Chris C. Wilson, and Dylan J. Fraser. Anthropogenic and habitat correlates of hybridization between hatchery and wild brook trout. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 71(5):688–697, May 2014. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/cjfas-2013-0460.

Larson:2014:SNP

[778] Wesley A. Larson, James E. Seeb, Carita E. Pascal, William D. Templin, and Lisa W. Seeb. Single-nucleotide polymorphisms (SNPs) identified through genotyping-by-sequencing improve genetic stock identification of Chinook salmon (Oncorhynchus tshawytscha) from western Alaska. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 71(5):698–708, May 2014. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2013-0502.

Bez:2014:IVR

[779] Nicolas Bez and Cheikh-Baye Braham. Indicator variables for a robust estimation of an acoustic index of abundance. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 71(5):709–718, May 2014. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/cjfas-2013-0437.

Anas:2014:ZCG

[780] M. U. Mohamed Anas, Kenneth A. Scott, Ryan N. Cooper, and Björn Wissel. Zooplankton communities are good indicators of potential impacts of athabasca oil sands operations on downwind boreal lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 71(5):719–732, May 2014. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2013-0472.

Ellis:2014:SET

[781] Nick Ellis, Francis Pantus, and C. Roland Pitcher. Scaling up experimental trawl impact results to fishery management scales — a modelling approach for a “hot time”. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 71 (5):733–746, May 2014. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2013-0426.

Arciszewski:2014:DRM

[782] T. J. Arciszewski, A. J. Farwell, M. R. Servos, T. D. Jardine, and K. R. Munkittrick. Differential recovery of δ ^{13}C in multiple tissues of white sucker across age classes after the closure of a pulp mill. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 71(5):747–755, May 2014. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2013-0643.

Elvidge:2014:PCI

[783] Chris K. Elvidge and Grant E. Brown. Predation costs of impaired chemosensory risk assessment on acid-impacted juvenile Atlantic salmon (Salmo salar). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 71(5):756– 762, May 2014. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
cjfas-2013-0633.

Roy:2014:ESP

[784] Denis Roy, David C. Hardie, Margaret A. Treble, James D. Reist, and Daniel E. Ruzzante. Evidence supporting panmixia in Greenland halibut (Reinhardtius hippoglossoides) in the Northwest Atlantic. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 71(5):763–774, May 2014. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2014-0004.

Ludsin:2014:PBC

[785] Stuart A. Ludsin, Kristen M. DeVanna, and Ralph E. H. Smith. Physical–biological coupling and the challenge of understanding fish recruitment in freshwater lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 71 (5):775–794, May 2014. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2013-0512.

Poisson:2014:MRS

[786] François Poisson, John David Filmalter, Anne-Lise Vernet, and Laurent Dagorn. Mortality rate of silky sharks (Carcharhinus falciformis) caught in the tropical tuna purse seine fishery in the Indian Ocean. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 71(6):795–798, June 2014. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2013-0561.

Kovach:2014:VEC

[787] Ryan P. Kovach, John E. Joyce, Scott C. Vulstek, Evan M. Barrientos, and David A. Tallmon. Variable effects of climate and density on the juvenile ecology of two salmonids in an Alaskan lake. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 71(6):799–807, June 2014. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2013-0577.

Tlusty:2014:MSD

[788] Michael F. Tlusty, Anita Kim, and Kathleen M. Castro. Modeling shell disease in American lobster (Homarus americanus) as individual-based health trajectories. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 71(6):808– 813, June 2014. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
cjfas-2013-0373.

Carney:2014:UMS

[789] Justin M. Carney and Milo D. Adkison. Using model simulations to compare performance of two commercial salmon management strategies in Bristol Bay, Alaska. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 71 (6):814–823, June 2014. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2013-0351.

DiRocco:2014:DAC

[790] Richard T. Di Rocco, Cowan F. Belanger, István Imre, Grant E. Brown, and Nicholas S. Johnson. Daytime avoidance of chemosensory alarm cues by adult sea lamprey (Petromyzon marinus). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 71(6):824–830, June 2014. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/cjfas-2013-0381.

Melnychuk:2014:MSD

[791] Michael C. Melnychuk, Josh Korman, Stephen Hausch, David W. Welch, Don J. F. McCubbing, and Carl J. Walters. Marine survival difference between wild and hatchery-reared steelhead trout determined during early downstream migration. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 71 (6):831–846, June 2014. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2013-0165.

Sagarese:2014:AGA

[792] Skyler R. Sagarese, Michael G. Frisk, Robert M. Cerrato, Kathy A. Sosebee, John A. Musick, and Paul J. Rago. Application of generalized additive models to examine ontogenetic and seasonal distributions of spiny dogfish (Squalus acanthias) in the northeast (US) shelf large marine ecosystem. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 71(6):847– 877, June 2014. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
cjfas-2013-0342.

Lyon:2014:EET

[793] Jarod P. Lyon, Tomas Bird, Simon Nicol, Joanne Kearns, Justin O’Mahony, Charles R. Todd, Ian G. Cowx, and Corey J. A. Bradshaw. Efficiency of electrofishing in turbid lowland rivers: implications for measuring temporal change in fish populations. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 71(6):878–886, June 2014. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/cjfas-2013-0287.

Tremblay:2014:RPL

[794] Roxane Tremblay, Reinhard Pienitz, and Pierre Legendre. Reconstructing phosphorus levels using models based on the modern diatom assemblages of 55 lakes in southern Quebec. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 71(6):887–914, June 2014. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/cjfas-2013-0469.

Homola:2014:WEM

[795] Jared J. Homola, Carl R. Ruetz III, Steven L. Kohler, and Ryan A. Thum. Weak effects of a microsporidian parasite on mottled sculpin in Michigan streams. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 71 (6):915–926, June 2014. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2013-0545.

Mittelbach:2014:FBT

[796] Gary G. Mittelbach, Nicholas G. Ballew, and Melissa K. Kjelvik. Fish behavioral types and their ecological consequences. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 71(6):927–944, June 2014. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2013-0558.

Johnson:2014:APG

[797] Matthew F. Johnson and Stephen P. Rice. Animal perception in gravelbed rivers: scales of sensing and environmental controls on sensory information. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 71(6):945–957, June 2014. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
cjfas-2013-0474.

Guill:2014:ECD

[798] Christian Guill, Eddy Carmack, and Barbara Drossel. Exploring cyclic dominance of sockeye salmon with a predator–prey model. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 71(7):959–972, July 2014. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2013-0441.

Thorson:2014:HVR

[799] James T. Thorson, Olaf P. Jensen, and Elise F. Zipkin. How variable is recruitment for exploited marine fishes? A hierarchical model for testing life history theory. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 71 (7):973–983, July 2014. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2013-0645. See corrigendum [749].

Meynier:2014:FDL

[800] Laureline Meynier, Patrick C. H. Morel, B. Louise Chilvers, Duncan D. S. Mackenzie, and Pádraig J. Duignan. Foraging diversity in lactating New Zealand sea lions: insights from qualitative and quantitative fatty acid analysis. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 71 (7):984–991, July 2014. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2013-0479.

Wittmann:2014:GCG

[801] Marion E. Wittmann, Christopher L. Jerde, Jennifer G. Howeth, Sean P. Maher, Andrew M. Deines, Jill A. Jenkins, Gregory W. Whitledge, Sarah R. Burbank, William L. Chadderton, Andrew R. Mahon, Jeffrey T. Tyson, Crysta A. Gantz, Reuben P. Keller, John M. Drake, and David M. Lodge. Grass carp in the Great Lakes region: establishment potential, expert perceptions, and re-evaluation of experimental evidence of ecological impact. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 71 (7):992–999, July 2014. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2013-0537.

Whitcomb:2014:IIR

[802] Amelia C. Whitcomb, Michael A. Banks, and Kathleen G. O’Malley. Influence of immune-relevant genes on mate choice and reproductive success in wild-spawning hatchery-reared and wild-born Coho salmon (Oncorhynchus kisutch). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 71 (7):1000–1009, July 2014. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2013-0501.

Kanno:2014:SVA

[803] Yoichiro Kanno, Benjamin H. Letcher, Jason C. Vokoun, and Elise F. Zipkin. Spatial variability in adult brook trout (Salvelinus fontinalis) survival within two intensively surveyed headwater stream networks. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 71(7):1010–1019, July 2014. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2013-0358.

Harvey:2014:MFP

[804] Chris J. Harvey. Mediation functions in Ecopath with Ecosim: handle with care. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 71(7):1020– 1029, July 2014. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
cjfas-2013-0594.

Poletto:2014:JGS

[805] Jamilynn B. Poletto, Dennis E. Cocherell, Natalie Ho, Joseph J. Cech, A. Peter Klimley, and Nann A. Fangue. Juvenile green sturgeon (Acipenser medirostris) and white sturgeon (Acipenser transmontanus) behavior near water-diversion fish screens: experiments in a laboratory swimming flume. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 71(7):1030– 1038, July 2014. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
cjfas-2013-0556.

McDonnell:2014:EEP

[806] Kevin N. McDonnell and Brian M. Roth. Evaluating the effect of pelagic zooplankton community composition and density on larval walleye (Sander vitreus) growth with a bioenergetic-based foraging model. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 71(7):1039–1049, July 2014. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2013-0498.

Loxterman:2014:EIS

[807] Janet L. Loxterman, Ernest R. Keeley, and Zacharia M. Njoroge. Evaluating the influence of stocking history and barriers to movement on the spatial extent of hybridization between westslope cutthroat trout and rainbow trout. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 71 (7):1050–1058, July 2014. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2013-0424.

Bowerman:2014:EFS

[808] Tracy Bowerman, Bethany T. Neilson, and Phaedra Budy. Effects of fine sediment, hyporheic flow, and spawning site characteristics on survival and development of bull trout embryos. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 71(7):1059–1071, July 2014. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/cjfas-2013-0372.

Rogers:2014:LMO

[809] Mark W. Rogers, David B. Bunnell, Charles P. Madenjian, and David M. Warner. Lake Michigan offshore ecosystem structure and food web changes from 1987 to 2008. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 71 (7):1072–1086, July 2014. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2013-0514.

Kring:2014:RSM

[810] Stefanie A. Kring, Stephanie E. Figary, Gregory L. Boyer, Susan B. Watson, and Michael R. Twiss. Rapid in situ measures of phytoplankton communities using the bbe FluoroProbe: evaluation of spectral calibration, instrument intercompatibility, and performance range. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 71(7):1087–1095, July 2014. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2013-0599.

Penney:2014:CRS

[811] Chantelle M. Penney, Gordon W. Nash, and A. Kurt Gamperl. Cardiorespiratory responses of seawater-acclimated adult Arctic char (Salvelinus alpinus) and Atlantic salmon (Salmo salar) to an acute temperature increase. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 71(7):1096–1105, July 2014. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
cjfas-2013-0569.

Folkvord:2014:TOB

[812] Arild Folkvord, Christian Jørgensen, Knut Korsbrekke, Richard D. M. Nash, Trygve Nilsen, and Jon Egil Skjæraasen. Trade-offs between growth and reproduction in wild Atlantic cod. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 71(7):1106–1112, July 2014. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2013-0600.

Calduch-Verdiell:2014:LHE

[813] Núria Calduch-Verdiell, Brian R. MacKenzie, James W. Vaupel, and Ken H. Andersen. A life-history evaluation of the impact of maternal effects on recruitment and fisheries reference points. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 71(7):1113–1120, July 2014. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2014-0034.

Swain:2014:CES

[814] Noel R. Swain, Morgan D. Hocking, Jennifer N. Harding, and John D. Reynolds. Corrigendum: Effects of salmon on the diet and condition of stream-resident sculpins. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 71(7): 1121, July 2014. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
cjfas-2014-0179. See [763].

Thompson:2014:ETP

[815] Kevin A. Thompson, Selina S. Heppell, and Grant G. Thompson. The effects of temperature and predator densities on the consumption of walleye pollock (Gadus chalcogrammus) by three groundfish in the Gulf of Alaska. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 71(8):1123–1133, August 2014. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
cjfas-2013-0260.

Whitney:2014:FAS

[816] James E. Whitney, Keith B. Gido, and David L. Propst. Factors associated with the success of native and nonnative species in an unfragmented arid-land riverscape. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 71(8): 1134–1145, August 2014. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2014-0153.

Strayer:2014:HRF

[817] David L. Strayer, Kathryn A. Hattala, Andrew W. Kahnle, and Robert D. Adams. Has the Hudson River fish community recovered from the zebra mussel invasion along with its forage base? Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 71(8):1146–1157, August 2014. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2013-0549.

Weber:2014:ESH

[818] Nicholas Weber, Nicolaas Bouwes, and Chris E. Jordan. Estimation of salmonid habitat growth potential through measurements of invertebrate food abundance and temperature. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 71(8):1158–1170, August 2014. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/cjfas-2013-0390.

Dolinsek:2014:DNH

[819] Ivan J. Dolinsek, Robert L. McLaughlin, James W. A. Grant, Lisa M. O’Connor, and Thomas C. Pratt. Do natural history data predict the movement ecology of fishes in Lake Ontario streams? Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 71(8):1171–1185, August 2014. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2013-0294.

Pena:2014:ADM

[820] Marian Peña, M. Pilar Olivar, Rosa Balbín, Jose Luis López-Jurado, Magdalena Iglesias, and Joan Miquel. Acoustic detection of mesopelagic fishes in scattering layers of the Balearic Sea (western Mediterranean). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 71(8):1186–1197, August 2014. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2013-0331.

Braun:2014:LHE

[821] Douglas C. Braun and John D. Reynolds. Life history and environmental influences on population dynamics in sockeye salmon. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 71(8):1198–1208, August 2014. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2013-0326.

Mussen:2014:CBF

[822] Timothy D. Mussen, Oliver Patton, Dennis Cocherell, Ali Ercan, Hossein Bandeh, M. Levent Kavvas, Joseph J. Cech, and Nann A. Fangue. Can behavioral fish-guidance devices protect juvenile Chinook salmon (Oncorhynchus tshawytscha) from entrainment into unscreened waterdiversion pipes? Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 71(8):1209– 1219, August 2014. CODEN CJFSDX. ISSN 0706-652X (print), 1205-

7533 (electronic). URL https://cdnsciencepub.com/doi/full/10.
1139/cjfas-2013-0601.

Chambers:2014:SBT

[823] Mark S. Chambers, Leesa A. Sidhu, and Ben O’Neill. Southern bluefin tuna (Thunnus maccoyii) shed tags at a higher rate in tuna farms than in the open ocean — two-stage tag retention models. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 71(8):1220–1228, August 2014. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2013-0325.

Conn:2014:EUF

[824] Paul B. Conn, Devin S. Johnson, Lowell W. Fritz, and Brian S. Fadely. Examining the utility of fishery and survey data to detect prey removal effects on Steller sea lions (Eumetopias jubatus). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 71(8):1229–1242, August 2014. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2013-0602.

Fernandez-Cebrian:2014:GRS

[825] Raquel Fernández-Cebrián, Rosa Maria Araguas, Nuria Sanz, and Jose Luis García-Marín. Genetic risks of supplementing trout populations with native stocks: a simulation case study from current Pyrenean populations. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 71(8):1243–1255, August 2014. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
cjfas-2013-0362.

Muir:2014:RLH

[826] Andrew M. Muir, Michael T. Arts, Marten A. Koops, Timothy B. Johnson, Charles C. Krueger, and Trent M. Sutton. Reproductive life-history strategies in lake whitefish (Coregonus clupeaformis) from the Laurentian Great Lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 71(8):1256– 1269, August 2014. CODEN CJFSDX. ISSN 0706-652X (print), 1205-

7533 (electronic). URL https://cdnsciencepub.com/doi/full/10.
1139/cjfas-2013-0254.

Railsback:2014:FAS

[827] Steven F. Railsback, Bret C. Harvey, and Jason L. White. Facultative anadromy in salmonids: linking habitat, individual life history decisions, and population-level consequences. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 71(8):1270–1278, August 2014. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/cjfas-2014-0091.

Winton:2014:ISV

[828] Megan V. Winton, Mark J. Wuenschel, and Richard S. McBride. Investigating spatial variation and temperature effects on maturity of female winter flounder (Pseudopleuronectes americanus) using generalized additive models. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 71(9):1279– 1290, September 2014. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2013-0617.

Garnett:2014:LVC

[829] Jeffrey A. Garnett and Darold P. Batzer. Longitudinal variation in community structure of floodplain fishes along two rivers of the southeastern USA. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 71(9):1291–1302, September 2014. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
cjfas-2013-0422.

Chezik:2014:FGDb

[830] Kyle A. Chezik, Nigel P. Lester, and Paul A. Venturelli. Fish growth and degree-days II: selecting a base temperature for an among-population study. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 71(9):1303–1311, September 2014. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
cjfas-2013-0615.

Longmore:2014:OSC

[831] Craig Longmore, Clive N. Trueman, Francis Neat, Per Erik Jorde, Halvor Knutsen, Sergio Stefanni, Diana Catarino, James A. Milton, and Stefano Mariani. Ocean-scale connectivity and life cycle reconstruction in a deep-sea fish. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 71(9):1312– 1323, September 2014. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2013-0343.

Yurista:2014:NLL

[832] Peder M. Yurista, Daniel L. Yule, Matt Balge, Jon D. VanAlstine, Jo A. Thompson, Allison E. Gamble, Tom R. Hrabik, John R. Kelly, Jason D. Stockwell, and Mark R. Vinson. A new look at the Lake Superior biomass size spectrum. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 71(9):1324– 1333, September 2014. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2013-0596.

Herb:2014:PCW

[833] William R. Herb, Lucinda B. Johnson, Peter C. Jacobson, and Heinz G. Stefan. Projecting cold-water fish habitat in lakes of the glacial lakes region under changing land use and climate regimes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 71(9):1334–1348, September 2014. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2013-0535.

Richardson:2014:IFF

[834] David E. Richardson, Michael C. Palmer, and Brian E. Smith. The influence of forage fish abundance on the aggregation of Gulf of Maine Atlantic cod (Gadus morhua) and their catchability in the fishery. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 71(9):1349–1362, September 2014. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2013-0489.

Chang:2014:DHR

[835] Shui-Kai Chang and Tzu-Lun Yuan. Deriving high-resolution spatiotemporal fishing effort of large-scale longline fishery from vessel monitoring system (VMS) data and validated by observer data. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 71(9):1363–1370, September 2014. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2013-0552.

Debes:2014:EDP

[836] Paul V. Debes and Jeffrey A. Hutchings. Effects of domestication on parr maturity, growth, and vulnerability to predation in Atlantic salmon. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 71(9):1371–1384, September 2014. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
cjfas-2013-0618.

Lloyd-Jones:2014:IES

[837] Luke R. Lloyd-Jones, Hien D. Nguyen, You-Gan Wang, and Michael F. O’Neill. Improved estimation of size-transition matrices using tag– recapture data. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 71(9):1385– 1394, September 2014. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2014-0080.

Rosenfeld:2014:MEH

[838] Jordan S. Rosenfeld. Modelling the effects of habitat on self-thinning, energy equivalence, and optimal habitat structure for juvenile trout. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 71(9):1395–1406, September 2014. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2013-0603.

Rudd:2014:ESE

[839] Merrill B. Rudd, Robert N. M. Ahrens, William E. Pine, and Stephania K. Bolden. Empirical, spatially explicit natural mortality and movement rate estimates for the threatened Gulf sturgeon (Acipenser oxyrinchus desotoi). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 71(9): 1407–1417, September 2014. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2014-0010.

Parkos:2014:OFC

[840] Joseph J. Parkos and Joel C. Trexler. Origins of functional connectivity in a human-modified wetland landscape. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 71(9):1418–1429, September 2014. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/cjfas-2013-0553.

Wiper:2014:EER

[841] Mallory L. Wiper, Stephanie Britton, and Dennis M. Higgs. Early experience and reproductive morph both affect brain morphology in adult male Chinook salmon (Oncorhynchus tshawytscha). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 71(9):1430–1436, September 2014. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2013-0624.

Stocks:2014:SMA

[842] Allison P. Stocks, Evgeny A. Pakhomov, and Brian P. V. Hunt. A simple method to assess the marine environment residence duration of juvenile sockeye salmon (Oncorhynchus nerka) using laser ablation. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 71(10):1437–1446, October 2014. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2014-0073.

Bar:2014:PHC

[843] Nadav Bar. Physiological and hormonal changes during prolonged starvation in fish. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 71(10):1447–1458, October 2014. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
cjfas-2013-0175.

Freon:2014:EEC

[844] Pierre Fréon, Angel Avadí, Wilbert Marin Soto, and Richard Negrón. Environmentally extended comparison table of large- versus small- and medium-scale fisheries: the case of the Peruvian anchoveta fleet. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 71(10):1459–1474, October 2014. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2013-0542.

Hasegawa:2014:EHC

[845] Koh Hasegawa, Kentaro Morita, Kazumasa Ohkuma, Tsutomu Ohnuki, and Yasutaka Okamoto. Effects of hatchery chum salmon fry on densitydependent intra- and interspecific competition between wild chum and masu salmon fry. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 71(10):1475– 1482, October 2014. CODEN CJFSDX. ISSN 0706-652X (print), 1205-

7533 (electronic). URL https://cdnsciencepub.com/doi/full/10.
1139/cjfas-2014-0167.

Hegeman:2014:MFM

[846] Ericka E. Hegeman, Scott W. Miller, and Karen E. Mock. Modeling freshwater mussel distribution in relation to biotic and abiotic habitat variables at multiple spatial scales. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 71(10):1483–1497, October 2014. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/cjfas-2014-0110.

Clark:2014:BSS

[847] Steven M. Clark, Jason B. Dunham, Jeffrey R. McEnroe, and Scott W. Lightcap. Breeding site selection by Coho salmon (Oncorhynchus kisutch) in relation to large wood additions and factors that influence reproductive success. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 71(10): 1498–1507, October 2014. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2014-0020.

Huhn:2014:NRD

[848] Daniel Hühn, Kay Lübke, Christian Skov, and Robert Arlinghaus. Natural recruitment, density-dependent juvenile survival, and the potential for additive effects of stock enhancement: an experimental evaluation of stocking northern pike (Esox lucius) fry. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 71(10):1508–1519, October 2014. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2013-0636.

Gamboa-Delgado:2014:ASI

[849] Julián Gamboa-Delgado, César Molina-Poveda, Daniel Enrique Godínez-Siordia, David Villarreal-Cavazos, Denis Ricque-Marie, and Lucía Elizabeth Cruz-Suárez. Application of stable isotope analysis to differentiate shrimp extracted by industrial fishing or produced through aquaculture practices. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 71(10):1520–1528, October 2014. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
cjfas-2014-0005.

Lam:2014:MOA

[850] Chi Hin Lam, Benjamin Galuardi, and Molly E. Lutcavage. Movements and oceanographic associations of bigeye tuna (Thunnus obesus) in the Northwest Atlantic. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 71(10): 1529–1543, October 2014. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2013-0511.

Manning:2014:MTT

[851] Nathan F. Manning, Jonathan M. Bossenbroek, Christine M. Mayer, David B. Bunnell, Jeff T. Tyson, Lars G. Rudstam, and James R.

Jackson. Modeling turbidity type and intensity effects on the growth and starvation mortality of age-0 yellow perch. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 71(10):1544–1553, October 2014. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2013-0528.

Hunt:2014:CFS

[852] Taylor L. Hunt, Khageswor Giri, Paul Brown, Brett A. Ingram, Paul L. Jones, Laurie J. B. Laurenson, and Anne M. Wallis. Consequences of fish stocking density in a recreational fishery. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 71(10):1554–1560, October 2014. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2013-0517.

Drouineau:2014:EEE

[853] Hilaire Drouineau, Christian Rigaud, Françoise Daverat, and Patrick Lambert. EvEel (evolutionary ecology-based model for eel): a model to explore the role of phenotypic plasticity as an adaptive response of three temperate eels to spatially structured environments. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 71(10):1561–1571, October 2014. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2014-0090.

Kirchhoff:2014:IAI

[854] Nicole T. Kirchhoff, Justin Nelligan, David Ellis, Karine Cadoret, Melanie Leef, and Barbara Nowak. Inter-annual and intra-annual variability in blood variables and parasitic loads of wild Thunnus maccoyii. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 71(10):1572–1578, October 2014. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2013-0457.

Beguer-Pon:2014:LSM

[855] Mélanie Béguer-Pon, Martin Castonguay, José Benchetrit, Daniel Hatin, Guy Verreault, Yves Mailhot, Valérie Tremblay, Denis Lefaivre, Michel Legault, David Stanley, and Julian J. Dodson. Large-scale migration patterns of silver American eels from the St. Lawrence River to the Gulf of St. Lawrence using acoustic telemetry. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 71(10):1579–1592, October 2014. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2013-0217.

Cooke:2014:WWM

[856] S. J. Cooke, R. Arlinghaus, D. M. Bartley, T. D. Beard, I. G. Cowx, T. E. Essington, O. P. Jensen, A. Lynch, W. W. Taylor, and R. Watson. Where the waters meet: sharing ideas and experiences between inland and marine realms to promote sustainable fisheries management. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 71(10):1593–1601, October 2014. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2014-0176.

Bylak:2014:BIS

[857] Aneta Bylak, Krzysztof Kukuła, and Józef Mitka. Beaver impact on stream fish life histories: the role of landscape and local attributes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 71(11):1603–1615, November 2014. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2014-0105.

Jones:2014:CDN

[858] Nicholas E. Jones, Bastian J. Schmidt, and Stephanie J. Melles. Characteristics and distribution of natural flow regimes in Canada: a habitat template approach. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 71(11): 1616–1624, November 2014. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2014-0040.

Avery-Gomm:2014:HDR

[859] Stephanie Avery-Gomm, Jordan S. Rosenfeld, John S. Richardson, and Michael Pearson. Hydrological drought and the role of refugia in an endangered riffle-dwelling fish, nooksack dace (Rhinichthys cataractae ssp.). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 71(11):1625–1634, November 2014. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
cjfas-2013-0585.

Grasset:2014:TOD

[860] Julie Grasset, Bérénice Bougas, Peter G. C. Campbell, Louis Bernatchez, and Patrice Couture. Temperature, oxygen, and diet modulate gene transcription and metabolic capacities in yellow perch. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 71(11):1635–1641, November 2014. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2014-0059.

Millie:2014:UAI

[861] David F. Millie, Gary R. Weckman, Gary L. Fahnenstiel, Hunter J. Carrick, Ehsan Ardjmand, William A. Young, Michael J. Sayers, and Robert A. Shuchman. Using artificial intelligence for CyanoHAB niche modeling: discovery and visualization of Microcystis — environmental associations within western Lake Erie. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 71(11):1642–1654, November 2014. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/cjfas-2013-0654.

Shelton:2014:SSM

[862] Andrew Olaf Shelton, James T. Thorson, Eric J. Ward, and Blake E. Feist. Spatial semiparametric models improve estimates of species abundance and distribution. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 71(11): 1655–1666, November 2014. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2013-0508.

Whiteley:2014:SEA

[863] Andrew R. Whiteley, Jason A. Coombs, Benjamin H. Letcher, and Keith H. Nislow. Simulation and empirical analysis of novel sibship-based genetic determination of fish passage. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 71(11):1667–1679, November 2014. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/cjfas-2014-0137.

Neville:2014:GMT

[864] Helen M. Neville and Douglas P. Peterson. Genetic monitoring of trout movement after culvert remediation: family matters. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 71(11):1680–1694, November 2014. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2014-0138.

Pinder:2014:IWD

[865] Kieran C. Pinder, M. Catherine Eimers, and Shaun A. Watmough. Impact of wetland disturbance on phosphorus loadings to lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 71(11):1695–1703, November 2014. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2014-0143.

Morris:2014:TSM

[866] Corey J. Morris, John M. Green, Paul V. R. Snelgrove, Curtis J. Pennell, and L. M. Neil Ollerhead. Temporal and spatial migration of Atlantic cod (Gadus morhua) inside and outside a marine protected area and evidence for the role of prior experience in homing. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 71(11):1704–1712, November 2014. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2014-0036.

Holbrook:2014:ERS

[867] Christopher M. Holbrook, Nicholas S. Johnson, Juan P. Steibel, Michael B. Twohey, Thomas R. Binder, Charles C. Krueger, and Michael L. Jones. Estimating reach-specific fish movement probabilities in rivers with a Bayesian state-space model: application to sea lamprey passage and capture at dams. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 71(11):1713–1729, November 2014. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/cjfas-2013-0581.

Richard:2014:TRH

[868] Antoine Richard, Louis Bernatchez, Eliane Valiquette, and Mélanie Dionne. Telemetry reveals how catch and release affects prespawning migration in Atlantic salmon (Salmo salar). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 71(11):1730–1739, November 2014. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2014-0072.

Jensen:2014:OLH

[869] Arne J. Jensen, Sten Karlsson, Peder Fiske, Lars Petter Hansen, Gunnel M. Østborg, and Kjetil Hindar. Origin and life history of Atlantic salmon (Salmo salar) near their northernmost oceanic limit. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 71(11):1740–1746, November 2014. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2014-0169.

Losee:2014:IOE

[870] James P. Losee, Jessica A. Miller, William T. Peterson, David J. Teel, and Kym C. Jacobson. Influence of ocean ecosystem variation on trophic interactions and survival of juvenile Coho and Chinook salmon. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 71(11):1747–1757, November 2014. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2014-0043.

LeMoine:2014:BUP

[871] Michael Thomas LeMoine and Leo Richard Bodensteiner. Barriers to upstream passage by two migratory sculpins, prickly sculpin (Cottus asper) and coastrange sculpin (Cottus aleuticus), in northern Puget Sound lowland streams. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 71(11):1758– 1765, November 2014. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2014-0029.

Castaneda:2014:MPS

[872] Rowshyra A. Castañeda, Suncica Avlijas, M. Anouk Simard, and Anthony Ricciardi. Microplastic pollution in St. Lawrence River sediments. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 71(12):1767–1771, December 2014. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
cjfas-2014-0281.

Yamamoto:2014:ARA

[873] Shoichiro Yamamoto, Kouji Mutou, Hidefumi Nakamura, Kouta Miyamoto, Kazuo Uchida, Kaori Takagi, Ken Fujimoto, Hideki Kaeriyama, and Tsuneo Ono. Assessment of radiocaesium accumulation by hatchery-reared salmonids after the Fukushima nuclear accident. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 71(12):1772–1775, December 2014. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
cjfas-2014-0170.

Jyvasjarvi:2014:SCC

[874] Jussi Jyväsjärvi, Marko Järvinen, and Heikki Hämäläinen. Spatial community concordance of summer phytoplankton and profundal macroinvertebrates in boreal lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 71 (12):1776–1783, December 2014. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/

doi/full/10.1139/cjfas-2014-0135.

Foley:2014:SIP

[875] Carolyn J. Foley, Gabriel J. Bowen, Thomas F. Nalepa, Marisol S. Sepúlveda, and Tomas O. Höök. Stable isotope patterns of benthic organisms from the Great Lakes region indicate variable dietary overlap of Diporeia spp. and dreissenid mussels. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 71(12):1784–1795, December 2014. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/cjfas-2013-0620.

Ebel:2014:BNL

[876] Jonathan D. Ebel, Amy M. Marcarelli, and André E. Kohler. Biofilm nutrient limitation, metabolism, and standing crop responses to experimental application of salmon carcass analog in Idaho streams. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 71(12):1796–1804, December 2014. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2014-0266.

Favaro:2014:PLR

[877] Corinna Favaro, Jonathan W. Moore, John D. Reynolds, and Michael P. Beakes. Potential loss and rehabilitation of stream longitudinal connectivity: fish populations in urban streams with culverts. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 71(12):1805–1816, December 2014. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2014-0199.

Charles:2014:UHM

[878] Colin Charles, Darren Gillis, and Elmer Wade. Using hidden Markov models to infer vessel activities in the snow crab (Chionoecetes opilio) fixed gear fishery and their application to catch standardization. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 71(12):1817–1829, December 2014. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2013-0572.

Beaulieu:2014:CPC

[879] Marieke Beaulieu, Frances Pick, Michelle Palmer, Sue Watson, Jenny Winter, Ron Zurawell, and Irene Gregory-Eaves. Comparing predictive cyanobacterial models from temperate regions. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 71(12):1830–1839, December 2014. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2014-0168.

Majarreis:2014:NSA

[880] Joanna M. Majarreis, Sue B. Watson, and Ralph E. H. Smith. Nutrient status and its assessment by pulse amplitude modulated (PAM) fluorometry of phytoplankton at sites in Lakes Erie and Ontario. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 71(12):1840–1851, December 2014. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2014-0017.

Mahlum:2014:ABR

[881] Shad Mahlum, Dan Kehler, David Cote, Yolanda F. Wiersma, and Les Stanfield. Assessing the biological relevance of aquatic connectivity to stream fish communities. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 71(12): 1852–1863, December 2014. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2013-0646.

Anlauf-Dunn:2014:HCC

[882] Kara J. Anlauf-Dunn, Eric J. Ward, Matt Strickland, and Kim Jones. Habitat connectivity, complexity, and quality: predicting adult Coho salmon occupancy and abundance. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 71(12):1864–1876, December 2014. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/cjfas-2014-0162.

Osterback:2014:LIS

[883] Ann-Marie K. Osterback, Danielle M. Frechette, Sean A. Hayes, Morgan H. Bond, Scott A. Shaffer, and Jonathan W. Moore. Linking individual size and wild and hatchery ancestry to survival and predation risk of threatened steelhead (Oncorhynchus mykiss). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 71(12):1877–1887, December 2014. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2014-0097.

Tonkin:2014:RRE

[884] Zeb Tonkin, Jarod Lyon, David S. L. Ramsey, Nick R. Bond, Graeme Hackett, Kyne Krusic-Golub, Brett A. Ingram, and Stephen R. Balcombe. Reservoir refilling enhances growth and recruitment of an endangered remnant riverine fish. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 71(12):1888–1899, December 2014. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/cjfas-2014-0081.

Harkonen:2014:EBI

[885] Laura Härkönen, Pekka Hyvärinen, Juuso Paappanen, and Anssi Vainikka. Explorative behavior increases vulnerability to angling in hatchery-reared brown trout (Salmo trutta). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 71(12):1900–1909, December 2014. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2014-0221.

vanEsch:2014:VMP

[886] B. P. M. van Esch and I. L. Y. Spierts. Validation of a model to predict fish passage mortality in pumping stations. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 71(12):1910–1923, December 2014. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2014-0035.

Hardy:2014:SOL

[887] Robert F. Hardy, Anton D. Tucker, Allen M. Foley, Barbara A. Schroeder, Robin J. Giove, and Anne B. Meylan. Spatiotemporal occurrence of loggerhead turtles (Caretta caretta) on the West Florida Shelf and apparent overlap with a commercial fishery. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 71(12):1924–1933, December 2014. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2014-0128.

Andrew:2015:UIW

[888] Ross G. Andrew and Kyle J. Hartman. Uneven inputs of woody debris to Appalachian streams from superstorm Sandy. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 72(1):1–6, January 2015. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/cjfas-2014-0208.

He:2015:CAS

[889] Ji X. He, James R. Bence, Charles P. Madenjian, Steven A. Pothoven, Norine E. Dobiesz, David G. Fielder, James E. Johnson, Mark P. Ebener, R. Adam Cottrill, Lloyd C. Mohr, and Scott R. Koproski. Coupling agestructured stock assessment and fish bioenergetics models: a system of time-varying models for quantifying piscivory patterns during the rapid trophic shift in the main basin of Lake Huron. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 72(1):7–23, January 2015. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/cjfas-2014-0161. See comment [1139] and re-

ply [1140].

Matabos:2015:VCO

[890] Marjolaine Matabos, Nils Piechaud, François De Montigny, Pierre-Marie Sarradin, and Jozée Sarrazin. The VENUS cabled observatory as a method to observe fish behaviour and species assemblages in a hypoxic fjord, Saanich Inlet (British Columbia, Canada). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 72(1):24–36, January 2015. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2013-0611.

Johnston:2015:OMR

[891] Fiona D. Johnston, Ben Beardmore, and Robert Arlinghaus. Optimal management of recreational fisheries in the presence of hooking mortality and noncompliance — predictions from a bioeconomic model incorporating a mechanistic model of angler behavior. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 72(1):37–53, January 2015. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/cjfas-2013-0650.

Burgess:2015:CFD

[892] Matthew G. Burgess. Consequences of fleet diversification in managed and unmanaged fisheries. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 72 (1):54–70, January 2015. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2014-0116.

Hamel:2015:RWA

[893] Martin J. Hamel, Mark A. Pegg, Reuben R. Goforth, Quinton E. Phelps, Kirk D. Steffensen, Jeremy J. Hammen, and Mathew L. Rugg. Rangewide age and growth characteristics of shovelnose sturgeon from mark– recapture data: implications for conservation and management. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 72(1):71–82, January 2015. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2014-0238.

Bradbury:2015:GEL

[894] Ian R. Bradbury, Lorraine C. Hamilton, Sara Rafferty, David Meerburg, Rebecca Poole, J. Brian Dempson, Martha J. Robertson, David G. Reddin, Vincent Bourret, Mélanie Dionne, Gerald Chaput, Timothy F. Sheehan, Timothy L. King, John R. Candy, and Louis Bernatchez. Genetic evidence of local exploitation of Atlantic salmon in a coastal subsistence fishery in the Northwest Atlantic. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 72(1):83–95, January 2015. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/cjfas-2014-0058.

Gordeeva:2015:PED

[895] Natalia V. Gordeeva, Sergey S. Alekseyev, Arkadii N. Matveev, and Vitalii P. Samusenok. Parallel evolutionary divergence in Arctic char Salvelinus alpinus complex from transbaikalia: variation in differentiation degree and segregation of genetic diversity among sympatric forms. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 72(1):96–115, January 2015. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2014-0014.

Mehner:2015:PDV

[896] Thomas Mehner. Partial diel vertical migration of sympatric vendace (Coregonus albula) and Fontane cisco (Coregonus fontanae) is driven by density dependence. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 72(1): 116–124, January 2015. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2014-0009.

Scott:2015:PPW

[897] Mark A. Scott, Rashpal S. Dhillon, Patricia M. Schulte, and Jeffrey G. Richards. Physiology and performance of wild and domestic strains of diploid and triploid rainbow trout (Oncorhynchus mykiss) in response to environmental challenges. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 72 (1):125–134, January 2015. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2013-0450.

Chu:2015:UAH

[898] Cindy Chu, Charles K. Minns, Nigel P. Lester, and Nicholas E. Mandrak. An updated assessment of human activities, the environment, and freshwater fish biodiversity in Canada. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 72(1):135–148, January 2015. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/cjfas-2013-0609.

Wurtsbaugh:2015:ASF

[899] Wayne A. Wurtsbaugh, Nicholas A. Heredia, Brian G. Laub, Christy S. Meredith, Harrison E. Mohn, Sarah E. Null, David A. Pluth, Brett B. Roper, W. Carl Saunders, David K. Stevens, Richard H. Walker, and Kit Wheeler. Approaches for studying fish production: Do river and lake researchers have different perspectives? Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 72(1):149–160, January 2015. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/cjfas-2014-0210.

Aalto:2015:SRM

[900] Emilius A. Aalto, E. J. Dick, and Alec D. MacCall. Separating recruitment and mortality time lags for a delay-difference production model. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 72(2):161–165, February 2015. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2013-0415.

Polivka:2015:JSS

[901] Karl M. Polivka, E. Ashley Steel, and Jenni L. Novak. Juvenile salmon and steelhead occupancy of stream pools treated and not treated with restoration structures, Entiat River, Washington. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 72(2):166–174, February 2015. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2014-0228.

Sweeney:2015:NAC

[902] Kathryn L. Sweeney, Kyle W. Shertzer, Lowell W. Fritz, and Andrew J. Read. A novel approach to compare pinniped populations across a broad geographic range. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 72(2):175–185, February 2015. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
cjfas-2014-0070.

Cosandey-Godin:2015:ABS

[903] Aurelie Cosandey-Godin, Elias Teixeira Krainski, Boris Worm, and Joanna Mills Flemming. Applying Bayesian spatiotemporal models to fisheries bycatch in the Canadian Arctic. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 72(2):186–197, February 2015. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2014-0159.

Segura:2015:GEL

[904] Angel Manuel Segura, Valentina Franco-Trecu, Paula Franco-Fraguas, and Matías Arim. Gape and energy limitation determine a humped relationship between trophic position and body size. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 72(2):198–205, February 2015. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2014-0093.

Ferreira:2015:EEW

[905] Verónica Ferreira, Eric Chauvet, and Cristina Canhoto. Effects of experimental warming, litter species, and presence of macroinvertebrates on litter decomposition and associated decomposers in a temperate mountain stream. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 72(2):206–216, February 2015. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
cjfas-2014-0119.

Alos:2015:RAI

[906] Josep Alós, Miquel Palmer, Pedro Trías, Carlos Díaz-Gil, and Robert Arlinghaus. Recreational angling intensity correlates with alteration of vulnerability to fishing in a carnivorous coastal fish species. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 72(2):217–225, February 2015. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2014-0183.

Samarasin:2015:FDB

[907] Pasan Samarasin, Charles K. Minns, Brian J. Shuter, William M. Tonn, and Michael D. Rennie. Fish diversity and biomass in northern Canadian lakes: northern lakes are more diverse and have greater biomass than expected based on species–energy theory. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 72(2):226–237, February 2015. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2014-0104. See cor-

rigendum [1057].

Hanselman:2015:MIL

[908] Dana H. Hanselman, Jonathan Heifetz, Katy B. Echave, and Sherri C. Dressel. Move it or lose it: movement and mortality of sablefish tagged in Alaska. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 72(2):238–251, February 2015. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
cjfas-2014-0251.

Ferter:2015:CRA

[909] Keno Ferter, Klaas Hartmann, Alf Ring Kleiven, Even Moland, and Esben Moland Olsen. Catch-and-release of Atlantic cod (Gadus morhua): post-release behaviour of acoustically pretagged fish in a natural marine environment. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 72(2):252–261, February 2015. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
cjfas-2014-0290.

McGilliard:2015:AMR

[910] Carey R. McGilliard, André E. Punt, Richard D. Methot, Jr., and Ray Hilborn. Accounting for marine reserves using spatial stock assessments. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 72(2):262–280, February 2015. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2013-0364.

Schaefer:2015:LSV

[911] Wayne F. Schaefer, Mark H. Schmitz, Vicki S. Blazer, Timothy J. Ehlinger, and John A. Berges. Localization and seasonal variation of blue pigment (sandercyanin) in walleye (Sander vitreus). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 72(2):281–289, February 2015. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2014-0139.

Boyd:2015:BPP

[912] Charlotte Boyd, Mathieu Woillez, Sophie Bertrand, Ramiro Castillo, Arnaud Bertrand, and André E. Punt. Bayesian posterior prediction of the patchy spatial distributions of small pelagic fish in regions of suitable habitat. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 72(2):290–303, February 2015. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
cjfas-2014-0234.

Falke:2015:CCV

[913] Jeffrey A. Falke, Rebecca L. Flitcroft, Jason B. Dunham, Kristina M. Mc-Nyset, Paul F. Hessburg, and Gordon H. Reeves. Climate change and vulnerability of bull trout (Salvelinus confluentus) in a fire-prone landscape. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 72(2):304–318, February 2015. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2014-0098.

Kendall:2015:ARS

[914] Neala W. Kendall, John R. McMillan, Matthew R. Sloat, Thomas W. Buehrens, Thomas P. Quinn, George R. Pess, Kirill V. Kuzishchin, Michelle M. McClure, and Richard W. Zabel. Anadromy and residency in steelhead and rainbow trout (Oncorhynchus mykiss): a review of the processes and patterns. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 72 (3):319–342, March 2015. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2014-0192.

Munroe:2015:GTV

[915] Samantha E. M. Munroe, Michelle R. Heupel, Aaron T. Fisk, and Colin A. Simpfendorfer. Geographic and temporal variation in the trophic ecology of a small-bodied shark: evidence of resilience to environmental change. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 72 (3):343–351, March 2015. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2014-0230.

Koopman:2015:DFF

[916] Heather N. Koopman, Andrew J. Westgate, and Zachary A. Siders. Declining fecundity and factors affecting embryo quality in the American lobster (Homarus americanus) from the Bay of Fundy. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 72(3):352–363, March 2015. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2014-0277.

Millar:2015:BEM

[917] Russell B. Millar. A better estimator of mortality rate from agefrequency data. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 72(3):364–375, March 2015. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
cjfas-2014-0193.

Barnett:2015:QPM

[918] Lewis A. K. Barnett, Marissa L. Baskett, and Louis W. Botsford. Quantifying the potential for marine reserves or harvest reductions to buffer temporal mismatches caused by climate change. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 72(3):376–389, March 2015. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2014-0243.

Oliver:2015:NEC

[919] Samantha K. Oliver, Donn K. Branstrator, Thomas R. Hrabik, Stephanie J. Guildford, and Robert E. Hecky. Nutrient excretion by crustacean zooplankton in the deep chlorophyll layer of Lake Superior. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 72(3):390–399, March 2015. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2014-0209.

Arkhipkin:2015:ATI

[920] Alexander Arkhipkin, Juan Argüelles, Zhanna Shcherbich, and Carmen Yamashiro. Ambient temperature influences adult size and life span in jumbo squid (Dosidicus gigas). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 72(3):400–409, March 2015. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/cjfas-2014-0386.

Mulazzani:2015:RAC

[921] Luca Mulazzani, Rosa Manrique, Giovanna Trevisan, Corrado Piccinetti, and Giulio Malorgio. The relationship among catch, fishing effort, and measures of fish stock abundance: implications in the Adriatic Sea. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 72(3):410–421, March 2015. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2014-0337.

Ramsay:2015:BTF

[922] Alice L. Ramsay, Roger N. Hughes, Simon R. Chenery, and Ian D. Mc-Carthy. Biogeochemical tags in fish: predicting spatial variations in strontium and manganese in Salmo trutta scales using stream water geochemistry. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 72(3):422–433, March 2015. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
cjfas-2014-0055.

Liermann:2015:SDS

[923] Martin C. Liermann, Dan Rawding, George R. Pess, and Bryce Glaser. The spatial distribution of salmon and steelhead redds and optimal sampling design. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 72(3):434–446, March 2015. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
cjfas-2014-0181.

Wyatt:2015:NTI

[924] Kevin H. Wyatt, Jill S. Bange, Andrea S. Fitzgibbon, Melody J. Bernot, and Allison R. Rober. Nutrients and temperature interact to regulate algae and heterotrophic bacteria in an Alaskan poor fen peatland. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 72(3):447–453, March 2015. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2014-0425.

Anderson:2015:DPC

[925] Joseph H. Anderson, Paul L. Faulds, Karl D. Burton, Michele E. Koehler, William I. Atlas, and Thomas P. Quinn. Dispersal and productivity of Chinook (Oncorhynchus tshawytscha) and Coho (Oncorhynchus kisutch) salmon colonizing newly accessible habitat. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 72(3):454–465, March 2015. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/cjfas-2014-0180.

Roni:2015:WPR

[926] Philip Roni, Tim Beechie, George Pess, and Karrie Hanson. Wood placement in river restoration: fact, fiction, and future direction. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 72(3):466–478, March 2015. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2014-0344.

Ogston:2015:WSE

[927] Lindsey Ogston, Sam Gidora, Matthew Foy, and Jordan Rosenfeld. Watershed-scale effectiveness of floodplain habitat restoration for juvenile Coho salmon in the Chilliwack River, British Columbia. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 72(4):479–490, April 2015. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2014-0189.

Handegard:2015:RCH

[928] Nils Olav Handegard, Alex De Robertis, Guillaume Rieucau, Kevin Boswell, and Gavin J. Macaulay. The reaction of a captive herring school to playbacks of a noise-reduced and a conventional research vessel. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 72(4):491–499, April 2015. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2014-0257.

Beardmore:2015:EMA

[929] Ben Beardmore, Len M. Hunt, Wolfgang Haider, Malte Dorow, and Robert Arlinghaus. Effectively managing angler satisfaction in recreational fisheries requires understanding the fish species and the anglers. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 72(4):500–513, April 2015. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2014-0177.

Bean:2015:MHI

[930] Jared R. Bean, Andrew C. Wilcox, William W. Woessner, and Clint C. Muhlfeld. Multiscale hydrogeomorphic influences on bull trout (Salvelinus confluentus) spawning habitat. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 72(4):514–526, April 2015. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/

doi/full/10.1139/cjfas-2013-0534.

Hoagstrom:2015:IRC

[931] Christopher W. Hoagstrom, Thomas P. Archdeacon, Stephen R. Davenport, David L. Propst, and James E. Brooks. Intrafragment riverscape conservation for an imperiled, small-bodied, pelagic-broadcast spawning minnow: speckled chub (Macrhybopsis aestivalis). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 72(4):527–537, April 2015. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2014-0191.

Musyl:2015:PMI

[932] Michael K. Musyl, Christopher D. Moyes, Richard W. Brill, Bruno L. Mourato, Andrew West, Lianne M. McNaughton, Wei-Chuan Chiang, and Chi-Lu Sun. Postrelease mortality in istiophorid billfish. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 72(4):538–556, April 2015. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2014-0323.

Winship:2015:EFP

[933] Arliss J. Winship, Michael R. O’Farrell, William H. Satterthwaite, Brian K. Wells, and Michael S. Mohr. Expected future performance of salmon abundance forecast models with varying complexity. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 72(4):557–569, April 2015. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2014-0247.

Raabe:2015:IWW

[934] Joshua K. Raabe and Michael A. Bozek. Influence of wind, wave, and water level dynamics on walleye eggs in a north temperate lake. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 72(4):570–581, April 2015. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2014-0320.

Goethel:2015:TPS

[935] Daniel R. Goethel, Christopher M. Legault, and Steven X. Cadrin. Testing the performance of a spatially explicit tag-integrated stock assessment model of yellowtail flounder (Limanda ferruginea) through simulation analysis. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 72(4):582– 601, April 2015. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
cjfas-2014-0244.

Winegardner:2015:RFO

[936] Amanda K. Winegardner, Emma E. Hodgson, and Adrienne M. Davidson. Reductions in federal oversight of aquatic systems in Canada: implications of the new navigation protection act. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 72(4):602–611, April 2015. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/cjfas-2014-0385.

Zimmermann:2015:BCF

[937] Fabian Zimmermann and Christian Jørgensen. Bioeconomic consequences of fishing-induced evolution: a model predicts limited impact on net present value. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 72 (4):612–624, April 2015. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2014-0006.

Graves:2015:TMJ

[938] John E. Graves, Andrew S. Wozniak, Rebecca M. Dickhut, Michele A. Cochran, Elizabeth H. MacDonald, Elizabeth Bush, Haritz Arrizabalaga, and Nicolas Goñi. Transatlantic movements of juvenile Atlantic bluefin tuna inferred from analyses of organochlorine tracers. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 72(4):625–633, April 2015. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2014-0305.

Brooks:2015:WDD

[939] Elizabeth N. Brooks and Jonathan J. Deroba. When “data” are not data: the pitfalls of post hoc analyses that use stock assessment model output. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 72(4):634–641, April 2015. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2014-0231.

Heathcote:2015:PBF

[940] Adam J. Heathcote, Paul A. del Giorgio, and Yves T. Prairie. Predicting bathymetric features of lakes from the topography of their surrounding landscape. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 72 (5):643–650, May 2015. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2014-0392.

Silsbe:2015:QEP

[941] Greg M. Silsbe, Ralph E. H. Smith, and Michael R. Twiss. Quantum efficiency of phytoplankton photochemistry measured continuously across gradients of nutrients and biomass in Lake Erie (Canada and USA) is strongly regulated by light but not by nutrient deficiency. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 72(5):651–660, May 2015. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2014-0365.

Hansen:2015:PWR

[942] Gretchen J. A. Hansen, Stephen R. Carpenter, Jereme W. Gaeta, Joseph M. Hennessy, and M. Jake Vander Zanden. Predicting walleye recruitment as a tool for prioritizing management actions. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 72(5):661–672, May 2015. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2014-0513.

Madison:2015:EER

[943] Barry N. Madison, John W. Heath, Daniel D. Heath, and Nicholas J. Bernier. Effects of early rearing environment and breeding strategy on social interactions and the hormonal response to stressors in juvenile Chinook salmon. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 72(5):673– 683, May 2015. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
cjfas-2014-0409.

Johnson:2015:RTJ

[944] Gary E. Johnson, Gene R. Ploskey, Nichole K. Sather, and David J. Teel. Residence times of juvenile salmon and steelhead in off-channel tidal freshwater habitats, Columbia River, USA. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 72(5):684–696, May 2015. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/cjfas-2014-0085.

Malick:2015:LPP

[945] Michael J. Malick, Sean P. Cox, Franz J. Mueter, and Randall M. Peterman. Linking phytoplankton phenology to salmon productivity along a north–south gradient in the Northeast Pacific Ocean. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 72(5):697–708, May 2015. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2014-0298.

Feyrer:2015:MSS

[946] Frederick Feyrer, James Hobbs, Shawn Acuna, Brian Mahardja, Lenny Grimaldo, Melinda Baerwald, Rachel C. Johnson, and Swee Teh. Metapopulation structure of a semi-anadromous fish in a dynamic environment. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 72(5):709–721, May 2015. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
cjfas-2014-0433.

Warren-Myers:2015:IDE

[947] Fletcher Warren-Myers, Tim Dempster, Per Gunnar Fjelldal, Tom Hansen, and Stephen E. Swearer. Immersion during egg swelling results in rapid uptake of stable isotope markers in salmonid otoliths. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 72(5):722–727, May 2015. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2014-0390.

Gibson:2015:EPT

[948] A. Jamie F. Gibson, Edmund A. Halfyard, Rod G. Bradford, Michael J. W. Stokesbury, and Anna M. Redden. Effects of predation on telemetry-based survival estimates: insights from a study on endangered Atlantic salmon smolts. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 72 (5):728–741, May 2015. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2014-0245.

Rodtka:2015:EOD

[949] Mike C. Rodtka, Chad S. Judd, Peter K. M. Aku, and Kevin M. Fitzsimmons. Estimating occupancy and detection probability of juvenile bull trout using backpack electrofishing gear in a west-central Alberta watershed. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 72 (5):742–750, May 2015. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2014-0175.

Houde:2015:GME

[950] Aimee Lee S. Houde, Craig A. Black, Chris C. Wilson, Trevor E. Pitcher, and Bryan D. Neff. Genetic and maternal effects on juvenile survival and fitness-related traits in three populations of Atlantic salmon. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 72(5):751–758, May 2015. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2014-0472.

Hawkshaw:2015:HCR

[951] Michael Hawkshaw and Carl Walters. Harvest control rules for mixedstock fisheries coping with autocorrelated recruitment variation, conservation of weak stocks, and economic well-being. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 72(5):759–766, May 2015. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2014-0212.

Au:2015:NAC

[952] David W. Au, Susan E. Smith, and Christina Show. New abbreviated calculation for measuring intrinsic rebound potential in exploited fish populations — example for sharks. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 72 (5):767–773, May 2015. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2014-0360.

Valentin:2015:URP

[953] Alexandra E. Valentin, Don Power, and Jean-Marie Sévigny. Understanding recruitment patterns of historically strong juvenile year classes in redfish (Sebastes spp.): the importance of species identity, population structure, and juvenile migration. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 72(5):774–784, May 2015. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/cjfas-2014-0149.

Moore:2015:BCR

[954] Jonathan W. Moore. Bidirectional connectivity in rivers and implications for watershed stability and management. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 72(5):785–795, May 2015. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/cjfas-2014-0478.

Kraus:2015:DHZ

[955] Richard T. Kraus, Carey T. Knight, Troy M. Farmer, Ann Marie Gorman, Paris D. Collingsworth, Glenn J. Warren, Patrick M. Kocovsky, and Joseph D. Conroy. Dynamic hypoxic zones in Lake Erie compress fish habitat, altering vulnerability to fishing gears. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 72(6):797–806, June 2015. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/cjfas-2014-0517.

Barnes:2015:ESP

[956] Brian D. Barnes and Wayne A. Wurtsbaugh. The effects of salinity on plankton and benthic communities in the Great Salt Lake, Utah, USA: a microcosm experiment. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 72 (6):807–817, June 2015. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2014-0396.

Ruggerone:2015:PLH

[957] Gregory T. Ruggerone and Brendan M. Connors. Productivity and life history of sockeye salmon in relation to competition with pink and sockeye salmon in the North Pacific Ocean. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 72(6):818–833, June 2015. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/cjfas-2014-0134.

Rypel:2015:PRW

[958] Andrew L. Rypel, Daisuke Goto, Greg G. Sass, and M. Jake Vander Zanden. Production rates of walleye and their relationship to exploitation in Escanaba Lake, Wisconsin, 1965–2009. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 72(6):834–844, June 2015. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/cjfas-2014-0394. See corrigendum [1011].

Turner:2015:CDC

[959] Sara M. Turner, Karin E. Limburg, and Eric P. Palkovacs. Can different combinations of natural tags identify river herring natal origin at different levels of stock structure? Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 72 (6):845–854, June 2015. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2014-0403.

Jensen:2015:BWM

[960] Arne Johan Jensen, Ola Håvard Diserud, Bengt Finstad, Peder Fiske, and Audun Håvard Rikardsen. Between-watershed movements of two anadromous salmonids in the Arctic. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 72(6):855–863, June 2015. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/cjfas-2015-0015.

Dragon:2015:EDM

[961] Anne-Cécile Dragon, Inna Senina, Olivier Titaud, Beatriz Calmettes, Anna Conchon, Haritz Arrizabalaga, and Patrick Lehodey. An ecosystem-driven model for spatial dynamics and stock assessment of North Atlantic albacore. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 72 (6):864–878, June 2015. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2014-0338.

Wischniowski:2015:IBP

[962] Stephen G. Wischniowski, Craig R. Kastelle, Timothy Loher, and Thomas E. Helser. Incorporation of bomb-produced^{14}C into fish otoliths. an example of basin-specific rates from the North Pacific Ocean. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 72(6):879–892, June 2015. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2014-0225.

Penaluna:2015:RGT

[963] Brooke E. Penaluna, Steve F. Railsback, Jason B. Dunham, Sherri Johnson, Robert E. Bilby, and Arne E. Skaugset. The role of the geophysical template and environmental regimes in controlling stream-living trout populations. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 72(6):893–901, June 2015. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
cjfas-2014-0377.

Cullis:2015:HCB

[964] James Duncan Sydney Cullis, Diane Marie McKnight, and Sarah A. Spaulding. Hydrodynamic control of benthic mats of Didymosphenia geminata at the reach scale. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 72 (6):902–914, June 2015. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2014-0314.

Mortensen:2015:DSH

[965] Rebecca A. Mortensen, Stephen A. Arnott, William J. Jones, and Dianne I. Greenfield. Development of a sandwich hybridization assay for the identification and quantification of red drum (Sciaenops ocellatus) eggs: a novel tool for fishery research and management. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 72(6):915–925, June 2015. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2014-0526.

Lawson:2015:EMN

[966] Zach J. Lawson, M. Jake Vander Zanden, Colin A. Smith, Emily Heald, Thomas R. Hrabik, and Stephen R. Carpenter. Experimental mixing of a north-temperate lake: testing the thermal limits of a cold-water invasive fish. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 72(6):926–937, June 2015. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
cjfas-2014-0346.

Aanes:2015:ESE

[967] Sondre Aanes and Jon Helge Vølstad. Efficient statistical estimators and sampling strategies for estimating the age composition of fish. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 72(6):938–953, June 2015. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2014-0408.

Kuchapski:2015:FCT

[968] Kathryn A. Kuchapski and Joseph B. Rasmussen. Food chain transfer and exposure effects of selenium in salmonid fish communities in two watersheds in the Canadian Rocky Mountains. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 72(7):955–967, July 2015. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/cjfas-2014-0484.

Spies:2015:WDW

[969] Ingrid Spies, Paul D. Spencer, and André E. Punt. Where do we draw the line? A simulation approach for evaluating management of marine fish stocks with isolation-by-distance stock structure. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 72(7):968–982, July 2015. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2014-0366.

Colborne:2015:FEN

[970] S. F. Colborne, A. D. M. Clapp, F. J. Longstaffe, and B. D. Neff. Foraging ecology of native pumpkinseed (Lepomis gibbosus) following the invasion of zebra mussels (Dreissena polymorpha). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 72(7):983–990, July 2015. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/cjfas-2014-0372.

Martin:2015:DCA

[971] Jean Martin, Quentin Rougemont, Hilaire Drouineau, Sophie Launey, Philippe Jatteau, Gilles Bareille, Sylvain Berail, Christophe Pécheyran, Eric Feunteun, Séverine Roques, David Clavé, David Jose Nachón, Carlos Antunes, Micaela Mota, Elodie Réveillac, and Françoise Daverat. Dispersal capacities of anadromous Allis shad population inferred from a coupled genetic and otolith approach. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 72(7):991–1003, July 2015. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/cjfas-2014-0510.

Ridgway:2015:CZO

[972] Mark S. Ridgway and Trevor A. Middel. Coastal zone occupancy by double-crested cormorants on a Laurentian Great Lake before, during, and after a food web regime shift. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 72(7):1004–1014, July 2015. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/cjfas-2014-0443.

Colman:2015:MPB

[973] John A. Colman, Jacob I. Nogueira, Oscar C. Pancorbo, Carol A. Batdorf, and Barbara A. Block. Mercury in Pacific bluefin tuna (Thunnus orientalis): bioaccumulation and trans-Pacific Ocean migration. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 72(7):1015–1023, July 2015. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2014-0476.

Cahalan:2015:CEF

[974] Jennifer A. Cahalan, Jason Gasper, and Jennifer Mondragon. Catch estimation in the federal trawl fisheries off Alaska: a simulation approach to compare the statistical properties of three trip-specific catch estimators. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 72(7):1024–1036, July 2015. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2014-0347.

Tommasi:2015:EEC

[975] Désirée Tommasi, Janet Nye, Charles Stock, Jonathan A. Hare, Michael Alexander, and Katie Drew. Effect of environmental conditions on juvenile recruitment of alewife (Alosa pseudoharengus) and blueback herring (Alosa aestivalis) in fresh water: a coastwide perspective. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 72(7):1037–1047, July 2015. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2014-0259.

Taylor:2015:GAD

[976] Eric B. Taylor and Shannan L. May-McNally. Genetic analysis of Dolly Varden (Salvelinus malma) across its North American range: evidence for a contact zone in southcentral Alaska. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 72(7):1048–1057, July 2015. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/cjfas-2015-0003.

Balasubramaniam:2015:SWI

[977] Ann M. Balasubramaniam, Roland I. Hall, Brent B. Wolfe, Jon N. Sweetman, and Xiaowa Wang. Source water inputs and catchment characteristics regulate limnological conditions of shallow subarctic lakes (Old Crow Flats, Yukon, Canada). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 72 (7):1058–1072, July 2015. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2014-0340.

Wynne:2015:REH

[978] Molly L. Payne Wynne, Karen A. Wilson, and Karin E. Limburg. Retrospective examination of habitat use by blueback herring (Alosa aestivalis) using otolith microchemical methods. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 72(7):1073–1086, July 2015. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2014-0206.

Guethjonsson:2015:MFA

[979] Sigurdur Gudjónsson, Sigurdur Már Einarsson, Ingi Rúnar Jónsson, and Jóhannes Gudbrandsson. Marine feeding areas and vertical movements of Atlantic salmon (Salmo salar) as inferred from recoveries of data storage tags. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 72(7):1087– 1098, July 2015. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
cjfas-2014-0562.

Roach:2015:HRT

[980] Katherine A. Roach and Kirk O. Winemiller. Hydrologic regime and turbidity influence entrance of terrestrial material into river food webs. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 72(7):1099–1112, July 2015. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2014-0459.

Godwin:2015:SLS

[981] Sean C. Godwin, Lawrence M. Dill, John D. Reynolds, and Martin Krkošek. Sea lice, sockeye salmon, and foraging competition: lousy fish are lousy competitors. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 72 (7):1113–1120, July 2015. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2014-0284.

Carr:2015:GHN

[982] Steven M. Carr and William S. Davidson. Greenland halibut in the North Atlantic are genetically homogeneous: a response to. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 72(7):1121, July 2015. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2014-0537.

Roy:2015:PIL

[983] Denis Roy, Margaret A. Treble, David C. Hardie, James D. Reist, and Daniel E. Ruzzante. Panmixia and incomplete lineage sorting can give similar mtDNA signals: response to comment by Carr and Davidson (2015). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 72(7):1122–1123, July 2015. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
cjfas-2015-0128.

Neuswanger:2015:LPC

[984] Jason R. Neuswanger, Mark S. Wipfli, Matthew J. Evenson, Nicholas F. Hughes, and Amanda E. Rosenberger. Low productivity of Chinook salmon strongly correlates with high summer stream discharge in two Alaskan rivers in the Yukon drainage. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 72(8):1125–1137, August 2015. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/cjfas-2014-0498.

Yates:2015:IHD

[985] Matthew C. Yates, Paul V. Debes, Dylan J. Fraser, and Jeffrey A. Hutchings. The influence of hybridization with domesticated conspecifics on alternative reproductive phenotypes in male Atlantic salmon in multiple temperature regimes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 72(8): 1138–1145, August 2015. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2014-0527.

Buxton:2015:MIS

[986] Todd H. Buxton, John M. Buffington, Daniele Tonina, Alexander K. Fremier, and Elowyn M. Yager. Modeling the influence of salmon spawning on hyporheic exchange of marine-derived nutrients in gravel stream beds. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 72(8):1146–1158, August 2015. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
cjfas-2014-0413.

Carruthers:2015:MAD

[987] Thomas R. Carruthers, John F. Walter, Murdoch K. McAllister, and Meaghan D. Bryan. Modelling age-dependent movement: an application to red and gag groupers in the Gulf of Mexico. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 72(8):1159–1176, August 2015. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2014-0471.

Ono:2015:HDM

[988] Kotaro Ono, André E. Punt, and Ray Hilborn. How do marine closures affect the analysis of catch and effort data? Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 72(8):1177–1190, August 2015. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2014-0146.

Vilela:2015:FSM

[989] Raúl Vilela and Jose María Bellido. Fishing suitability maps: helping fishermen reduce discards. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 72 (8):1191–1201, August 2015. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2013-0522.

Sarnelle:2015:TIF

[990] Orlando Sarnelle, Jeffrey D. White, Theresa E. Geelhoed, and Carrie L. Kozel. Type III functional response in the zebra mussel, Dreissena polymorpha. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 72(8):1202–1207, August 2015. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
cjfas-2015-0076.

McAdam:2015:RWE

[991] D. Steven O. McAdam. Retrospective weight-of-evidence analysis identifies substrate change as the apparent cause of recruitment failure in the upper Columbia River white sturgeon (Acipenser transmontanus). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 72(8):1208–1220, August 2015. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2014-0423.

White:2015:HIM

[992] Jeffrey D. White, Stephen K. Hamilton, and Orlando Sarnelle. Heatinduced mass mortality of invasive zebra mussels (Dreissena polymorpha) at sublethal water temperatures. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 72(8):1221–1229, August 2015. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/cjfas-2015-0064.

Chan:2015:RIV

[993] Farrah T. Chan, Hugh J. MacIsaac, and Sarah A. Bailey. Relative importance of vessel hull fouling and ballast water as transport vectors of nonindigenous species to the Canadian Arctic. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 72(8):1230–1242, August 2015. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2014-0473.

Ivan:2015:EAS

[994] Lori N. Ivan and Tomas O. Höök. Energy allocation strategies of young temperate fish: an eco-genetic modeling approach. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 72(8):1243–1258, August 2015. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2014-0197.

Panikian:2015:IQH

[995] Garo Panikian, James Cussens, and Jonathan W. Pitchford. Identification and quantification of heteroscedasticity in stock–recruitment relationships. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 72(8):1259–1271, August 2015. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
cjfas-2014-0549.

Creed:2015:RCF

[996] Irena F. Creed, Diane M. McKnight, Brian A. Pellerin, Mark B. Green, Brian A. Bergamaschi, George R. Aiken, Douglas A. Burns, Stuart E. G. Findlay, Jamie B. Shanley, Rob G. Striegl, Brent T. Aulenbach, David W. Clow, Hjalmar Laudon, Brian L. McGlynn, Kevin J. McGuire, Richard A. Smith, and Sarah M. Stackpoole. The river as a chemostat: fresh perspectives on dissolved organic matter flowing down the river continuum. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 72(8):1272–1285, August 2015. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
cjfas-2014-0400.

Kessel:2015:TIA

[997] Steven T. Kessel and Nigel E. Hussey. Tonic immobility as an anaesthetic for elasmobranchs during surgical implantation procedures. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 72(9):1287–1291, September 2015. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2015-0136.

McGarvey:2015:CSL

[998] Richard McGarvey, André E. Punt, Janet M. Matthews, John E. Feenstra, Caleb Gardner, Paul Burch, Klaas Hartmann, and Adrian Linnane. Comparing size-limit and quota policies to increase economic yield in a lobster fishery. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 72(9):1292– 1305, September 2015. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2014-0405.

Oligny-Hebert:2015:EDT

[999] Hélène Oligny-Hébert, Caroline Senay, Eva C. Enders, and Daniel Boisclair. Effects of diel temperature fluctuation on the standard metabolic rate of juvenile Atlantic salmon (Salmo salar): influence of acclimation temperature and provenience. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 72(9):1306–1315, September 2015. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/cjfas-2014-0345.

Stawitz:2015:SSA

[1000] Christine C. Stawitz, Timothy E. Essington, Trevor A. Branch, Melissa A. Haltuch, Anne B. Hollowed, and Paul D. Spencer. A statespace approach for detecting growth variation and application to North Pacific groundfish. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 72(9):1316– 1328, September 2015. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2014-0558.

Jacobson:2015:SZA

[1001] Larry D. Jacobson, Lisa C. Hendrickson, and Jiashen Tang. Solar zenith angles for biological research and an expected catch model for diel vertical migration patterns that affect stock size estimates for longfin inshore squid (Doryteuthis pealeii). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 72(9):1329–1338, September 2015. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/cjfas-2014-0436.

Stich:2015:IMM

[1002] Daniel S. Stich, Michael T. Kinnison, John F. Kocik, and Joseph D. Zydlewski. Initiation of migration and movement rates of Atlantic salmon smolts in fresh water. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 72(9): 1339–1351, September 2015. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2014-0570.

Stich:2015:CWS

[1003] Daniel S. Stich, Michael M. Bailey, Christopher M. Holbrook, Michael T. Kinnison, and Joseph D. Zydlewski. Catchment-wide survival of wildand hatchery-reared Atlantic salmon smolts in a changing system. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 72(9):1352–1365, September 2015. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2014-0573.

Eldoy:2015:MMH

[1004] Sindre Håvarstein Eldøy, Jan Grimsrud Davidsen, Eva Bonsak Thorstad, Fred Whoriskey, Kim Aarestrup, Tor Fredrik Næsje, Lars Rønning, Aslak Darre Sjursen, Audun Håvard Rikardsen, and Jo Vegar Arnekleiv. Marine migration and habitat use of anadromous brown trout (Salmo trutta). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 72(9):1366– 1378, September 2015. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2014-0560.

Hartill:2015:CRH

[1005] Bruce William Hartill and Charles Thomas Thesen Edwards. Comparison of recreational harvest estimates provided by onsite and offsite surveys: detecting bias and corroborating estimates. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 72(9):1379–1389, September 2015. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2014-0451.

Sard:2015:FIS

[1006] Nicholas M. Sard, Kathleen G. O’Malley, Dave P. Jacobson, Michael J. Hogansen, Marc A. Johnson, and Michael A. Banks. Factors influencing spawner success in a spring Chinook salmon (Oncorhynchus tshawytscha) reintroduction program. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 72(9): 1390–1397, September 2015. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2015-0007.

Forget:2015:BVT

[1007] Fabien G. Forget, Manuela Capello, John D. Filmalter, Rodney Govinden, Marc Soria, Paul D. Cowley, and Laurent Dagorn. Behaviour and vulnerability of target and non-target species at drifting fish aggregating devices (FADs) in the tropical tuna purse seine fishery determined by acoustic telemetry. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 72(9): 1398–1405, September 2015. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2014-0458.

Samways:2015:CPB

[1008] Kurt M. Samways, Peter R. Leavitt, Pierre Magnan, Marco A. Rodríguez, and Pedro R. Peres-Neto. Convergent polymorphism between stream and lake habitats: the case of brook char. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 72(9):1406–1414, September 2015. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2015-0116.

Spies:2015:UGM

[1009] Ingrid Spies and André E. Punt. The utility of genetics in marine fisheries management: a simulation study based on Pacific cod off Alaska. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 72(9):1415–1432, September 2015. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2014-0050.

Hilborn:2015:WFS

[1010] Ray Hilborn, Elizabeth A. Fulton, Bridget S. Green, Klaas Hartmann, Sean R. Tracey, and Reg A. Watson. When is a fishery sustainable? Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 72(9):1433–1441, September 2015. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
cjfas-2015-0062.

Rypel:2015:CPR

[1011] Andrew L. Rypel, Daisuke Goto, Greg G. Sass, and M. Jake Vander Zanden. Corrigendum: Production rates of walleye and their relationship to exploitation in Escanaba Lake, Wisconsin, 1965–2009. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 72(9):1442, September 2015. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:/

/cdnsciencepub.com/doi/full/10.1139/cjfas-2015-0308. See [958].

Weber:2015:IVT

[1012] Gregory M. Weber, Mark A. Hostuttler, Kenneth J. Semmens, and Brian A. Beers. Induction and viability of tetraploids in brook trout (Salvelinus fontinalis). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 72(10): 1443–1449, October 2015. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2014-0536.

Fritts:2015:EMA

[1013] Andrea K. Fritts, James T. Peterson, Peter D. Hazelton, and Robert B. Bringolf. Evaluation of methods for assessing physiological biomarkers of stress in freshwater mussels. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 72(10):1450–1459, October 2015. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/cjfas-2014-0564.

Fritts:2015:NAF

[1014] Andrea K. Fritts, James T. Peterson, Jason M. Wisniewski, and Robert B. Bringolf. Nonlethal assessment of freshwater mussel physiological response to changes in environmental factors. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 72(10):1460–1468, October 2015. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2014-0565.

Nelson:2015:QES

[1015] Michelle C. Nelson, Morgan D. Hocking, Jennifer N. Harding, Joel M. S. Harding, and John D. Reynolds. Quantifying the effects of stream habitat on populations of breeding Pacific salmon. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 72(10):1469–1476, October 2015. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2014-0253.

Harris:2015:LHC

[1016] Les N. Harris, Robert Bajno, Colin P. Gallagher, Itsuro Koizumi, Lucy K. Johnson, Kimberly L. Howland, Eric B. Taylor, and James D. Reist. Life-history characteristics and landscape attributes as drivers of genetic variation, gene flow, and fine-scale population structure in northern Dolly Varden (Salvelinus malma malma) in Canada. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 72(10):1477–1493, October 2015. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2015-0016.

Zhang:2015:CPS

[1017] Fan Zhang, Kevin B. Reid, and Thomas D. Nudds. Counterintuitive patterns of spawning stock age structure and recruitment in Lake Erie yellow perch (Perca flavescens). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 72(10): 1494–1502, October 2015. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2014-0489.

Rossetto:2015:NTM

[1018] Marisa Rossetto, Fiorenza Micheli, Andrea Saenz-Arroyo, Jose Antonio Espinoza Montes, and Giulio Alessandro De Leo. No-take marine reserves can enhance population persistence and support the fishery of abalone. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 72(10):1503–1517, October 2015. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
cjfas-2013-0623.

Adkison:2015:ESA

[1019] Milo D. Adkison and Curry J. Cunningham. The effects of salmon abundance and run timing on the performance of management by emergency order. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 72(10):1518–1526, October 2015. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
cjfas-2015-0174.

Eldridge:2015:FGP

[1020] William H. Eldridge, Bernard W. Sweeney, and J. Mac Law. Fish growth, physiological stress, and tissue condition in response to rate of temperature change during cool or warm diel thermal cycles. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 72(10):1527–1537, October 2015. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2014-0350.

Jones:2015:RES

[1021] Michael L. Jones, Travis O. Brenden, and Brian J. Irwin. Re-examination of sea lamprey control policies for the St. Marys River: completion of an adaptive management cycle. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 72(10):1538–1551, October 2015. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/cjfas-2014-0567.

Malick:2015:AMP

[1022] Michael J. Malick, Sean P. Cox, Randall M. Peterman, Thomas C. Wainwright, and William T. Peterson. Accounting for multiple pathways in the connections among climate variability, ocean processes, and Coho salmon recruitment in the Northern California Current. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 72(10):1552–1564, October 2015. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2014-0509.

Stowell:2015:GLM

[1023] Sierra M. Love Stowell, Christopher M. Kennedy, Stower C. Beals, Jessica L. Metcalf, and Andrew P. Martin. The genetic legacy of more than a century of stocking trout: a case study in Rocky Mountain National Park, Colorado, USA. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 72(10):

1565–1574, October 2015. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2015-0009.

Chase:2015:MPD

[1024] Nathan M. Chase, Colleen A. Caldwell, Scott A. Carleton, William R. Gould, and James A. Hobbs. Movement patterns and dispersal potential of pecos bluntnose shiner (Notropis simus pecosensis) revealed using otolith microchemistry. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 72(10): 1575–1583, October 2015. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2014-0574.

Moreira:2015:OGD

[1025] Amanda L. Moreira and Eric B. Taylor. The origin and genetic divergence of “black” Kokanee, a novel reproductive ecotype of Oncorhynchus nerka. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 72(10):1584–1595, October 2015. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2015-0145.

McDevitt-Irwin:2015:MSN

[1026] Jamie Marie McDevitt-Irwin, Susanna Drake Fuller, Catharine Grant, and Julia Kathleen Baum. Missing the safety net: evidence for inconsistent and insufficient management of at-risk marine fishes in Canada. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 72(10):1596–1608, October 2015. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2015-0030.

Green:2015:CSE

[1027] Corey P. Green, Simon G. Robertson, Paul A. Hamer, Patti Virtue, George D. Jackson, and Natalie A. Moltschaniwskyj. Combining statolith element composition and Fourier shape data allows discrimination of spatial and temporal stock structure of arrow squid (Nototodarus gouldi). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 72(11):1609–1618, November 2015. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
cjfas-2014-0559.

Perala:2015:DRS

[1028] Tommi Perälä and Anna Kuparinen. Detecting regime shifts in fish stock dynamics. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 72(11):1619–1628, November 2015. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
cjfas-2014-0406.

Narimatsu:2015:IGE

[1029] Yoji Narimatsu, Shigeho Kakehi, Shin ichi Ito, Yuji Okazaki, Ryo Inagawa, and Toshikazu Yano. Impact of the Great East Japan Earthquake tsunami on growth and survival of Pacific cod (Gadus macrocephalus). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 72(11):1629–1638, November 2015. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
cjfas-2014-0467.

Malison:2015:JSG

[1030] Rachel L. Malison, Lisa A. Eby, and Jack A. Stanford. Juvenile salmonid growth, survival, and production in a large river floodplain modified by beavers (Castor canadensis). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 72(11):1639–1651, November 2015. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/cjfas-2015-0147.

deKerckhove:2015:WLB

[1031] Derrick Tupper de Kerckhove, Edyta Agnes Blukacz-Richards, Brian John Shuter, Liset Cruz-Font, and Peter Arnold Abrams. Wind on lakes brings predator and prey together in the pelagic zone. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 72(11):1652–1662, November 2015. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2015-0119.

Seekell:2015:TOB

[1032] David A. Seekell, Jean-François Lapierre, and Jan Karlsson. Trade-offs between light and nutrient availability across gradients of dissolved organic carbon concentration in Swedish lakes: implications for patterns in primary production. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 72(11):

1663–1671, November 2015. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2015-0187.

Friedland:2015:LEP

[1033] Kevin D. Friedland, Robert T. Leaf, Trond Kristiansen, and Scott I. Large. Layered effects of parental condition and larval survival on the recruitment of neighboring haddock stocks. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 72(11):1672–1681, November 2015. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2015-0084.

Atlas:2015:ISC

[1034] William I. Atlas, Thomas W. Buehrens, Don J. F. McCubbing, Robert Bison, and Jonathan W. Moore. Implications of spatial contraction for density dependence and conservation in a depressed population of anadromous fish. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 72(11):1682– 1693, November 2015. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2014-0532.

Jones:2015:FAS

[1035] Douglas A. Jones, Eva Bergman, and Larry Greenberg. Food availability in spring affects smolting in brown trout (Salmo trutta). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 72(11):1694–1699, November 2015. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2015-0106.

Wilson:2015:TFA

[1036] Steven G. Wilson, Ian D. Jonsen, Robert J. Schallert, James E. Ganong, Michael R. Castleton, Aaron D. Spares, André M. Boustany, Michael J. W. Stokesbury, and Barbara A. Block. Tracking the fidelity of Atlantic bluefin tuna released in Canadian waters to the Gulf of Mexico spawning grounds. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 72(11):1700– 1717, November 2015. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2015-0110.

Shedd:2015:ERL

[1037] Kyle R. Shedd, Frank A. von Hippel, James J. Willacker, Troy R. Hamon, Ora L. Schlei, John K. Wenburg, Joe L. Miller, and Scott A. Pavey. Ecological release leads to novel ontogenetic diet shift in Kokanee (Oncorhynchus nerka). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 72(11): 1718–1730, November 2015. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2015-0146.

McManamay:2015:ICP

[1038] Ryan A. McManamay, Brandon K. Peoples, Donald J. Orth, Charles A. Dolloff, and David C. Matthews. Isolating causal pathways between flow and fish in the regulated river hierarchy. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 72(11):1731–1748, November 2015. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2015-0227.

Michel:2015:CSO

[1039] Cyril J. Michel, Arnold J. Ammann, Steven T. Lindley, Philip T. Sandstrom, Eric D. Chapman, Michael J. Thomas, Gabriel P. Singer, A. Peter Klimley, and R. Bruce MacFarlane. Chinook salmon outmigration survival in wet and dry years in California’s Sacramento River. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 72(11):1749–1759, November 2015. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2014-0528.

Vivancos:2015:QCI

[1040] Aurélien Vivancos and Gerard P. Closs. Quantification and comparison of individual space-use strategies in foraging drift-feeding fish using fine-scale, multidimensional movement analysis. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 72(11):1760–1768, November 2015. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2015-0006.

Raoult:2015:ASE

[1041] Vincent Raoult, Troy F. Gaston, and Jane E. Williamson. Not all sawsharks are equal: species of co-existing sawsharks show plasticity in trophic consumption both within and between species. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 72(11):1769–1775, November 2015. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2015-0307.

Katz:2015:EPR

[1042] Rachel A. Katz and Mary C. Freeman. Evidence of population resistance to extreme low flows in a fluvial-dependent fish species. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 72(11):1776–1787, November 2015. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2015-0173.

Anonymous:2015:NAN

[1043] Anonymous. Note of appreciation / note de reconnaissance. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 72(12), December 2015. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2015-0437.

Rose:2015:NCC

[1044] George A. Rose and Sherrylynn Rowe. Northern cod comeback. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 72(12):1789–1798, December 2015. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2015-0346.

Hume:2015:APA

[1045] John B. Hume, Trevor D. Meckley, Nicholas S. Johnson, Thomas M. Luhring, Michael J. Siefkes, and C. Michael Wagner. Application of a putative alarm cue hastens the arrival of invasive sea lamprey (Petromyzon marinus) at a trapping location. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 72(12):1799–1806, December 2015. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/cjfas-2014-0535.

Walsworth:2015:CSE

[1046] Timothy E. Walsworth and Daniel E. Schindler. Coho salmon escapement and trends in migration timing to a data-poor river: estimates from a Bayesian hierarchical model. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 72(12):1807–1816, December 2015. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/cjfas-2014-0554.

Goodman:2015:PSM

[1047] Damon H. Goodman, Stewart B. Reid, Nicholas A. Som, and William R. Poytress. The punctuated seaward migration of Pacific lamprey (Entosphenus tridentatus): environmental cues and implications for streamflow management. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 72(12):1817– 1828, December 2015. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2015-0063.

Thompson:2015:EHR

[1048] Neil F. Thompson and Michael S. Blouin. The effects of high rearing density on the potential for domestication selection in hatchery culture of steelhead (Oncorhynchus mykiss). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 72(12):1829–1834, December 2015. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/cjfas-2015-0233.

Bjorkland:2015:SPR

[1049] Rhema Bjorkland, Daniel C. Dunn, Michelle McClure, Jason Jannot, Marlene A. Bellman, Mary Gleason, and Katja Schiffers. Spatiotemporal patterns of rockfish bycatch in US west coast groundfish fisheries: opportunities for reducing incidental catch of depleted species. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 72(12):1835–1846, December 2015. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2014-0242.

Buchinger:2015:BER

[1050] Tyler J. Buchinger, Weiming Li, and Nicholas S. Johnson. Behavioral evidence for a role of chemoreception during reproduction in lake trout. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 72(12):1847–1852, December 2015. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
cjfas-2015-0351.

Barbini:2015:UOR

[1051] Santiago A. Barbini, Luis O. Lucifora, and Daniel E. Figueroa. Using opportunistic records from a recreational fishing magazine to assess population trends of sharks. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 72(12): 1853–1859, December 2015. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2015-0087.

Satterthwaite:2015:WPE

[1052] William H. Satterthwaite and Stephanie M. Carlson. Weakening portfolio effect strength in a hatchery-supplemented Chinook salmon population complex. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 72(12):1860–1875, December 2015. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
cjfas-2015-0169.

McLean:2015:ULS

[1053] Adrienne R. McLean, Jessica Barber, Gale Bravener, Andrew M. Rous, and Robert L. McLaughlin. Understanding low success trapping invasive sea lampreys: an entry-level analysis. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 72(12):1876–1885, December 2015. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/cjfas-2015-0140.

Perez-Fuentetaja:2015:TTF

[1054] Alicia Pérez-Fuentetaja, Susan A. Mackintosh, Lisa R. Zimmerman, Mark D. Clapsadl, Mehran Alaee, and Diana S. Aga. Trophic transfer of flame retardants (PBDEs) in the food web of Lake Erie. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 72(12):1886–1896, December 2015. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2015-0088.

Thorson:2015:SDD

[1055] James T. Thorson, James N. Ianelli, Stephan B. Munch, Kotaro Ono, and Paul D. Spencer. Spatial delay-difference models for estimating spatiotemporal variation in juvenile production and population abundance. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 72(12):1897–1915, December 2015. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
cjfas-2014-0543.

vanPoorten:2015:DFR

[1056] Brett T. van Poorten and Carl J. Walters. Dependence of feeding rates on body mass when food density is limiting to growth. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 72(12):1916–1922, December 2015. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2015-0204.

Samarasin:2015:CFD

[1057] Pasan Samarasin, Charles K. Minns, Brian J. Shuter, William M. Tonn, and Michael D. Rennie. Corrigendum: Fish diversity and biomass in northern Canadian lakes: northern lakes are more diverse and have greater biomass than expected based on species–energy theory. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 72(12):1923, December 2015. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:/

/cdnsciencepub.com/doi/full/10.1139/cjfas-2015-0467. See [907].

Bickel:2016:PDA

[1058] Tobias Oliver Bickel and Christine Perrett. Precise determination of aquatic plant wet mass using a salad spinner. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 73(1):1–4, January 2016. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/cjfas-2015-0274.

Buchheister:2016:DTL

[1059] André Buchheister and Robert J. Latour. Dynamic trophic linkages in a large estuarine system — support for supply-driven dietary changes using delta generalized additive mixed models. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 73(1):5–17, January 2016. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/cjfas-2014-0441.

Binder:2016:SSF

[1060] Thomas R. Binder, Stephen C. Riley, Christopher M. Holbrook, Michael J. Hansen, Roger A. Bergstedt, Charles R. Bronte, Ji He, and Charles C. Krueger. Spawning site fidelity of wild and hatchery lake trout (Salvelinus namaycush) in northern Lake Huron. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 73(1):18–34, January 2016. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2015-0175.

Deshpande:2016:CBD

[1061] Ashok D. Deshpande, Bruce W. Dockum, and Andrew F. J. Draxler. Contaminant bioaccumulation dynamics in young-of-the-year bluefish subpopulations in New York Bight with a special reference to the condition and nursery area fidelity subsequent to recruitment. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 73(1):35–52, January 2016. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2015-0369.

Klobucar:2016:CSV

[1062] Stephen L. Klobucar and Phaedra Budy. Consequences of seasonal variation in reservoir water level for predatory fishes: linking visual foraging and prey densities. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 73 (1):53–64, January 2016. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2015-0008.

Watras:2016:CDC

[1063] Carl J. Watras, Kenneth A. Morrison, Noah R. Lottig, and Timothy K. Kratz. Comparing the diel cycles of dissolved organic matter fluorescence in a clear-water and two dark-water Wisconsin lakes: potential insights into lake metabolism. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 73 (1):65–75, January 2016. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2015-0172.

Simmons:2016:APE

[1064] Megan Simmons, Andrew Tucker, W. Lindsay Chadderton, Christopher L. Jerde, and Andrew R. Mahon. Active and passive environmental DNA surveillance of aquatic invasive species. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 73(1):76–83, January 2016. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/cjfas-2015-0262.

Heikinheimo:2016:EMC

[1065] Outi Heikinheimo, Pekka Rusanen, and Katja Korhonen. Estimating the mortality caused by great cormorant predation on fish stocks: pikeperch in the Archipelago Sea, northern Baltic Sea, as an example. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 73(1):84–93, January 2016. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2015-0033.

Goerig:2016:BTP

[1066] Elsa Goerig, Theodore Castro-Santos, and Normand Émile Bergeron. Brook trout passage performance through culverts. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 73(1):94–104, January 2016. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2015-0089.

Korman:2016:FCA

[1067] Josh Korman, Michael D. Yard, and Charles B. Yackulic. Factors controlling the abundance of rainbow trout in the Colorado River in Grand Canyon in a reach utilized by endangered humpback chub. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 73(1):105–124, January 2016. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2015-0101.

Yard:2016:SSP

[1068] Michael D. Yard, Josh Korman, Carl J. Walters, and Theodore A. Kennedy. Seasonal and spatial patterns of growth of rainbow trout in the Colorado River in Grand Canyon, Arizona. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 73(1):125–139, January 2016. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2015-0102.

MacKinnon:2016:IHC

[1069] Brett D. MacKinnon, Jay Sagin, Helen M. Baulch, Karl-Erich Lindenschmidt, and Timothy D. Jardine. Influence of hydrological connectivity on winter limnology in floodplain lakes of the Saskatchewan River Delta, Saskatchewan. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 73(1): 140–152, January 2016. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2015-0210.

Devlin:2016:CVC

[1070] R. H. Devlin, L. Park, D. M. Sakhrani, J. D. Baker, A. R. Marshall, E. LaHood, S. E. Kolesar, M. R. Mayo, C. A. Biagi, and M. Uh. Corrigendum: Variation of Y-chromosome DNA markers in Chinook salmon (Oncorhynchus tshawytscha) populations. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 73(1):153, January 2016. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/cjfas-2015-0499. See [3].

Anonymous:2016:SIJ

[1071] Anonymous. Special issue from the Johan Hjort Symposium on Recruitment Dynamics and Stock Variability held in Bergen, Norway, on 7–9 October 2014 / Numéro spécial tiré du Symposium Johan Hjort sur la dynamique du recrutement et la variabilité des stocks qui s’est tenu à Bergen (Norvège) du 7 au 9 octobre 2014. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 73(2), February 2016. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/cjfas-2015-0492.

Kjesbu:2016:JHS

[1072] Olav Sigurd Kjesbu, C. Tara Marshall, Richard D. M. Nash, Svein Sundby, Brian J. Rothschild, and Michael Sinclair. Johan Hjort Symposium on Recruitment Dynamics and Stock Variability. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 73(2), February 2016. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2015-0491.

Takahashi:2016:IVR

[1073] Motomitsu Takahashi, Chiyuki Sassa, Kou Nishiuchi, and Youichi Tsukamoto. Interannual variations in rates of larval growth and development of jack mackerel (Trachurus japonicus) in the East China Sea: implications for juvenile survival. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 73(2):155–162, February 2016. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/cjfas-2015-0077.

Folkvord:2016:IHD

[1074] Arild Folkvord, Geir Gundersen, Jon Albretsen, Lars Asplin, Stein Kaartvedt, and Jarl Giske. Impact of hatch date on early life growth and survival of Mueller’s pearlside (Maurolicus muelleri) larvae and lifehistory consequences. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 73(2): 163–176, February 2016. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2015-0040.

Huse:2016:SAU

[1075] Geir Huse. A spatial approach to understanding herring population dynamics. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 73(2):177–188, February 2016. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
cjfas-2015-0095.

Opdal:2016:LTS

[1076] Anders Frugård Opdal and Frode B. Vikebø. Long-term stability in modelled zooplankton influx could uphold major fish spawning grounds on the Norwegian continental shelf. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 73 (2):189–196, February 2016. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2014-0524.

Volkenandt:2016:FSS

[1077] Mareike Volkenandt, Ian O’Connor, Jean-Marc Guarini, Simon Berrow, and Ciaran O’Donnell. Fine-scale spatial association between baleen whales and forage fish in the Celtic Sea. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 73(2):197–204, February 2016. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://

cdnsciencepub.com/doi/full/10.1139/cjfas-2015-0073.

Pepin:2016:RIL

[1078] Pierre Pepin. Reconsidering the impossible — linking environmental drivers to growth, mortality, and recruitment of fish. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 73(2):205–215, February 2016. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2015-0091.

Uriarte:2016:ANM

[1079] Andrés Uriarte, Leire Ibaibarriaga, Lionel Pawlowski, Jacques Massé, Pierre Petitgas, María Santos, and Dankert Skagen. Assessing natural mortality of Bay of Biscay anchovy from survey population and biomass estimates. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 73(2): 216–234, February 2016. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2015-0096.

Johannesen:2016:SVC

[1080] Edda Johannesen, Geir Odd Johansen, and Knut Korsbrekke. Seasonal variation in cod feeding and growth in a changing sea. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 73(2):235–245, February 2016. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2015-0052.

Bogstad:2016:ELH

[1081] Bjarte Bogstad, Natalia A. Yaragina, and Richard D. M. Nash. The early life-history dynamics of Northeast Arctic cod: levels of natural mortality and abundance during the first 3 years of life. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 73(2):246–256, February 2016. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2015-0093.

Morgan:2016:CPS

[1082] M. Joanne Morgan, Peter A. Shelton, Fernando González-Costas, and Diana González-Troncoso. Compensation potential in six depleted groundfish stocks from the Northwest Atlantic. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 73(2):257–269, February 2016. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2015-0056.

Castillo-Jordan:2016:CRP

[1083] Claudio Castillo-Jordán, Neil L. Klaer, Geoffrey N. Tuck, Stewart D. Frusher, Luis A. Cubillos, Sean R. Tracey, and Michael J. Salinger. Coincident recruitment patterns of Southern Hemisphere fishes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 73(2):270–278, February 2016. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2015-0069.

Hermansen:2016:RAS

[1084] Gudmund H. Hermansen, Nils Lid Hjort, and Olav S. Kjesbu. Recent advances in statistical methodology applied to the hjort liver index time series (1859–2012) and associated influential factors. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 73(2):279–295, February 2016. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2015-0086.

Cadigan:2016:SSS

[1085] Noel G. Cadigan. A state-space stock assessment model for northern cod, including under-reported catches and variable natural mortality rates. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 73(2):296–308, February 2016. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2015-0047.

Dankel:2016:CUQ

[1086] Dorothy Jane Dankel, Jon Helge Vølstad, and Sondre Aanes. Communicating uncertainty in quota advice: a case for confidence interval harvest control rules (CI-HCRs) for fisheries. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 73(2):309–317, February 2016. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/cjfas-2015-0078.

Anonymous:2016:ACa

[1087] Anonymous. Announcement / Communiqué. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 73(3):iii, March 2016. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/cjfas-2016-0058.

Bret:2016:IDH

[1088] Victor Bret, Benjamin Bergerot, Hervé Capra, Véronique Gouraud, and Nicolas Lamouroux. Influence of discharge, hydraulics, water temperature, and dispersal on density synchrony in brown trout populations (Salmo trutta). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 73(3):319–329, March 2016. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
cjfas-2015-0209.

Herbst:2016:EWS

[1089] Seth J. Herbst, Bryan S. Stevens, Daniel B. Hayes, and Patrick A. Hanchin. Estimating walleye (Sander vitreus) movement and fishing mortality using state-space models: implications for management of spatially structured populations. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 73 (3):330–348, March 2016. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2015-0021.

Legault:2016:WDS

[1090] Christopher M. Legault and Michael C. Palmer. In what direction should the fishing mortality target change when natural mortality increases within an assessment? Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 73 (3):349–357, March 2016. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2015-0232.

VanLeeuwen:2016:DDI

[1091] Travis E. Van Leeuwen, Oliver E. Hooker, Neil B. Metcalfe, and Colin E. Adams. Differences in diet-induced flexibility in morphology and growth in a partially migratory species. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 73(3):358–365, March 2016. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/cjfas-2015-0300.

Pazhayamadom:2016:SSC

[1092] Deepak George Pazhayamadom, Ciarán J. Kelly, Emer Rogan, and Edward A. Codling. Self-starting cumulative sum harvest control rule (SS-CUSUM-HCR) for status-quo management of data-limited fisheries. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 73(3):366–381, March 2016. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2015-0039.

Reid:2016:SPT

[1093] Carolyn R. Reid and Shaun A. Watmough. Spatial patterns, trends, and the potential long-term impacts of tree harvesting on lake calcium levels in the Muskoka River Watershed, Ontario, Canada. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 73(3):382–393, March 2016. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2015-0231.

Troia:2016:TDP

[1094] Matthew J. Troia, Michael A. Denk, and Keith B. Gido. Temperaturedependent performance as a driver of warm-water fish species replacement along the river continuum. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 73(3):394–405, March 2016. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/cjfas-2015-0094.

Corrigan:2016:MCS

[1095] Shannon Corrigan, Charlie Huveneers, Adam Stow, and Luciano B. Beheregaray. A multilocus comparative study of dispersal in three codistributed demersal sharks from eastern Australia. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 73(3):406–415, March 2016. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2015-0085.

Brodnik:2016:LDU

[1096] Reed M. Brodnik, Michael E. Fraker, Eric J. Anderson, Lucia Carreon-Martinez, Kristen M. DeVanna, Daniel D. Heath, Julie M. Reichert, Edward F. Roseman, and Stuart A. Ludsin. Larval dispersal underlies demographically important intersystem connectivity in a Great Lakes yellow perch (Perca flavescens) population. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 73(3):416–426, March 2016. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/cjfas-2015-0161.

Sahashi:2016:HSR

[1097] Genki Sahashi and Taku Yoshiyama. A hump-shaped relationship between migration distance and adult pink salmon morphology suggests interactive effects of migration costs and bear predation. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 73(3):427–435, March 2016. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2015-0176.

Lyle:2016:UOS

[1098] Jeremy M. Lyle, Simon T. Willcox, and Klaas Hartmann. Underwater observations of seal–fishery interactions and the effectiveness of an exclusion device in reducing bycatch in a midwater trawl fishery. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 73(3):436–444, March 2016. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2015-0273.

Robertson:2016:ILT

[1099] Dale M. Robertson, William J. Rose, and Paul C. Reneau. Interannual and long-term changes in the trophic state of a multibasin lake: effects of morphology, climate, winter aeration, and beaver activity. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 73(3):445–460, March 2016. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2015-0249.

Tsuboi:2016:BMA

[1100] Jun ichi Tsuboi, Kentaro Morita, Thomas Klefoth, Shinsuke Endou, and Robert Arlinghaus. Behaviour-mediated alteration of positively sizedependent vulnerability to angling in response to historical fishing pressure in a freshwater salmonid. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 73 (3):461–468, March 2016. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2014-0571.

Bastardie:2016:CDD

[1101] Francois Bastardie, J. Rasmus Nielsen, and Tanja Miethe. Corrigendum: DISPLACE: a dynamic, individual-based model for spatial fishing planning and effort displacement — integrating underlying fish population models. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 73(3):469, March 2016. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
cjfas-2015-0569. See [749].

Anonymous:2016:SBA

[1102] Anonymous. Size-based approaches in aquatic ecology and fisheries science: a symposium in honour of Rob Peters / ^{≪} approches basées sur la taille en écologie aquatique et sciences halieutiques : un symposium en l’honneur de Rob Peters ^{≫}. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 73 (4), April 2016. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
cjfas-2016-0078.

Cyr:2016:RHP

[1103] Hélène Cyr. Robert Henry Peters (1946–1996). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 73(4), April 2016. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/cjfas-2015-0543.

Giacomini:2016:SBA

[1104] Henrique C. Giacomini, Brian J. Shuter, and Julia K. Baum. Size-based approaches to aquatic ecosystems and fisheries science: a symposium in honour of Rob Peters. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 73 (4):471–476, April 2016. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2016-0100.

Sprules:2016:SBS

[1105] William Gary Sprules and Lauren Emily Barth. Surfing the biomass size spectrum: some remarks on history, theory, and application. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 73(4):477–495, April 2016. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2015-0115.

Robinson:2016:TRD

[1106] James P. W. Robinson and Julia K. Baum. Trophic roles determine coral reef fish community size structure. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 73(4):496–505, April 2016. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/cjfas-2015-0178.

Mehner:2016:EFP

[1107] Thomas Mehner, Caroline Keeling, Matthias Emmrich, Kerstin Holmgren, Christine Argillier, Pietro Volta, Ian J. Winfield, and Sandra Brucet. Effects of fish predation on density and size spectra of prey fish communities in lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 73 (4):506–518, April 2016. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2015-0034.

Arranz:2016:DDE

[1108] Ignasi Arranz, Thomas Mehner, Lluís Benejam, Christine Argillier, Kerstin Holmgren, Erik Jeppesen, Torben L. Lauridsen, Pietro Volta, Ian J. Winfield, and Sandra Brucet. Density-dependent effects as key drivers of intraspecific size structure of six abundant fish species in lakes across Europe. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 73 (4):519–534, April 2016. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2014-0508.

Chu:2016:CPU

[1109] Cindy Chu, Nigel P. Lester, Henrique C. Giacomini, Brian J. Shuter, and Donald A. Jackson. Catch-per-unit-effort and size spectra of lake fish assemblages reflect underlying patterns in ecological conditions and anthropogenic activities across regional and local scales. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 73(4):535–546, April 2016. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2015-0150.

Benejam:2016:AEC

[1110] Lluís Benejam, Franco Teixeira de Mello, Mariana Meerhoff, Marcelo Loureiro, Erik Jeppesen, and Sandra Brucet. Assessing effects of change in land use on size-related variables of fish in subtropical streams. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 73(4):547–556, April 2016. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2015-0025.

Wheeland:2016:AML

[1111] Laura J. Wheeland and George A. Rose. Acoustic measures of lake community size spectra. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 73 (4):557–564, April 2016. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2014-0446.

deKerckhove:2016:ADF

[1112] Derrick Tupper de Kerckhove, Brian John Shuter, and Scott Milne. Acoustically derived fish size spectra within a lake and the statistical power to detect environmental change. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 73(4):565–574, April 2016. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/cjfas-2015-0222.

Andersen:2016:TFS

[1113] Ken H. Andersen, Nis S. Jacobsen, and K. D. Farnsworth. The theoretical foundations for size spectrum models of fish communities. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 73(4):575–588, April 2016. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2015-0230.

Spence:2016:PUD

[1114] Michael A. Spence, Paul G. Blackwell, and Julia L. Blanchard. Parameter uncertainty of a dynamic multispecies size spectrum model. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 73(4):589–597, April 2016. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2015-0022.

Datta:2016:ESP

[1115] Samik Datta and Julia L. Blanchard. The effects of seasonal processes on size spectrum dynamics. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 73 (4):598–610, April 2016. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2015-0468.

Canales:2016:SPS

[1116] T. Mariella Canales, Richard Law, and Julia L. Blanchard. Shifts in plankton size spectra modulate growth and coexistence of anchovy and sardine in upwelling systems. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 73 (4):611–621, April 2016. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2015-0181.

vanZwieten:2016:NPI

[1117] Paul A. M. van Zwieten, Jeppe Kolding, Michael J. Plank, Robert E. Hecky, Thomas B. Bridgeman, Sally MacIntyre, Ole Seehausen, and Greg M. Silsbe. The Nile perch invasion in Lake Victoria: cause or consequence of the haplochromine decline? Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 73(4):622–643, April 2016. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/cjfas-2015-0130.

Kolding:2016:MFY

[1118] Jeppe Kolding, Nis S. Jacobsen, Ken H. Andersen, and Paul A. M. van Zwieten. Maximizing fisheries yields while maintaining community structure. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 73(4):644–655, April 2016. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
cjfas-2015-0098.

McCann:2016:FWS

[1119] Kevin S. McCann, Gabriel Gellner, Bailey C. McMeans, Tina Deenik, Gordon Holtgrieve, Neil Rooney, Lee Hannah, Michael Cooperman, and So Nam. Food webs and the sustainability of indiscriminate fisheries. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 73(4):656–665, April 2016. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2015-0044.

Jacobsen:2016:CMP

[1120] Nis S. Jacobsen, Timothy E. Essington, and Ken H. Andersen. Comparing model predictions for ecosystem-based management. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 73(4):666–676, April 2016. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2014-0561.

Juanes:2016:LBA

[1121] Francis Juanes. A length-based approach to predator–prey relationships in marine predators. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 73 (4):677–684, April 2016. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2015-0159.

Arim:2016:ECF

[1122] Matías Arim, Ana I. Borthagaray, and Henrique Corrêa Giacomini. Energetic constraints to food chain length in a metacommunity framework. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 73(4):685–692, April 2016. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2015-0156.

Shuter:2016:FLH

[1123] Brian J. Shuter, Henrique C. Giacomini, Derrick de Kerckhove, and Kris Vascotto. Fish life history dynamics: shifts in prey size structure evoke shifts in predator maturation traits. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 73(4):693–708, April 2016. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/cjfas-2015-0190.

DeLong:2016:IBR

[1124] John P. DeLong and Matthew Walsh. The interplay between resource supply and demand determines the influence of predation on prey body size. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 73(4):709– 715, April 2016. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
cjfas-2015-0029.

vanderLee:2016:SFM

[1125] Adam S. van der Lee and Marten A. Koops. Are small fishes more sensitive to habitat loss? A generic size-based model. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 73(4):716–726, April 2016. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2015-0026.

Shi:2016:RPC

[1126] Xiaoli Shi, Xuhui Zhao, Min Zhang, Zhou Yang, Ping Xu, and Fanxiang Kong. The responses of phytoplankton communities to elevated CO_{2} show seasonal variations in the highly eutrophic Lake Taihu. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 73(5):727–736, May 2016. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2015-0151.

Haring:2016:DEQ

[1127] Michaela W. Haring, Tom A. Johnston, Murray D. Wiegand, Aaron T. Fisk, and Trevor E. Pitcher. Differences in egg quantity and quality among hatchery- and wild-origin Chinook salmon (Oncorhynchus tshawytscha). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 73(5):737–746, May 2016. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
cjfas-2015-0083.

Connelly:2016:ACL

[1128] Tara L. Connelly, Tara N. Businski, Don Deibel, Christopher C. Parrish, and Piotr Trela. Annual cycle of lipid content and lipid class composition in zooplankton from the Beaufort Sea shelf, Canadian Arctic. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 73(5):747–758, May 2016. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2015-0333.

Dodrill:2016:PSA

[1129] Michael J. Dodrill, Charles B. Yackulic, Theodore A. Kennedy, and John W. Hayes. Prey size and availability limits maximum size of rainbow trout in a large tailwater: insights from a drift-foraging bioenergetics model. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 73(5):759– 772, May 2016. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
cjfas-2015-0268.

Railsback:2016:MPR

[1130] Steven F. Railsback, Bret C. Harvey, Sarah J. Kupferberg, Margaret M. Lang, Scott McBain, and Hart H. Welsh. Modeling potential river management conflicts between frogs and salmonids. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 73(5):773–784, May 2016. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2015-0267.

Bohman:2016:PHN

[1131] Patrik Bohman, Lennart Edsman, Alfred Sandström, Per Nyström, Marika Stenberg, Pia Hertonsson, and Jacob Johansson. Predicting harvest of non-native signal crayfish in lakes — a role for changing climate?

Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 73(5):785–792, May 2016. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2015-0241.

Sinclair:2016:QDC

[1132] James S. Sinclair, Shelley E. Arnott, and Amelia Cox. The quick and the dead: copepods dominate as cladocerans decline following invasion by Hemimysis anomala. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 73 (5):793–803, May 2016. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2015-0111.

Larsen:2016:RRD

[1133] Martin H. Larsen, Jörgen I. Johnsson, Joacim Näslund, Søren T. Thomassen, and Kim Aarestrup. Reduced rearing density increases postrelease migration success of Atlantic salmon (Salmo salar) smolts. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 73(5):804–810, May 2016. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2014-0563.

Patin:2016:USS

[1134] Rémi Patin, Lauren A. Rogers, and Jan Ohlberger. Using a state-space population model to detect age-dependent species interactions. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 73(5):811–818, May 2016. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2015-0004.

Keeley:2016:BCE

[1135] Ernest R. Keeley, Steven O. Campbell, and André E. Kohler. Bioenergetic calculations evaluate changes to habitat quality for salmonid fishes in streams treated with salmon carcass analog. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 73(5):819–831, May 2016. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/cjfas-2015-0265.

Caldwell:2016:NNP

[1136] Timothy J. Caldwell, Frank M. Wilhelm, and Andy Dux. Non-native pelagic macroinvertebrate alters population dynamics of herbivorous zooplankton in a large deep lake. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 73 (5):832–843, May 2016. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2015-0144.

Evans:2016:EMG

[1137] Melissa L. Evans, Marc A. Johnson, Dave Jacobson, Jinliang Wang, Michael Hogansen, and Kathleen G. O’Malley. Evaluating a multigenerational reintroduction program for threatened salmon using genetic parentage analysis. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 73 (5):844–852, May 2016. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2015-0317.

Hershberger:2016:PUE

[1138] Paul K. Hershberger, Kyle A. Garver, and James R. Winton. Principles underlying the epizootiology of viral hemorrhagic septicemia in Pacific herring and other fishes throughout the North Pacific Ocean. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 73(5):853–859, May 2016. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2015-0417.

Riley:2016:MSD

[1139] Stephen C. Riley and Erin S. Dunlop. Misapplied survey data and model uncertainty result in incorrect conclusions about the role of predation on alewife population dynamics in Lake Huron: a comment on. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 73(5):860–864, May 2016. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:/

/cdnsciencepub.com/doi/full/10.1139/cjfas-2015-0237. See [889]

and reply [1140].

Bence:2016:RCR

[1140] James R. Bence, Charles P. Madenjian, Ji X. He, David G. Fielder, Steven A. Pothoven, Norine E. Dobiesz, James E. Johnson, Mark P. Ebener, R. Adam Cottrill, Lloyd C. Mohr, and Scott R. Koproski. Reply to comments by Riley and Dunlop on He et al. (2015). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 73(5):865–868, May 2016. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:/

/cdnsciencepub.com/doi/full/10.1139/cjfas-2015-0522. See [889,

1139].

Pink:2016:TIP

[1141] Melissa Pink and Mark V. Abrahams. Temperature and its impact on predation risk within aquatic ecosystems. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 73(6):869–876, June 2016. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/cjfas-2015-0302.

Porreca:2016:EER

[1142] Anthony P. Porreca, William D. Hintz, Gregory W. Whitledge, Neil P. Rude, Edward J. Heist, and James E. Garvey. Establishing ecologically relevant management boundaries: linking movement ecology with the conservation of Scaphirhynchus sturgeon. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 73(6):877–884, June 2016. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/cjfas-2015-0352.

Chang:2016:RSS

[1143] Jui-Han Chang, Yong Chen, William Halteman, and Carl Wilson. Roles of spatial scale in quantifying stock–recruitment relationships for American lobsters in the inshore Gulf of Maine. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 73(6):885–909, June 2016. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/cjfas-2015-0018.

LeRoy:2016:SCI

[1144] Carri J. LeRoy, Dylan G. Fischer, Walton M. Andrews, Lisa Belleveau, Clyde H. Barlow, Jennifer A. Schweitzer, Joseph K. Bailey, Jane C. Marks, and Jeff C. Kallestad. Salmon carcasses influence genetic linkages between forests and streams. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 73 (6):910–920, June 2016. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2015-0439.

Schroeder:2016:JLH

[1145] R. Kirk Schroeder, Luke D. Whitman, Brian Cannon, and Paul Olmsted. Juvenile life-history diversity and population stability of spring Chinook salmon in the Willamette River basin, Oregon. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 73(6):921–934, June 2016. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2015-0314.

Brooks:2016:RFE

[1146] Elizabeth N. Brooks and Christopher M. Legault. Retrospective forecasting — evaluating performance of stock projections for New England groundfish stocks. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 73(6):935– 950, June 2016. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
cjfas-2015-0163.

Hasselman:2016:GSC

[1147] Daniel J. Hasselman, Eric C. Anderson, Emily E. Argo, N. David Bethoney, Stephen R. Gephard, David M. Post, Bradley P. Schondelmeier, Thomas F. Schultz, Theodore V. Willis, and Eric P. Palkovacs. Genetic stock composition of marine bycatch reveals disproportional impacts on depleted river herring genetic stocks. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 73(6):951–963, June 2016. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/cjfas-2015-0402.

Jackson:2016:ECD

[1148] Leland J. Jackson, Lisa Corbett, and Garry Scrimgeour. Environmental constraints on Didymosphenia geminata occurrence and bloom formation in Canadian Rocky Mountain lotic systems. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 73(6):964–972, June 2016. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/cjfas-2015-0361.

Leadley:2016:UAV

[1149] Todd A. Leadley, Anne M. McLeod, Timothy B. Johnson, Daniel Heath, and Ken G. Drouillard. Uncovering adaptive versus acclimatized alterations in standard metabolic rate in brown bullhead (Ameiurus nebulosus). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 73(6):973–981, June 2016. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
cjfas-2015-0375.

OConnell-Milne:2016:ICH

[1150] Sorrel O’Connell-Milne, Candida Savage, and William Rayment. The influence of commercial harvesting on parasite infection in the bivalve Austrovenus stutchburyi. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 73 (6):982–989, June 2016. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2015-0401.

Nicola:2016:TFB

[1151] Graciela G. Nicola, Daniel Ayllón, Benigno Elvira, and Ana Almodóvar. Territorial and foraging behaviour of juvenile Mediterranean trout under changing conditions of food and competitors. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 73(6):990–998, June 2016. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/cjfas-2015-0304.

Kuczynski:2016:CRL

[1152] Anika Kuczynski, Martin T. Auer, Colin N. Brooks, and Amanda G. Grimm. The Cladophora resurgence in Lake Ontario: characterization and implications for management. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 73(6):999–1013, June 2016. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/cjfas-2015-0460.

Thorson:2016:CHV

[1153] James T. Thorson, Olaf P. Jensen, and Elise F. Zipkin. Corrigendum: How variable is recruitment for exploited marine fishes? A hierarchical model for testing life history theory. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 73(6):1014, June 2016. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/cjfas-2016-0167. See [799].

Quinn:2016:PPM

[1154] Thomas P. Quinn, Philip McGinnity, and Thomas E. Reed. The paradox of “premature migration” by adult anadromous salmonid fishes: patterns and hypotheses. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 73 (7):1015–1030, July 2016. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2015-0345.

Cao:2016:MMF

[1155] Yong Cao, Leon Hinz, Brian Metzke, Jeffrey Stein, and Ann Holtrop. Modeling and mapping fish abundance across wadeable streams of Illinois, USA, based on landscape-level environmental variables. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 73(7):1031–1046, July 2016. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2015-0343.

Roni:2016:IVE

[1156] Philip Roni, Christopher Johnson, Trenton De Boer, George Pess, Andrew Dittman, and David Sear. Interannual variability in the effects of physical habitat and parentage on Chinook salmon egg-to-fry survival. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 73(7):1047–1059, July 2016. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2015-0372.

Parker:2016:FDA

[1157] Aaron D. Parker, David C. Glover, Samuel T. Finney, P. Bradley Rogers, Jeffrey G. Stewart, and Robert L. Simmonds. Fish distribution, abundance, and behavioral interactions within a large electric dispersal barrier designed to prevent Asian carp movement. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 73(7):1060–1071, July 2016. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/cjfas-2015-0309.

Eberts:2016:CDR

[1158] Rebecca L. Eberts, Björn Wissel, Richard G. Manzon, Joanna Y. Wilson, Douglas R. Boreham, and Christopher M. Somers. Consistent differential resource use by sympatric lake (Coregonus clupeaformis) and round (Prosopium cylindraceum) whitefish in Lake Huron: a multi-time scale isotopic niche analysis. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 73 (7):1072–1080, July 2016. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2015-0324.

Wall:2016:NRE

[1159] C. Eric Wall, Nicolaas Bouwes, Joseph M. Wheaton, W. Carl Saunders, and Stephen N. Bennett. Net rate of energy intake predicts reachlevel steelhead (Oncorhynchus mykiss) densities in diverse basins from a large monitoring program. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 73 (7):1081–1091, July 2016. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2015-0290.

He:2016:UTV

[1160] Ji X. He, James R. Bence, Edward F. Roseman, David G. Fielder, and Mark P. Ebener. Using time-varying asymptotic length and body condition of top piscivores to indicate ecosystem regime shift in the main basin of Lake Huron: a Bayesian hierarchical modeling approach. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 73(7):1092–1103, July 2016. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2015-0235.

Bilodeau:2016:ANM

[1161] François Bilodeau, Roger Schetagne, Jean Therrien, and Richard Verdon. Absence of noticeable mercury effects on fish populations in boreal reservoirs despite threefold to sevenfold increases in mercury concentrations. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 73(7):1104–1125, July 2016. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2015-0035.

Mollet:2016:EIA

[1162] Fabian M. Mollet, Jan Jaap Poos, Ulf Dieckmann, and Adriaan D. Rijnsdorp. Evolutionary impact assessment of the North Sea plaice fishery. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 73(7):1126–1137, July 2016. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2014-0568.

Scott:2016:FCS

[1163] David C. Scott, Michael Arbeider, Jennifer Gordon, and Jonathan W. Moore. Flood control structures in tidal creeks associated with reduction in nursery potential for native fishes and creation of hotspots for invasive species. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 73(7):1138– 1148, July 2016. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
cjfas-2015-0311.

Pick:2016:BAC

[1164] Frances R. Pick. Blooming algae: a Canadian perspective on the rise of toxic cyanobacteria. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 73 (7):1149–1158, July 2016. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2015-0470.

Pardo:2016:MIR

[1165] Sebastián A. Pardo, Holly K. Kindsvater, John D. Reynolds, and Nicholas K. Dulvy. Maximum intrinsic rate of population increase in sharks, rays, and chimaeras: the importance of survival to maturity. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 73(8):1159–1163, August 2016. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2016-0069.

Bateman:2016:RFC

[1166] Andrew W. Bateman, Stephanie J. Peacock, Brendan Connors, Zephyr Polk, Dana Berg, Martin Krkošek, and Alexandra Morton. Recent failure to control sea louse outbreaks on salmon in the Broughton Archipelago, British Columbia. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 73(8):1164– 1172, August 2016. CODEN CJFSDX. ISSN 0706-652X (print), 1205-

7533 (electronic). URL https://cdnsciencepub.com/doi/full/10.
1139/cjfas-2016-0122.

Polymeropoulos:2016:MED

[1167] Elias. T. Polymeropoulos, Nicholas G. Elliott, and Peter B. Frappell. The maternal effect of differences in egg size influence metabolic rate and hypoxia induced hatching in Atlantic salmon eggs: implications for respiratory gas exchange across the egg capsule. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 73(8):1173–1181, August 2016. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2015-0358.

Hojesjo:2016:SRH

[1168] Johan Höjesjö, Rasmus Kaspersson, and John D. Armstrong. Sizerelated habitat use in juvenile Atlantic salmon: the importance of intercohort competition. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 73(8): 1182–1189, August 2016. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2015-0446.

Freshwater:2016:DIP

[1169] Cameron Freshwater, Marc Trudel, Terry D. Beacham, Lyse Godbout, Chrys-Ellen M. Neville, Strahan Tucker, and Francis Juanes. Disentangling individual- and population-scale processes within a latitudinal size gradient in sockeye salmon. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 73 (8):1190–1201, August 2016. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2015-0344.

Crossman:2016:CRD

[1170] Jill Crossman, M. Catherine Eimers, Shaun A. Watmough, Martyn N. Futter, Jason Kerr, Scott R. Baker, and Peter J. Dillon. Can recovery from disturbance explain observed declines in total phosphorus in Precambrian Shield catchments? Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 73 (8):1202–1212, August 2016. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2015-0312.

Thome:2016:ILW

[1171] Christopher Thome, Charles Mitz, Christopher M. Somers, Richard G. Manzon, Douglas R. Boreham, and Joanna Y. Wilson. Incubation of lake whitefish (Coregonus clupeaformis) embryos in cooling water discharge and the impacts of fluctuating thermal regimes on development. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 73(8):1213–1221, August 2016. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2015-0286.

Happel:2016:EQF

[1172] Austin Happel, Logan Stratton, Colleen Kolb, Chris Hays, Jacques Rinchard, and Sergiusz Czesny. Evaluating quantitative fatty acid signature analysis (QFASA) in fish using controlled feeding experiments. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 73(8):1222–1229, August 2016. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2015-0328.

Benoit:2016:GRA

[1173] Pierre-Olivier Benoît, Beatrix E. Beisner, and Christopher T. Solomon. Growth rate and abundance of common fishes is negatively related to dissolved organic carbon concentration in lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 73(8):1230–1236, August 2016. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2015-0340.

Cruz-Font:2016:ECA

[1174] Liset Cruz-Font, Brian J. Shuter, and Paul J. Blanchfield. Energetic costs of activity in wild lake trout: a calibration study using acceleration transmitters and positional telemetry. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 73(8):1237–1250, August 2016. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/cjfas-2015-0323.

Sethi:2016:SAM

[1175] Suresh Andrew Sethi and Catherine Bradley. Statistical arrival models to estimate missed passage counts at fish weirs. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 73(8):1251–1260, August 2016. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2015-0318.

Miller:2016:SSA

[1176] Timothy J. Miller, Jonathan A. Hare, and Larry A. Alade. A statespace approach to incorporating environmental effects on recruitment in an age-structured assessment model with an application to southern New England yellowtail flounder. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 73 (8):1261–1270, August 2016. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2015-0339.

Breivik:2016:LGM

[1177] Olav Nikolai Breivik, Geir Storvik, and Kjell Nedreaas. Latent Gaussian models to decide on spatial closures for bycatch management in the Barents Sea shrimp fishery. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 73 (8):1271–1280, August 2016. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2015-0322.

Groner:2016:QIS

[1178] Maya L. Groner, Gregor F. McEwan, Erin E. Rees, George Gettinby, and Crawford W. Revie. Quantifying the influence of salinity and temperature on the population dynamics of a marine ectoparasite. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 73(8):1281–1291, August 2016. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2015-0444.

Naman:2016:CCI

[1179] Sean M. Naman, Jordan S. Rosenfeld, and John S. Richardson. Causes and consequences of invertebrate drift in running waters: from individuals to populations and trophic fluxes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 73(8):1292–1305, August 2016. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/cjfas-2015-0363.

Boswell:2016:STP

[1180] Kevin M. Boswell, Guillaume Rieucau, Johanna J. Vollenweider, John R. Moran, Ron A. Heintz, Jason K. Blackburn, and David J. Csepp. Are spatial and temporal patterns in Lynn Canal overwintering Pacific herring related to top predator activity? Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 73(9):1307–1318, September 2016. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/cjfas-2015-0192.

Cook:2016:EGS

[1181] Robin M. Cook and Vanessa Trijoulet. The effects of grey seal predation and commercial fishing on the recovery of a depleted cod stock. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 73(9):1319–1329, September 2016. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2015-0423.

Fetzer:2016:YCV

[1182] William W. Fetzer, Collin J. Farrell, James R. Jackson, and Lars G. Rudstam. Year-class variation drives interactions between warm-water predators and yellow perch. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 73(9): 1330–1341, September 2016. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2015-0275.

Ou:2016:SHS

[1183] Chouly Ou and Kirk O. Winemiller. Seasonal hydrology shifts production sources supporting fishes in rivers of the Lower Mekong Basin. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 73(9):1342–1362, September 2016. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2015-0214.

Mahe:2016:IBW

[1184] Kélig Mahe, Clémence Oudard, Tiphaine Mille, James Keating, Patricia Gonçalves, Lotte Worsoe Clausen, Gróa Petursdottir, Helle Rasmussen, Elna Meland, Eugene Mullins, John K. Pinnegar, Åge Hoines, and Verena M. Trenkel. Identifying blue whiting (Micromesistius poutassou) stock structure in the Northeast Atlantic by otolith shape analysis. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 73(9):1363–1371, September 2016. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2015-0332.

Kurota:2016:EIP

[1185] Hiroyuki Kurota, Murdoch K. McAllister, Eric A. Parkinson, and N. T. Johnston. Evaluating the influence of predator–prey interactions on stock assessment and management reference points for a large lake ecosystem. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 73(9):1372–1388, September 2016. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
cjfas-2014-0414.

Laberge:2016:RBO

[1186] Frédéric Laberge, Nicholas Edmunds, Irene Yin-Liao, and Kevin S. Mc-Cann. Relationship between ontogenetic changes in foraging ecology and muscle lactate dehydrogenase activity in wild smallmouth bass (Micropterus dolomieu). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 73(9): 1389–1394, September 2016. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2015-0537.

Al-Chokhachy:2016:BTR

[1187] Robert Al-Chokhachy, David Schmetterling, Chris Clancy, Pat Saffel, Ryan Kovach, Leslie Nyce, Brad Liermann, Wade Fredenberg, and Ron Pierce. Are brown trout replacing or displacing bull trout populations in a changing climate? Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 73(9):1395– 1404, September 2016. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2015-0293.

Tsoukali:2016:ALR

[1188] Stavroula Tsoukali, Karin H. Olsson, André W. Visser, and Brian R. MacKenzie. Adult lifetime reproductive value in fish depends on size and fecundity type. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 73(9):1405– 1412, September 2016. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2015-0378.

Tsehaye:2016:EOW

[1189] Iyob Tsehaye, Brian M. Roth, and Greg G. Sass. Exploring optimal walleye exploitation rates for northern Wisconsin Ceded Territory lakes using a hierarchical Bayesian age-structured model. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 73(9):1413–1433, September 2016. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2015-0191.

Moore:2016:PNE

[1190] Jean-Sébastien Moore, Les N. Harris, Steven T. Kessel, Louis Bernatchez, Ross F. Tallman, and Aaron T. Fisk. Preference for nearshore and estuarine habitats in anadromous Arctic char (Salvelinus alpinus) from the Canadian high Arctic (Victoria Island, Nunavut) revealed by acoustic telemetry. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 73 (9):1434–1445, September 2016. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/

doi/full/10.1139/cjfas-2015-0436.

Depestele:2016:FLM

[1191] Jochen Depestele, Marie-Joëlle Rochet, Ghislain Dorémus, Pascal Laffargue, and Eric Willem Maria Stienen. Favorites and leftovers on the menu of scavenging seabirds: modelling spatiotemporal variation in discard consumption. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 73(9):1446– 1459, September 2016. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2015-0326.

Tentelier:2016:SUE

[1192] Cédric Tentelier, Nicolas Larranaga, Olivier Lepais, Aurélie Manicki, Jacques Rives, and Frédéric Lange. Space use and its effects on reproductive success of anadromous Atlantic salmon. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 73(10):1461–1471, October 2016. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2015-0518.

Harrison:2016:IEP

[1193] Joel W. Harrison, E. Todd Howell, Susan B. Watson, and Ralph E. H. Smith. Improved estimates of phytoplankton community composition based on in situ spectral fluorescence: use of ordination and field-derived norm spectra for the bbe FluoroProbe. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 73(10):1472–1482, October 2016. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2015-0360.

Miller:2016:UTD

[1194] Erin Miller, Marthe Larsen Haarr, and Rémy Rochette. Using temperature-dependent embryonic growth models to predict time of hatch of American lobster (Homarus americanus) in nature. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 73(10):1483–1492, October 2016. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2015-0355.

Allen:2016:PBM

[1195] David Allen, Jordan Rosenfeld, and Jeffrey Richards. Physiological basis of metabolic trade-offs between growth and performance among different strains of rainbow trout. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 73(10): 1493–1506, October 2016. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2015-0429.

Stanley:2016:RVO

[1196] Ryan R. E. Stanley, Claudio DiBacco, Simon R. Thorrold, Paul V. R. Snelgrove, Corey J. Morris, Robert S. Gregory, Steven E. Campana, and Ian R. Bradbury. Regional variation in otolith geochemistry of juvenile Atlantic cod (Gadus morhua) in coastal Newfoundland. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 73(10):1507–1519, October 2016. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2015-0353.

Myrvold:2016:JSM

[1197] Knut Marius Myrvold and Brian Patrick Kennedy. Juvenile steelhead movements in relation to stream habitat, population density, and body size: consequences for individual growth rates. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 73(10):1520–1529, October 2016. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2016-0007.

Lennox:2016:LTE

[1198] Preston A. Lennox and Joseph B. Rasmussen. Long-term effects of channelization on a cold-water stream community. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 73(10):1530–1537, October 2016. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2015-0561.

Hannan:2016:PES

[1199] Kelly D. Hannan, Jennifer D. Jeffrey, Caleb T. Hasler, and Cory D. Suski. Physiological effects of short- and long-term exposure to elevated carbon dioxide on a freshwater mussel, Fusconaia flava. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 73(10):1538–1546, October 2016. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2016-0083.

Tracey:2016:CIP

[1200] Sean R. Tracey, Klaas Hartmann, Melanie Leef, and Jaime McAllister. Capture-induced physiological stress and postrelease mortality for southern bluefin tuna (Thunnus maccoyii) from a recreational fishery. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 73(10):1547–1556, October 2016. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2015-0516.

Caron:2016:OIC

[1201] Antoine Caron, Pauline Pannetier, Maikel Rosabal, Hélène Budzinski, Mathilde Lauzent, Pierre Labadie, Bouchra Nasri, Fabien Pierron, Magalie Baudrimont, and Patrice Couture. Organic and inorganic contamination impacts on metabolic capacities in American and European yellow eels. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 73(10):1557–1566, October 2016. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
cjfas-2015-0473.

Dembkowski:2016:SLY

[1202] Daniel J. Dembkowski, David W. Willis, and Melissa R. Wuellner. Synchrony in larval yellow perch abundance: the influence of the Moran Effect during early life history. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 73(10):1567–1574, October 2016. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/cjfas-2015-0310.

Andrews:2016:AGB

[1203] Allen H. Andrews, Edward E. DeMartini, Jeff A. Eble, Brett M. Taylor, Dong Chun Lou, and Robert L. Humphreys. Age and growth of bluespine unicornfish (Naso unicornis): a half-century life-span for a keystone browser, with a novel approach to bomb radiocarbon dating in the Hawaiian Islands. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 73(10): 1575–1586, October 2016. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2016-0019.

Haggarty:2016:LRF

[1204] Dana R. Haggarty, Steve J. D. Martell, and Jonathan B. Shurin. Lack of recreational fishing compliance may compromise effectiveness of rockfish conservation areas in British Columbia. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 73(10):1587–1598, October 2016. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2015-0205.

Campana:2016:TMU

[1205] Steven E. Campana. Transboundary movements, unmonitored fishing mortality, and ineffective international fisheries management pose risks for pelagic sharks in the Northwest Atlantic. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 73(10):1599–1607, October 2016. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2015-0502.

Anonymous:2016:CER

[1206] Anonymous. Cross-ecosystem resource subsidies: from land to water and back again / Apports de ressources entre écosystèmes : de la terre à l’eau et vice versa. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 73(11), November 2016. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
cjfas-2016-0441.

Richardson:2016:GQA

[1207] John S. Richardson and Mark S. Wipfli. Getting quantitative about consequences of cross-ecosystem resource subsidies on recipient consumers. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 73(11):1609–1615, November 2016. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
cjfas-2016-0242.

Weaver:2016:SLC

[1208] Daniel M. Weaver, Stephen M. Coghlan, and Joseph Zydlewski. Sea lamprey carcasses exert local and variable food web effects in a nutrientlimited Atlantic coastal stream. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 73(11):1616–1625, November 2016. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/cjfas-2015-0506.

Rine:2016:TPS

[1209] Kristin M. Rine, Mark S. Wipfli, Erik R. Schoen, Timothy L. Nightengale, and Craig A. Stricker. Trophic pathways supporting juvenile Chinook and Coho salmon in the glacial Susitna River, Alaska: patterns of freshwater, marine, and terrestrial food resource use across a seasonally dynamic habitat mosaic. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 73(11): 1626–1641, November 2016. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2015-0555.

Li:2016:SSF

[1210] Judith L. Li, William J. Gerth, Richard P. Van Driesche, Doug S. Bateman, and Alan T. Herlihy. Seasonal and spatial fluctuations in Oncorhynchus trout diet in a temperate mixed-forest watershed. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 73(11):1642–1649, November 2016. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2015-0520.

Mejia:2016:LGS

[1211] Francine H. Mejia, Colden V. Baxter, Eric K. Berntsen, and Alexander K. Fremier. Linking groundwater–surface water exchange to food production and salmonid growth. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 73(11):1650–1660, November 2016. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/cjfas-2015-0535.

Smits:2016:TCS

[1212] Adrianne P. Smits, Daniel E. Schindler, Jonathan B. Armstrong, Michael T. Brett, Jackie L. Carter, and Bianca S. Santos. Thermal constraints on stream consumer responses to a marine resource subsidy. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 73(11):1661–1671, November 2016. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
cjfas-2015-0420.

Boucek:2016:OYD

[1213] Ross E. Boucek, Mariajesus Soula, Felipe Tamayo, and Jennifer S. Rehage. A once in 10 year drought alters the magnitude and quality of a floodplain prey subsidy to coastal river fishes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 73(11):1672–1678, November 2016. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2015-0507.

Roon:2016:IEB

[1214] David A. Roon, Mark S. Wipfli, Tricia L. Wurtz, and Arny L. Blanchard. Invasive European bird cherry (Prunus padus) reduces terrestrial prey subsidies to urban Alaskan salmon streams. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 73(11):1679–1690, November 2016. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2015-0548.

Piovia-Scott:2016:VRS

[1215] Jonah Piovia-Scott, Steven Sadro, Roland A. Knapp, James Sickman, Karen L. Pope, and Sudeep Chandra. Variation in reciprocal subsidies between lakes and land: perspectives from the mountains of California. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 73(11):1691–1701, November 2016. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
cjfas-2015-0549.

Anonymous:2016:NAN

[1216] Anonymous. Note of appreciation / note de reconnaissance. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 73(12), December 2016. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2016-0403.

Anonymous:2016:ACb

[1217] Anonymous. Announcement / Communiqué. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 73(12), December 2016. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/cjfas-2016-0471.

Siegfried:2016:ISA

[1218] Kate I. Siegfried, Erik H. Williams, Kyle W. Shertzer, and Lewis G. Coggins. Improving stock assessments through data prioritization. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 73(12):1703–1711, December 2016. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2015-0398.

Roley:2016:PWP

[1219] Sarah S. Roley and Jennifer L. Tank. Pore water physicochemical constraints on the endangered clubshell mussel (Pleurobema clava). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 73(12):1712–1722, December 2016. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2015-0442.

Freshwater:2016:DMB

[1220] Cameron Freshwater, Marc Trudel, Terry D. Beacham, Lyse Godbout, Chrys-Ellen M. Neville, Strahan Tucker, and Francis Juanes. Divergent migratory behaviours associated with body size and ocean entry phenology in juvenile sockeye salmon. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 73(12):1723–1732, December 2016. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/cjfas-2015-0425.

Louhi:2016:LTM

[1221] Pauliina Louhi, Teppo Vehanen, Ari Huusko, Aki Mäki-Petäys, and Timo Muotka. Long-term monitoring reveals the success of salmonid habitat restoration. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 73(12):1733–1741, December 2016. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
cjfas-2015-0546.

Langangen:2016:RSM

[1222] Øystein Langangen, Geir Ottersen, Lorenzo Ciannelli, Frode B. Vikebø, and Leif Christian Stige. Reproductive strategy of a migratory fish stock: implications of spatial variations in natural mortality. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 73(12):1742–1749, December 2016. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2015-0321.

Dunmall:2016:UTL

[1223] Karen M. Dunmall, Neil J. Mochnacz, Christian E. Zimmerman, Charles Lean, and James D. Reist. Using thermal limits to assess establishment of fish dispersing to high-latitude and high-elevation watersheds. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 73(12):1750–1758, December 2016. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2016-0051.

deEyto:2016:RNA

[1224] Elvira de Eyto, Catherine Dalton, Mary Dillane, Eleanor Jennings, Philip McGinnity, Barry O’Dwyer, Russell Poole, Ger Rogan, and David Taylor. The response of North Atlantic diadromous fish to multiple stressors, including land use change: a multidecadal study. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 73(12):1759–1769, December 2016. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2015-0450.

Monroy:2016:EAC

[1225] Silvia Monroy, Jon Miranda-Apodaca, Usue Pérez-López, Amaia Mena-Petite, Arturo Elosegi, Alberto Muñoz-Rueda, and Jesús Pozo. Elevated atmospheric CO_{2} interacts with drought and competition to produce complex results in plant quality and subsequent microbial aquatic decomposition. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 73(12):1770– 1776, December 2016. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2016-0034.

Keyser:2016:WPT

[1226] Freya M. Keyser, Jeremy E. Broome, Rodney G. Bradford, Brian Sanderson, and Anna M. Redden. Winter presence and temperature-related diel vertical migration of striped bass (Morone saxatilis) in an extreme high-flow passage in the inner Bay of Fundy. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 73(12):1777–1786, December 2016. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2016-0002.

Hordyk:2016:SLS

[1227] Adrian R. Hordyk, Kotaro Ono, Jeremy D. Prince, and Carl J. Walters. A simple length-structured model based on life history ratios and incorporating size-dependent selectivity: application to spawning potential ratios for data-poor stocks. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 73(12): 1787–1799, December 2016. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2015-0422.

Webb:2016:CFE

[1228] Joel B. Webb, Laura M. Slater, Ginny L. Eckert, and Gordon H. Kruse. The contribution of fecundity and embryo quality to reproductive potential of eastern Bering Sea snow crab (Chionoecetes opilio). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 73(12):1800–1814, December 2016. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2015-0485.

Smith:2016:DPR

[1229] Sara E. Smith, Shawn P. Sitar, Frederick W. Goetz, Mar Huertas, Brandon M. Armstrong, and Cheryl A. Murphy. Differential physiological response to sea lamprey parasitism between lake trout (Salvelinus namaycush) morphotypes from Lake Superior. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 73(12):1815–1829, December 2016. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2015-0448.

Bloomer:2016:EOD

[1230] Jack Bloomer, David Sear, Peter Dutey-Magni, and Paul Kemp. The effects of oxygen depletion due to upwelling groundwater on the posthatch fitness of Atlantic salmon (Salmo salar). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 73(12):1830–1840, December 2016. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2016-0189.

Samsing:2016:SLL

[1231] Francisca Samsing, Frode Oppedal, Sussie Dalvin, Ingrid Johnsen, Tone Vågseth, and Tim Dempster. Salmon lice (Lepeophtheirus salmonis) development times, body size, and reproductive outputs follow universal models of temperature dependence. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 73(12):1841–1851, December 2016. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/cjfas-2016-0050. See corrigendum [1453].

Willmes:2016:IRA

[1232] Malte Willmes, Justin J. G. Glessner, Scott A. Carleton, Paul C. Gerrity, and James A. Hobbs. ^{87}Sr/^{86}Sr isotope ratio analysis by laser ablation MC-ICP-MS in scales, spines, and fin rays as a nonlethal alternative to otoliths for reconstructing fish life history. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 73(12):1852–1860, December 2016. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2016-0103.

Neuswanger:2016:MFT

[1233] Jason R. Neuswanger, Mark S. Wipfli, Amanda E. Rosenberger, and Nicholas F. Hughes. Measuring fish and their physical habitats: versatile 2D and 3D video techniques with user-friendly software. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 73(12):1861–1873, December 2016. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2016-0010.

Nadon:2016:SSS

[1234] Marc O. Nadon and Jerald S. Ault. A stepwise stochastic simulation approach to estimate life history parameters for data-poor fisheries. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 73(12):1874–1884, December 2016. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2015-0303.

Maheu:2016:RAI

[1235] Audrey Maheu, André St-Hilaire, Daniel Caissie, Nassir El-Jabi, Guillaume Bourque, and Daniel Boisclair. A regional analysis of the impact of dams on water temperature in medium-size rivers in eastern Canada. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 73(12):1885–1897, December 2016. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
cjfas-2015-0486.

Bateman:2016:ESA

[1236] Douglas S. Bateman, Matthew R. Sloat, Robert E. Gresswell, Aaron M. Berger, David P. Hockman-Wert, David W. Leer, and Arne E. Skaugset.

Effects of stream-adjacent logging in fishless headwaters on downstream coastal cutthroat trout. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 73(12): 1898–1913, December 2016. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2015-0455.

Jech:2016:TBP

[1237] J. Michael Jech and Ian H. McQuinn. Towards a balanced presentation and objective interpretation of acoustic and trawl survey data, with specific reference to the eastern Scotian Shelf. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 73(12):1914–1921, December 2016. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2016-0113.

Gallagher:2017:BRF

[1238] Austin J. Gallagher, Erica R. Staaterman, Steven J. Cooke, and Neil Hammerschlag. Behavioural responses to fisheries capture among sharks caught using experimental fishery gear. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 74(1):1–7, January 2017. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/cjfas-2016-0165.

Gibson-Reinemer:2017:DAN

[1239] Daniel K. Gibson-Reinemer, Brian S. Ickes, and John H. Chick. Development and assessment of a new method for combining catch per unit effort data from different fish sampling gears: multigear mean standardization (MGMS). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 74(1):8–14, January 2017. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
cjfas-2016-0003.

Katinic:2017:CDM

[1240] Peter J. Katinić, David A. Patterson, and Ronald C. Ydenberg. Condition dependence in the marine exit timing of sockeye salmon (Oncorhynchus nerka) returning to Copper Creek, Haida Gwaii. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 74(1):15–22, January 2017. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2015-0533.

Izzo:2017:SRE

[1241] Christopher Izzo, Zoë A. Doubleday, Gretchen L. Grammer, Morgan C. F. Disspain, Qifeng Ye, and Bronwyn M. Gillanders. Seasonally resolved environmental reconstructions using fish otoliths. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 74(1):23–31, January 2017. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2016-0055.

Louhi:2017:SAR

[1242] Pauliina Louhi, John S. Richardson, and Timo Muotka. Sediment addition reduces the importance of predation on ecosystem functions in experimental stream channels. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 74 (1):32–40, January 2017. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2015-0530.

Richardson:2017:FMD

[1243] David P. Richardson, André E. Kohler, Million Hailemichael, and Bruce P. Finney. The fate of marine-derived nutrients: tracing δ ^{13}C and δ ^{15}N through oligotrophic freshwater and linked riparian ecosystems following salmon carcass analog additions. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 74(1):41–55, January 2017. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/cjfas-2015-0500.

French:2017:BTS

[1244] W. E. French, B. Vondracek, L. C. Ferrington, J. C. Finlay, and D. J. Dieterman. Brown trout (Salmo trutta) growth and condition along a winter thermal gradient in temperate streams. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 74(1):56–64, January 2017. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/cjfas-2016-0005.

Cisneros-Montemayor:2017:TID

[1245] Andrés Miguel Cisneros-Montemayor, William Wai Lung Cheung, Karin Bodtker, Louise Teh, Nadja Steiner, Megan Bailey, Carie Hoover, and Ussif Rashid Sumaila. Towards an integrated database on Canadian ocean resources: benefits, current states, and research gaps. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 74(1):65–74, January 2017. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2015-0573.

Plichard:2017:CES

[1246] Laura Plichard, Hervé Capra, Raphaël Mons, Hervé Pella, and Nicolas Lamouroux. Comparing electrofishing and snorkelling for characterizing fish assemblages over time and space. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 74(1):75–86, January 2017. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/cjfas-2015-0578.

Xu:2017:TPL

[1247] Meng Xu, Jeppe Kolding, and Joel E. Cohen. Taylor’s power law and fixed-precision sampling: application to abundance of fish sampled by gillnets in an African lake. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 74 (1):87–100, January 2017. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2016-0009.

Dala-Corte:2017:IMP

[1248] Renato Bolson Dala-Corte, Fernando Gertum Becker, and Adriano Sanches Melo. The importance of metacommunity processes for long-term turnover of riffle-dwelling fish assemblages depends on spatial position within a dendritic network. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 74(1):101–115, January 2017. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/cjfas-2016-0049.

Fernandez-Chacon:2017:CMD

[1249] Albert Fernández-Chacón, Even Moland, Sigurd Heiberg Espeland, Alf Ring Kleiven, and Esben Moland Olsen. Causes of mortality in depleted populations of Atlantic cod estimated from multi-event modelling of mark–recapture and recovery data. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 74(1):116–126, January 2017. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/cjfas-2015-0313.

Tunnah:2017:DPD

[1250] Louise Tunnah, Suzanne Currie, and Tyson J. MacCormack. Do prior diel thermal cycles influence the physiological response of Atlantic salmon (Salmo salar) to subsequent heat stress? Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 74(1):127–139, January 2017. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/cjfas-2016-0157.

Lambert:2017:IER

[1251] Max R. Lambert, Aaron B. Stoler, Meredith S. Smylie, Rick A. Relyea, and David K. Skelly. Interactive effects of road salt and leaf litter on wood frog sex ratios and sexual size dimorphism. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 74(2):141–146, February 2017. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2016-0324.

Isaak:2017:SPE

[1252] Daniel J. Isaak, Jay M. Ver Hoef, Erin E. Peterson, Dona L. Horan, and David E. Nagel. Scalable population estimates using spatial-stream-network (SSN) models, fish density surveys, and national geospatial database frameworks for streams. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 74(2):147–156, February 2017. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/cjfas-2016-0247.

Lamperth:2017:LSE

[1253] James S. Lamperth, Thomas P. Quinn, and Mara S. Zimmerman. Levels of stored energy but not marine foraging patterns differentiate seasonal ecotypes of wild and hatchery steelhead (Oncorhynchus mykiss) returning to the Kalama River, Washington. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 74(2):157–167, February 2017. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/cjfas-2016-0018.

Moriarty:2017:NME

[1254] Pamela E. Moriarty, Timothy E. Essington, and Eric J. Ward. A novel method to estimate prey contributions to predator diets. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 74(2):168–177, February 2017. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2016-0149.

Zydlewski:2017:HCA

[1255] Joseph Zydlewski, Daniel Stich, and Douglas Sigourney. Hard choices in assessing survival past dams — a comparison of single- and pairedrelease strategies. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 74(2):178–190, February 2017. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
cjfas-2015-0480.

Jeffrey:2017:CBS

[1256] Kyla M. Jeffrey, Isabelle M. Côté, James R. Irvine, and John D. Reynolds. Changes in body size of Canadian Pacific salmon over six decades. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 74(2):191–201, February 2017. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
cjfas-2015-0600.

Foldvik:2017:SDM

[1257] A. Foldvik, S. Einum, and A. G. Finstad. Spatial diffusion modelling of juvenile Atlantic salmon (Salmo salar) shows ontogenetic increase in movement rates. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 74(2): 202–207, February 2017. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2015-0315.

Bennett:2017:NMN

[1258] Joseph R. Bennett, Kathleen M. Rühland, and John P. Smol. No magic number: determining cost-effective sample size and enumeration effort for diatom-based environmental assessment analyses. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 74(2):208–215, February 2017. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2016-0066.

Trystram:2017:FPT

[1259] Clément Trystram, Karyne M. Rogers, Marc Soria, and Sébastien Jaquemet. Feeding patterns of two sympatric shark predators in coastal ecosystems of an oceanic island. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 74 (2):216–227, February 2017. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2016-0105.

Perkin:2017:LHT

[1260] Joshuah S. Perkin, Natalie E. Knorp, Thomas C. Boersig, Amy E. Gebhard, Lucas A. Hix, and Thomas C. Johnson. Life history theory predicts long-term fish assemblage response to stream impoundment. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 74(2):228–239, February 2017. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2015-0593.

Grothues:2017:HFS

[1261] Thomas M. Grothues, Arthur E. Newhall, James F. Lynch, Kaela S. Vogel, and Glen G. Gawarkiewicz. High-frequency side-scan sonar fish reconnaissance by autonomous underwater vehicles. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 74(2):240–255, February 2017. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2015-0301.

Lauffenburger:2017:CBT

[1262] Nathan Lauffenburger, Alex De Robertis, and Stan Kotwicki. Combining bottom trawls and acoustics in a diverse semipelagic environment: What is the contribution of walleye pollock (Gadus chalcogrammus) to near-bottom acoustic backscatter in the eastern Bering Sea? Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 74(2):256–264, February 2017. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2015-0481.

Franssen:2017:ITT

[1263] Nathan R. Franssen, Eliza I. Gilbert, Angela P. James, and Jason E. Davis. Isotopic tissue turnover and discrimination factors following a laboratory diet switch in Colorado pikeminnow (Ptychocheilus lucius). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 74(2):265–272, February 2017. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2015-0531.

Berejikian:2017:RSA

[1264] Barry A. Berejikian, Jeffrey J. Hard, Christopher P. Tatara, Donald M. Van Doornik, Penny Swanson, and Donald A. Larsen. Rearing strategies alter patterns of size-selective mortality and heritable size variation in steelhead trout (Oncorhynchus mykiss). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 74(2):273–283, February 2017. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2016-0175.

Favaro:2017:ANL

[1265] Brett Favaro and Martin Olszynski. Authorized net losses of fish habitat demonstrate need for improved habitat protection in Canada. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 74(3):285–291, March 2017. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2016-0480.

Kessel:2017:FDL

[1266] S. T. Kessel, N. E. Hussey, R. E. Crawford, D. J. Yurkowski, D. M. Webber, T. A. Dick, and A. T. Fisk. First documented large-scale horizontal movements of individual Arctic cod (Boreogadus saida). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 74(3):292–296, March 2017. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2016-0196.

Cupp:2017:RIS

[1267] Aaron R. Cupp, Richard A. Erickson, Kim T. Fredricks, Nicholas M. Swyers, Tyson W. Hatton, and Jon J. Amberg. Responses of invasive silver and bighead carp to a carbon dioxide barrier in outdoor ponds. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 74(3):297–305, March 2017. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2015-0472.

Bunnell:2017:TSR

[1268] David B. Bunnell, Tomas O. Höök, Cary D. Troy, Wentao Liu, Charles P. Madenjian, and Jean V. Adams. Testing for synchrony in recruitment among four Lake Michigan fish species. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 74(3):306–315, March 2017. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/cjfas-2015-0534.

Dzul:2017:ITH

[1269] M. C. Dzul, C. B. Yackulic, J. Korman, M. D. Yard, and J. D. Muehlbauer. Incorporating temporal heterogeneity in environmental conditions into a somatic growth model. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 74(3):316–326, March 2017. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/cjfas-2016-0056.

Vaha:2017:GSI

[1270] Juha-Pekka Vähä, Jaakko Erkinaro, Morten Falkegård, Panu Orell, and Eero Niemelä. Genetic stock identification of Atlantic salmon and its evaluation in a large population complex. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 74(3):327–338, March 2017. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/cjfas-2015-0606.

Sala:2017:ASS

[1271] Antonello Sala, Jure Brčić, Bent Herrmann, Alessandro Lucchetti, and Massimo Virgili. Assessment of size selectivity in hydraulic clam dredge fisheries. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 74(3):339–348, March 2017. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
cjfas-2015-0199.

Cao:2017:IAPa

[1272] Jie Cao, Yong Chen, and R. Anne Richards. Improving assessment of Pandalus stocks using a seasonal, size-structured assessment model with environmental variables. part i: Model description and application. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 74(3):349–362, March 2017. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2016-0020.

Cao:2017:IAPb

[1273] Jie Cao, Yong Chen, and R. Anne Richards. Improving assessment of Pandalus stocks using a seasonal, size-structured assessment model with environmental variables. Part II: Model evaluation and simulation. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 74(3):363–376, March 2017. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2016-0021.

Jasonowicz:2017:LOY

[1274] Andrew J. Jasonowicz, Frederick W. Goetz, Giles W. Goetz, and Krista M. Nichols. Love the one you’re with: genomic evidence of panmixia in the sablefish (Anoplopoma fimbria). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 74(3):377–387, March 2017. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/cjfas-2016-0012.

Scheld:2017:SFS

[1275] Andrew M. Scheld and Christopher M. Anderson. Selective fishing and shifting production in multispecies fisheries. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 74(3):388–395, March 2017. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/cjfas-2015-0494.

Rosengren:2017:BWE

[1276] Malin Rosengren, Eli Kvingedal, Joacim Näslund, Jörgen I. Johnsson, and Kristina Sundell. Born to be wild: effects of rearing density and environmental enrichment on stress, welfare, and smolt migration in hatchery-reared Atlantic salmon. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 74(3):396–405, March 2017. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/cjfas-2015-0515.

Hitt:2017:BTU

[1277] Nathaniel P. Hitt, Erin L. Snook, and Danielle L. Massie. Brook trout use of thermal refugia and foraging habitat influenced by brown trout. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 74(3):406–418, March 2017. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2016-0255.

Carlisle:2017:IEU

[1278] Aaron B. Carlisle, Steven Y. Litvin, Daniel J. Madigan, Kady Lyons, Jennifer S. Bigman, Melissa Ibarra, and Joseph J. Bizzarro. Interactive effects of urea and lipid content confound stable isotope analysis in elasmobranch fishes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 74 (3):419–428, March 2017. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2015-0584.

McKinney:2017:MMS

[1279] Garrett J. McKinney, James E. Seeb, and Lisa W. Seeb. Managing mixed-stock fisheries: genotyping multi-SNP haplotypes increases power for genetic stock identification. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 74(4):429–434, April 2017. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/cjfas-2016-0443.

Zhu:2017:CTA

[1280] Feng-Jiao Zhu, Kun Hu, Zong-Ying Yang, and Xian-Le Yang. Comparative transcriptome analysis of the hepatopancreas of Eriocheir sinensis following oral gavage with enrofloxacin. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 74(4):435–444, April 2017. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/cjfas-2016-0041.

Podgorniak:2017:SPF

[1281] T. Podgorniak, M. Angelini, E. De Oliveira, F. Daverat, and F. Pierron. Selective pressure of fishways upon morphological and muscle enzymatic traits of migrating glass eels. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 74 (4):445–451, April 2017. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2016-0110.

Lima:2017:ITT

[1282] Ana Carolina Lima, Diogo Sayanda, Amadeu M. V. M. Soares, Frederick J. Wrona, and Kieran A. Monaghan. Integrating taxonomic and trait analyses to assess the impact of damming on fish communities in a northern cold region river. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 74 (4):452–463, April 2017. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2016-0074.

Ward:2017:EEP

[1283] Hillary G. M. Ward, John R. Post, Nigel P. Lester, Paul J. Askey, and Theresa Godin. Empirical evidence of plasticity in life-history characteristics across climatic and fish density gradients. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 74(4):464–474, April 2017. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2016-0023.

Westgaard:2017:GPS

[1284] Jon-Ivar Westgaard, Atal Saha, Matthew Kent, Hanne Hellerud Hansen, Halvor Knutsen, Lorenz Hauser, Steven X. Cadrin, Ole Thomas Albert, and Torild Johansen. Genetic population structure in Greenland halibut (Reinhardtius hippoglossoides) and its relevance to fishery management. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 74(4):475–485, April 2017. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2015-0430.

Krag:2017:IEP

[1285] Ludvig A. Krag, Bent Herrmann, Jordan Feekings, Henrik S. Lund, and Junita D. Karlsen. Improving escape panel selectivity in Nephropsdirected fisheries by actively stimulating fish behavior. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 74(4):486–493, April 2017. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2015-0568.

Lanier:2017:UYU

[1286] Wendy E. Lanier, Kevin R. Bestgen, W. Chris Funk, and Larissa L. Bailey. Unpalatable, yet unprotected: trout reduce survival and development rate of rare toad tadpoles despite chemical defense. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 74(4):494–502, April 2017. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2016-0231.

Jones:2017:TER

[1287] Nicholas E. Jones and Bastian J. Schmidt. Tributary effects in rivers: interactions of spatial scale, network structure, and landscape characteristics. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 74(4):503–510, April 2017. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
cjfas-2015-0493.

Laurel:2017:RSJ

[1288] Benjamin J. Laurel, David Cote, Robert S. Gregory, Lauren Rogers, Halvor Knutsen, and Esben Moland Olsen. Recruitment signals in juvenile cod surveys depend on thermal growth conditions. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 74(4):511–523, April 2017. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2016-0035.

Edwards:2017:SOW

[1289] Kristin M. Edwards, Gary W. Rice, and S. Laurie Sanderson. Separating oil from water: suspension-feeding goldfish ingest liquid vegetable oil. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 74(4):524–532, April 2017. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2016-0197.

Crook:2017:TSV

[1290] David A. Crook, Katherine Lacksen, Alison J. King, Duncan J. Buckle, Steven J. Tickell, Jonathon D. Woodhead, Roland Maas, Simon A. Townsend, and Michael M. Douglas. Temporal and spatial variation in strontium in a tropical river: implications for otolith chemistry analyses of fish migration. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 74 (4):533–545, April 2017. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2016-0153.

Steel:2017:CLL

[1291] E. Ashley Steel, Ariel Muldoon, Rebecca L. Flitcroft, Julie C. Firman, Kara J. Anlauf-Dunn, Kelly M. Burnett, and Robert J. Danehy. Current landscapes and legacies of land-use past: understanding the distribution of juvenile Coho salmon (Oncorhynchus kisutch) and their habitats along the Oregon Coast, USA. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 74 (4):546–561, April 2017. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2015-0589.

Wennerstrom:2017:TSW

[1292] Lovisa Wennerström, Jens Olsson, Nils Ryman, and Linda Laikre. Temporally stable, weak genetic structuring in brackish water northern pike (Esox lucius) in the Baltic Sea indicates a contrasting divergence pattern relative to freshwater populations. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 74(4):562–571, April 2017. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/cjfas-2016-0039.

Warburton:2017:VDP

[1293] Manna L. Warburton, Malcolm R. Reid, Claudine H. Stirling, and Gerry Closs. Validation of depth-profiling LA-ICP-MS in otolith applications. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 74(4):572–581, April 2017. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2016-0063.

Perez-Rodriguez:2017:DFC

[1294] A. Pérez-Rodríguez, D. Howell, M. Casas, F. Saborido-Rey, and A. Ávila de Melo. Dynamic of the Flemish Cap commercial stocks: use of a gadget multispecies model to determine the relevance and synergies among predation, recruitment, and fishing. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 74(4):582–597, April 2017. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/cjfas-2016-0111.

Holmes:2017:LEN

[1295] Teslin G. Holmes, William M. Tonn, Cynthia A. Paszkowski, and Garry J. Scrimgeour. Limited effects of non-native trout on microcrustacean zooplankton in boreal foothills lakes: comparisons of stocked, unstocked, and fishless lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 74 (4):598–608, April 2017. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2015-0495.

Twining:2017:NLA

[1296] Cornelia W. Twining, Eric P. Palkovacs, Maya A. Friedman, Daniel J. Hasselman, and David M. Post. Nutrient loading by anadromous fishes: species-specific contributions and the effects of diversity. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 74(4):609–619, April 2017. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2016-0136.

Krkosek:2017:PBI

[1297] Martin Krkošek. Population biology of infectious diseases shared by wild and farmed fish. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 74(4):620– 628, April 2017. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
cjfas-2016-0379.

Pepino:2017:ILS

[1298] Marc Pépino, Marco A. Rodríguez, and Pierre Magnan. Incorporating lakes in stream fish habitat models: are we missing a key landscape attribute? Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 74(5):629– 635, May 2017. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
cjfas-2016-0221.

Kallasvuo:2017:MSD

[1299] Meri Kallasvuo, Jarno Vanhatalo, and Lari Veneranta. Modeling the spatial distribution of larval fish abundance provides essential information for management. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 74 (5):636–649, May 2017. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2016-0008.

Smith:2017:SBR

[1300] Stephen J. Smith, Jessica A. Sameoto, and Craig J. Brown. Setting biological reference points for sea scallops (Placopecten magellanicus) allowing for the spatial distribution of productivity and fishing effort. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 74(5):650–667, May 2017. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2015-0595.

Reed:2017:AHH

[1301] Thomas E. Reed, Elvira de Eyto, Kealan O’Higgins, Paddy Gargan, William Roche, Jonathan White, Niall O’Maoileidigh, Thomas P. Quinn, and Phil McGinnity. Availability of holding habitat in lakes and rivers affects the incidence of spring (premature) upriver migration by Atlantic salmon. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 74(5):668– 679, May 2017. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
cjfas-2016-0191.

Ashworth:2017:AGR

[1302] Eloïse C. Ashworth, Norman G. Hall, S. Alex Hesp, Peter G. Coulson, and Ian C. Potter. Age and growth rate variation influence the functional relationship between somatic and otolith size. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 74(5):680–692, May 2017. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/cjfas-2015-0471.

Chen:2017:EFE

[1303] Kuan-Yu Chen, Stuart A. Ludsin, Morgan M. Corey, Paris D. Collingsworth, Megan K. Nims, John W. Olesik, Konrad Dabrowski, Jason J. van Tassell, and Elizabeth A. Marschall. Experimental and field evaluation of otolith strontium as a marker to discriminate between riverspawning populations of walleye in Lake Erie. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 74(5):693–701, May 2017. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/cjfas-2015-0565.

Mauger:2017:STR

[1304] Sue Mauger, Rebecca Shaftel, Jason C. Leppi, and Daniel J. Rinella. Summer temperature regimes in southcentral Alaska streams: watershed drivers of variation and potential implications for Pacific salmon. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 74(5):702–715, May 2017. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2016-0076.

Quick:2017:ESE

[1305] Nicola Quick, Lindesay Scott-Hayward, Dina Sadykova, Doug Nowacek, and Andrew Read. Effects of a scientific echo sounder on the behavior of short-finned pilot whales (Globicephala macrorhynchus). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 74(5):716–726, May 2017. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2016-0293.

Wall:2017:DMW

[1306] C. Eric Wall, Nicolaas Bouwes, Joseph M. Wheaton, Stephen N. Bennett, W. Carl Saunders, Pete A. McHugh, and Chris E. Jordan. Design and monitoring of woody structures and their benefits to juvenile steelhead (Oncorhynchus mykiss) using a net rate of energy intake model. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 74(5):727–738, May 2017. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2016-0131.

Yochum:2017:URI

[1307] Noëlle Yochum, Allan W. Stoner, David B. Sampson, Craig Rose, Alan Pazar, and Robert Eder. Utilizing reflex impairment to assess the role of discard mortality in “Size, sex, and season” management for Oregon Dungeness crab (Cancer magister) fisheries. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 74(5):739–750, May 2017. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/cjfas-2016-0029.

Perala:2017:ENR

[1308] Tommi A. Perälä, Douglas P. Swain, and Anna Kuparinen. Examining nonstationarity in the recruitment dynamics of fishes using Bayesian change point analysis. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 74 (5):751–765, May 2017. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2016-0177.

Eger:2017:TSS

[1309] Aaron M. Eger, Janelle M. R. Curtis, Marie-Josée Fortin, Isabelle M. Côté, and Frédéric Guichard. Transferability and scalability of species distribution models: a test with sedentary marine invertebrates. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 74(5):766–778, May 2017. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2016-0129.

Albertsen:2017:COL

[1310] Christoffer Moesgaard Albertsen, Anders Nielsen, and Uffe Høgsbro Thygesen. Choosing the observational likelihood in state-space stock assessment models. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 74 (5):779–789, May 2017. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2015-0532.

Castro-Santos:2017:MBA

[1311] Theodore Castro-Santos, Xiaotao Shi, and Alex Haro. Migratory behavior of adult sea lamprey and cumulative passage performance through four fishways. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 74(5):790–800, May 2017. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
cjfas-2016-0089.

Fellman:2017:LLS

[1312] Jason B. Fellman, Brian Buma, Eran Hood, Richard T. Edwards, and David V. D’Amore. Linking LiDAR with streamwater biogeochemistry in coastal temperate rainforest watersheds. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 74(6):801–811, June 2017. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/cjfas-2016-0130.

Takasuka:2017:DGB

[1313] Akinori Takasuka, Ayumu Sakai, and Ichiro Aoki. Dynamics of growthbased survival mechanisms in Japanese anchovy (Engraulis japonicus) larvae. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 74(6):812–823, June 2017. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
cjfas-2016-0120.

Sorensen:2017:NAI

[1314] Michael L. Sorensen and Donn K. Branstrator. The North American invasive zooplanktivore Bythotrephes longimanus is less hypoxia-tolerant than the native Leptodora kindtii. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 74(6):824–832, June 2017. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/

doi/full/10.1139/cjfas-2016-0188.

Neumann:2017:QPB

[1315] Viola Neumann, Matthias Schaber, Margit Eero, Uwe Böttcher, and Friedrich W. Köster. Quantifying predation on Baltic cod early life stages. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 74(6):833–842, June 2017. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
cjfas-2016-0215.

Harris:2017:EOD

[1316] Julianne E. Harris and Jeffrey C. Jolley. Estimation of occupancy, density, and abundance of larval lampreys in tributary river mouths upstream of dams on the Columbia River, Washington and Oregon. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 74(6):843–852, June 2017. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2016-0212.

Haworth:2017:FWT

[1317] Matthew R. Haworth and Kevin R. Bestgen. Flow and water temperature affect reproduction and recruitment of a Great Plains cyprinid. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 74(6):853–863, June 2017. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2016-0238.

Horbowy:2017:EMD

[1318] Jan Horbowy and Anna Luzeńczyk. Effects of multispecies and densitydependent factors on MSY reference points: example of the Baltic Sea sprat. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 74(6):864–870, June 2017. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
cjfas-2016-0220.

Guzzo:2017:CCA

[1319] Matthew M. Guzzo and Paul J. Blanchfield. Climate change alters the quantity and phenology of habitat for lake trout (Salvelinus namaycush) in small Boreal Shield lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 74 (6):871–884, June 2017. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2016-0190.

Goerig:2017:MIB

[1320] Elsa Goerig and Theodore Castro-Santos. Is motivation important to brook trout passage through culverts? Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 74(6):885–893, June 2017. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/cjfas-2016-0237.

Camaclang:2017:MIP

[1321] Abbey E. Camaclang, Janelle M. R. Curtis, Ilona Naujokaitis-Lewis, Mark S. Poesch, and Marten A. Koops. Modelling the impact of poaching on metapopulation viability for data-limited species. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 74(6):894–906, June 2017. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2015-0508.

Thomas:2017:HST

[1322] Austen C. Thomas, Benjamin W. Nelson, Monique M. Lance, Bruce E. Deagle, and Andrew W. Trites. Harbour seals target juvenile salmon of conservation concern. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 74 (6):907–921, June 2017. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2015-0558.

Froehlich:2017:WDH

[1323] Halley E. Froehlich, Timothy E. Essington, and P. Sean McDonald. When does hypoxia affect management performance of a fishery? A management strategy evaluation of Dungeness crab (Metacarcinus magister) fisheries in Hood Canal, Washington, USA. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 74(6):922–932, June 2017. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/cjfas-2016-0269.

Kai:2017:EHS

[1324] Mikihiko Kai and Hiroki Yokoi. Evaluation of harvest strategies for pelagic sharks taking ecological characteristics into consideration: an example for North Pacific blue shark. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 74(6):933–947, June 2017. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/cjfas-2016-0170.

Lagarde:2017:TVL

[1325] Raphaël Lagarde, Nils Teichert, Henri Grondin, Hélène Magalon, Agathe Pirog, and Dominique Ponton. Temporal variability of larval drift of tropical amphidromous gobies along a watershed in Réunion Island. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 74(6):948–957, June 2017. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2016-0101.

Miller:2017:EOD

[1326] Michael J. Miller and Katsumi Tsukamoto. The ecology of oceanic dispersal and survival of anguillid leptocephali. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 74(6):958–971, June 2017. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/cjfas-2016-0281.

Record:2017:OSS

[1327] Nicholas R. Record, John D. O’Brien, Karen Stamieszkin, and Jeffrey A. Runge. Omic-style statistical clustering reveals old and new patterns in the Gulf of Maine ecosystem. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 74 (7):973–979, July 2017. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2016-0151.

Studinski:2017:ERD

[1328] Jered M. Studinski, Andrew W. Hafs, Jonathan M. Niles, and Kyle J. Hartman. The effects of riparian disturbance on the condition and summer diets of age-0 brook trout (Salvelinus fontinalis) in three central Appalachian streams. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 74 (7):980–986, July 2017. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2016-0301.

Mateo:2017:CCE

[1329] Maria Mateo, Patrick Lambert, Stéphane Tetard, Martin Castonguay, Bruno Ernande, and Hilaire Drouineau. Cause or consequence? Exploring the role of phenotypic plasticity and genetic polymorphism in the emergence of phenotypic spatial patterns of the European eel. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 74(7):987–999, July 2017. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2016-0214.

Krieger:2017:DEH

[1330] Joseph R. Krieger and James S. Diana. Development and evaluation of a habitat suitability model for young lake sturgeon (Acipenser fulvescens) in the North Channel of the St. Clair River, Michigan. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 74(7):1000–1008, July 2017. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2016-0357.

Baumann:2017:MBI

[1331] Zofia Baumann, Robert P. Mason, David O. Conover, Prentiss Balcom, Celia Y. Chen, Kate L. Buckman, Nicholas S. Fisher, and Hannes Baumann. Mercury bioaccumulation increases with latitude in a coastal marine fish (Atlantic silverside, Menidia menidia). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 74(7):1009–1015, July 2017. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2016-0396. See cor-

rigendum [1561].

Schmiing:2017:PMR

[1332] Mara Schmiing, Jorge Fontes, and Pedro Afonso. Predictive mapping of reproductive fish habitats to aid marine conservation planning. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 74(7):1016–1027, July 2017. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2015-0538.

Wiedenmann:2017:EAB

[1333] John Wiedenmann, Michael Wilberg, Andrea Sylvia, and Thomas Miller. An evaluation of acceptable biological catch (ABC) harvest control rules designed to limit overfishing. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 74 (7):1028–1040, July 2017. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2016-0381.

Muehlbauer:2017:DEN

[1334] Jeffrey D. Muehlbauer, Theodore A. Kennedy, Adam J. Copp, and Thomas A. Sabol. Deleterious effects of net clogging on the quantification of stream drift. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 74 (7):1041–1048, July 2017. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2016-0365.

Pierce:2017:RDC

[1335] Andrew L. Pierce, Joseph W. Blodgett, Timothy D. Cavileer, Lea R. Medeiros, Josh Boyce, Lucius K. Caldwell, William J. Bosch, Ryan Branstetter, David E. Fast, Douglas R. Hatch, and James J. Nagler. Reproductive development in captive reconditioned female steelhead kelts: evidence for consecutive and skip spawning life histories. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 74(7):1049–1060, July 2017. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2016-0065.

Harris:2017:ITM

[1336] Julianne E. Harris and Joseph E. Hightower. An integrated tagging model to estimate mortality rates of Albemarle Sound–Roanoke River striped bass. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 74(7):1061– 1076, July 2017. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
cjfas-2016-0141.

Debertin:2017:MGP

[1337] Allan J. Debertin, James R. Irvine, Carrie A. Holt, Gladys Oka, and Marc Trudel. Marine growth patterns of southern British Columbia chum salmon explained by interactions between density-dependent competition and changing climate. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 74 (7):1077–1087, July 2017. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2016-0265.

Ashworth:2017:NPB

[1338] Eloïse C. Ashworth, S. Alex Hesp, and Norman G. Hall. A new proportionality-based back-calculation approach, which employs traditional forms of growth equations, improves estimates of length at age. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 74(7):1088–1099, July 2017. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2016-0409.

Cortes:2017:HPS

[1339] Federico Cortés and Juan A. Waessle. Hotspots for porbeagle shark (Lamna nasus) bycatch in the southwestern Atlantic (51^{o}S–57^{o}S). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 74(7):1100–1110, July 2017. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2016-0114.

Dauphin:2017:SWV

[1340] Guillaume J. R. Dauphin, Catherine Brugel, Marion Legrand, and Etienne Prévost. Separating wild versus stocking components in fish recruitment without identification data: a hierarchical modelling approach. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 74(7):1111–1124, July 2017. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2015-0443.

Zhang:2017:REB

[1341] Fan Zhang, Kevin B. Reid, and Thomas D. Nudds. Relative effects of biotic and abiotic factors during early life history on recruitment dynamics: a case study. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 74 (7):1125–1134, July 2017. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2016-0155.

Gibeau:2017:RRH

[1342] P. Gibeau, B. M. Connors, and W. J. Palen. Run-of-river hydropower and salmonids: potential effects and perspective on future research. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 74(7):1135–1149, July 2017. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2016-0253.

Cooke:2017:DCM

[1343] Emma L. L. Cooke, Alexander D. M. Wilson, Chris K. Elvidge, and Steven J. Cooke. Does capture method or the presence of aquatic protected areas influence the selective harvest of behavioural types in largemouth bass? Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 74(8): 1151–1157, August 2017. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2016-0192.

Crane:2017:TBC

[1344] Derek P. Crane and John M. Farrell. Trends in body condition of smallmouth bass and northern pike (1982–2013) following multiple ecological perturbations in the St. Lawrence River. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 74(8):1158–1172, August 2017. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2016-0160.

Chasco:2017:ECS

[1345] Brandon Chasco, Isaac C. Kaplan, Austen Thomas, Alejandro Acevedo-Gutiérrez, Dawn Noren, Michael J. Ford, M. Bradley Hanson, Jonathan Scordino, Steve Jeffries, Scott Pearson, Kristin N. Marshall, and Eric J. Ward. Estimates of Chinook salmon consumption in Washington State inland waters by four marine mammal predators from 1970 to 2015. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 74(8):1173–1194, August 2017. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2016-0203.

Jones:2017:COF

[1346] William A. Jones and David M. Checkley. Classification of otoliths of fishes common in the Santa Barbara Basin based on morphology and chemical composition. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 74(8): 1195–1207, August 2017. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2015-0566.

Naman:2017:HSP

[1347] Sean M. Naman, Jordan S. Rosenfeld, Laura C. Third, and John S. Richardson. Habitat-specific production of aquatic and terrestrial invertebrate drift in small forest streams: implications for drift-feeding fish.

Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 74(8):1208–1217, August 2017. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2016-0406.

Rahikainen:2017:IEO

[1348] Mika Rahikainen, Kirsi-Maaria Hoviniemi, Samu Mäntyniemi, Jarno Vanhatalo, Inari Helle, Maiju Lehtiniemi, Jukka Pönni, and Sakari Kuikka. Impacts of eutrophication and oil spills on the Gulf of Finland herring stock. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 74(8): 1218–1232, August 2017. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2016-0108.

Amoroso:2017:MNB

[1349] Ricardo O. Amoroso, Michael D. Tillotson, and Ray Hilborn. Measuring the net biological impact of fisheries enhancement: pink salmon hatcheries can increase yield, but with apparent costs to wild populations. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 74(8):1233–1242, August 2017. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
cjfas-2016-0334.

Fryer:2017:MAV

[1350] R. J. Fryer, K. Summerbell, and F. G. O’Neill. A meta-analysis of vertical stratification in demersal trawl gears. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 74(8):1243–1250, August 2017. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/cjfas-2016-0391.

Samarasin:2017:OSE

[1351] Pasan Samarasin, Scott M. Reid, and Nicholas E. Mandrak. Optimal sampling effort required to characterize wetland fish communities. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 74(8):1251–1259, August 2017. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2016-0424.

Nguyen:2017:SSE

[1352] Vivian M. Nguyen, Jill L. Brooks, Nathan Young, Robert J. Lennox, Neal Haddaway, Frederick G. Whoriskey, Robert Harcourt, and Steven J. Cooke. To share or not to share in the emerging era of big data: perspectives from fish telemetry researchers on data sharing. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 74(8):1260–1274, August 2017. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2016-0261.

Kendall:2017:DPP

[1353] Neala W. Kendall, Gary W. Marston, and Matthew M. Klungle. Declining patterns of Pacific Northwest steelhead trout (Oncorhynchus mykiss) adult abundance and smolt survival in the ocean. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 74(8):1275–1290, August 2017. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2016-0486.

Blanke:2017:CCS

[1354] Chelsey M. Blanke, Yoshito Chikaraishi, Yuko Takizawa, Shawn A. Steffan, Prarthana S. Dharampal, and M. Jake Vander Zanden. Comparing compound-specific and bulk stable nitrogen isotope trophic discrimination factors across multiple freshwater fish species and diets. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 74(8):1291–1297, August 2017. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2016-0420.

Remmal:2017:FMC

[1355] Y. Remmal, C. Hudon, P. B. Hamilton, M. Rondeau, and P. Gagnon. Forecasting the magnitude and composition of phytoplankton blooms in a eutrophic lowland river (Rivière Yamaska, Que., Canada). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 74(8):1298–1311, August 2017. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2016-0305.

Pavlov:2017:DMM

[1356] Dmitrii S. Pavlov and Victor N. Mikheev. Downstream migration and mechanisms of dispersal of young fish in rivers. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 74(8):1312–1323, August 2017. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2016-0298.

Manhard:2017:EFI

[1357] Christopher V. Manhard, John E. Joyce, William W. Smoker, and Anthony J. Gharrett. Ecological factors influencing lifetime productivity of pink salmon (Oncorhynchus gorbuscha) in an Alaskan stream. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 74(9):1325–1336, September 2017. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2016-0335.

Brunsdon:2017:DDD

[1358] Eric Blake Brunsdon, Dylan John Fraser, William Rundle Ardren, and James William Angus Grant. Dispersal and density-dependent growth of Atlantic salmon (Salmo salar) juveniles: clumped versus dispersed stocking. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 74(9):1337–1347, September 2017. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
cjfas-2015-0488.

Marriott:2017:ESI

[1359] Ross J. Marriott, Berwin A. Turlach, Kevin Murray, and David V. Fairclough. Evaluation of spatiotemporal imputations for fishing catch rate standardization. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 74(9):1348– 1361, September 2017. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2016-0182.

Evans:2017:FCA

[1360] Nathan T. Evans, Yiyuan Li, Mark A. Renshaw, Brett P. Olds, Kristy Deiner, Cameron R. Turner, Christopher L. Jerde, David M. Lodge, Gary A. Lamberti, and Michael E. Pfrender. Fish community assessment with eDNA metabarcoding: effects of sampling design and bioinformatic filtering. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 74(9):1362– 1374, September 2017. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2016-0306.

Phillips:2017:PPI

[1361] Elizabeth M. Phillips, John K. Horne, and Jeannette E. Zamon. Predator–prey interactions influenced by a dynamic river plume. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 74(9):1375–1390, September 2017. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2016-0302.

Beacham:2017:PII

[1362] Terry D. Beacham, Colin Wallace, Cathy MacConnachie, Kim Jonsen, Brenda McIntosh, John R. Candy, Robert H. Devlin, and Ruth E. Withler. Population and individual identification of Coho salmon in British Columbia through parentage-based tagging and genetic stock identification: an alternative to coded-wire tags. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 74(9):1391–1410, September 2017. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2016-0452.

Weldrick:2017:SDN

[1363] Christine K. Weldrick and Dennis E. Jelinski. Seasonal dynamics in a nearshore isotopic niche and spatial subsidies from multi-trophic aquaculture. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 74(9):1411–1421, September 2017. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
cjfas-2016-0372.

VanBeveren:2017:PPI

[1364] Elisabeth Van Beveren, Jean-Marc Fromentin, Sylvain Bonhommeau, Anne-Elise Nieblas, Luisa Metral, Blandine Brisset, Marko Jusup, Robert Klaus Bauer, Pablo Brosset, and Claire Saraux. Predator– prey interactions in the face of management regulations: changes in Mediterranean small pelagic species are not due to increased tuna predation. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 74(9):1422–1430, September 2017. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
cjfas-2016-0152.

Bell:2017:CIT

[1365] Donovan A. Bell, Ryan P. Kovach, Scott C. Vulstek, John E. Joyce, and David A. Tallmon. Climate-induced trends in predator–prey synchrony differ across life-history stages of an anadromous salmonid. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 74(9):1431–1438, September 2017. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2016-0309.

Verschoor:2017:IIL

[1366] Mark J. Verschoor, Chistopher R. Powe, Eric McQuay, Sherry L. Schiff, Jason J. Venkiteswaran, Jiahua Li, and Lewis A. Molot. Internal iron loading and warm temperatures are preconditions for cyanobacterial dominance in embayments along Georgian Bay, Great Lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 74(9):1439–1453, September 2017. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2016-0377.

Li:2017:TPA

[1367] Yan Li, Hong Zhu Wang, Xiao Min Liang, Qing Yu, Xu Cheng Xiao, Jian Chun Shao, and Hai Jun Wang. Total phytoplankton abundance is determined by phosphorus input: evidence from an 18-month fertilization experiment in four subtropical ponds. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 74(9):1454–1461, September 2017. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/cjfas-2016-0057.

Ramos:2017:LFD

[1368] S. Ramos, C. B. Paris, and M. M. Angélico. Larval fish dispersal along an estuarine–ocean gradient. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 74 (9):1462–1473, September 2017. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/

doi/full/10.1139/cjfas-2016-0325.

Glas:2017:MDR

[1369] Martin Glas, Michael Tritthart, Bernhard Zens, Hubert Keckeis, Aaron Lechner, Timothy Kaminskas, and Helmut Habersack. Modelling the dispersal of riverine fish larvae: from a raster-based analysis of movement patterns within a racetrack flume to a rheoreaction-based correlated random walk (RCRW) model approach. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 74(9):1474–1489, September 2017. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/cjfas-2016-0287.

Kincaid:2017:ECB

[1370] Kate Kincaid and George Rose. Effects of closing bottom trawling on fisheries, biodiversity, and fishing communities in a boreal marine ecosystem: the Hawke Box off Labrador, Canada. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 74(9):1490–1502, September 2017. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/cjfas-2016-0343.

Oresland:2017:RCA

[1371] Vidar Øresland, Mats Ulmestrand, Ann-Lisbeth Agnalt, and Gert Oxby. Recorded captures of American lobster (Homarus americanus) in Swedish waters and an observation of predation on the European lobster (Homarus gammarus). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 74(10): 1503–1506, October 2017. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2016-0532.

Price:2017:CWS

[1372] Michael H. H. Price, Karl K. English, Andrew G. Rosenberger, Misty MacDuffee, and John D. Reynolds. Canada’s wild salmon policy: an assessment of conservation progress in British Columbia. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 74(10):1507–1518, October 2017. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2017-0127.

Manhard:2017:EPS

[1373] Christopher V. Manhard, John E. Joyce, and Anthony J. Gharrett. Evolution of phenology in a salmonid population: a potential adaptive response to climate change. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 74(10):1519–1527, October 2017. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/cjfas-2017-0028.

Eshaghzadeh:2017:CBM

[1374] Hamid Eshaghzadeh, Carles Alcaraz, Arash Akbarzadeh, and Enric Gisbert. The combination of bivariate and multivariate methods to analyze character synchronization and early allometric growth patterns in the stellate sturgeon (Acipenser stellatus) as tools for better understanding larval behavior. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 74(10): 1528–1537, October 2017. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2016-0288.

Ward-Campbell:2017:FAA

[1375] Belinda Ward-Campbell, Karl Cottenie, Nicholas E. Mandrak, and Robert McLaughlin. Fish assemblages in agricultural drains are resilient to habitat change caused by drain maintenance. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 74(10):1538–1548, October 2017. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2016-0361.

Quinn:2017:PCA

[1376] Brady K. Quinn, Joël Chassé, and Rémy Rochette. Potential connectivity among American lobster fisheries as a result of larval drift across the species’ range in eastern North America. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 74(10):1549–1563, October 2017. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2016-0416.

Baker:2017:OSM

[1377] Leanne F. Baker, Kyle J. Artym, and Heidi K. Swanson. Optimal sampling methods for modelling the occupancy of Arctic grayling (Thymallus arcticus) in the Canadian barrenlands. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 74(10):1564–1574, October 2017. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2016-0429.

Davison:2017:LHE

[1378] Raziel J. Davison and William H. Satterthwaite. Life history effects on hatchery contributions to ocean harvest and natural-area spawning. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 74(10):1575–1587, October 2017. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2016-0457.

Brewitt:2017:HEC

[1379] Kimberly S. Brewitt, Eric M. Danner, and Jonathan W. Moore. Hot eats and cool creeks: juvenile Pacific salmonids use mainstem prey while in thermal refuges. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 74(10): 1588–1602, October 2017. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2016-0395.

Baker:2017:FOS

[1380] Cindy F. Baker, Don J. Jellyman, Kathryn Reeve, Shannan Crow, Michael Stewart, Tyler Buchinger, and Weiming Li. First observations of spawning nests in the pouched lamprey (Geotria australis). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 74(10):1603–1611, October 2017. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2016-0292.

Bret:2017:UIR

[1381] Victor Bret, Hervé Capra, Véronique Gouraud, Nicolas Lamouroux, Jérémy Piffady, Laurence Tissot, and Etienne Rivot. Understanding inter-reach variation in brown trout (Salmo trutta) mortality rates using a hierarchical Bayesian state-space model. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 74(10):1612–1627, October 2017. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2016-0240.

Chen:2017:FPN

[1382] Bingyao Chen, Lin Wang, Hui Wang, Shanshan Li, Thomas A. Jefferson, Lian Wang, Erin K. Tang, Xinrong Xu, and Guang Yang. Finless porpoises (Neophocaena asiaeorientalis) in the East China Sea: insights into feeding habits using morphological, molecular, and stable isotopic techniques. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 74(10):1628–1645, October 2017. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
cjfas-2016-0119.

Whitlock:2017:UGP

[1383] S. L. Whitlock, L. D. Schultz, C. B. Schreck, and J. E. Hess. Using genetic pedigree reconstruction to estimate effective spawner abundance from redd surveys: an example involving Pacific lamprey (Entosphenus tridentatus). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 74(10): 1646–1653, October 2017. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2016-0154.

Winegardner:2017:CDD

[1384] Amanda K. Winegardner, Natasha Salter, Stéphane Aebischer, Reinhard Pienitz, Alison M. Derry, Boswell Wing, Beatrix E. Beisner, and Irene Gregory-Eaves. Cladoceran diversity dynamics in lakes from a northern mining region: responses to multiple stressors characterized by alpha and beta diversity. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 74(10): 1654–1667, October 2017. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2016-0449.

Deslauriers:2017:DFM

[1385] David Deslauriers, Alex J. Rosburg, and Steven R. Chipps. Development of a foraging model framework to reliably estimate daily food consumption by young fishes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 74(10): 1668–1681, October 2017. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2016-0331.

Weber:2017:QSS

[1386] Michael J. Weber, Michael L. Brown, David H. Wahl, and Daniel E. Shoup. Quantifying the spatial scale of common carp (Cyprinus carpio) recruitment synchrony. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 74(10): 1682–1691, October 2017. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2016-0254.

Anonymous:2017:SOR

[1387] Anonymous. Space oddity: Recent advances incorporating spatial processes in the fishery stock assessment and management interface / Odyssée spatiale : avancées récentes intégrant des processus spatiaux dans l’interface d’évaluation et de gestion des stocks halieutiques. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 74(11), November 2017. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2017-0401.

Berger:2017:ISO

[1388] Aaron M. Berger, Daniel R. Goethel, and Patrick D. Lynch. Introduction to “Space Oddity: Recent Advances Incorporating Spatial Processes in the Fishery Stock Assessment and Management Interface”. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 74(11):1693–1697, November 2017. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2017-0296.

Berger:2017:SOM

[1389] Aaron M. Berger, Daniel R. Goethel, Patrick D. Lynch, Terrance Quinn, Sophie Mormede, Jeremy McKenzie, and Alistair Dunn. Space oddity: The mission for spatial integration. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 74(11):1698–1716, November 2017. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/cjfas-2017-0150.

Turner:2017:ESD

[1390] Sara M. Turner, Jonathan A. Hare, John P. Manderson, David E. Richardson, and John J. Hoey. Evaluation of species distribution forecasts: a potential predictive tool for reducing incidental catch in pelagic fisheries. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 74(11):1717–1731, November 2017. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
cjfas-2016-0274.

Miller:2017:SVC

[1391] Rebecca R. Miller, John C. Field, Jarrod A. Santora, Melissa H. Monk, Rosemary Kosaka, and Cynthia Thomson. Spatial valuation of California marine fisheries as an ecosystem service. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 74(11):1732–1748, November 2017. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2016-0228.

Ducharme-Barth:2017:CAV

[1392] Nicholas D. Ducharme-Barth and Robert N. M. Ahrens. Classification and analysis of VMS data in vertical line fisheries: incorporating uncertainty into spatial distributions. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 74(11):1749–1764, November 2017. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/cjfas-2016-0181.

Kai:2017:SVS

[1393] Mikihiko Kai, James T. Thorson, Kevin R. Piner, and Mark N. Maunder. Spatiotemporal variation in size-structured populations using fishery data: an application to shortfin mako (Isurus oxyrinchus) in the Pacific Ocean. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 74(11):1765– 1780, November 2017. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2016-0327.

Cao:2017:SIS

[1394] Jie Cao, James T. Thorson, R. Anne Richards, and Yong Chen. Spatiotemporal index standardization improves the stock assessment of northern shrimp in the Gulf of Maine. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 74(11):1781–1793, November 2017. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/cjfas-2016-0137.

Thorson:2017:ASV

[1395] James T. Thorson, Robert Fonner, Melissa A. Haltuch, Kotaro Ono, and Henning Winker. Accounting for spatiotemporal variation and fisher targeting when estimating abundance from multispecies fishery data. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 74(11):1794–1807, November 2017. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2015-0598.

Cadigan:2017:SMS

[1396] Noel G. Cadigan, Elmer Wade, and Anders Nielsen. A spatiotemporal model for snow crab (Chionoecetes opilio) stock size in the southern Gulf of St. Lawrence. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 74(11): 1808–1820, November 2017. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2016-0260.

Langseth:2017:CPL

[1397] Brian J. Langseth and Amy M. Schueller. Calculation of populationlevel fishing mortality for single- versus multi-area models: application to models with spatial structure. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 74(11):1821–1831, November 2017. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/cjfas-2016-0295.

Lee:2017:EAM

[1398] Hui-Hua Lee, Kevin R. Piner, Mark N. Maunder, Ian G. Taylor, and Richard D. Methot, Jr. Evaluation of alternative modelling approaches to account for spatial effects due to age-based movement. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 74(11):1832–1844, November 2017. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2016-0294.

Lauretta:2017:RBT

[1399] Matthew V. Lauretta and Daniel R. Goethel. The robustness of Brownie tag return models to complex spatiotemporal dynamics evaluated through simulation analysis. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 74(11):1845–1861, November 2017. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/cjfas-2016-0291.

Liu:2017:VHM

[1400] Chang Liu, Geoffrey W. Cowles, Douglas R. Zemeckis, Steven X. Cadrin, and Micah J. Dean. Validation of a hidden Markov model for the geolocation of Atlantic cod. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 74(11): 1862–1877, November 2017. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2016-0376.

Goethel:2017:ASC

[1401] Daniel R. Goethel and Aaron M. Berger. Accounting for spatial complexities in the calculation of biological reference points: effects of misdiagnosing population structure for stock status indicators. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 74(11):1878–1894, November 2017. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2016-0290.

deMoor:2017:QUP

[1402] Carryn L. de Moor, Douglas S. Butterworth, and Carl D. van der Lingen. The quantitative use of parasite data in multistock modelling of South African sardine (Sardinops sagax). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 74(11):1895–1903, November 2017. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/cjfas-2016-0280.

Truesdell:2017:EUC

[1403] Samuel B. Truesdell, Deborah R. Hart, and Yong Chen. Effects of unequal capture probability on stock assessment abundance and mortality estimates: an example using the US Atlantic sea scallop fishery. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 74(11):1904–1917, November 2017. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2016-0296.

Denson:2017:DNS

[1404] LaTreese S. Denson, David B. Sampson, and Andi Stephens. Data needs and spatial structure considerations in stock assessments with regional differences in recruitment and exploitation. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 74(11):1918–1929, November 2017. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2016-0277.

Vincent:2017:STR

[1405] Matthew T. Vincent, Travis O. Brenden, and James R. Bence. Simulation testing the robustness of a multiregion, tag-integrated assessment model that exhibits natal homing and estimates natural mortality and reporting rate. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 74(11):1930– 1949, November 2017. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2016-0297.

Little:2017:EES

[1406] L. Richard Little, André E. Punt, Geoffrey N. Tuck, and Bruce D. Mapstone. Exploring the effect of sampling, protogyny, and larval advection on stock estimates subject to no-take closures in a spatially complex coral reef line fishery on the Great Barrier Reef, Australia. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 74(11):1950–1959, November 2017. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2016-0349.

Punt:2017:EMC

[1407] André E. Punt, Malcolm Haddon, L. Richard Little, and Geoffrey N. Tuck. The effect of marine closures on a feedback control management strategy used in a spatially aggregated stock assessment: a case study based on pink ling in Australia. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 74(11):1960–1973, November 2017. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/cjfas-2016-0017.

Kapur:2017:SFC

[1408] Maia R. Kapur and Erik C. Franklin. Simulating future climate impacts on tropical fisheries: are contemporary spatial fishery management strategies sufficient? Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 74(11): 1974–1989, November 2017. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2016-0200.

Kerr:2017:MIS

[1409] Lisa A. Kerr, Steven X. Cadrin, David H. Secor, and Nathan G. Taylor. Modeling the implications of stock mixing and life history uncertainty of Atlantic bluefin tuna. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 74(11): 1990–2004, November 2017. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2016-0067.

Anonymous:2017:NAN

[1410] Anonymous. Note of appreciation / note de reconnaissance. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 74(12), December 2017. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2017-0384.

Orihel:2017:IPL

[1411] Diane M. Orihel, Helen M. Baulch, Nora J. Casson, Rebecca L. North, Chris T. Parsons, Dalila C. M. Seckar, and Jason J. Venkiteswaran. Internal phosphorus loading in Canadian fresh waters: a critical review and data analysis. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 74(12):2005– 2029, December 2017. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2016-0500.

Klobucar:2017:FEA

[1412] Stephen L. Klobucar, Torrey W. Rodgers, and Phaedra Budy. At the forefront: evidence of the applicability of using environmental DNA to quantify the abundance of fish populations in natural lentic waters with additional sampling considerations. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 74(12):2030–2034, December 2017. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/cjfas-2017-0114.

Dockery:2017:SPS

[1413] David R. Dockery, Thomas E. McMahon, Kevin M. Kappenman, and Matthew Blank. Swimming performance of sauger (Sander canadensis) in relation to fish passage. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 74(12):2035–2044, December 2017. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/cjfas-2016-0410.

Meckley:2017:HFF

[1414] Trevor D. Meckley, Eliezer Gurarie, James R. Miller, and C. Michael Wagner. How fishes find the shore: evidence for orientation to bathymetry from the non-homing sea lamprey. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 74(12):2045–2058, December 2017. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2016-0412.

Feiner:2017:SBT

[1415] Zachary S. Feiner, Stephen C. Chong, David G. Fielder, James A. Hoyle, Carey Knight, Thomas E. Lauer, Michael V. Thomas, Jeff T. Tyson, and Tomas O. Höök. Sex-based trade-offs among growth, mortality, and maturation in Great Lakes yellow perch stocks. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 74(12):2059–2072, December 2017. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2016-0173.

Novaczek:2017:HRS

[1416] Emilie Novaczek, Rodolphe Devillers, Evan Edinger, and Luiz Mello. High-resolution seafloor mapping to describe coastal denning habitat of a Canadian species at risk: Atlantic wolffish (Anarhichas lupus). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 74(12):2073–2084, December 2017. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2016-0414.

Rous:2017:SMB

[1417] Andrew M. Rous, Adrienne R. McLean, Jessica Barber, Gale Bravener, Theodore Castro-Santos, Christopher M. Holbrook, Istvan Imre, Thomas C. Pratt, and Robert L. McLaughlin. Spatial mismatch between sea lamprey behaviour and trap location explains low success at trapping for control. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 74(12): 2085–2097, December 2017. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2016-0445.

Stanistreet:2017:UPA

[1418] Joy E. Stanistreet, Douglas P. Nowacek, Simone Baumann-Pickering, Joel T. Bell, Danielle M. Cholewiak, John A. Hildebrand, Lynne E. W. Hodge, Hilary B. Moors-Murphy, Sofie M. Van Parijs, and Andrew J. Read. Using passive acoustic monitoring to document the distribution of beaked whale species in the western North Atlantic Ocean. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 74(12):2098–2109, December 2017. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2016-0503.

Page:2017:CSI

[1419] Henry M. Page, Scott D. Cooper, Sheila W. Wiseman, Danuta Bennett, Kristie Klose, Steven Sadro, Craig Nelson, and Thomas Even. Comparisons of stable isotope (C, H, N) signatures for revealing organic matter sources and trophic relationships in headwater streams. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 74(12):2110–2121, December 2017. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2016-0322.

Bardarson:2017:OSD

[1420] Hlynur Bardarson, Bruce. J. McAdam, Vilhjalmur Thorsteinsson, Einar Hjorleifsson, and Gudrun Marteinsdottir. Otolith shape differences between ecotypes of Icelandic cod (Gadus morhua) with known migratory behaviour inferred from data storage tags. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 74(12):2122–2130, December 2017. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2016-0307.

Budy:2017:QLT

[1421] Phaedra E. Budy, Tracy Bowerman, Robert Al-Chokhachy, Mary Conner, and Howard Schaller. Quantifying long-term population growth rates of threatened bull trout: challenges, lessons learned, and opportunities. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 74(12):2131–2143, December 2017. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
cjfas-2016-0336.

Cambie:2017:DFM

[1422] Giulia Cambiè, Julia R. Pantin, Harriet Lincoln, Jan G. Hiddink, Gwladys T. Lambert, and Michel J. Kaiser. Diversity of fishing métier use can affect incomes and costs in small-scale fisheries. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 74(12):2144–2152, December 2017. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2016-0367.

Teichert:2017:IBL

[1423] Maxim A. K. Teichert, Anders Foldvik, Sigurd Einum, Anders G. Finstad, Torbjørn Forseth, and Ola Ugedal. Interactions between local population density and limited habitat resources determine movements of juvenile Atlantic salmon. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 74(12): 2153–2160, December 2017. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2016-0047.

Link:2018:SLO

[1424] Jason S. Link. System-level optimal yield: increased value, less risk, improved stability, and better fisheries. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 75(1):1–16, January 2018. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/cjfas-2017-0250.

Andrews:2018:BMM

[1425] Allen H. Andrews, Robert L. Humphreys, and Jeffery D. Sampaga. Blue marlin (Makaira nigricans) longevity estimates confirmed with bomb radiocarbon dating. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 75(1):17–25, January 2018. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
cjfas-2017-0031.

St-Gelais:2018:LFR

[1426] Nicolas F. St-Gelais, Anneli Jokela, and Beatrix E. Beisner. Limited functional responses of plankton food webs in northern lakes following diamond mining. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 75(1):26–35, January 2018. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
cjfas-2016-0418.

Gallagher:2018:IED

[1427] Brian K. Gallagher and David H. Secor. Intensified environmental and density-dependent regulation of white perch recruitment after an ecosystem shift in the Hudson River Estuary. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 75(1):36–46, January 2018. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/cjfas-2016-0455.

Vogt:2018:DIE

[1428] R. J. Vogt, S. Sharma, and P. R. Leavitt. Direct and interactive effects of climate, meteorology, river hydrology, and lake characteristics on water quality in productive lakes of the Canadian prairies. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 75(1):47–59, January 2018. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2016-0520.

Kotwicki:2018:CDB

[1429] Stan Kotwicki, Patrick H. Ressler, James N. Ianelli, André E. Punt, and John K. Horne. Combining data from bottom-trawl and acoustic-trawl surveys to estimate an index of abundance for semipelagic species. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 75(1):60–71, January 2018. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2016-0362.

Chizinski:2018:BNI

[1430] Christopher J. Chizinski, Dustin R. Martin, Daizaburo Shizuka, and Kevin L. Pope. Bipartite networks improve understanding of effects of waterbody size and angling method on angler–fish interactions. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 75(1):72–81, January 2018. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2016-0176.

Jarrin:2018:DMP

[1431] Jose R. Marin Jarrin, Timothy B. Johnson, Stuart A. Ludsin, Julie M. Reichert, and Kevin L. Pangle. Do models parameterized with observations from the system predict larval yellow perch (Perca flavescens) growth performance better in Lake Erie? Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 75(1):82–94, January 2018. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/cjfas-2016-0392.

Happel:2018:SVL

[1432] Austin Happel, Jory L. Jonas, Paul R. McKenna, Jacques Rinchard, Ji Xiang He, and Sergiusz J. Czesny. Spatial variability of lake trout diets in Lakes Huron and Michigan revealed by stomach content and fatty acid profiles. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 75(1):95–105, January 2018. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
cjfas-2016-0202.

Hansen:2018:WRS

[1433] Gretchen J. A. Hansen, Stephen R. Midway, and Tyler Wagner. Walleye recruitment success is less resilient to warming water temperatures in lakes with abundant largemouth bass populations. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 75(1):106–115, January 2018. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2016-0249.

Lush:2018:EAP

[1434] Lynn Lush, Isabel Costa, Kimberly Marshall, Juan Carlos Pérez-Casanova, Rénald Belley, and Dounia Hamoutene. Effects of acidic pH on wild, farmed, and F1 hybrid Atlantic salmon (Salmo salar) parr. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 75(1):116–120, January 2018. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2017-0006.

Pennock:2018:CFM

[1435] Casey A. Pennock, David Bender, Jordan Hofmeier, Jessica A. Mounts, Ryan Waters, Vaughn D. Weaver, and Keith B. Gido. Can fishways mitigate fragmentation effects on Great Plains fish communities? Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 75(1):121–130, January 2018. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2016-0466.

Iacarella:2018:PBI

[1436] Josephine C. Iacarella, Emma J. Hudgins, Jaimie T. A. Dick, and Anthony Ricciardi. Predatory behaviour of an invasive amphipod in response to varying conspecific densities under higher-order predation risk. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 75(1):131–140, January 2018. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2016-0417.

Song:2018:DET

[1437] Yiqing Song, Fei Cheng, Brian R. Murphy, and Songguang Xie. Downstream effects of the Three Gorges Dam on larval dispersal, spatial distribution, and growth of the four major Chinese carps call for reprioritizing conservation measures. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 75(1): 141–151, January 2018. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2016-0278.

Dunlop:2018:REM

[1438] Erin S. Dunlop, Rob McLaughlin, Jean V. Adams, Michael Jones, Oana Birceanu, Mark R. Christie, Lori A. Criger, Julia L. M. Hinderer, Robert M. Hollingworth, Nicholas S. Johnson, Stephen R. Lantz, Weiming Li, James Miller, Bruce J. Morrison, David Mota-Sanchez, Andrew Muir, Maria S. Sepúlveda, Todd Steeves, Lisa Walter, Erin Westman, Isaac Wirgin, and Michael P. Wilkie. Rapid evolution meets invasive species control: the potential for pesticide resistance in sea lamprey. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 75(1):152–168, January 2018. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2017-0015.

Avlijas:2018:ETT

[1439] Sunčica Avlijaš, Anthony Ricciardi, and Nicholas E. Mandrak. Eurasian tench (Tinca tinca): the next Great Lakes invader. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 75(2):169–179, February 2018. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2017-0171.

Shulzitski:2018:PLR

[1440] K. Shulzitski, S. Sponaugle, M. Hauff, K. D. Walter, E. K. D’Alessandro, and R. K. Cowen. Patterns in larval reef fish distributions and assemblages, with implications for local retention in mesoscale eddies. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 75(2):180–192, February 2018. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2016-0304.

Zens:2018:MPR

[1441] Bernhard Zens, Martin Glas, Michael Tritthart, Helmut Habersack, and Hubert Keckeis. Movement patterns and rheoreaction of larvae of a fluvial specialist (nase, Chondrostoma nasus): the role of active versus passive components of behaviour in dispersal. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 75(2):193–200, February 2018. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2016-0276.

Winger:2018:CFE

[1442] Paul D. Winger, Jenna G. Munden, Truong X. Nguyen, Scott M. Grant, and George Legge. Comparative fishing to evaluate the viability of an aligned footgear designed to reduce seabed contact in northern shrimp bottom trawl fisheries. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 75(2): 201–210, February 2018. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2016-0461.

Gislason:2018:RCA

[1443] Davíd Gíslason, Robert L. McLaughlin, Beren W. Robinson, Andy Cook, and Erin S. Dunlop. Rapid changes in age and size at maturity in Lake Erie yellow perch (Perca flavescens) are not explained by harvest. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 75(2):211–223, February 2018. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2016-0211.

Westerberg:2018:MDE

[1444] Håkan Westerberg, Selma Pacariz, Lasse Marohn, Vilhelm Fagerström, Klaus Wysujack, Michael J. Miller, Marko Freese, Jan-Dag Pohlmann, and Reinhold Hanel. Modeling the drift of European (Anguilla anguilla) and American (Anguilla rostrata) eel larvae during the year of spawning. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 75(2):224–234, February 2018. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2016-0256.

Rousseau:2018:SDL

[1445] Yann Rousseau, Fabian Blanchard, and Antoine Gardel. Spatiotemporal dynamics of larval fish in a tropical estuarine mangrove: example of the Mahury River Estuary (French Guiana). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 75(2):235–246, February 2018. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2016-0267.

Lechner:2018:IDC

[1446] Aaron Lechner, Hubert Keckeis, Martin Glas, Michael Tritthart, Helmut Habersack, Laurin Andorfer, and Paul Humphries. The influence of discharge, current speed, and development on the downstream dispersal of larval nase (Chondrostoma nasus) in the River Danube. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 75(2):247–259, February 2018. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2016-0340.

Madigan:2018:IIM

[1447] Daniel J. Madigan, Zofia Baumann, Aaron B. Carlisle, Owyn Snodgrass, Heidi Dewar, and Nicholas S. Fisher. Isotopic insights into migration patterns of Pacific bluefin tuna in the eastern Pacific Ocean. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 75(2):260–270, February 2018. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2016-0504.

Andersen:2018:EBL

[1448] Ken H. Andersen, Lise Marty, and Robert Arlinghaus. Evolution of boldness and life history in response to selective harvesting. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 75(2):271–281, February 2018. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2016-0350.

Minto:2018:IUI

[1449] Cóilín Minto, John Hinde, and Rui Coelho. Including unsexed individuals in sex-specific growth models. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 75 (2):282–292, February 2018. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2016-0450.

Brignon:2018:ETO

[1450] William R. Brignon, James T. Peterson, Jason B. Dunham, Howard A. Schaller, and Carl B. Schreck. Evaluating trade-offs in bull trout reintroduction strategies using structured decision making. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 75(2):293–307, February 2018. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2016-0516.

Jackson:2018:ESF

[1451] Breeanne K. Jackson and S. Mažeika P. Sullivan. Ecosystem size and flooding drive trophic dynamics of riparian spiders in a fire-prone Sierra Nevada river system. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 75(2): 308–318, February 2018. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2017-0045.

dAraujoCouto:2018:LVL

[1452] Thiago Belisario d’Araújo Couto, Jansen Zuanon, Julian D. Olden, and Gonçalo Ferraz. Longitudinal variability in lateral hydrologic connectivity shapes fish occurrence in temporary floodplain ponds. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 75(2):319–328, February 2018. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2016-0388.

Samsing:2018:CSL

[1453] Francisca Samsing, Frode Oppedal, Sussie Dalvin, Ingrid Johnsen, Tone Vågseth, and Tim Dempster. Corrigendum: Salmon lice (Lepeophtheirus salmonis) development times, body size, and reproductive outputs follow universal models of temperature dependence. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 75(2):329, February 2018. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/cjfas-2017-0447. See [1231].

Keefer:2018:SSS

[1454] Matthew L. Keefer, Michael A. Jepson, Tami S. Clabough, Eric L. Johnson, Shawn R. Narum, Jon E. Hess, and Christopher C. Caudill. Sea-to-sea survival of late-run adult steelhead (Oncorhynchus mykiss) from the Columbia and Snake Rivers. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 75 (3):331–341, March 2018. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2016-0430.

Wiedenmann:2018:USA

[1455] John Wiedenmann and Olaf P. Jensen. Uncertainty in stock assessment estimates for New England groundfish and its impact on achieving target harvest rates. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 75 (3):342–356, March 2018. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2016-0484.

Hoffle:2018:LSG

[1456] Hannes Höffle, Cindy J. G. Van Damme, Clive Fox, Stéphanie Lelièvre, Christophe Loots, Richard D. M. Nash, Sandrine Vaz, Peter J. Wright, and Peter Munk. Linking spawning ground extent to environmental factors — patterns and dispersal during the egg phase of four North Sea fishes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 75(3):357–374, March 2018. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
cjfas-2016-0310.

Farmer:2018:ADG

[1457] Nicholas A. Farmer and Jerald S. Ault. Accounting for detection gaps when evaluating reef fish habitat use in an acoustic array. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 75(3):375–388, March 2018. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2016-0494.

Wheeler:2018:ASE

[1458] Kit Wheeler, Scott W. Miller, and Todd A. Crowl. Adfluvial salmonids have engineering but not fertilization impacts in tributaries of a central Utah reservoir. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 75(3):389–401, March 2018. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
cjfas-2016-0462.

Grimaldo:2018:IRE

[1459] Eduardo Grimaldo, Manu Sistiaga, Bent Herrmann, Roger B. Larsen, Jesse Brinkhof, and Ivan Tatone. Improving release efficiency of cod (Gadus morhua) and haddock (Melanogrammus aeglefinus) in the Barents Sea demersal trawl fishery by stimulating escape behaviour. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 75(3):402–416, March 2018. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2017-0002.

Monk:2018:EPP

[1460] Christopher T. Monk and Robert Arlinghaus. Eurasian perch, Perca fluviatilis, spatial behaviour determines vulnerability independent of angler skill in a whole-lake reality mining experiment. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 75(3):417–428, March 2018. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2017-0029.

Johnson:2018:ESP

[1461] Nicholas S. Johnson, Dennis Higgs, Thomas R. Binder, J. Ellen Marsden, Tyler Buchinger, Linnea Brege, Tyler Bruning, Steven Farha, and Charles C. Krueger. Evidence of sound production by spawning lake trout (Salvelinus namaycush) in lakes Huron and Champlain. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 75(3):429–438, March 2018. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2016-0511.

Cunningham:2018:GMS

[1462] Curry J. Cunningham, Trevor A. Branch, Tyler H. Dann, Matt Smith, James E. Seeb, Lisa W. Seeb, and Ray Hilborn. A general model for salmon run reconstruction that accounts for interception and differences in availability to harvest. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 75 (3):439–451, March 2018. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2016-0360.

Okamura:2018:TBC

[1463] Hiroshi Okamura, Shoko H. Morita, Tetsuichiro Funamoto, Momoko Ichinokawa, and Shinto Eguchi. Target-based catch-per-unit-effort standardization in multispecies fisheries. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 75(3):452–463, March 2018. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/cjfas-2016-0460.

Altenritter:2018:UBS

[1464] Matthew E. Altenritter, Gayle Barbin Zydlewski, Michael T. Kinnison, Joseph D. Zydlewski, and Gail S. Wippelhauser. Understanding the basis of shortnose sturgeon (Acipenser brevirostrum) partial migration in the Gulf of Maine. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 75 (3):464–473, March 2018. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2017-0083.

Lazartigues:2018:OMI

[1465] Angélique Lazartigues, Chantale Girard, Philippe Brodeur, Frédéric Lecomte, Marc Mingelbier, and Pascal Sirois. Otolith microchemistry to identify sources of larval yellow perch in a fluvial lake: an approach towards freshwater fish management. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 75(3):474–487, March 2018. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/cjfas-2016-0289.

Raby:2018:DBT

[1466] Graham D. Raby, Christopher S. Vandergoot, Todd A. Hayden, Matthew D. Faust, Richard T. Kraus, John M. Dettmers, Steven J. Cooke, Yingming Zhao, Aaron T. Fisk, and Charles C. Krueger. Does behavioural thermoregulation underlie seasonal movements in Lake Erie walleye? Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 75(3):488–496, March 2018. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
cjfas-2017-0145.

Schumann:2018:ICN

[1467] David A. Schumann, Christopher S. Uphoff, Casey W. Schoenebeck, and Katie N. B. Graeb. Incorporation of carbon and nitrogen isotopes in age-0 walleye (Sander vitreus) tissues following a laboratory diet switch experiment. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 75 (4):497–505, April 2018. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2017-0301.

Keyser:2018:PIE

[1468] Freya Keyser, Brendan F. Wringe, Nicholas W. Jeffery, J. Brian Dempson, Steven Duffy, and Ian R. Bradbury. Predicting the impacts of escaped farmed Atlantic salmon on wild salmon populations. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 75(4):506–512, April 2018. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2017-0386.

Kaylor:2018:CCA

[1469] Matthew J. Kaylor and Dana R. Warren. Canopy closure after four decades of postlogging riparian forest regeneration reduces cutthroat trout biomass in headwater streams through bottom-up pathways. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 75(4):513–524, April 2018. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2016-0519.

Fraley:2018:RTM

[1470] Kevin M. Fraley, Jeffrey A. Falke, Megan V. McPhee, and Anupma Prakash. Rainbow trout movement behavior and habitat occupancy are influenced by sex and Pacific salmon presence in an Alaska river system. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 75(4):525–537, April 2018. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2016-0459.

Gerig:2018:IEI

[1471] Brandon S. Gerig, David N. Weber, Dominic T. Chaloner, Lillian M. McGill, and Gary A. Lamberti. Interactive effects of introduced Pacific salmon and brown trout on native brook trout: an experimental and modeling approach. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 75 (4):538–548, April 2018. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2016-0502.

Yates:2018:VSM

[1472] Adam G. Yates, Robert B. Brua, Joseph M. Culp, Roger G. Young, and Patricia A. Chambers. Variation in stream metabolism and benthic invertebrate composition along longitudinal profiles of two contrasting river systems. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 75(4):549– 559, April 2018. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
cjfas-2016-0198.

Piercy:2018:HTR

[1473] Julius J. B. Piercy, David J. Smith, Jamaluddin Jompa, Stephen D. Simpson, and Edward A. Codling. High temporal resolution sampling reveals reef fish settlement is highly clustered. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 75(4):560–568, April 2018. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/cjfas-2016-0318.

Ward:2018:NUD

[1474] Peter R. B. Ward, Paul J. Anders, G. Wayne Minshall, Charlie Holderman, Genevieve M. Hoyle, and Hassen Yassien. Nutrient uptake during low-level fertilization of a large, seventh-order oligotrophic river. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 75(4):569–579, April 2018. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2017-0062.

DeMartini:2018:CGA

[1475] Edward E. DeMartini, Allen H. Andrews, Kathrine G. Howard, Brett M. Taylor, Dong-Chun Lou, and Mary K. Donovan. Comparative growth, age at maturity and sex change, and longevity of Hawaiian parrotfishes, with bomb radiocarbon validation. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 75(4):580–589, April 2018. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/cjfas-2016-0523.

Sanchez-Hernandez:2018:MFI

[1476] Javier Sánchez-Hernández and Fernando Cobo. Modelling the factors influencing ontogenetic dietary shifts in stream-dwelling brown trout (Salmo trutta). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 75(4):590– 599, April 2018. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
cjfas-2017-0021.

Lyons:2018:EOM

[1477] John Lyons, Timothy P. Parks, Kristi L. Minahan, and Aaron S. Ruesch. Evaluation of oxythermal metrics and benchmarks for the protection of cisco (Coregonus artedi) habitat quality and quantity in Wisconsin lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 75(4):600–608, April 2018. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2017-0043.

Yoshimura:2018:ERC

[1478] Mayumi Yoshimura and Akio Akama. Elevated radioactive contamination from the Fukushima nuclear power plant accident in aquatic biota from a river with a lake in its upper reaches. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 75(4):609–620, April 2018. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/cjfas-2016-0333.

Jensen:2018:RBM

[1479] Arne Johan Jensen, Bengt Finstad, Peder Fiske, Torbjørn Forseth, Audun Håvard Rikardsen, and Ola Ugedal. Relationship between marine growth and sea survival of two anadromous salmonid fish species. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 75(4):621–628, April 2018. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2016-0408.

DeRobertis:2018:CBM

[1480] Alex De Robertis, Robert Levine, and Christopher D. Wilson. Can a bottom-moored echo sounder array provide a survey-comparable index of abundance? Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 75(4):629– 640, April 2018. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
cjfas-2017-0013.

Holgerson:2018:SAS

[1481] Meredith A. Holgerson, Max R. Lambert, L. Kealoha Freidenburg, and David K. Skelly. Suburbanization alters small pond ecosystems: shifts in nitrogen and food web dynamics. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 75(4):641–652, April 2018. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/cjfas-2016-0526.

Wiwchar:2018:SIC

[1482] Logan D. Wiwchar, Matthew J. H. Gilbert, Ashley V. Kasurak, and Keith B. Tierney. Schooling improves critical swimming performance in zebrafish (Danio rerio). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 75 (4):653–661, April 2018. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2017-0141.

Jensen:2018:ELS

[1483] Arne Johan Jensen, Bengt Finstad, and Peder Fiske. Evidence for the linkage of survival of anadromous Arctic char and brown trout during winter to marine growth during the previous summer. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 75(5):663–672, May 2018. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2017-0077.

Purchase:2018:LTS

[1484] Craig F. Purchase. Low tolerance of salt water in a marine fish: new and historical evidence for surprising local adaption in the well-studied commercially exploited capelin. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 75 (5):673–681, May 2018. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2017-0058.

Manhard:2018:LCM

[1485] Christopher V. Manhard, Nicholas A. Som, Russell W. Perry, and John M. Plumb. A laboratory-calibrated model of Coho salmon growth with utility for ecological analyses. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 75(5):682–690, May 2018. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/

doi/full/10.1139/cjfas-2016-0506.

Miller:2018:EEA

[1486] Timothy J. Miller and Saang-Yoon Hyun. Evaluating evidence for alternative natural mortality and process error assumptions using a statespace, age-structured assessment model. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 75(5):691–703, May 2018. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/cjfas-2017-0035.

Brcic:2018:CSM

[1487] Jure Brčić, Bent Herrmann, and Antonello Sala. Can a square-mesh panel inserted in front of the cod end improve size and species selectivity in Mediterranean trawl fisheries? Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 75(5):704–713, May 2018. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/cjfas-2017-0123.

Realis-Doyelle:2018:TAG

[1488] Emilie Réalis-Doyelle, Enric Gisbert, Carles Alcaraz, Fabrice Teletchea, and Alain Pasquet. Temperature affects growth allometry and development patterns in brown trout (Salmo trutta) fry: a multitrait approach. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 75(5):714–722, May 2018. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2017-0037.

Trijoulet:2018:GSP

[1489] Vanessa Trijoulet, Steven J. Holmes, and Robin M. Cook. Grey seal predation mortality on three depleted stocks in the West of Scotland: What are the implications for stock assessments? Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 75(5):723–732, May 2018. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/cjfas-2016-0521.

Pedersen:2018:LTG

[1490] Eric J. Pedersen, Daisuke Goto, Jereme W. Gaeta, Gretchen J. A. Hansen, Greg G. Sass, M. Jake Vander Zanden, Thomas A. Cichosz, and Andrew L. Rypel. Long-term growth trends in northern Wisconsin walleye populations under changing biotic and abiotic conditions. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 75(5):733–745, May 2018. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2017-0084.

Hwan:2018:DSS

[1491] Jason L. Hwan, Albert Fernández-Chacón, Mathieu Buoro, and Stephanie M. Carlson. Dry season survival of juvenile salmonids in an intermittent coastal stream. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 75 (5):746–758, May 2018. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2017-0026.

Mumby:2018:FEN

[1492] James A. Mumby, Timothy B. Johnson, Thomas J. Stewart, Edmund A. Halfyard, Brian C. Weidel, Maureen G. Walsh, Jana R. Lantry, and Aaron T. Fisk. Feeding ecology and niche overlap of Lake Ontario offshore forage fish assessed with stable isotopes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 75(5):759–771, May 2018. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/cjfas-2016-0150. See corrigendum [1560].

McCann:2018:CLT

[1493] Erin L. McCann, Nicholas S. Johnson, and Kevin L. Pangle. Corresponding long-term shifts in stream temperature and invasive fish migration. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 75(5):772–778, May 2018. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2017-0195.

Wyman:2018:FSH

[1494] M. T. Wyman, M. J. Thomas, R. R. McDonald, A. R. Hearn, R. D. Battleson, E. D. Chapman, P. Kinzel, J. T. Minear, E. A. Mora, J. M. Nelson, M. D. Pagel, and A. P. Klimley. Fine-scale habitat selection of green sturgeon (Acipenser medirostris) within three spawning locations in the Sacramento River, California. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 75(5):779–791, May 2018. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/cjfas-2017-0072.

Keckeis:2018:PUD

[1495] Hubert Keckeis, Catriona Clemmesen, Paul Humphries, and Su Sponaugle. Paths to the unknown: dispersal during the early life of fishes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 75(5):792–796, May 2018. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2017-0536.

Doherty:2018:UAV

[1496] Beau Doherty, Samuel D. N. Johnson, and Sean P. Cox. Using autonomous video to estimate the bottom-contact area of longline trap gear and presence–absence of sensitive benthic habitat. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 75(5):797–812, May 2018. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2016-0483.

Rochette:2018:LNC

[1497] Rémy Rochette, Bernard Sainte-Marie, Marc Allain, Jackie Baker, Louis Bernatchez, Virginia Boudreau, Michel Comeau, Jean Côté, Gilles Miron, Laura Ramsay, Kevin Squires, and M. John Tremblay. The lobster node of the CFRN: co-constructed and collaborative research on productivity, stock structure, and connectivity in the American lobster (Homarus americanus). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 75 (5):813–824, May 2018. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2016-0426.

Rose:2018:DRL

[1498] George A. Rose and Sherrylynn Rowe. Does redistribution or local growth underpin rebuilding of Canada’s Northern cod? Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 75(6):825–835, June 2018. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2017-0421.

Morgan:2018:GPS

[1499] Thomas D. Morgan, Carly F. Graham, Andrew G. McArthur, Amogelang R. Raphenya, Douglas R. Boreham, Richard G. Manzon, Joanna Y. Wilson, Stacey L. Lance, Kimberly L. Howland, Paul H. Patrick, and Christopher M. Somers. Genetic population structure of the round whitefish (Prosopium cylindraceum) in North America: multiple markers reveal glacial refugia and regional subdivision. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 75(6):836–849, June 2018. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/cjfas-2016-0528.

Melli:2018:FCH

[1500] Valentina Melli, Junita D. Karlsen, Jordan P. Feekings, Bent Herrmann, and Ludvig A. Krag. FLEXSELECT: counter-herding device to reduce bycatch in crustacean trawl fisheries. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 75(6):850–860, June 2018. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/cjfas-2017-0226.

Takahashi:2018:NCT

[1501] Kohji Takahashi and Reiji Masuda. Net-chasing training improves the behavioral characteristics of hatchery-reared red sea bream (Pagrus major) juveniles. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 75(6):861–867, June 2018. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
cjfas-2017-0073.

Debertin:2018:LZA

[1502] Allan J. Debertin, J. Mark Hanson, and Simon C. Courtenay. Linking zooplankton assemblages with oceanographic zones in an Atlantic coastal ecosystem. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 75(6):868– 882, June 2018. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
cjfas-2016-0342.

Carruthers:2018:AUD

[1503] Thomas Carruthers, Laurence Kell, and Carlos Palma. Accounting for uncertainty due to data processing in virtual population analysis using Bayesian multiple imputation. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 75(6):883–896, June 2018. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/

doi/full/10.1139/cjfas-2017-0165.

Myrvold:2018:IWT

[1504] Knut Marius Myrvold and Brian Patrick Kennedy. Increasing water temperatures exacerbate the potential for density dependence in juvenile steelhead. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 75(6):897–907, June 2018. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
cjfas-2016-0497.

Haarr:2018:FAL

[1505] Marthe Larsen Haarr, Bernard Sainte-Marie, Michel Comeau, M. John Tremblay, and Rémy Rochette. Female American lobster (Homarus americanus) size-at-maturity declined in Canada during the 20th and early 21st centuries. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 75 (6):908–924, June 2018. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2016-0434.

Baril:2018:LSA

[1506] André-Marcel Baril, John T. Buszkiewicz, Pascale M. Biron, Quinton E. Phelps, and James W. A. Grant. Lake sturgeon (Acipenser fulvescens) spawning habitat: a quantitative review. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 75(6):925–933, June 2018. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/cjfas-2017-0100.

Brenden:2018:IRS

[1507] Travis O. Brenden, Iyob Tsehaye, James R. Bence, Jeannette Kanefsky, and Kim T. Scribner. Indexing recruitment for source populations contributing to mixed fisheries by incorporating age in genetic stock identification models. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 75(6):934– 954, June 2018. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
cjfas-2016-0525.

Andersson:2018:HFM

[1508] Luke C. Andersson and John D. Reynolds. Habitat features mediate selective consumption of salmon by bears. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 75(6):955–963, June 2018. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/cjfas-2017-0055.

Richard:2018:DCF

[1509] Gaëtan Richard, Christophe Guinet, Julien Bonnel, Nicolas Gasco, and Paul Tixier. Do commercial fisheries display optimal foraging? The case of longline fishers in competition with odontocetes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 75(6):964–976, June 2018. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2016-0498.

Pizzochero:2018:UMS

[1510] Ana Carolina Pizzochero, Loïc N. Michel, Simon R. Chenery, Ian D. McCarthy, Marcelo Vianna, Olaf Malm, Gilles Lepoint, Krishna Das, and Paulo R. Dorneles. Use of multielement stable isotope ratios to investigate ontogenetic movements of Micropogonias furnieri in a tropical Brazilian estuary. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 75 (6):977–986, June 2018. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2017-0148.

Tarpey:2018:SNP

[1511] Carolyn M. Tarpey, James E. Seeb, Garrett J. McKinney, William D. Templin, Alexander Bugaev, Shunpei Sato, and Lisa W. Seeb. Singlenucleotide polymorphism data describe contemporary population structure and diversity in allochronic lineages of pink salmon (Oncorhynchus gorbuscha). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 75(6):987–997, June 2018. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
cjfas-2017-0023.

Hamilton:2018:SBD

[1512] Trevor J. Hamilton, David I. Kline, and Martin Tresguerres. Shoaling behaviour is differentially altered by ethanol and dopamine D1 receptor modulators in tropical marine forage fish. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 75(7):999–1004, July 2018. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/cjfas-2018-0084.

Minns:2018:FIP

[1513] Charles K. Minns, Brian J. Shuter, Andrew Davidson, and Shusen Wang. Factors influencing peak summer surface water temperature in Canada’s large lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 75(7):1005–1018, July 2018. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
cjfas-2017-0061.

Rudd:2018:AVR

[1514] Merrill B. Rudd and James T. Thorson. Accounting for variable recruitment and fishing mortality in length-based stock assessments for data-limited fisheries. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 75 (7):1019–1035, July 2018. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2017-0143.

Muench:2018:PPF

[1515] Angela Muench, Geret Sean DePiper, and Chad Demarest. On the precision of predicting fishing location using data from the vessel monitoring system (VMS). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 75(7):1036– 1047, July 2018. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
cjfas-2016-0446.

Som:2018:MFE

[1516] Nicholas A. Som, Russell W. Perry, Edward C. Jones, Kyle De Juilio, Paul Petros, William D. Pinnix, and Derek L. Rupert. N-mix for fish: estimating riverine salmonid habitat selection via N-mixture models. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 75(7):1048–1058, July 2018. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2017-0027.

Klinard:2018:HUS

[1517] Natalie V. Klinard, Aaron T. Fisk, Steven T. Kessel, Edmund A. Halfyard, and Scott F. Colborne. Habitat use and small-scale residence patterns of sympatric sunfish species in a large temperate river. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 75(7):1059–1069, July 2018. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2017-0125.

Kissinger:2018:FSP

[1518] Benjamin C. Kissinger, Les N. Harris, Danny Swainson, W. Gary Anderson, Margaret F. Docker, and James D. Reist. Fine-scale population structure in lake trout (Salvelinus namaycush) influenced by life history variation in the Husky Lakes drainage basin, Northwest Territories, Canada. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 75(7):1070– 1081, July 2018. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
cjfas-2016-0524.

Dorner:2018:STP

[1519] Brigitte Dorner, Matthew J. Catalano, and Randall M. Peterman. Spatial and temporal patterns of covariation in productivity of Chinook salmon populations of the northeastern Pacific Ocean. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 75(7):1082–1095, July 2018. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2017-0197.

Beacham:2018:PII

[1520] Terry D. Beacham, Colin Wallace, Cathy MacConnachie, Kim Jonsen, Brenda McIntosh, John R. Candy, and Ruth E. Withler. Population and individual identification of Chinook salmon in British Columbia through parentage-based tagging and genetic stock identification with single nucleotide polymorphisms. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 75 (7):1096–1105, July 2018. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2017-0168.

Oskarsson:2018:IEA

[1521] Gudmundur J. Óskarsson, Jónbjörn Pálsson, and Asta Gudmundsdottir. An ichthyophoniasis epizootic in Atlantic herring in marine waters around Iceland. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 75(7):1106– 1116, July 2018. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
cjfas-2017-0219.

Hostetter:2018:WDA

[1522] Nathan J. Hostetter, Beth Gardner, Allen F. Evans, Bradley M. Cramer, Quinn Payton, Ken Collis, and Daniel D. Roby. Wanted dead or alive: a state-space mark–recapture–recovery model incorporating multiple recovery types and state uncertainty. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 75(7):1117–1127, July 2018. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/cjfas-2016-0246.

Mucci:2018:TIV

[1523] Alfonso Mucci, Maurice Levasseur, Yves Gratton, Chloé Martias, Michael Scarratt, Denis Gilbert, Jean-Éric Tremblay, Gustavo Ferreyra, and Bruno Lansard. Tidally induced variations of pH at the head of the Laurentian Channel. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 75 (7):1128–1141, July 2018. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2017-0007.

Weisser:2018:VMP

[1524] Manuel Weisser, Hilmar Hofmann, Jorge Encinas Fernández, and Frank Peeters. Vertical migration patterns of the different larval instars of Chaoborus flavicans and the influence of dissolved oxygen concentrations. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 75(7):1142–1150, July 2018. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2017-0157.

Jasper:2018:HBE

[1525] James R. Jasper, Margaret Short, Chris Shelden, and W. Stewart Grant. Hierarchical Bayesian estimation of unobserved salmon passage through weirs. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 75(7):1151–1159, July 2018. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
cjfas-2016-0398.

Larson:2018:RDS

[1526] Wesley A. Larson, Yniv Palti, Gunagtu Gao, Kenneth I. Warheit, and James E. Seeb. Rapid discovery of SNPs that differentiate hatchery steelhead trout from ESA-listed natural-origin steelhead trout using a 57K SNP array. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 75(7):1160– 1168, July 2018. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
cjfas-2017-0116.

Olson:2018:LSR

[1527] Mark H. Olson, Janet M. Fischer, Craig E. Williamson, Erin P. Overholt, and Nora Theodore. Landscape-scale regulators of water transparency in mountain lakes: implications of projected glacial loss. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 75(7):1169–1176, July 2018. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2017-0215.

Huusko:2018:SMS

[1528] Riina Huusko, Pekka Hyvärinen, Mikko Jaukkuri, Aki Mäki-Petäys, Panu Orell, and Jaakko Erkinaro. Survival and migration speed of radiotagged Atlantic salmon (Salmo salar) smolts in two large rivers: one without and one with dams. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 75 (8):1177–1184, August 2018. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2017-0134.

Ono:2018:MSA

[1529] Kotaro Ono, Alan C. Haynie, Anne B. Hollowed, James N. Ianelli, Carey R. McGilliard, and André E. Punt. Management strategy analysis for multispecies fisheries, including technical interactions and human behavior in modelling management decisions and fishing. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 75(8):1185–1202, August 2018. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2017-0135.

Wor:2018:LAM

[1530] Catarina Wor, Murdoch K. McAllister, Steven J. D. Martell, Nathan G. Taylor, and Carl J. Walters. A Lagrangian approach to model movement of migratory species. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 75 (8):1203–1214, August 2018. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2017-0093.

Scopel:2018:SDB

[1531] Lauren C. Scopel, Antony W. Diamond, Stephen W. Kress, Adrian R. Hards, and Paula Shannon. Seabird diets as bioindicators of Atlantic herring recruitment and stock size: a new tool for ecosystem-based fisheries management. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 75(8): 1215–1229, August 2018. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2017-0140.

Laman:2018:USD

[1532] Edward A. Laman, Christopher N. Rooper, Kali Turner, Sean Rooney, Dan W. Cooper, and Mark Zimmermann. Using species distribution models to describe essential fish habitat in Alaska. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 75(8):1230–1255, August 2018. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2017-0181.

Salisbury:2018:GDA

[1533] Sarah J. Salisbury, Connor Booker, Gregory R. McCracken, Tom Knight, Donald Keefe, Robert Perry, and Daniel E. Ruzzante. Genetic divergence among and within Arctic char (Salvelinus alpinus) populations inhabiting landlocked and sea-accessible sites in Labrador, Canada. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 75(8):1256–1269, August 2018. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2017-0163.

Patey:2018:EBC

[1534] Géraldine Patey, Catherine M. Couillard, Hilaire Drouineau, Guy Verreault, Fabien Pierron, Patrick Lambert, Magalie Baudrimont, and Patrice Couture. Early back-calculated size-at-age of Atlantic yellow eels sampled along ecological gradients in the Gironde and St. Lawrence hydrographical systems. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 75(8): 1270–1279, August 2018. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2017-0025.

Duplisea:2018:EIF

[1535] Daniel E. Duplisea. Eliminating implausible fisheries assessment models using fishers’ knowledge. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 75 (8):1280–1290, August 2018. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2017-0178.

Marandel:2018:IGD

[1536] Florianne Marandel, Pascal Lorance, Marco Andrello, Grégory Charrier, Sabrina Le Cam, Sigrid Lehuta, and Verena M. Trenkel. Insights from genetic and demographic connectivity for the management of rays and skates. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 75(8):1291–1302, August 2018. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
cjfas-2017-0291.

Magee:2018:MOS

[1537] Madeline R. Magee, Peter B. McIntyre, and Chin H. Wu. Modeling oxythermal stress for cool-water fishes in lakes using a cumulative dosage approach. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 75(8):1303–1312, August 2018. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
cjfas-2017-0260.

Simmons:2018:CIB

[1538] Carrie M. Simmons and Stephen T. Szedlmayer. Competitive interactions between gray triggerfish (Balistes capriscus) and red snapper (Lutjanus campechanus) in laboratory and field studies in the northern Gulf of Mexico. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 75(8):1313–1318, August 2018. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
cjfas-2017-0039.

Regish:2018:SKA

[1539] Amy M. Regish, John T. Kelly, Michael F. O’Dea, and Stephen D. Mc-Cormick. Sensitivity of Na^{+} / K^{+}-ATPase isoforms to acid and aluminum explains differential effects on Atlantic salmon osmoregulation in fresh water and seawater. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 75 (8):1319–1328, August 2018. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2017-0198.

George:2018:CGV

[1540] Ellen M. George, Matthew P. Hare, Darran L. Crabtree, Brian F. Lantry, and Lars G. Rudstam. Comparison of genetic and visual identification of cisco and lake whitefish larvae from Chaumont Bay, Lake Ontario. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 75(8):1329–1336, August 2018. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2017-0186.

Feiner:2018:IEI

[1541] Zachary S. Feiner, Timothy D. Malinich, and Tomas O. Höök. Interacting effects of identity, size, and winter severity determine temporal consistency of offspring phenotype. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 75 (8):1337–1345, August 2018. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2017-0220.

Freshwater:2018:CPD

[1542] Cameron Freshwater, Brian J. Burke, Mark D. Scheuerell, Sue C. H. Grant, Marc Trudel, and Francis Juanes. Coherent population dynamics associated with sockeye salmon juvenile life history strategies. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 75(8):1346–1356, August 2018. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2017-0251.

Cahill:2018:MCC

[1543] Christopher L. Cahill, Stephanie Mogensen, Kyle L. Wilson, Ariane Cantin, R. Nilo Sinnatamby, Andrew J. Paul, Paul Christensen, Jessica R. Reilly, Linda Winkel, Anne Farineau, and John R. Post. Multiple challenges confront a high-effort inland recreational fishery in decline. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 75(9):1357–1368, September 2018. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
cjfas-2018-0086.

Thorson:2018:TPC

[1544] James T. Thorson. Three problems with the conventional delta-model for biomass sampling data, and a computationally efficient alternative. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 75(9):1369–1382, September 2018. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
cjfas-2017-0266.

Hemstrom:2018:FLI

[1545] William Hemstrom, Stan van de Wetering, and Michael Banks. Fish ladder installation across a historical barrier asymmetrically increased conspecific introgressive hybridization between wild winter and summer run steelhead salmon in the Siletz River, Oregon. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 75(9):1383–1392, September 2018. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2016-0411.

Dzul:2018:EDA

[1546] M. C. Dzul, C. B. Yackulic, and J. Korman. Estimating disperser abundance using open population models that incorporate data from continuous detection passive integrated transponder arrays. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 75(9):1393–1404, September 2018. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2017-0304.

Bell:2018:RFC

[1547] Richard J. Bell, Anthony Wood, Jonathan Hare, David Richardson, John Manderson, and Timothy Miller. Rebuilding in the face of climate change. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 75(9):1405– 1414, September 2018. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2017-0085.

Jensen:2018:FRG

[1548] Alexander J. Jensen and Michael L. Jones. Forecasting the response of Great Lakes sea lamprey (Petromyzon marinus) to barrier removals. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 75(9):1415–1426, September 2018. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
cjfas-2017-0243.

Cauwelier:2018:ISG

[1549] Eef Cauwelier, John Gilbey, James Sampayo, Lee Stradmeyer, and Stuart J. Middlemas. Identification of a single genomic region associated with seasonal river return timing in adult Scottish Atlantic salmon (Salmo salar), using a genome-wide association study. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 75(9):1427–1435, September 2018. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2017-0293.

Li:2018:EST

[1550] Yan Li, Tyler Wagner, Yan Jiao, Robert Lorantas, and Cheryl A. Murphy. Evaluating spatial and temporal variability in growth and mortality for recreational fisheries with limited catch data. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 75(9):1436–1452, September 2018. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2017-0052.

Rolla:2018:PIF

[1551] Matteo Rolla, Geordie Biffoni, Stefano Brighenti, Rocco Iacobuzio, Kevin Liautaud, Cristian Pasquaretta, and Rocco Tiberti. Predation by introduced fish can magnify the terrestrial arthropod subsidies in mountain lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 75(9):1453–1464, September 2018. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
cjfas-2017-0121.

Mushet:2018:IRA

[1552] Graham R. Mushet, Karlee Flear, Brendan Wiltse, Andrew M. Paterson, and Brian F. Cumming. Increased relative abundance of colonial scaled chrysophytes since pre-industrial times in minimally disturbed lakes from the Experimental Lakes Area, Ontario. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 75(9):1465–1476, September 2018. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2017-0074.

Edmunds:2018:RHS

[1553] Nicholas B. Edmunds, Kevin S. McCann, and Frédéric Laberge. Relative heart size and fish foraging ecology in a lake food web. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 75(9):1477–1484, September 2018. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2017-0412.

Zhang:2018:SEE

[1554] Chongliang Zhang, Yong Chen, Binduo Xu, Ying Xue, and Yiping Ren. Sampling effects on the effectiveness of ecological indicators in detecting fishery-induced community changes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 75(9):1485–1494, September 2018. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/cjfas-2017-0199.

Venditti:2018:EHS

[1555] David A. Venditti, Ryan N. Kinzer, Kimberly A. Apperson, Bruce Barnett, Matthew Belnap, Timothy Copeland, Matthew P. Corsi, and Kurt Tardy. Effects of hatchery supplementation on abundance and productivity of natural-origin Chinook salmon: two decades of evaluation and implications for conservation programs. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 75(9):1495–1510, September 2018. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/cjfas-2016-0344.

Drenner:2018:TPB

[1556] S. Matthew Drenner, Scott G. Hinch, Nathan B. Furey, Timothy D. Clark, Shaorong Li, Tobi Ming, Ken M. Jeffries, David A. Patterson, Steven J. Cooke, David Robichaud, David W. Welch, Anthony P. Farrell, and Kristina M. Miller. Transcriptome patterns and blood physiology associated with homing success of sockeye salmon during their final stage of marine migration. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 75(9):1511– 1524, September 2018. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2017-0391.

Braccini:2018:DSC

[1557] Matias Braccini, Simon de Lestang, and Rory McAuley. Dusky sharks (Carcharhinus obscurus) undertake large-scale migrations between tropical and temperate ecosystems. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 75(9):1525–1533, September 2018. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/cjfas-2017-0313.

Mickle:2018:ITR

[1558] Megan F. Mickle and Dennis M. Higgs. Integrating techniques: a review of the effects of anthropogenic noise on freshwater fish. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 75(9):1534–1541, September 2018. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2017-0245.

Marshall:2018:GCN

[1559] Brian E. Marshall. Guilty as charged: Nile perch was the cause of the haplochromine decline in Lake Victoria. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 75(9):1542–1559, September 2018. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2017-0056.

Mumby:2018:CFE

[1560] James A. Mumby, Timothy B. Johnson, Thomas J. Stewart, Edmund A. Halfyard, Brian C. Weidel, Maureen G. Walsh, Jana R. Lantry, and Aaron T. Fisk. Corrigendum: Feeding ecology and niche overlap of Lake Ontario offshore forage fish assessed with stable isotopes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 75(9):1560–1561, September 2018. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
cjfas-2018-0218. See [1492].

Baumann:2018:CMB

[1561] Zofia Baumann, Robert P. Mason, David O. Conover, Prentiss Balcom, Celia Y. Chen, Kate L. Buckman, Nicholas S. Fisher, and Hannes Baumann. Corrigendum: Mercury bioaccumulation increases with latitude in a coastal marine fish (Atlantic silverside, Menidia menidia). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 75(9):1562, September 2018. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
cjfas-2018-0248. See [1331].

Sun:2018:ASD

[1562] Ming Sun, Chongliang Zhang, Yong Chen, Binduo Xu, Ying Xue, and Yiping Ren. Assessing the sensitivity of data-limited methods (DLMs) to the estimation of life-history parameters from length–frequency data. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 75(10):1563–1572, October 2018. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2017-0325.

Madenjian:2018:TRG

[1563] Charles P. Madenjian, Todd A. Hayden, Tyler B. Peat, Christopher S. Vandergoot, David G. Fielder, Ann Marie Gorman, Steven A. Pothoven, John M. Dettmers, Steven J. Cooke, Yingming Zhao, and Charles C. Krueger. Temperature regimes, growth, and food consumption for female and male adult walleye in Lake Huron and Lake Erie: a bioenergetics analysis. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 75(10):1573–1586, October 2018. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
cjfas-2017-0280.

Godwin:2018:HSL

[1564] Sean C. Godwin, Martin Krkošek, John D. Reynolds, Luke A. Rogers, and Lawrence M. Dill. Heavy sea louse infection is associated with decreased stomach fullness in wild juvenile sockeye salmon. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 75(10):1587–1595, October 2018. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2017-0267.

Czarnecka:2018:DPW

[1565] Magdalena Czarnecka and Oliver Miler. Decay processes in woody debris influence the taxonomic and functional composition of littoral macroinvertebrates. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 75(10):1596–1605, October 2018. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
cjfas-2017-0364.

Yokota:2018:PAA

[1566] Masashi Yokota, Jungo Takeda, Naoki Suzuki, and Kazumi Sakuramoto. A protocol for the analysis of aquatic biodiversity by multiple β diversities. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 75(10):1606–1613, October 2018. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
cjfas-2017-0231.

Mollenhauer:2018:SSL

[1567] Robert Mollenhauer, Joshua B. Mouser, and Shannon K. Brewer. Sampling the stream landscape: improving the applicability of an ecoregionlevel capture probability model for stream fishes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 75(10):1614–1625, October 2018. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2016-0422.

Neumann:2018:FEP

[1568] Viola Neumann, Friedrich W. Köster, and Margit Eero. Fish egg predation by Baltic sprat and herring: do species characteristics and development stage matter? Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 75(10): 1626–1634, October 2018. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2017-0105.

Wildermuth:2018:SUQ

[1569] Robert P. Wildermuth, Gavin Fay, and Sarah Gaichas. Structural uncertainty in qualitative models for ecosystem-based management of Georges Bank. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 75(10):1635–1643, October 2018. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
cjfas-2017-0149.

Allen:2018:LSG

[1570] Brandon E. Allen, Ella Bowles, Matthew R. J. Morris, and Sean M. Rogers. Loss of SNP genetic diversity following population collapse in a recreational walleye (Sander vitreus) fishery. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 75(10):1644–1651, October 2018. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2017-0164.

Lucchetti:2018:MUS

[1571] Alessandro Lucchetti, Emilio Notti, Antonello Sala, and Massimo Virgili. Multipurpose use of side-scan sonar technology for fisheries science. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 75(10):1652–1662, October 2018. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2017-0359.

Bellanger:2018:ITO

[1572] Manuel Bellanger, Claire Macher, Mathieu Merzéréaud, Olivier Guyader, and Christelle Le Grand. Investigating trade-offs in alternative catch share systems: an individual-based bio-economic model applied to the Bay of Biscay sole fishery. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 75 (10):1663–1679, October 2018. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/

doi/full/10.1139/cjfas-2017-0075.

Lemopoulos:2018:GWD

[1573] Alexandre Lemopoulos, Silva Uusi-Heikkilä, Anti Vasemägi, Ari Huusko, Harri Kokko, and Anssi Vainikka. Genome-wide divergence patterns support fine-scaled genetic structuring associated with migration tendency in brown trout. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 75(10): 1680–1692, October 2018. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2017-0014.

Riseng:2018:ECM

[1574] Catherine M. Riseng, Kevin E. Wehrly, Li Wang, Edward S. Rutherford, James E. McKenna, Lucinda B. Johnson, Lacey A. Mason, Christian Castiglione, Thomas P. Hollenhorst, Beth L. Sparks-Jackson, and Scott P. Sowa. Ecosystem classification and mapping of the Laurentian Great Lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 75(10):1693– 1712, October 2018. CODEN CJFSDX. ISSN 0706-652X (print), 1205-

7533 (electronic). URL https://cdnsciencepub.com/doi/full/10.
1139/cjfas-2017-0242.

Vignon:2018:STS

[1575] Matthias Vignon. Short-term stress for long-lasting otolith morphology — brief embryological stress disturbance can reorient otolith ontogenetic trajectory. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 75(10):1713–1722, October 2018. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
cjfas-2017-0110.

Perry:2018:TFC

[1576] Diana Perry, Thomas A. B. Staveley, Linus Hammar, Alyssa Meyers, Regina Lindborg, and Martin Gullström. Temperate fish community variation over seasons in relation to large-scale geographic seascape variables. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 75(10):1723–1732, October 2018. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
cjfas-2017-0032.

Nguyen:2018:AKA

[1577] Vivian M. Nguyen, Nathan Young, and Steven J. Cooke. Applying a knowledge–action framework for navigating barriers to incorporating telemetry science into fisheries management and conservation: a qualitative study. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 75(10):1733–1743, October 2018. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
cjfas-2017-0303.

Bordeleau:2018:NCS

[1578] Xavier Bordeleau, Jan G. Davidsen, Sindre H. Eldøy, Aslak D. Sjursen, Fred G. Whoriskey, and Glenn T. Crossin. Nutritional correlates of spatiotemporal variations in the marine habitat use of brown trout (Salmo trutta) veteran migrants. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 75(10): 1744–1754, October 2018. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2017-0350.

Krueger:2018:ATO

[1579] Charles C. Krueger, Christopher M. Holbrook, Thomas R. Binder, Christopher S. Vandergoot, Todd A. Hayden, Darryl W. Hondorp, Nancy Nate, Kelli Paige, Stephen C. Riley, Aaron T. Fisk, and Steven J. Cooke. Acoustic telemetry observation systems: challenges encountered and overcome in the Laurentian Great Lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 75(10):1755–1763, October 2018. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2017-0406.

Zhang:2018:FDE

[1580] Fan Zhang, Davíd Gíslason, Kevin B. Reid, Allan J. Debertin, Katrine Turgeon, and Thomas D. Nudds. Failure to detect ecological and evolutionary effects of harvest on exploited fish populations in a managed fisheries ecosystem. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 75(10): 1764–1771, October 2018. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2017-0217.

Anonymous:2018:AND

[1581] Anonymous. Appointment of a new deputy executive Editor-in-Chief. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 75(11), November 2018. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2018-deeic.

Bowlby:2018:PRH

[1582] Heather Dawn Bowlby. Productivity of riverine habitats may be changing for American eel. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 75(11): 1773–1777, November 2018. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2018-0188.

Bourret:2018:UFG

[1583] Samuel L. Bourret and Niall G. Clancy. Using forensic geochemistry via fish otoliths to investigate an illegal fish introduction. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 75(11):1778–1783, November 2018. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2018-0082.

Runnebaum:2018:HSM

[1584] Jocelyn Runnebaum, Lisha Guan, Jie Cao, Loretta O’Brien, and Yong Chen. Habitat suitability modeling based on a spatiotemporal model: an example for cusk in the Gulf of Maine. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 75(11):1784–1797, November 2018. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/cjfas-2017-0316.

Campbell:2018:MPT

[1585] Stuart D. Campbell and Patricia Chow-Fraser. Models to predict total phosphorus concentrations in coastal embayments of eastern Georgian Bay, Lake Huron. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 75(11): 1798–1810, November 2018. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2017-0095.

Daniels:2018:ECR

[1586] Jason Daniels, Gérald Chaput, and Jonathan Carr. Estimating consumption rate of Atlantic salmon smolts (Salmo salar) by striped bass (Morone saxatilis) in the Miramichi River estuary using acoustic telemetry. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 75(11):1811–1822, November 2018. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2017-0373.

Cienciala:2018:SLB

[1587] Piotr Cienciala and Marwan A. Hassan. Spatial linkages between geomorphic and hydraulic conditions and invertebrate drift characteristics in a small mountain stream. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 75(11): 1823–1835, November 2018. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2017-0170.

Willmes:2018:FCR

[1588] Malte Willmes, James A. Hobbs, Anna M. Sturrock, Zachary Bess, Levi S. Lewis, Justin J. G. Glessner, Rachel C. Johnson, Ryon Kurth, and Jason Kindopp. Fishery collapse, recovery, and the cryptic decline of wild salmon on a major California river. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 75(11):1836–1848, November 2018. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/cjfas-2017-0273.

Snouck-Hurgronje:2018:FFO

[1589] Julia E. Snouck-Hurgronje, David M. Kaplan, Emmanuel Chassot, Alexandra Maufroy, and Daniel Gaertner. Fishing on floating objects (FOBs): how French tropical tuna purse seiners split fishing effort between GPS-monitored and unmonitored FOBs. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 75(11):1849–1858, November 2018. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2017-0152.

Koizumi:2018:EWL

[1590] Shuntaro Koizumi, Nicola Craig, Jacob A. Zwart, Patrick T. Kelly, Jacob P. Ziegler, Brian C. Weidel, Stuart E. Jones, and Christopher T.

Solomon. Experimental whole-lake dissolved organic carbon increase alters fish diet and density but not growth or productivity. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 75(11):1859–1867, November 2018. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2017-0283.

Bellinger:2018:CRA

[1591] Kristy L. Bellinger, Gary H. Thorgaard, and Patrick A. Carter. Complex relationships among rearing temperature, growth, and sprint speed in clonal lines of rainbow trout. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 75(11):1868–1877, November 2018. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/cjfas-2017-0218.

Hasnain:2018:ETR

[1592] Sarah S. Hasnain, Michael D. Escobar, and Brian J. Shuter. Estimating thermal response metrics for North American freshwater fish using Bayesian phylogenetic regression. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 75(11):1878–1885, November 2018. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/cjfas-2017-0278.

Perry:2018:FME

[1593] Russell W. Perry, Adam C. Pope, Jason G. Romine, Patricia L. Brandes, Jon R. Burau, Aaron R. Blake, Arnold J. Ammann, and Cyril J. Michel. Flow-mediated effects on travel time, routing, and survival of juvenile Chinook salmon in a spatially complex, tidally forced river delta. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 75(11):1886–1901, November 2018. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2017-0310.

Guan:2018:CSS

[1594] Lu Guan, John F. Dower, and Pierre Pepin. Characterizing spatial structures of larval fish assemblages at multiple scales in relation to environmental heterogeneity in the Strait of Georgia (British Columbia, Canada). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 75(11):1902–1914, November 2018. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
cjfas-2017-0194.

Sinclair-Waters:2018:GTM

[1595] Marion Sinclair-Waters, Paul Bentzen, Corey J. Morris, Daniel E. Ruzzante, Matthew P. Kent, Sigbjørn Lien, and Ian R. Bradbury. Genomic tools for management and conservation of Atlantic cod in a coastal marine protected area. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 75(11): 1915–1925, November 2018. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2017-0254.

Lamothe:2018:LTD

[1596] Karl A. Lamothe, Donald A. Jackson, and Keith M. Somers. Long-term directional trajectories among lake crustacean zooplankton communities and water chemistry. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 75(11): 1926–1939, November 2018. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2017-0518.

Long:2018:HDS

[1597] William Christopher Long, Peter A. Cummiskey, and J. Eric Munk. How does stocking density affect enhancement success for hatchery-reared red king crab? Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 75(11):1940–1948, November 2018. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
cjfas-2017-0330.

Roloson:2018:RTO

[1598] Scott D. Roloson, Kyle M. Knysh, Michael R. S. Coffin, Karen L. Gormley, Christina C. Pater, and Michael R. van den Heuvel. Rainbow trout (Oncorhynchus mykiss) habitat overlap with wild Atlantic salmon (Salmo salar) and brook trout (Salvelinus fontinalis) in natural streams: do habitat and landscape factors override competitive interactions? Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 75(11):1949–1959, November 2018. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2017-0342.

Kovach:2018:LTP

[1599] Ryan P. Kovach, Jonathan B. Armstrong, David A. Schmetterling, Robert Al-Chokhachy, and Clint C. Muhlfeld. Long-term population dynamics and conservation risk of migratory bull trout in the upper Columbia River basin. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 75(11): 1960–1968, November 2018. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2017-0466.

Parks:2018:PPD

[1600] Timothy P. Parks and Andrew L. Rypel. Predator–prey dynamics mediate long-term production trends of cisco (Coregonus artedi) in a northern Wisconsin lake. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 75(11):1969– 1976, November 2018. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2017-0302.

Watson:2018:PRW

[1601] Jordan T. Watson and Alan C. Haynie. Paths to resilience: the walleye pollock fleet uses multiple fishing strategies to buffer against environmental change in the Bering Sea. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 75(11):1977–1989, November 2018. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/cjfas-2017-0315.

Kenworthy:2018:MEM

[1602] Matthew D. Kenworthy, Jonathan H. Grabowski, Craig A. Layman, Graham D. Sherwood, Sean P. Powers, Charles H. Peterson, Rachel K. Gittman, Danielle A. Keller, and F. Joel Fodrie. Movement ecology of a mobile predatory fish reveals limited habitat linkages within a temperate estuarine seascape. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 75(11): 1990–1998, November 2018. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2017-0308.

Frechette:2018:UST

[1603] Danielle M. Frechette, Stephen J. Dugdale, Julian J. Dodson, and Normand E. Bergeron. Understanding summertime thermal refuge use by adult Atlantic salmon using remote sensing, river temperature monitoring, and acoustic telemetry. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 75 (11):1999–2010, November 2018. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/

doi/full/10.1139/cjfas-2017-0422.

Bass:2018:LSC

[1604] Arthur L. Bass, Scott G. Hinch, David A. Patterson, Steven J. Cooke, and Anthony P. Farrell. Location-specific consequences of beach seine and gillnet capture on upriver-migrating sockeye salmon migration behavior and fate. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 75(11):2011– 2023, November 2018. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2017-0474.

Beguer-Pon:2018:MST

[1605] Mélanie Béguer-Pon, Guy Verreault, David Stanley, Martin Castonguay, and Julian J. Dodson. The migration of stocked, trapped and transported, and wild female American silver eels through the Gulf of St. Lawrence. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 75(11):2024–2037, November 2018. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
cjfas-2017-0356.

Murray:2018:FMC

[1606] Taryn S. Murray, Paul D. Cowley, Rhett H. Bennett, and Amber-Robyn Childs. Fish on the move: connectivity of an estuary-dependent fishery species evaluated using a large-scale acoustic telemetry array. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 75(11):2038–2052, November 2018. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2017-0361.

Healy:2018:TPR

[1607] Stephen J. Healy, Scott G. Hinch, Arthur L. Bass, Nathan B. Furey, David W. Welch, Erin L. Rechisky, Erika J. Eliason, Andrew G. Lotto, and Kristina M. Miller. Transcriptome profiles relate to migration fate in hatchery steelhead (Oncorhynchus mykiss) smolts. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 75(11):2053–2068, November 2018. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2017-0424.

Teffer:2018:IFC

[1608] Amy K. Teffer, Arthur L. Bass, Kristi M. Miller, David A. Patterson, Francis Juanes, and Scott G. Hinch. Infections, fisheries capture, temperature, and host responses: multistressor influences on survival and behaviour of adult Chinook salmon. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 75(11):2069–2083, November 2018. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/cjfas-2017-0491.

Lewandoski:2018:EPC

[1609] Sean A. Lewandoski, Mary Anne Bishop, and Megan K. McKinzie. Evaluating Pacific cod migratory behavior and site fidelity in a fjord environment using acoustic telemetry. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 75(11):2084–2095, November 2018. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/cjfas-2017-0432.

Morse:2018:SSC

[1610] Bryan L. Morse, Brady K. Quinn, Michel Comeau, and Rémy Rochette. Stock structure and connectivity of the American lobster (Homarus americanus) in the southern Gulf of St. Lawrence: Do benthic movements matter? Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 75(11):2096– 2108, November 2018. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2017-0346.

Anonymous:2018:NAN

[1611] Anonymous. Note of appreciation / note de reconnaissance. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 75(12), December 2018. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2018-0362.

Stewart:2018:ISE

[1612] Ian J. Stewart and Allan C. Hicks. Interannual stability from ensemble modelling. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 75(12):2109– 2113, December 2018. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2018-0238.

Aylesworth:2018:RTP

[1613] Lindsay Aylesworth and Ting-Chun Kuo. Reporting time period matters: quantifying catch rates and exploring recall bias from fisher interviews in Thailand. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 75(12): 2114–2122, December 2018. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2017-0169.

Johnston:2018:ESL

[1614] Thomas A. Johnston. Egg size and lipid content of lake trout (Salvelinus namaycush) in the wild and in captivity. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 75(12):2123–2135, December 2018. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2017-0408.

Li:2018:CSO

[1615] Yang Li, James R. Bence, and Travis O. Brenden. Can spawning origin information of catch or a recruitment penalty improve assessment and fishery management performance for a spatially structured stock assessment model? Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 75(12): 2136–2148, December 2018. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2017-0523.

Murdoch:2018:EOE

[1616] Andrew R. Murdoch, Chad J. Herring, Charles H. Frady, Kevin See, and Chris E. Jordan. Estimating observer error and steelhead redd abundance using a modified Gaussian area-under-the-curve framework. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 75(12):2149–2158, December 2018. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2017-0335.

Miller:2018:TEV

[1617] Timothy J. Miller, Loretta O’Brien, and Paula S. Fratantoni. Temporal and environmental variation in growth and maturity and effects on management reference points of Georges Bank Atlantic cod. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 75(12):2159–2171, December 2018. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2017-0124.

Sinnatamby:2018:STR

[1618] R. Niloshini Sinnatamby, Madalena C. Pinto, Fiona D. Johnston, Andrew J. Paul, Craig J. Mushens, Jim D. Stelfox, Hillary G. M. Ward, and John R. Post. Seasonal timing of reproductive migrations in adfluvial bull trout: an assessment of sex, spawning experience, population density, and environmental factors. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 75(12):2172–2183, December 2018. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/cjfas-2017-0542.

Osterback:2018:EES

[1619] Ann-Marie K. Osterback, Cynthia H. Kern, Emerson A. Kanawi, Jeffrey M. Perez, and Joseph D. Kiernan. The effects of early sandbar formation on the abundance and ecology of Coho salmon (Oncorhynchus kisutch) and steelhead trout (Oncorhynchus mykiss) in a central California coastal lagoon. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 75(12): 2184–2197, December 2018. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2017-0455.

Harbicht:2018:EAD

[1620] Andrew B. Harbicht, Theodore Castro-Santos, Dimitry Gorsky, David M. Hand, Dylan J. Fraser, and William R. Ardren. Environmental, anthropogenic, and dietary influences on fine-scale movement patterns of Atlantic salmon through challenging waters. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 75(12):2198–2210, December 2018. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2017-0476.

Heaston:2018:AFW

[1621] Emily D. Heaston, Matthew J. Kaylor, and Dana R. Warren. Aquatic food web response to patchy shading along forested headwater streams. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 75(12):2211–2220, December 2018. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
cjfas-2017-0464.

Griffiths:2018:ICR

[1622] Katherine Griffiths, Joshua Thienpont, Adam Jeziorski, and John P. Smol. The impact of calcium-rich diamond mining effluent on downstream cladoceran communities in softwater lakes of the Northwest Territories, Canada. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 75(12):2221– 2232, December 2018. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2017-0469.

Santana-Garcon:2018:RVR

[1623] Julia Santana-Garcon, Corey B. Wakefield, Stacey R. Dorman, Ainslie Denham, Stuart Blight, Brett W. Molony, and Stephen J. Newman. Risk versus reward: interactions, depredation rates, and bycatch mitigation of dolphins in demersal fish trawls. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 75(12):2233–2240, December 2018. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/cjfas-2017-0203.

Kiyama:2018:ISC

[1624] Shoichi Kiyama and Satoshi Yamazaki. The impact of stock collapse on small-scale fishers’ behavior: evidence from Japan. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 75(12):2241–2254, December 2018. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2017-0091.

Tucker:2018:LTA

[1625] Taaja R. Tucker, Edward F. Roseman, Robin L. DeBruyne, Jeremy J. Pritt, David H. Bennion, Darryl W. Hondorp, and James C. Boase. Longterm assessment of ichthyoplankton in a large North American river system reveals changes in fish community dynamics. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 75(12):2255–2270, December 2018. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2017-0511.

Dender:2018:PIB

[1626] Mitchel G. E. Dender, Pauline M. Capelle, Oliver P. Love, Daniel D. Heath, John W. Heath, and Christina A. D. Semeniuk. Phenotypic integration of behavioural and physiological traits is related to variation in growth among stocks of Chinook salmon. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 75(12):2271–2279, December 2018. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2017-0367.

vanPoorten:2018:EBS

[1627] Brett T. van Poorten, Shannon Harris, and Allison Hebert. Evaluating benefits of stocking on sockeye recovery projections in a nutrientenhanced mixed life history population. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 75(12):2280–2290, December 2018. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2017-0438.

Rypel:2018:EPW

[1628] Andrew L. Rypel, Daisuke Goto, Greg G. Sass, and M. Jake Vander Zanden. Eroding productivity of walleye populations in northern Wisconsin lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 75(12):2291–2301, December 2018. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
cjfas-2017-0311.

Neville:2018:OCR

[1629] Victoria Neville, George Rose, Sherrylynn Rowe, Robyn Jamieson, and Glenn Piercey. Otolith chemistry and redistributions of Northern cod: evidence of Smith Sound–Bonavista Corridor connectivity. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 75(12):2302–2312, December 2018. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2017-0357.

Haraldstad:2018:CDS

[1630] Tormod Haraldstad, Erik Höglund, Frode Kroglund, Anders Lamberg, Esben Moland Olsen, and Thrond Oddvar Haugen. Condition-dependent skipped spawning in anadromous brown trout (Salmo trutta). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 75(12):2313–2319, December 2018. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2017-0076.

Shaw:2018:MEB

[1631] Stephanie L. Shaw, Greg G. Sass, and Justin A. VanDeHey. Maternal effects better predict walleye recruitment in Escanaba Lake, Wisconsin, 1957–2015: implications for regulations. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 75(12):2320–2331, December 2018. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2017-0318.

Whitmore:2018:FSM

[1632] Margaret M. Whitmore and Matthew K. Litvak. Fine-scale movement of juvenile Atlantic sturgeon (Acipenser oxyrinchus oxyrinchus) during aggregations in the lower Saint John River Basin, New Brunswick, Canada. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 75(12):2332–2342, December 2018. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
cjfas-2017-0430.

Mulder:2018:OTH

[1633] Ingeborg M. Mulder, Corey J. Morris, J. Brian Dempson, Ian A. Fleming, and Michael Power. Overwinter thermal habitat use in lakes by anadromous Arctic char. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 75(12): 2343–2353, December 2018. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2017-0420.

Whitmore:2018:SDM

[1634] Margaret M. Whitmore and Matthew K. Litvak. Seasonal distribution and movement of juvenile Atlantic sturgeon (Acipenser oxyrinchus oxyrinchus) in the lower Saint John River Basin, New Brunswick, Canada. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 75(12):2354–2363, December 2018. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
cjfas-2017-0429.

Potts:2018:CCR

[1635] Warren Mason Potts, Alexander Claus Winkler, Matthew Cameron Parkinson, Carmen Van Dunem Santos, Warwick Hans Hugo Sauer, and Amber-Robyn Childs. Comparing catch rate, conventional tagging, and acoustic telemetry data for understanding the migration patterns of coastal fishes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 75(12):2364– 2374, December 2018. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2017-0428.

Middleton:2018:ENW

[1636] Collin T. Middleton, Scott G. Hinch, Eduardo G. Martins, Douglas C. Braun, David A. Patterson, Nicholas J. Burnett, Vanessa Minke-Martin, and Matthew T. Casselman. Effects of natal water concentration and temperature on the behaviour of up-river migrating sockeye salmon. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 75(12):2375–2389, December 2018. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
cjfas-2017-0490.

Ivanova:2018:IVT

[1637] Silviya V. Ivanova, Steven T. Kessel, Justin Landry, Caitlin O’Neill, Montana F. McLean, Mario Espinoza, Svein Vagle, Nigel E. Hussey, and Aaron T. Fisk. Impact of vessel traffic on the home ranges and movement of shorthorn sculpin (Myoxocephalus scorpius) in the nearshore environment of the high Arctic. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 75(12): 2390–2400, December 2018. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2017-0418.

Comeau:2018:ALP

[1638] Michel Comeau and J. Mark Hanson. American lobster: persistence in the face of high, size-selective, fishing mortality — a perspective from the southern Gulf of St. Lawrence. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 75(12):2401–2411, December 2018. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/cjfas-2017-0374.

Anttila:2019:CSA

[1639] Katja Anttila, Anthony P. Farrell, David A. Patterson, Scott G. Hinch, and Erika J. Eliason. Cardiac SERCA activity in sockeye salmon populations: an adaptive response to migration conditions. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 76(1):1–5, January 2019. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2018-0334.

Bruestle:2019:NTI

[1640] Eric L. Bruestle, Curtis Karboski, Anna Hussey, Aaron T. Fisk, Knut Mehler, Christopher Pennuto, and Dimitry Gorsky. Novel trophic interaction between lake sturgeon (Acipenser fulvescens) and non-native species in an altered food web. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 76(1):6–14, January 2019. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/

doi/full/10.1139/cjfas-2017-0282.

Smolinski:2019:IOE

[1641] Szymon Smoliński. Incorporation of optimal environmental signals in the prediction of fish recruitment using random forest algorithms. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 76(1):15–27, January 2019. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2017-0554.

Duguay:2019:IBU

[1642] Jason M. Duguay, R. W. Jay Lacey, and Theodore Castro-Santos. Influence of baffles on upstream passage of brook trout and brown trout in an experimental box culvert. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 76 (1):28–41, January 2019. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2017-0453.

Erkinaro:2019:LHV

[1643] Jaakko Erkinaro, Yann Czorlich, Panu Orell, Jorma Kuusela, Morten Falkegård, Maija Länsman, Henni Pulkkinen, Craig R. Primmer, and Eero Niemelä. Life history variation across four decades in a diverse population complex of Atlantic salmon in a large subarctic river. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 76(1):42–55, January 2019. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2017-0343.

Addicott:2019:IPC

[1644] Ethan T. Addicott, Kailin Kroetz, Matthew N. Reimer, James N. Sanchirico, Daniel K. Lew, and Justine Huetteman. Identifying the potential for cross-fishery spillovers: a network analysis of Alaskan permitting patterns. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 76(1):56–68, January 2019. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
cjfas-2017-0550.

Sardenne:2019:SHL

[1645] Fany Sardenne, N’Guessan Constance Diaha, Monin Justin Amandé, Iker Zudaire, Lydie I. E. Couturier, Luisa Metral, Fabienne Le Grand, and Nathalie Bodin. Seasonal habitat and length influence on the trophic niche of co-occurring tropical tunas in the eastern Atlantic Ocean. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 76(1):69–80, January 2019. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2017-0368.

Hegg:2019:WDY

[1646] Jens C. Hegg, Brian P. Kennedy, and Paul Chittaro. What did you say about my mother? The complexities of maternally derived chemical signatures in otoliths. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 76 (1):81–94, January 2019. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2017-0341.

Shelton:2019:UHM

[1647] Andrew Olaf Shelton, William H. Satterthwaite, Eric J. Ward, Blake E. Feist, and Brian Burke. Using hierarchical models to estimate stockspecific and seasonal variation in ocean distribution, survivorship, and aggregate abundance of fall run Chinook salmon. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 76(1):95–108, January 2019. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2017-0204.

Kopf:2019:MFC

[1648] R. Keller Kopf, Paul Humphries, Nick R. Bond, Neil C. Sims, Robyn J. Watts, Ross M. Thompson, Sally Hladyz, John D. Koehn, Alison J. King, Nicole McCasker, and Simon McDonald. Macroecology of fish community biomass–size structure: effects of invasive species and river regulation. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 76(1):109–122, January 2019. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2017-0544.

Sparks:2019:IST

[1649] Morgan M. Sparks, Jeffrey A. Falke, Thomas P. Quinn, Milo D. Adkison, Daniel E. Schindler, Krista Bartz, Dan Young, and Peter A. H. Westley. Influences of spawning timing, water temperature, and climatic warming on early life history phenology in western Alaska sockeye salmon. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 76(1):123–135, January 2019. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2017-0468.

Rechisky:2019:QSA

[1650] Erin L. Rechisky, Aswea D. Porter, Timothy D. Clark, Nathan B. Furey, Marika Kirstin Gale, Scott G. Hinch, and David W. Welch. Quantifying survival of age-2 Chilko Lake sockeye salmon during the first 50 days of migration. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 76(1): 136–152, January 2019. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2017-0425.

Wang:2019:DMF

[1651] Jocelyn Yun-Ling Wang, Christopher M. Anderson, Curry J. Cunningham, Ray Hilborn, and Michael R. Link. Does more fish mean more money? Evaluating alternative escapement goals in the Bristol Bay salmon fishery. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 76(1):153–167, January 2019. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
cjfas-2017-0336.

Neuenhoff:2019:CDC

[1652] Rachel D. Neuenhoff, Douglas P. Swain, Sean P. Cox, Murdoch K. McAllister, Andrew W. Trites, Carl J. Walters, and Mike O. Hammill. Continued decline of a collapsed population of Atlantic cod (Gadus morhua) due to predation-driven Allee effects. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 76(1):168–184, January 2019. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/cjfas-2017-0190.

Campbell:2019:PHE

[1653] Emily Y. Campbell, Jason B. Dunham, Gordon H. Reeves, and Steve M. Wondzell. Phenology of hatching, emergence, and end-of-season body size in young-of-year Coho salmon in thermally contrasting streams draining the Copper River Delta, Alaska. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 76(2):185–191, February 2019. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2018-0003.

Kuriyama:2019:ITS

[1654] Peter T. Kuriyama, Trevor A. Branch, Allan C. Hicks, John H. Harms, and Owen S. Hamel. Investigating three sources of bias in hook-and-line surveys: survey design, gear saturation, and multispecies interactions. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 76(2):192–207, February 2019. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2017-0286.

Nogueira:2019:UMS

[1655] Adriana Nogueira, Nick Tolimieri, and Diana González-Troncoso. Using multivariate state-space models to examine commercial stocks of redfish (Sebastes spp.) on the Flemish Cap. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 76(2):208–216, February 2019. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/cjfas-2017-0435.

Bi:2019:BSA

[1656] Rujia Bi, Yan Jiao, Can Zhou, and Eric Hallerman. A Bayesian spatiotemporal approach to inform management unit appropriateness. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 76(2):217–237, February 2019. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2017-0526.

Waller:2019:LSR

[1657] Diane L. Waller, Michelle R. Bartsch, Lynn A. Bartsch, and Craig A. Jackson. Lethal and sublethal responses of native mussels (Unionidae: Lampsilis siliquoidea and Lampsilis higginsii) to elevated carbon dioxide. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 76(2):238–248, February 2019. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2017-0543.

Binion-Rock:2019:SKD

[1658] Samantha M. Binion-Rock, Brian J. Reich, and Jeffrey A. Buckel. A spatial kernel density method to estimate the diet composition of fish. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 76(2):249–267, February 2019. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2017-0306.

Xu:2019:NSP

[1659] Haikun Xu, James T. Thorson, Richard D. Methot, Jr., and Ian G. Taylor. A new semi-parametric method for autocorrelated age- and time-varying selectivity in age-structured assessment models. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 76(2):268–285, February 2019. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2017-0446.

Bond:2019:OFC

[1660] Rosealea M. Bond, Colin L. Nicol, Joseph D. Kiernan, and Brian C. Spence. Occurrence, fate, and confounding influence of ghost passive integrated transponder tags in an intensively monitored watershed. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 76(2):286–298, February 2019. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2017-0409.

Ohlberger:2019:BLC

[1661] Jan Ohlberger, Samuel J. Brenkman, Patrick Crain, George R. Pess, Jeffrey J. Duda, Thomas W. Buehrens, Thomas P. Quinn, and Ray Hilborn. A Bayesian life-cycle model to estimate escapement at maximum sustained yield in salmon based on limited information. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 76(2):299–307, February 2019. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2017-0382.

Larsen:2019:ENG

[1662] Roger B. Larsen, Manu Sistiaga, Bent Herrmann, Jesse Brinkhof, Ivan Tatone, and Juan Santos. The effect of Nordmøre grid length and angle on codend entry of bycatch fish species and shrimp catches. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 76(2):308–319, February 2019. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2018-0069.

Koeck:2019:ASA

[1663] Barbara Koeck, Libor Závorka, David Aldvén, Joacim Näslund, Robert Arlinghaus, Per-Ove Thörnqvist, Svante Winberg, Björn Thrandur Björnsson, and Jörgen I. Johnsson. Angling selects against active and stress-resilient phenotypes in rainbow trout. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 76(2):320–333, February 2019. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2018-0085.

Cote:2019:CSC

[1664] David Cote, Jean-Marc Nicolas, Frederick Whoriskey, Adam. M. Cook, Jeremy Broome, Paul M. Regular, and Darrin Baker. Characterizing snow crab (Chionoecetes opilio) movements in the Sydney Bight (Nova Scotia, Canada): a collaborative approach using multiscale acoustic telemetry. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 76(2):334–346, February 2019. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
cjfas-2017-0472.

Andersen:2019:LIB

[1665] Ken H. Andersen, Nis Sand Jacobsen, and P. Daniël van Denderen. Limited impact of big fish mothers for population replenishment. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 76(3):347–349, March 2019. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2018-0354.

Gendron:2019:LEP

[1666] Louise Gendron, Denis Lefaivre, and Bernard Sainte-Marie. Local egg production and larval losses to advection contribute to interannual and long-term variability of American lobster (Homarus americanus) settlement intensity. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 76(3):350–363, March 2019. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
cjfas-2017-0565.

Turschak:2019:SVT

[1667] Benjamin A. Turschak, Sergiusz Czesny, Jason C. Doll, Brice K. Grunert, Tomas O. Höök, John Janssen, and Harvey A. Bootsma. Spatial variation in trophic structure of nearshore fishes in Lake Michigan as it relates to water clarity. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 76(3):364–377, March 2019. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
cjfas-2017-0390.

Warkentin:2019:UPF

[1668] Luke Warkentin, Corinna Favaro, David C. Scott, Rebecca Seifert, and Jonathan W. Moore. Urban planning for fishes: untangling a new project’s effects from old infrastructure and regional patterns. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 76(3):378–389, March 2019. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2017-0354.

Abrial:2019:IFV

[1669] E. Abrial, L. A. Espínola, A. P. Rabuffetti, M. L. Amsler, and K. M. Wantzen. Interannual flow variability in a large subtropical–temperate floodplain: a challenge for fish reproduction. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 76(3):390–400, March 2019. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/cjfas-2017-0579.

Thorson:2019:SAC

[1670] James T. Thorson and Melissa A. Haltuch. Spatiotemporal analysis of compositional data: increased precision and improved workflow using model-based inputs to stock assessment. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 76(3):401–414, March 2019. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/cjfas-2018-0015.

Bouska:2019:DUF

[1671] Kristen L. Bouska, Gregory W. Whitledge, Christopher Lant, and Justin Schoof. Drivers and uncertainties of forecasted range shifts for warmwater fishes under climate and land cover change. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 76(3):415–425, March 2019. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2018-0002.

Campbell:2019:PCA

[1672] Teresa Campbell, James Simmons, Jessica Sáenz, Christopher L. Jerde, William Cowan, Sudeep Chandra, and Zeb Hogan. Population connectivity of adfluvial and stream-resident Lahontan cutthroat trout: implications for resilience, management, and restoration. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 76(3):426–437, March 2019. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2017-0483.

Guernon:2019:CAB

[1673] Stéphanie Guernon, Matthew C. Yates, Dylan J. Fraser, and Alison M. Derry. The coevolution of adult body mass and excretion rate between genetically size-divergent brook trout populations. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 76(3):438–446, March 2019. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2017-0508.

Nelson:2019:WCS

[1674] Benjamin W. Nelson, Carl J. Walters, Andrew W. Trites, and Murdoch K. McAllister. Wild Chinook salmon productivity is negatively related to seal density and not related to hatchery releases in the Pacific Northwest. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 76(3):447–462, March 2019. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
cjfas-2017-0481.

Nguyen:2019:WUK

[1675] Vivian M. Nguyen, Nathan Young, Marianne Corriveau, Scott G. Hinch, and Steven J. Cooke. What is “usable” knowledge? Perceived barriers for integrating new knowledge into management of an iconic Canadian fishery. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 76(3):463–474, March 2019. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2017-0305.

Morita:2019:EMT

[1676] Kentaro Morita. Earlier migration timing of salmonids: an adaptation to climate change or maladaptation to the fishery? Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 76(3):475–479, March 2019. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2018-0078.

Stephenson:2019:IDO

[1677] Robert L. Stephenson, Melanie Wiber, Stacey Paul, Eric Angel, Ashleen Benson, Anthony Charles, Omer Chouinard, Dan Edwards, Paul Foley, Dan Lane, Jim McIsaac, Barb Neis, Courtenay Parlee, Evelyn Pinkerton, Mark Saunders, Kevin Squires, and U. Rashid Sumaila. Integrating diverse objectives for sustainable fisheries in Canada. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 76(3):480–496, March 2019. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2017-0345.

Lennox:2019:OMS

[1678] Robert J. Lennox, Cecilia Engler-Palma, Katie Kowarski, Alexander Filous, Rebecca Whitlock, Steven J. Cooke, and Marie Auger-Méthé. Optimizing marine spatial plans with animal tracking data. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 76(3):497–509, March 2019. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2017-0495.

Puerta:2019:SPS

[1679] Patricia Puerta, Bethany Johnson, Lorenzo Ciannelli, Thomas Helser, and Robert Lauth. Subsampling populations with spatially structured traits: a field comparison of stratified and random strategies. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 76(4):511–522, April 2019. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2017-0248.

Dauphin:2019:HMD

[1680] Guillaume J. R. Dauphin, Gérald Chaput, Cindy Breau, and Richard A. Cunjak. Hierarchical model detects decadal changes in calibration relationships of single-pass electrofishing indices of abundance of Atlantic salmon in two large Canadian catchments. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 76(4):523–542, April 2019. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/cjfas-2017-0456.

Dorazio:2019:SSM

[1681] Robert M. Dorazio and Melissa Price. State-space models to infer movements and behavior of fish detected in a spatial array of acoustic receivers. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 76(4):543–550, April 2019. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2018-0067.

Moran:2019:BIC

[1682] Benjamin M. Moran and Eric C. Anderson. Bayesian inference from the conditional genetic stock identification model. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 76(4):551–560, April 2019. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/cjfas-2018-0016.

Weber:2019:ARD

[1683] Michael J. Weber and Michael L. Brown. Application of a robust design occupancy model for assessing fish recruitment. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 76(4):561–568, April 2019. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2018-0083.

Pohlmann:2019:ELN

[1684] Jan-Dag Pohlmann, Marko Freese, Stefan Reiser, and Reinhold Hanel. Evaluation of lethal and nonlethal assessment methods of muscle fat content in European eel (Anguilla anguilla). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 76(4):569–575, April 2019. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/cjfas-2018-0032.

Rajasilta:2019:IEC

[1685] Marjut Rajasilta, Jari Hänninen, Lea Laaksonen, Päivi Laine, Jukka-Pekka Suomela, Ilppo Vuorinen, and Katja Mäkinen. Influence of environmental conditions, population density, and prey type on the lipid content in Baltic herring (Clupea harengus membras) from the northern Baltic Sea. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 76(4):576– 585, April 2019. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
cjfas-2017-0504.

Bourdaud:2019:IIF

[1686] Pierre Bourdaud, Morgane Travers-Trolet, Youen Vermard, and Paul Marchal. Improving the interpretation of fishing effort and pressures in mixed fisheries using spatial overlap metrics. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 76(4):586–596, April 2019. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/cjfas-2017-0529.

Embke:2019:MFE

[1687] Holly S. Embke, Patrick M. Kocovsky, Tatiana Garcia, Christine M. Mayer, and Song S. Qian. Modeling framework to estimate spawning and hatching locations of pelagically spawned eggs. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 76(4):597–607, April 2019. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2018-0047.

Tanaka:2019:MCI

[1688] Kisei R. Tanaka, Jui-Han Chang, Ying Xue, Zengguang Li, Larry Jacobson, and Yong Chen. Mesoscale climatic impacts on the distribution of Homarus americanus in the US inshore Gulf of Maine. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 76(4):608–625, April 2019. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2018-0075.

Landry:2019:MTA

[1689] Justin J. Landry, Steve T. Kessel, Montana F. McLean, Silviya V. Ivanova, Nigel E. Hussey, Caitlin O’Neill, Svein Vagle, Terry A. Dick, and Aaron T. Fisk. Movement types of an Arctic benthic fish, shorthorn sculpin (Myoxocephalus scorpius), during open-water periods in response to biotic and abiotic factors. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 76 (4):626–635, April 2019. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2017-0389.

Kocan:2019:TMF

[1690] Richard M. Kocan. Transmission models for the fish pathogen Ichthyophonus: synthesis of field observations and empirical studies. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 76(4):636–642, April 2019. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2018-0166.

Miller:2019:ICK

[1691] Shana K. Miller, Alejandro Anganuzzi, Doug S. Butterworth, Campbell R. Davies, Greg P. Donovan, Amanda Nickson, Rebecca A. Rademeyer, and Victor Restrepo. Improving communication: the key to more effective MSE processes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 76 (4):643–656, April 2019. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2018-0134.

Honsey:2019:BLF

[1692] Andrew E. Honsey, Paul A. Venturelli, and Nigel P. Lester. Bioenergetic and limnological foundations for using degree-days derived from air temperatures to describe fish growth. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 76(4):657–669, April 2019. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/cjfas-2018-0051.

Thompson:2019:CFR

[1693] Susan A. Thompson, Robert L. Stephenson, George A. Rose, and Stacey D. Paul. Collaborative fisheries research: the Canadian fisheries research network experience. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 76 (5):671–681, May 2019. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2018-0450.

Carrasquilla-Henao:2019:PIR

[1694] Mauricio Carrasquilla-Henao, K. Lynne Yamanaka, Dana Haggarty, and Francis Juanes. Predicting important rockfish (Sebastes spp.) habitat from large-scale longline surveys for southern British Columbia, Canada. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 76(5):682–694, May 2019. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2017-0458.

Wringe:2019:DES

[1695] Brendan F. Wringe, Eric C. Anderson, Nicholas W. Jeffery, Ryan R. E. Stanley, and Ian R. Bradbury. Development and evaluation of SNP panels for the detection of hybridization between wild and escaped Atlantic salmon (Salmo salar) in the western Atlantic. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 76(5):695–704, May 2019. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/cjfas-2017-0394.

Hayes:2019:INF

[1696] John W. Hayes, Eric O. Goodwin, Joanne E. Clapcott, and Karen A. Shearer. The influence of natural flow and temperature and introduced brown trout on the temporal variation in native fish abundance in a “reference” stream. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 76 (5):705–722, May 2019. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2018-0033.

Finucci:2019:ULM

[1697] Brittany Finucci, Matthew R. Dunn, and Richard Arnold. Using length– mass relationships to estimate life history: an application to deepsea fishes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 76(5):723–739, May 2019. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
cjfas-2018-0017.

Turgeon:2019:BRI

[1698] Katrine Turgeon, Christian Turpin, and Irene Gregory-Eaves. Boreal river impoundments caused nearshore fish community assemblage shifts but little change in diversity: a multiscale analysis. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 76(5):740–752, May 2019. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2017-0561.

Kovach:2019:ELU

[1699] Ryan P. Kovach, Clint C. Muhlfeld, Robert Al-Chokhachy, Jeffrey V. Ojala, and Eric K. Archer. Effects of land use on summer thermal regimes in critical salmonid habitats of the Pacific Northwest. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 76(5):753–761, May 2019. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2018-0165.

Stich:2019:DPP

[1700] Daniel S. Stich, Timothy F. Sheehan, and Joseph D. Zydlewski. A dam passage performance standard model for American shad. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 76(5):762–779, May 2019. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2018-0008.

Helser:2019:TAA

[1701] Thomas E. Helser, Irina Benson, Jason Erickson, Jordan Healy, Craig Kastelle, and Jonathan A. Short. A transformative approach to ageing fish otoliths using Fourier transform near infrared spectroscopy: a case study of eastern Bering Sea walleye pollock (Gadus chalcogrammus). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 76(5):780–789, May 2019. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2018-0112.

Reed:2019:HEM

[1702] Thomas E. Reed, Paulo Prodöhl, Caroline Bradley, John Gilbey, Philip McGinnity, Craig R. Primmer, and Philip J. Bacon. Heritability estimation via molecular pedigree reconstruction in a wild fish population reveals substantial evolutionary potential for sea age at maturity, but not size within age classes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 76 (5):790–805, May 2019. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2018-0123.

Simonin:2019:EMF

[1703] Paul W. Simonin, Lars G. Rudstam, Patrick J. Sullivan, Donna L. Parrish, and Bernard Pientka. Early mortality and freshwater forage fish recruitment: nonnative alewife and native rainbow smelt interactions in Lake Champlain. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 76 (5):806–814, May 2019. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2017-0571.

Bozeman:2019:MFS

[1704] Bryan B. Bozeman and Gary D. Grossman. Mechanics of foraging success and optimal microhabitat selection in Alaskan Arctic grayling (Thymallus arcticus). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 76(5):815– 830, May 2019. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
cjfas-2018-0115.

Watras:2019:MTC

[1705] C. J. Watras, D. Grande, A. W. Latzka, and L. S. Tate. Mercury trends and cycling in northern Wisconsin related to atmospheric and hydrologic processes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 76 (5):831–846, May 2019. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2018-0157.

Miller:2019:ECE

[1706] Timothy J. Miller, Deborah R. Hart, Karen Hopkins, Norman H. Vine, Richard Taylor, Amber D. York, and Scott M. Gallager. Estimation of the capture efficiency and abundance of Atlantic sea scallops (Placopecten magellanicus) from paired photographic–dredge tows using hierarchical models. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 76(6):847– 855, June 2019. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
cjfas-2018-0024.

McKinnell:2019:MSM

[1707] Skip McKinnell. Modelling salmon migration as a mixture problem. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 76(6):856–870, June 2019. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2017-0546.

Hayes:2019:RBB

[1708] John W. Hayes, Eric O. Goodwin, Karen A. Shearer, and D. Murray Hicks. Relationship between background invertebrate drift concentration and flow over natural flow recession and prediction with a drift transport model. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 76(6):871– 885, June 2019. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
cjfas-2017-0340.

Wuenschel:2019:MFC

[1709] Mark J. Wuenschel, W. David McElroy, Kenneth Oliveira, and Richard S. McBride. Measuring fish condition: an evaluation of new and old metrics for three species with contrasting life histories. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 76(6):886–903, June 2019. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/cjfas-2018-0076.

Xu:2019:SAD

[1710] Meng Xu, Jeppe Kolding, and Joel E. Cohen. Sequential analysis and design of fixed-precision sampling of Lake Kariba fishes using Taylor’s power law. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 76(6):904–917, June 2019. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
cjfas-2018-0091.

Espinasse:2019:IHS

[1711] Boris Espinasse, Brian P. V. Hunt, Yago Doson Coll, and Evgeny A. Pakhomov. Investigating high seas foraging conditions for salmon in the North Pacific: insights from a 100-year scale archive for Rivers Inlet sockeye salmon. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 76 (6):918–927, June 2019. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2018-0010.

Ciepiela:2019:QTS

[1712] Lindsy R. Ciepiela and Annika W. Walters. Quantifying ^{87}Sr/^{86}Sr temporal stability and spatial heterogeneity for use in tracking fish movement. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 76(6):928–936, June 2019. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
cjfas-2018-0124.

Guan:2019:ISV

[1713] Lisha Guan, Yong Chen, James A. Wilson, Timothy Waring, Lisa A. Kerr, and Xiujuan Shan. The influence of spatially variable and connected recruitment on complex stock dynamics and its ecological and management implications. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 76 (6):937–949, June 2019. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2018-0151.

Schroeder:2019:SWV

[1714] Isaac D. Schroeder, Jarrod A. Santora, Steven J. Bograd, Elliott L. Hazen, Keith M. Sakuma, Andrew M. Moore, Christopher A. Edwards, Brian K. Wells, and John C. Field. Source water variability as a driver of rockfish recruitment in the California Current ecosystem: implications for climate change and fisheries management. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 76(6):950–960, June 2019. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/cjfas-2017-0480.

Hasler:2019:ACD

[1715] Caleb T. Hasler, Christa M. Woodley, Eric V. Schneider, Bryton K. Hixson, Cynthia J. Fowler, Stephen R. Midway, Cory D. Suski, and David L.

Smith. Avoidance of carbon dioxide in flowing water by bighead carp. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 76(6):961–969, June 2019. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2018-0026.

Carruthers:2019:LSS

[1716] Thomas R. Carruthers, Kornelia Dabrowska, Wolfgang Haider, Eric A. Parkinson, Divya A. Varkey, Hillary Ward, Murdoch K. McAllister, Theresa Godin, Brett Van Poorten, Paul J. Askey, Kyle L. Wilson, Len M. Hunt, Adrian Clarke, Eric Newton, Carl Walters, and John R. Post. Landscape-scale social and ecological outcomes of dynamic angler and fish behaviours: processes, data, and patterns. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 76(6):970–988, June 2019. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2018-0168.

Feddern:2019:NMM

[1717] Megan Feddern, Hannah R. Bassett, Katherine N. McElroy, Marta Ree, Marcus Gho, and Ray Hilborn. A novel method for modeling age and length selectivity of sockeye salmon as applied to the Bristol Bay Port Moller test fishery. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 76 (6):989–997, June 2019. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2018-0018.

Ziegler:2019:CWH

[1718] Jacob P. Ziegler, Colin J. Dassow, Stuart E. Jones, Alexander J. Ross, and Christopher T. Solomon. Coarse woody habitat does not predict largemouth bass young of year mortality during the open-water season. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 76(6):998–1005, June 2019. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2018-0050.

Wiedenmann:2019:CRO

[1719] John Wiedenmann and Olaf P. Jensen. Could recent overfishing of New England groundfish have been prevented? A retrospective evaluation of alternative management strategies. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 76(6):1006–1018, June 2019. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/cjfas-2018-0129.

Faust:2019:ATP

[1720] Matthew D. Faust, Christopher S. Vandergoot, Travis O. Brenden, Richard. T. Kraus, Travis Hartman, and Charles C. Krueger. Acoustic telemetry as a potential tool for mixed-stock analysis of fishery harvest: a feasibility study using Lake Erie walleye. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 76(6):1019–1030, June 2019. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/cjfas-2017-0522.

Pinkerton:2019:RSJ

[1721] Evelyn Pinkerton, Anne K. Salomon, and Frank Dragon. Reconciling social justice and ecosystem-based management in the wake of a successful predator reintroduction. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 76 (6):1031–1039, June 2019. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2018-0441.

Iverson:2019:OTN

[1722] Sara J. Iverson, Aaron T. Fisk, Scott G. Hinch, Joanna Mills Flemming, Steven J. Cooke, and Frederick G. Whoriskey. The ocean tracking network: Advancing frontiers in aquatic science and management. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 76(7):1041–1051, July 2019. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2018-0481.

Larranaga:2019:IBT

[1723] Nicolas Larranaga, Magnus L. Wallerius, Haoyu Guo, Julien Cucherousset, and Jörgen I. Johnsson. Invasive brook trout disrupt the diel activity and aggregation patterns of native brown trout. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 76(7):1052–1059, July 2019. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2018-0110.

Pitman:2019:LAF

[1724] Kara J. Pitman, Samantha M. Wilson, Elissa Sweeney-Bergen, Paddy Hirshfield, Mark C. Beere, and Jonathan W. Moore. Linking anglers, fish, and management in a catch-and-release steelhead trout fishery. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 76(7):1060–1072, July 2019. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2018-0080.

Michielsens:2019:SAM

[1725] Catherine G. J. Michielsens and Jim D. Cave. In-season assessment and management of salmon stocks using a Bayesian time–density model. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 76(7):1073–1085, July 2019. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2018-0213.

Wathen:2019:BAI

[1726] Gus Wathen, Jacob E. Allgeier, Nicolaas Bouwes, Michael M. Pollock, Daniel E. Schindler, and Chris E. Jordan. Beaver activity increases habitat complexity and spatial partitioning by steelhead trout. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 76(7):1086–1095, July 2019. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2018-0171.

Grant:2019:LTS

[1727] Jon Grant, Michelle Simone, and Tara Daggett. Long-term studies of lobster abundance at a salmon aquaculture site, eastern Canada. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 76(7):1096–1102, July 2019. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2017-0547.

Feeney:2019:IMS

[1728] R. G. Feeney, D. V. Boelke, J. J. Deroba, S. Gaichas, B. J. Irwin, and M. Lee. Integrating management strategy evaluation into fisheries management: advancing best practices for stakeholder inclusion based on an MSE for Northeast US Atlantic herring. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 76(7):1103–1111, July 2019. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/cjfas-2018-0125.

Deroba:2019:DRM

[1729] Jonathan J. Deroba, Sarah K. Gaichas, Min-Yang Lee, Rachel G. Feeney, Deirdre Boelke, and Brian J. Irwin. The dream and the reality: meeting decision-making time frames while incorporating ecosystem and economic models into management strategy evaluation. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 76(7):1112–1133, July 2019. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2018-0128.

Marjadi:2019:LBS

[1730] Meghna N. Marjadi, Allison H. Roy, Adrian Jordaan, Benjamin I. Gahagan, Michael P. Armstrong, and Andrew R. Whiteley. Larger body size and earlier run timing increase alewife reproductive success in a whole lake experiment. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 76 (7):1134–1146, July 2019. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2017-0451.

Dammerman:2019:RCA

[1731] Kari J. Dammerman, Molly A. H. Webb, and Kim T. Scribner. Riverine characteristics and adult demography influence female lake sturgeon (Acipenser fulvescens) spawning behavior, reproductive success, and ovarian quality. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 76 (7):1147–1160, July 2019. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2018-0141.

Puffer:2019:DIU

[1732] Michael Puffer, Ole Kristian Berg, Frøydis Bolme Hamnes, Vidar Bentsen, Wouter Koch, Ola Ugedal, Torbjørn Forseth, Jo Vegar Arnekleiv, and Sigurd Einum. Density-independent use of shallow riverine areas in juvenile Atlantic salmon. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 76 (7):1161–1170, July 2019. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2017-0500.

Crisafulli:2019:DSC

[1733] B. M. Crisafulli, D. V. Fairclough, I. S. Keay, P. Lewis, J. R. How, K. L. Ryan, S. M. Taylor, and C. B. Wakefield. Does a spatiotemporal closure to fishing Chrysophrys auratus (Sparidae) spawning aggregations also protect individuals during migration? Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 76(7):1171–1185, July 2019. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/cjfas-2017-0449.

Thresher:2019:EAG

[1734] Ronald E. Thresher, Michael Jones, and D. Andrew R. Drake. Evaluating active genetic options for the control of sea lamprey (Petromyzon marinus) in the Laurentian Great Lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 76(7):1186–1202, July 2019. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/cjfas-2018-0153.

Davis:2019:GSC

[1735] Matthew M. Davis, Pablo de Jesus Suárez-Moo, and Toby S. Daly-Engel. Genetic structure and congeneric range overlap among sharpnose sharks (genus Rhizoprionodon) in the Northwest Atlantic Ocean. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 76(7):1203–1211, July 2019. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2018-0019.

Bond:2019:EBW

[1736] Morgan H. Bond, Tyler G. Nodine, Tim J. Beechie, and Richard W. Zabel. Estimating the benefits of widespread floodplain reconnection for Columbia River Chinook salmon. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 76(7):1212–1226, July 2019. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/cjfas-2018-0108.

Lake:2019:USI

[1737] James L. Lake, Jonathan R. Serbst, Anne Kuhn, Nathan J. Smucker, Phillip Edwards, Alan Libby, Michael Charpentier, and Kenneth Miller. Use of stable isotopes in benthic organic material as a baseline for estimating fish trophic positions in lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 76(7):1227–1237, July 2019. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/cjfas-2017-0381.

Brooks:2019:BIM

[1738] J. L. Brooks, J. M. Chapman, A. N. Barkley, S. T. Kessel, N. E. Hussey, S. G. Hinch, D. A. Patterson, K. J. Hedges, S. J. Cooke, A. T. Fisk, S. H. Gruber, and V. M. Nguyen. Biotelemetry informing management: case studies exploring successful integration of biotelemetry data into fisheries and habitat management. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 76 (7):1238–1252, July 2019. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2017-0530.

Geiger:2019:MTQ

[1739] Hal Geiger. In memory of Terry Quinn (1952–2019). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 76(8), August 2019. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/cjfas-2019-0194.

Thorn:2019:TEE

[1740] Michael W. Thorn, Morag F. Dick, Leo Oviedo, Christopher G. Guglielmo, and Yolanda E. Morbey. Transgenerational effects of egg nutrients on the early development of Chinook salmon (Oncorhynchus tshawytscha) across a thermal gradient. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 76(8):1253–1262, August 2019. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2018-0013.

Sanchez-Hernandez:2019:DDP

[1741] Javier Sánchez-Hernández, Anders G. Finstad, Jo Vegar Arnekleiv, Gaute Kjærstad, and Per-Arne Amundsen. Drivers of diet patterns in a globally distributed freshwater fish species. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 76(8):1263–1274, August 2019. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2018-0207.

OLeary:2019:UHS

[1742] Cecilia A. O’Leary, Timothy J. Miller, James T. Thorson, and Janet A. Nye. Understanding historical summer flounder (Paralichthys dentatus) abundance patterns through the incorporation of oceanographydependent vital rates in Bayesian hierarchical models. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 76(8):1275–1294, August 2019. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2018-0092.

Targett:2019:GEF

[1743] Timothy E. Targett, Paul A. Grecay, and Rachel L. Dixon. Growth of the estuarine fish Fundulus heteroclitus in response to diel-cycling hypoxia and acidification: interaction with temperature. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 76(8):1295–1304, August 2019. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2018-0216.

Millar:2019:IAE

[1744] Russell B. Millar and Christopher D. Nottingham. Improved approximations for estimation of size-transition probabilities within size-structured models. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 76(8):1305–1313, August 2019. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
cjfas-2017-0444.

Riedl:2019:IPA

[1745] Hannah L. Riedl, William H. Clements, and Liba Pejchar. An introduced plant is associated with declines in terrestrial arthropods, but no change in stream invertebrates. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 76(8): 1314–1325, August 2019. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2018-0098.

McQueen:2019:GCG

[1746] Kate McQueen, J. Paige Eveson, Bodo Dolk, Thomas Lorenz, Thomas Mohr, Franziska M. Schade, and Uwe Krumme. Growth of cod (Gadus morhua) in the western Baltic Sea: estimating improved growth parameters from tag–recapture data. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 76 (8):1326–1337, August 2019. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2018-0081.

Pinto:2019:CMD

[1747] Cecilia Pinto, Morgane Travers-Trolet, Jed I. Macdonald, Etienne Rivot, and Youen Vermard. Combining multiple data sets to unravel the spatiotemporal dynamics of a data-limited fish stock. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 76(8):1338–1349, August 2019. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2018-0149.

Herrmann:2019:SSL

[1748] Bent Herrmann, Manu Sistiaga, Eduardo Grimaldo, Roger B. Larsen, Leonore Olsen, Jesse Brinkhof, and Ivan Tatone. Size selectivity and length-dependent escape behaviour of haddock in a sorting device combining a grid and a square mesh panel. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 76(8):1350–1361, August 2019. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/cjfas-2017-0461.

Adelfio:2019:WWR

[1749] Luca A. Adelfio, Steven M. Wondzell, Nathan J. Mantua, and Gordon H. Reeves. Warm winters reduce landscape-scale variability in the duration of egg incubation for Coho salmon (Oncorhynchus kisutch) on the Copper River Delta, Alaska. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 76 (8):1362–1375, August 2019. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2018-0152.

Kennedy:2019:CCE

[1750] Patrick J. Kennedy, Paul J. Blanchfield, Karen A. Kidd, Michael J. Paterson, Cheryl L. Podemski, and Michael D. Rennie. Changes in the condition, early growth, and trophic position of lake trout (Salvelinus namaycush) in response to an experimental aquaculture operation. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 76(8):1376–1387, August 2019. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2017-0578.

Hanks:2019:EID

[1751] R. Daniel Hanks and Kyle J. Hartman. Evaluation of the influences of dam release types, land use, and habitat affecting abundance, richness, diversity, and community structure of larval and juvenile fish. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 76(8):1388–1397, August 2019. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2017-0107.

Michel:2019:DOM

[1752] Cyril J. Michel. Decoupling outmigration from marine survival indicates outsized influence of streamflow on cohort success for California’s Chinook salmon populations. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 76 (8):1398–1410, August 2019. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2018-0140.

Johnston:2019:PRB

[1753] T. A. Johnston, A. D. Ehrman, G. L. Hamilton, B. K. Nugent, P. A. Cott, and J. M. Gunn. Plenty of room at the bottom: niche variation and segregation in large-bodied benthivores of boreal lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 76(8):1411–1422, August 2019. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2018-0180.

Lopes:2019:PSD

[1754] Priscila F. M. Lopes, Júlia T. Verba, Alpina Begossi, and Maria Grazia Pennino. Predicting species distribution from fishers’ local ecological knowledge: a new alternative for data-poor management. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 76(8):1423–1431, August 2019. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2018-0148.

Watson:2019:DDS

[1755] Breanna M. Watson, Carlo A. Biagi, Sara L. Northrup, Michael L. A. Ohata, Colin Charles, Paul J. Blanchfield, Sam V. Johnston, Paul J. Askey, Brett T. van Poorten, and Robert H. Devlin. Distinct diel and seasonal behaviours in rainbow trout detected by fine-scale acoustic telemetry in a lake environment. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 76 (8):1432–1445, August 2019. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2018-0293.

Childress:2019:DER

[1756] Evan S. Childress, Keith H. Nislow, Andrew R. Whiteley, Matthew J. O’Donnell, and Benjamin H. Letcher. Daily estimates reveal fine-scale temporal and spatial variation in fish survival across a stream network. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 76(8):1446–1458, August 2019. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2018-0191.

Paterson:2019:LTC

[1757] Michael J. Paterson, Kenneth G. Beaty, David L. Findlay, Willie J. Findlay, Sherry L. Schiff, Vincent L. St. Louis, and Jason J. Venkiteswaran. Long-term changes in nutrient dynamics and plankton communities following the creation of a new reservoir. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 76(8):1459–1469, August 2019. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/cjfas-2018-0197.

Patterson:2019:DAA

[1758] Toby A. Patterson and Richard D. Pillans. Designing acoustic arrays for estimation of mortality rates in riverine and estuarine systems. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 76(9):1471–1479, September 2019. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2018-0090.

Lento:2019:EBB

[1759] Jennifer Lento, Michelle A. Gray, Allison J. Ferguson, and R. Allen Curry. Establishing baseline biological conditions and monitoring metrics for stream benthic macroinvertebrates and fish in an area of potential shale gas development. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 76(9): 1480–1494, September 2019. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2017-0574.

Miller:2019:OHR

[1760] Steve Miller, Andrew Rassweiler, Laura Dee, Kristin M. Kleisner, Tracey Mangin, Ricardo Oliveros-Ramos, Jorge Tam, Francisco P. Chavez, Miguel Ñiquen, Sarah E. Lester, Merrick Burden, Steven Gaines, and Christopher Costello. Optimal harvest responses to environmental forecasts depend on resource knowledge and how it can be used. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 76(9):1495–1502, September 2019. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2018-0283.

Ziegler:2019:MDS

[1761] Catherine M. Ziegler, Joshua P. Zacharias, and Michael G. Frisk. Migration diversity, spawning behavior, and habitat utilization of winter flounder. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 76(9):1503–1514, September 2019. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
cjfas-2018-0053.

Bjornsson:2019:TBC

[1762] Björn Björnsson. Thermoregulatory behaviour in cod: Is the thermal preference in free-ranging adult Atlantic cod affected by food abundance? Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 76(9):1515–1527, September 2019. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
cjfas-2018-0305.

Tableau:2019:DCP

[1763] Adrien Tableau, Jeremy S. Collie, Richard J. Bell, and Cóilín Minto. Decadal changes in the productivity of New England fish populations. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 76(9):1528–1540, September 2019. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
cjfas-2018-0255.

Mickle:2019:HCB

[1764] Megan F. Mickle, Scott M. Miehls, Nicholas S. Johnson, and Dennis M. Higgs. Hearing capabilities and behavioural response of sea lamprey (Petromyzon marinus) to low-frequency sounds. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 76(9):1541–1548, September 2019. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2018-0359.

Henderson:2019:EST

[1765] Mark J. Henderson, Ilysa S. Iglesias, Cyril J. Michel, Arnold J. Ammann, and David D. Huff. Estimating spatial–temporal differences in Chinook salmon outmigration survival with habitat- and predation-related covariates. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 76(9):1549–1561, September 2019. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
cjfas-2018-0212.

McCallum:2019:ART

[1766] Erin S. McCallum, Aneesh P. H. Bose, Naylor Lobban, Julie R. Marentette, Harri Pettitt-Wade, Marten A. Koops, Aaron T. Fisk, and Sigal Balshine. Alternative reproductive tactics, an overlooked source of life history variation in the invasive round goby. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 76(9):1562–1570, September 2019. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2018-0340.

Chudnow:2019:ECP

[1767] Rachel Chudnow, Brett van Poorten, and Murdoch McAllister. Estimating cross-population variation in juvenile compensation in survival for bull trout (Salvelinus confluentus): a Bayesian hierarchical approach. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 76(9):1571–1580, September 2019. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
cjfas-2017-0555.

Almodovar:2019:CDB

[1768] Ana Almodóvar, Daniel Ayllón, Graciela G. Nicola, Bror Jonsson, and Benigno Elvira. Climate-driven biophysical changes in feeding and breeding environments explain the decline of southernmost European Atlantic salmon populations. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 76(9): 1581–1595, September 2019. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2018-0297.

Yoon:2019:ETD

[1769] Gwangseok R. Yoon, David Deslauriers, Eva C. Enders, Jason R. Treberg, and W. Gary Anderson. Effects of temperature, dissolved oxygen, and substrate on the development of metabolic phenotypes in age-0 lake sturgeon (Acipenser fulvescens): implications for overwintering survival. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 76(9):1596–1607, September 2019. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
cjfas-2018-0399.

Seitz:2019:PSA

[1770] Andrew C. Seitz, Michael B. Courtney, Mark D. Evans, and Kaitlyn Manishin. Pop-up satellite archival tags reveal evidence of intense predation on large immature Chinook salmon (Oncorhynchus tshawytscha) in the North Pacific Ocean. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 76(9): 1608–1615, September 2019. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2018-0490.

Edmunds:2019:RBW

[1771] Nicholas B. Edmunds, Timothy J. Bartley, Amanda Caskenette, Frédéric Laberge, and Kevin S. McCann. Relationship between water transparency and walleye (Sander vitreus) muscle glycolytic potential in northwestern Ontario lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 76 (9):1616–1623, September 2019. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/

doi/full/10.1139/cjfas-2017-0556.

Sagarese:2019:LLD

[1772] Skyler R. Sagarese, William J. Harford, John F. Walter, Meaghan D. Bryan, J. Jeffery Isely, Matthew W. Smith, Daniel R. Goethel, Adyan B. Rios, Shannon L. Cass-Calay, Clay E. Porch, Thomas R. Carruthers, and Nancie J. Cummings. Lessons learned from data-limited evaluations of data-rich reef fish species in the Gulf of Mexico: implications for providing fisheries management advice for data-poor stocks. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 76(9):1624–1639, September 2019. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2017-0482.

Hutniczak:2019:VCF

[1773] Barbara Hutniczak, Douglas Lipton, John Wiedenmann, and Michael Wilberg. Valuing changes in frequency of fish stock assessments. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 76(9):1640–1652, September 2019. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2018-0130.

Carruthers:2019:UMS

[1774] T. R. Carruthers and A. R. Hordyk. Using management strategy evaluation to establish indicators of changing fisheries. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 76(9):1653–1668, September 2019. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2018-0223.

Cunningham:2019:MSE

[1775] Curry J. Cunningham, Christopher M. Anderson, Jocelyn Yun-Ling Wang, Michael Link, and Ray Hilborn. A management strategy evaluation of the commercial sockeye salmon fishery in Bristol Bay, Alaska. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 76(9):1669–1683, September 2019. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
cjfas-2018-0133.

Stephen:2019:AUC

[1776] Craig Stephen, Laura Griffith-Cochrane, and Joy Wade. Addressing the unique challenges of community-based capture–hold–release aquariums through a facility health program. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 76(9):1684–1687, September 2019. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/cjfas-2018-0425.

Goethel:2019:RAM

[1777] Daniel R. Goethel, Sean M. Lucey, Aaron M. Berger, Sarah K. Gaichas, Melissa A. Karp, Patrick D. Lynch, and John F. Walter. Recent advances in management strategy evaluation: introduction to the special issue “Under pressure: addressing fisheries challenges with Management Strategy Evaluation”. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 76(10): 1689–1696, October 2019. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2019-0084.

Hutchings:2019:LHV

[1778] Jeffrey A. Hutchings, William R. Ardren, Bjørn T. Barlaup, Eva Bergman, Keith D. Clarke, Larry A. Greenberg, Colin Lake, Jorma Piironen, Pascal Sirois, Line E. Sundt-Hansen, and Dylan J. Fraser. Life-history variability and conservation status of landlocked Atlantic salmon: an overview. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 76(10): 1697–1708, October 2019. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2018-0413.

Mazur:2019:UIB

[1779] Mackenzie Mazur, Bai Li, Jui-Han Chang, and Yong Chen. Using an individual-based model to simulate the Gulf of Maine American lobster (Homarus americanus) fishery and evaluate the robustness of current management regulations. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 76 (10):1709–1718, October 2019. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/

doi/full/10.1139/cjfas-2018-0122.

Staton:2019:BIU

[1780] Benjamin A. Staton and Matthew J. Catalano. Bayesian information updating procedures for Pacific salmon run size indicators: evaluation in the presence and absence of auxiliary migration timing information. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 76(10):1719–1727, October 2019. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2018-0176.

Du:2019:CBS

[1781] Xue L. Du, Irena F. Creed, Ryan J. Sorichetti, and Charles G. Trick. Cyanobacteria biomass in shallow eutrophic lakes is linked to the presence of iron-binding ligands. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 76 (10):1728–1739, October 2019. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/

doi/full/10.1139/cjfas-2018-0261.

Steel:2019:TLC

[1782] E. Ashley Steel, Amy Marsha, Aimee H. Fullerton, Julian D. Olden, Narasimhan K. Larkin, Se-Yeun Lee, and Akida Ferguson. Thermal landscapes in a changing climate: biological implications of water temperature patterns in an extreme year. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 76(10):1740–1756, October 2019. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/cjfas-2018-0244.

Tsuboi:2019:TOA

[1783] Jun ichi Tsuboi, Kohichi Kaji, Shinya Baba, and Robert Arlinghaus. Trade-offs in the adaptation towards hatchery and natural conditions drive survival, migration, and angling vulnerability in a territorial fish in the wild. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 76(10):1757–1767, October 2019. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
cjfas-2018-0256.

Casas-Monroy:2019:ETC

[1784] Oscar Casas-Monroy, Julie Vanden Byllaardt, Johanna Bradie, Andrea Sneekes, Klaas Kaag, and Sarah A. Bailey. Effect of temperature on chlorine treatment for elimination of freshwater phytoplankton in ballast water: bench-scale test. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 76(10): 1768–1780, October 2019. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2018-0179.

Karlsen:2019:UVD

[1785] Junita Diana Karlsen, Ludvig Ahm Krag, Bent Herrmann, and Henrik Skaarup Lund. Using vertical distribution to separate fish from crustaceans in a mixed species trawl fishery. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 76(10):1781–1794, October 2019. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2018-0313.

Stevens:2019:MID

[1786] Justin R. Stevens, John F. Kocik, and Timothy F. Sheehan. Modeling the impacts of dams and stocking practices on an endangered Atlantic salmon (Salmo salar) population in the Penobscot River, Maine, USA. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 76(10):1795–1807, October 2019.

CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2018-0225.

Warshafsky:2019:TSB

[1787] Zoemma T. Warshafsky, Troy D. Tuckey, Wolfgang K. Vogelbein, Robert J. Latour, and Andrew R. Wargo. Temporal, spatial, and biological variation of nematode epidemiology in American eels. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 76(10):1808–1818, October 2019. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2018-0136.

Johnson:2019:ERD

[1788] Samuel D. N. Johnson and Sean P. Cox. Evaluating the role of data quality when sharing information in hierarchical multistock assessment models, with an application to Dover sole. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 76(10):1819–1835, October 2019. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/cjfas-2018-0048.

Smeltz:2019:SSH

[1789] T. Scott Smeltz, Bradley P. Harris, John V. Olson, and Suresh A. Sethi. A seascape-scale habitat model to support management of fishing impacts on benthic ecosystems. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 76(10):1836–1844, October 2019. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/cjfas-2018-0243.

Mickle:2019:BMC

[1790] Megan F. Mickle, Christopher M. Harris, Oliver P. Love, and Dennis M. Higgs. Behavioural and morphological changes in fish exposed to ecologically relevant boat noises. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 76(10):1845–1853, October 2019. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/cjfas-2018-0258.

Vivancos:2019:UFS

[1791] Aurélien Vivancos and Gerry Closs. Using fine-scale spatial analysis to study behavioural strategies prevalence in wild groups of drift-feeding fish. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 76(10):1854–1861, October 2019. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
cjfas-2018-0221.

Rub:2019:CAC

[1792] A. Michelle Wargo Rub, Nicholas A. Som, Mark J. Henderson, Benjamin P. Sandford, Donald M. Van Doornik, David J. Teel, Matthew J. Tennis, Olaf P. Langness, Bjorn K. van der Leeuw, and David D. Huff. Changes in adult Chinook salmon (Oncorhynchus tshawytscha) survival within the lower Columbia River amid increasing pinniped abundance. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 76(10):1862–1873, October 2019. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2018-0290.

Andrews:2019:GLH

[1793] Allen H. Andrews, Edward E. DeMartini, Jon Brodziak, Ryan S. Nichols, and Robert L. Humphreys. Growth and longevity of Hawaiian grouper (Hyporthodus quernus) — input for management and conservation of a large, slow-growing grouper. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 76 (10):1874–1884, October 2019. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/

doi/full/10.1139/cjfas-2018-0170.

Harlyan:2019:VFH

[1794] Ledhyane Ika Harlyan, Dengke Wu, Ryosuke Kinashi, Methee Kaewnern, and Takashi Matsuishi. Validation of a feedback harvest control rule in data-limited conditions for managing multispecies fisheries. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 76(10):1885–1893, October 2019. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2018-0318.

Goethel:2019:CFL

[1795] Daniel R. Goethel, Sean M. Lucey, Aaron M. Berger, Sarah K. Gaichas, Melissa A. Karp, Patrick D. Lynch, John F. Walter, Jonathan J. Deroba, Shana Miller, and Michael J. Wilberg. Closing the feedback loop: on stakeholder participation in management strategy evaluation. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 76(10):1895–1913, October 2019. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2018-0162.

Lamothe:2019:MRE

[1796] Karl A. Lamothe and D. Andrew R. Drake. Moving repatriation efforts forward for imperilled Canadian freshwater fishes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 76(10):1914–1921, October 2019. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2018-0295.

Litzow:2019:NEO

[1797] Michael A. Litzow, Lorenzo Ciannelli, Curry J. Cunningham, Bethany Johnson, and Patricia Puerta. Nonstationary effects of ocean temperature on Pacific salmon productivity. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 76(11):1923–1928, November 2019. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/cjfas-2019-0120.

Prestie:2019:EOI

[1798] Kate Prestie, Iain D. Phillips, Douglas P. Chivers, and Timothy D. Jardine. Effects of ontogeny and invasive crayfish on feeding ecology and mercury concentrations of predatory fishes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 76(11):1929–1939, November 2019. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2018-0455.

Morrison:2019:DOS

[1799] Christie M. Morrison, Mélodie Kunegel-Lion, Colin P. Gallagher, Rick J. Wastle, Ellen V. Lea, Tracey N. Loewen, James D. Reist, Kimberly L. Howland, and Keith B. Tierney. Decoupling of otolith and somatic growth during anadromous migration in a northern salmonid. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 76(11):1940–1953, November 2019. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2018-0306.

Gebauer:2019:PEI

[1800] Radek Gebauer, Lukáš Veselý, Tatyana Vanina, Miloš Buřič, Antonín Kouba, and Bořek Drozd. Prediction of ecological impact of two alien gobiids in habitat structures of differing complexity. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 76(11):1954–1961, November 2019. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2018-0346.

Janisse:2019:LCV

[1801] Kevyn Janisse, Pauline M. Capelle, John W. Heath, Mitchel G. Dender, Daniel D. Heath, and Christina A. D. Semeniuk. Life in captivity: varied behavioural responses to novel setting and food types in first-generation hybrids of farmed and wild juvenile Chinook salmon (Oncorhynchus tshawytscha). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 76(11):1962– 1970, November 2019. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2018-0311.

Ciepiela:2019:LHV

[1802] Lindsy R. Ciepiela and Annika W. Walters. Life-history variation of two inland salmonids revealed through otolith microchemistry analysis. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 76(11):1971–1981, November 2019. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
cjfas-2018-0087.

Bunnell:2019:SVE

[1803] David B. Bunnell, Steven A. Pothoven, Patricia M. Armenio, Lauren Eaton, David M. Warner, Ashley K. Elgin, Lyuba E. Burlakova, and Alexander Y. Karatayev. Spatiotemporal variability in energetic condition of alewife and round goby in Lake Michigan. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 76(11):1982–1992, November 2019. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2018-0391.

Shink:2019:UDM

[1804] Katie G. Shink, Trent M. Sutton, James M. Murphy, and J. Andrés López. Utilizing DNA metabarcoding to characterize the diet of marinephase Arctic lamprey (Lethenteron camtschaticum) in the eastern Bering Sea. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 76(11):1993–2002, November 2019. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
cjfas-2018-0299.

Bade:2019:SBD

[1805] Andrew P. Bade, Thomas R. Binder, Matthew D. Faust, Christopher S. Vandergoot, Travis J. Hartman, Richard T. Kraus, Charles C. Krueger, and Stuart A. Ludsin. Sex-based differences in spawning behavior account for male-biased harvest in Lake Erie walleye (Sander vitreus). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 76(11):2003–2012, November 2019. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
cjfas-2018-0339.

Baldigo:2019:PRB

[1806] Barry P. Baldigo, Scott D. George, Timothy J. Sullivan, Charles T. Driscoll, Douglas A. Burns, Shuai Shao, and Gregory B. Lawrence. Probabilistic relations between acid–base chemistry and fish assemblages in streams of the western Adirondack Mountains, New York, USA. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 76(11):2013–2026, November 2019. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2018-0260.

Kanive:2019:SSA

[1807] Paul E. Kanive, Jay J. Rotella, Salvador J. Jorgensen, Taylor K. Chapple, James E. Hines, Scot D. Anderson, and Barbara A. Block. Size-specific apparent survival rate estimates of white sharks using mark–recapture models. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 76(11):2027–2034, November 2019. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
cjfas-2018-0142.

Maguffee:2019:EPO

[1808] Alexander C. Maguffee, Reneé Reilly, Richard Clark, and Michael L. Jones. Examining the potential of otolith chemistry to determine natal origins of wild Lake Michigan Chinook salmon. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 76(11):2035–2044, November 2019. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2018-0041.

Nishijima:2019:IMT

[1809] Shota Nishijima, Shigenori Suzuki, Momoko Ichinokawa, and Hiroshi Okamura. Integrated multi-timescale modeling untangles anthropogenic, environmental, and biological effects on catchability. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 76(11):2045–2056, November 2019. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2018-0114.

Mayfield:2019:AMM

[1810] Mariah P. Mayfield, Thomas E. McMahon, Jay J. Rotella, Robert E. Gresswell, Trevor Selch, Patrick Saffel, Jason Lindstrom, and Bradley Liermann. Application of multistate modeling to estimate salmonid survival and movement in relation to spatial and temporal variation in metal exposure in a mining-impacted river. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 76(11):2057–2068, November 2019. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/cjfas-2018-0280.

Mitz:2019:TDS

[1811] Charles Mitz, Christopher Thome, Mary Ellen Cybulski, Christopher M. Somers, Richard G. Manzon, Joanna Y. Wilson, and Douglas R. Boreham. Thermal dependence of size-at-hatch in the lake whitefish (Coregonus clupeaformis). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 76(11): 2069–2079, November 2019. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2018-0097.

Kelson:2019:DGS

[1812] Suzanne J. Kelson, Michael R. Miller, Tasha Q. Thompson, Sean M. O’Rourke, and Stephanie M. Carlson. Do genomics and sex predict migration in a partially migratory salmonid fish, Oncorhynchus mykiss? Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 76(11):2080–2088, November 2019. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
cjfas-2018-0394.

Li:2019:DIG

[1813] Yan Li and Tyler Wagner. Does incorporating gear selectivity during macroscale investigations of fish growth reduce size-selective sampling bias in parameter estimates? Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 76(11):2089–2101, November 2019. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/cjfas-2018-0355.

Stuparyk:2019:ECM

[1814] Blake R. Stuparyk, Mark Graham, Jenna Cook, Mitchell A. Johnsen, Karen K. Christensen-Dalsgaard, and Rolf D. Vinebrooke. Experimental culling of minnows suppresses cyanobacterial bloom under low-nutrient conditions. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 76(11):2102–2109, November 2019. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
cjfas-2018-0396.

Brinkhof:2019:EQI

[1815] Jesse Brinkhof, Bent Herrmann, Roger B. Larsen, and Tiago Veiga-Malta. Effect of a quality-improving cod end on size selectivity and catch patterns of cod in bottom trawl fishery. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 76(11):2110–2120, November 2019. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2018-0402.

Pennino:2019:MDT

[1816] Maria Grazia Pennino, Elena Guijarro-García, Raul Vilela, José Luis del Río, and Jose Maria Bellido. Modeling the distribution of thorny skate (Amblyraja radiata) in the southern Grand Banks (Newfoundland, Canada). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 76(11):2121–2130, November 2019. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
cjfas-2018-0302.

Feiner:2019:IFB

[1817] Zachary S. Feiner, Stephanie L. Shaw, and Greg G. Sass. Influences of female body condition on recruitment success of walleye (Sander vitreus) in Wisconsin lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 76(11): 2131–2144, November 2019. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2018-0364.

Pregler:2019:SSA

[1818] Kasey C. Pregler, R. Daniel Hanks, Evan S. Childress, Nathaniel P. Hitt, Daniel J. Hocking, Benjamin H. Letcher, Tyler Wagner, and Yoichiro Kanno. State-space analysis of power to detect regional brook trout population trends over time. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 76(11): 2145–2155, November 2019. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2018-0241.

Brookson:2019:MDN

[1819] Cole B. Brookson, Shane R. de Solla, Kim J. Fernie, Monina Cepeda, and Chelsea M. Rochman. Microplastics in the diet of nestling doublecrested cormorants (Phalacrocorax auritus), an obligate piscivore in a freshwater ecosystem. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 76(11): 2156–2163, November 2019. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2018-0388.

Anonymous:2019:NAN

[1820] Anonymous. Note of appreciation / note de reconnaissance. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 76(12), December 2019. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2019-0321.

Lescord:2019:MBR

[1821] Gretchen L. Lescord, Thomas Johnston, Brian A. Branfireun, and John M. Gunn. Mercury bioaccumulation in relation to changing physicochemical and ecological factors across a large and undisturbed boreal watershed. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 76(12):2165–2175, December 2019. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
cjfas-2018-0465.

Lennox:2019:BCB

[1822] Robert J. Lennox, Jacqueline M. Chapman, William M. Twardek, Franziska Broell, Kristin Bøe, Frederick G. Whoriskey, Ian A. Fleming, Martha Robertson, and Steven J. Cooke. Biologging in combination with biotelemetry reveals behavior of Atlantic salmon following exposure to capture and handling stressors. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 76(12):2176–2183, December 2019. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/cjfas-2018-0477.

Rooper:2019:FAC

[1823] Christopher N. Rooper, Pam Goddard, and Rachel Wilborn. Are fish associations with corals and sponges more than an affinity to structure? Evidence across two widely divergent ecosystems. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 76(12):2184–2198, December 2019. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2018-0264.

Li:2019:MPD

[1824] Yan Li, Laura M. Lee, and Jason Rock. Modeling population dynamics and nonstationary processes of difficult-to-age fishery species with a hierarchical Bayesian two-stage model. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 76(12):2199–2214, December 2019. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/cjfas-2018-0298.

Rose:2019:SOP

[1825] Craig S. Rose, Julie K. Nielsen, John R. Gauvin, Timothy Loher, Suresh A. Sethi, Andrew C. Seitz, Michael B. Courtney, and Paige Drobny. Survival outcome patterns revealed by deploying advanced tags in quantity: Pacific halibut (Hippoglossus stenolepis) survivals after release from trawl catches through expedited sorting. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 76(12):2215–2224, December 2019. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2018-0350.

Macnaughton:2019:UMA

[1826] Camille J. Macnaughton, David Deslauriers, Erinn L. Ipsen, Emily Corey, and Eva C. Enders. Using meta-analysis to derive a respiration model for Atlantic salmon (Salmo salar) to assess bioenergetics requirements of juveniles in two Canadian rivers. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 76(12):2225–2234, December 2019. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2018-0436.

Bayse:2019:HLC

[1827] Shannon M. Bayse, Stephen D. McCormick, and Theodore Castro-Santos. How lipid content and temperature affect American shad (Alosa sapidissima) attempt rate and sprint swimming: implications for overcoming migration barriers. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 76 (12):2235–2244, December 2019. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/

doi/full/10.1139/cjfas-2018-0406.

Nesslage:2019:PES

[1828] Geneviève M. Nesslage and Michael J. Wilberg. A performance evaluation of surplus production models with time-varying intrinsic growth in dynamic ecosystems. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 76(12): 2245–2255, December 2019. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2018-0292.

Song:2019:LRG

[1829] Yiqing Song, Fei Cheng, Peng Ren, Zhen Wang, and Songguang Xie. Longitudinal recovery gradients of drifting larval fish assemblages in the middle reach of the Yangtze River: impact of the Three Gorges Dam and conservation implementation. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 76(12):2256–2267, December 2019. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/cjfas-2018-0409.

Barth:2019:CZC

[1830] Lauren Emily Barth, Brian John Shuter, William Gary Sprules, Charles Kenneth Minns, and James Anthony Rusak. Calibration of the zooplankton community size spectrum as an indicator of change in Canadian Shield lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 76(12): 2268–2287, December 2019. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2018-0371.

Ruzzante:2019:LCL

[1831] Daniel E. Ruzzante, Gregory R. McCracken, Sarah J. Salisbury, Hilary T. Brewis, Donald Keefe, Oscar E. Gaggiotti, and Robert Perry. Landscape, colonization, and life history: their effects on genetic diversity in four sympatric species inhabiting a dendritic system. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 76(12):2288–2302, December 2019. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2018-0416.

Barouillet:2019:IGF

[1832] Cécilia Barouillet, Brian F. Cumming, Kathleen R. Laird, Christopher J. Perrin, and Daniel T. Selbie. Influence of glacial flour on the primary and secondary production of sockeye salmon nursery lakes: a comparative modern and paleolimnological study. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 76(12):2303–2314, December 2019. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/cjfas-2018-0372.

Nelligan:2019:LTT

[1833] Clare Nelligan, Adam Jeziorski, Kathleen M. Rühland, Andrew M. Paterson, and John P. Smol. Long-term trends in hypolimnetic volumes and dissolved oxygen concentrations in Boreal Shield lakes of southcentral Ontario, Canada. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 76(12): 2315–2325, December 2019. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2018-0278.

Wolf:2019:ARB

[1834] Nathan Wolf, Sarah R. Webster, Jeffery M. Welker, and Bradley P. Harris. Assessing the relationship between diet and size-at-age in Pacific halibut (Hippoglossus stenolepis) using δ ^{13}C and δ ^{15}N analysis. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 76(12):2326–2331, December 2019. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2018-0380.

Aljafary:2019:CPN

[1835] M. Aljafary, D. M. Gillis, and P. Comeau. Is catch proportional to nominal effort? Conceptual, fleet dynamic, and statistical considerations in catch standardization. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 76(12): 2332–2342, December 2019. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2018-0303.

Chalde:2019:EWD

[1836] Tomás Chalde, Cristina F. Nardi, and Daniel A. Fernández. Early warning: detection of exotic Coho salmon (Oncorhynchus kisutch) by environmental DNA and evidence of establishment at the extreme south of Patagonia. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 76(12):2343–2349, December 2019. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
cjfas-2018-0271.

Bevelhimer:2019:MIA

[1837] Mark S. Bevelhimer, Brenda M. Pracheil, Allison M. Fortner, Ryan Saylor, and Kendra L. Deck. Mortality and injury assessment for three species of fish exposed to simulated turbine blade strike. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 76(12):2350–2363, December 2019. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2018-0386.

Boe:2019:ITL

[1838] Kristin Bøe, Michael Power, Martha J. Robertson, Corey J. Morris, J. Brian Dempson, Curtis J. Pennell, and Ian A. Fleming. The influence of temperature and life stage in shaping migratory patterns during the early marine phase of two Newfoundland (Canada) Atlantic salmon (Salmo salar) populations. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 76 (12):2364–2376, December 2019. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/

doi/full/10.1139/cjfas-2018-0320.

Kuriyama:2019:CSD

[1839] Peter T. Kuriyama, Daniel S. Holland, Lewis A. K. Barnett, Trevor A. Branch, Robert L. Hicks, and Kurt E. Schnier. Catch shares drive fleet consolidation and increased targeting but not spatial effort concentration nor changes in location choice in a multispecies trawl fishery. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 76(12):2377–2389, December 2019. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2019-0005.

Smukall:2019:ESC

[1840] Matthew J. Smukall, Amy Shaw, and Donald C. Behringer. Effect of simulated catch-and-release angling on postrelease mortality and egg viability in sockeye salmon (Oncorhynchus nerka). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 76(12):2390–2395, December 2019. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2018-0426.

Bortolotti:2019:DEM

[1841] Lauren E. Bortolotti, Vincent L. St. Louis, and Rolf D. Vinebrooke. Drivers of ecosystem metabolism in restored and natural prairie wetlands. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 76(12):2396–2407, December 2019. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
cjfas-2018-0419.

Jensen:2019:CAM

[1842] Arne Johan Jensen, Bengt Finstad, and Peder Fiske. The cost of anadromy: marine and freshwater mortality rates in anadromous Arctic char and brown trout in the Arctic region of Norway. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 76(12):2408–2417, December 2019. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2018-0428.

Turley:2019:IWE

[1843] Brendan D. Turley and Ryan R. Rykaczewski. Influence of wind events on larval fish mortality rates in the southern California Current Ecosystem. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 76(12):2418–2432, December 2019. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
cjfas-2018-0458.